### v3

# Consumption Partial Insurance

Abstract
We measure the extent of consumption insurance to income shocks accounting for high-order moments of the income distribution. We derive a nonlinear consumption function, in which the extent of insurance varies with the sign and magnitude of income shocks. Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption compared to 9% of an equal-sized positive shock – and the pass-through increases as the shock worsens. We find similar patterns for transitory shocks among the least wealthy. Households are willing to sacrifice more than 1/8 of lifetime consumption to eliminate tail income risk. Our results are consistent with surveys of consumption responses to hypothetical events and suggest that tail risk matters substantially for consumption.
Keywords: Income risk, skewness, kurtosis, consumption, partial insurance, PSID
JEL Classification: D12, D15, D31, E21
1 Introduction
How does consumption respond to income shocks? The answer to this question is important for determining the welfare effects of shifts in the income distribution, i.e., the link between income and consumption inequality, for understanding how families cope after adverse events, for designing government insurance, and for the dynamics of business cycles. A consistent empirical finding is that consumption is partially insured against income shocks, so income fluctuations translate into consumption fluctuations less than one-to-one (e.g., Blundell et al., 2008; Heathcote et al., 2014). [^1] The study of consumption insurance dates back to tests of the permanent income (Deaton and Paxson, 1994) and complete markets hypotheses (Cochrane, 1991; Attanasio and Davis, 1996; Hayashi et al., 1996; Brav et al., 2002). Jappelli and Pistaferri (2010) and Crawley and Theloudis (2024) review this literature. With a few exceptions discussed subsequently, most measures of consumption insurance do not distinguish between good and bad shocks, or between small and large ones. However, recent evidence shows that the distribution of income risk has a long and fat left tail; that is, there is an asymmetrically large probability that households face some really bad income shocks. This paper measures the consumption response to income shocks precisely when the distribution of said shocks may exhibit long and fat tails.
In a seminal paper, Blundell et al. (2008) –henceforth, BPP– introduce an empirical methodology to measure the pass-through of income shocks of varying persistence into consumption. Using Panel Study of Income Dynamics (PSID) and Consumer Expenditure Survey (CEX) data over 1980–1992, BPP find evidence of partial insurance to permanent shocks and full insurance to transitory ones. BPP rely on a workhorse covariance between income and consumption growth, namely a second moment of their joint distribution. The idea is that the extent to which income growth (driven by income shocks) varies with consumption growth reflects the strength of the transmission of income shocks into consumption.
In a separate strand of literature, Guvenen et al. (2021) use data from the U.S. Social Security Administration and establish two new important facts about the distribution of unexplained income growth, namely the distribution of income shocks. First, income growth is very negatively skewed, i.e., it has a long left tail, so there are far more people in the data who experience large negative than large positive shocks. Second, income growth exhibits excess kurtosis, i.e., fat tails, so there are far more people in the data who face either small or large income shocks, than people who face moderate ones, relative to a Gaussian density. These are now benchmark facts about income dynamics, observed in many countries (e.g., De Nardi et al., 2021, compare the U.S. and the Netherlands).
If most shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks. In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution, respectively, may matter for the pass-through of shocks into consumption. This is precisely what we investigate in this paper. Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks). This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution. The second moment is a poor summary of the tails and, thus, abstracts from the modern higher-order facts that characterize income dynamics.
We offer four main contributions:
First, we re-estimate BPP using recent PSID data over 1999–2019, targeting the original income and consumption second moments. However, while BPP imputed consumption from the CEX (the PSID previously lacked comprehensive consumption data), such data are now internally available in the PSID. We subsequently use this data to study tail income risk, so it is natural to start with a ‘replication’ of BPP using the newer data. This ‘replication’ is insightful in its own right not only because of the different time period, but also because we no longer need to rely on imputed consumption from the CEX.
Second, we introduce into BPP the third and fourth moments of the joint distribution of income and consumption, in addition to the second moments that BPP originally use. This is our first attempt to go beyond the workhorse income-consumption covariance and measure partial insurance targeting additional information in the tails of the distribution.
Third, we generalize the method of measuring consumption insurance to explicitly permit a role for tail income risk. BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks. Our first and second contributions above revolve firmly around such a linear framework. We argue (as also does Carroll, 2001) that linearity is a poor approximation to behavior when income is subject to tail shocks. We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks. We characterize and empirically measure the consumption pass-through of income shocks; in this environment, the degree of partial insurance depends on the magnitude and sign of shocks. While a nonlinear consumption function has been estimated before, most notably in Arellano et al. (2017), identification in our case is far less data demanding relative to them: we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution. This has certain important advantages, which we discuss below.
Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk. De Nardi et al. (2019), Busch and Ludwig (2021), and Guvenen et al. (2024) provide similar calculations through structural models that fully specify the insurance channels at the nexus of income and consumption. By contrast, we remain agnostic about specific insurance mechanisms; we estimate an ‘aggregate’ of all partial insurance available to households, and then use a parametric utility function to estimate the welfare costs of tail risk.
Our baseline empirical implementation uses income and consumption panel data from the PSID over the years 1999–2019. We focus on a representative sample of stable married couples that is similar to the sample in BPP. We show that unexplained income growth in this sample features large negative skewness and excess kurtosis, qualitatively and numerically close to the skewness and kurtosis in the administrative data of Guvenen et al. (2021).
We then obtain five main findings:
First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively. Left skewness implies that a negative shock is typically more unsettling, more hurtful than a positive shock, because it is further away from the zero mean. Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders. Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad.
Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP. A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP. In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using the recent PSID data but a different method, Arellano et al. (2023) also find a low pass-through of permanent shocks. We discuss this point further when we present our empirical results. We attribute one third of the difference from BPP to the consumption imputation from the CEX that BPP use. The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth. Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute higher frequency shocks and their transmission into consumption. The last third reflects the different calendar times.
Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance. This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare implications of tail risk (e.g., Barro, 2009; De Nardi et al., 2019). One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find. This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of bad versus good shocks, and of small versus large ones. In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters. We draw three main conclusions. First, bad permanent shocks are more hurtful than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$. Second, the pass-through of bad shocks increases with the severity of the shock. Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also have increasingly larger transmission rates as their magnitude worsens. Third, transitory shocks provide mixed evidence. They are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021). Among low wealth households, however, transitory shocks shift consumption significantly, and negative shocks pass through at increasingly larger rates than positive ones. Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one. As there is clearly a limit to how much one can dissave or borrow, a large negative shock ends up affecting consumption by relatively more.
The asymmetric transmission of bad versus good permanent shocks intensifies with the age of the household. While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks. Severely bad permanent shocks thus cause more harm among older than younger households. Guvenen et al. (2021) show that income skewness becomes more negative with the age of the earner – there is more room for income to fall as older workers have higher incomes; we find a similar pattern in our data too. In turn, we document that there is also more room for consumption to fall, so bad shocks are double hurtful among the old. Overall, the results from our quadratic specification are indicative of limited insurance to left tail income risk. Moreover, if the econometrician assumed a linear consumption function, she would overestimate the true extent of insurance to it, i.e., she would infer excessive insurance.
[…]
There is a large literature in macro and labor economics that studies the transmission of income shocks into consumption, i.e., the link between income and consumption inequality. There are two distinct strands in this literature: a quasi-reduced form branch that measures the degree of partial insurance in the data without fully specifying the underlying insurance mechanisms, and a structural branch that fully models specific insurance channels.
The seminal paper in the first branch, namely BPP, sparked a large volume of research. [^3] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical distribution of consumption and income, so they are early precursors in this literature. Blundell et al. (2016) extend BPP to an environment with wage risk to measure the contribution of family labor supply to partial insurance. Theloudis (2021) extends this further to allow for household heterogeneity, which he estimates using information in the third moments. Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find. Hryshko and Manovskii (2022) carry out an interesting subsample analysis, identifying sets of households in the PSID with different degrees of insurance based on certain features of their incomes. Crawley and Kuchler (2023) offer another empirical extension, addressing neglected time aggregation in income and consumption in the PSID.
The previous papers restrict income shocks to transmit linearly into consumption. Arellano et al. (2017); Arellano et al. (2023) generalize the income and consumption process by letting permanent shocks feature nonlinear persistence and nonlinear transmission into consumption. [^4] The latter paper advances the former by allowing for flexible heterogeneity and unbalanced panels. Using recent PSID data, they measure the consumption response to income at about 0.3 on average (parameter closest to our definition of pass-through; specification with heterogeneity in the second paper), and the response varies over the distribution of shocks with evidence of asymmetries similar to ours. Identification is quite demanding as they require observing the entire consumption-income joint distribution; our case, by contrast, is far simpler: we need second moments of consumption and up to fourth moments of income. This is appealing in itself (we use more accessible information) and because, as we show later, higher-order features of the consumption distribution exhibit notable discrepancies across survey data, calling into question the conclusions drawn from them. The price of our econometrically more accessible approach is that we have to maintain a linear income process (though one that is subject to tail shocks), while Arellano et al. (2017) do not. Yet, we learn that nonlinearity in income is not crucial for obtaining asymmetric responses to income shocks.
The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models. [^5] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand. The early works focus on self-insurance, providing mixed evidence for how much insurance the model permits. Kaplan and Violante (2010) find less insurance compared to BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance. More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durable purchases (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021). Except Madera (2019), these papers abstract from higher income moments due to the computational difficulties involved. De Nardi et al. (2019) embed a non-Gaussian income process into a model of consumption/savings and show that tail income risk increases partial insurance due to stronger precautionary motives. In a similar model, Busch and Ludwig (2021) target income skewness and kurtosis, they distinguish between good and bad shocks, and they show that the latter are worse insured. Guvenen et al. (2024) find that tail shocks are worse insured than average shocks, but the benchmark methods of measuring this, effectively BPP, overstate the true extent of insurance. Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; this enables us to measure the overall degree of insurance from all mechanisms together keeping our model partially unspecified, while accounting for higher-order features of the income distribution.
The asymmetric response to bad and good shocks is recently also found in surveys that elicit responses to hypothetical events. Bunn et al. (2018) find that the marginal propensity to consume from bad shocks is larger than from good shocks; Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are larger than to gains, and the response increases with the severity of the loss. We document that these patterns exist also in survey data, where, unlike elicitation surveys, income shocks are realized but unobserved. [^6] Given our finding of heterogeneous responses to shocks, our paper is also broadly related to the literature that studies MPC heterogeneity (e.g., Lewis et al., 2022; Crawley and Kuchler, 2023).
The paper is also related to the newer income dynamics literature that examines the non-Gaussian features of the income distribution. Geweke and Keane (2000) and Bonhomme and Robin (2010) are early works that use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth. Guvenen et al. (2021) use data from the population of working males in the U.S. to provide further evidence of left skewness and excess kurtosis in earnings growth, both accentuated with age and the level of earnings. Guvenen et al. (2014) and Busch et al. (2018) study the cyclical nature of skewness in several countries. While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption. We thus contribute to a growing literature on the various economics/finance implications of tail risk, e.g. for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
In the remainder of the paper, section 2 presents the model while section 3 establishes identification of the parameters. Section 4 discusses the empirical implementation, section 5 presents the empirical results, and section 6 discusses welfare. Section 7 concludes. The appendix provides theoretical derivations and additional results.
2 Lifecycle model with tail income risk
The framework through which we study the extent of partial insurance to income shocks is a consumption lifecycle model with idiosyncratic income risk.
Household $i$ chooses lifetime consumption $\{C_{it}\}_{t=0}^{T}$ and savings $\{A_{it+1}\}_{t=0}^{T}$ to maximize its expected discounted utility over its lifecycle. The problem is formulated as
$\max_{\{C_{it},A_{it+1}\}_{t=0}^{T}}{\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\beta^{t}U(C_{it};\mathbf{Z}_{it}),$ (1)
subject to the lifetime budget constraint
$A_{i0}+{\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{Y_{it}}{(1+r)^{t}}={\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{C_{it}}{(1+r)^{t}},$ (2)
where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y$, and $T$ is the end of the horizon. We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.
2.1 Income process
Income is the only source of (idiosyncratic) uncertainty that the household faces. Its logarithm is given by
$\ln Y_{it}=\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}+P_{it}+v_{it},$
where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race); $\boldsymbol{\delta}$ is the loading factor of those variables on the logarithm of income. The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$. We assume that permanent income has a unit root, $P_{it}=P_{it-1}+\zeta_{it}$, where $\zeta_{it}$ is the permanent income shock. These modeling choices give rise to the familiar permanent-transitory formulation
$\Delta y_{it}=\zeta_{it}+\Delta v_{it},$ (3)
where $\Delta y_{it}=\Delta\ln Y_{it}-\Delta\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}$ is income growth net of the deterministic profile, i.e., unexplained income growth, and $\Delta$ is the first difference operator. [^7] See Meghir and Pistaferri (2011) for a comprehensive review of modern income processes.
We let the permanent and transitory income shocks, $\zeta_{it}$ and $v_{it}$, be non-Gaussian, with first four central moments given by
${\operatorname*{\mathbb{E}}}(\zeta_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{\zeta_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{\zeta_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{\zeta_{t}}&\text{for }m=4\\[-2.0pt] \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}(v_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{v_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{v_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{v_{t}}&\text{for }m=4.\\[-2.0pt] \end{cases}\vspace{-1ex}$
$\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction. $\gamma$ denotes the skewness/asymmetry of the distributions, i.e., the relative lengths of the upper and lower tails. Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distributions to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al. (2021) provide some evidence for that. Otherwise, shocks across households are identically distributed within $t$. Permanent and transitory shocks are mutually independent and independent over time.
As discussed in the introduction, several recent studies document that the cross-sectional distribution of income growth exhibits left skewness and excess kurtosis, i.e., a long and fat left tail. Our specification of the moments of income shocks enables our income process to be consistent with these facts, allowing us to subsequently decompose its second- and higher-order moments into the relative contributions of permanent and transitory shocks.
2.2 Linear consumption function
In line with BPP, we first consider a linear specification of the household consumption function, given by
$\Delta c_{it}=\xi_{it}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$ (4)
where $\Delta c_{it}$ is the residual from a regression of real consumption growth, $\Delta\ln C_{it}$, on the taste observables $\mathbf{Z}_{it}$ (i.e., consumption growth net of its predictable profile). The linear function (4) stems from a log-linearization (i.e., a first-order Taylor series expansion) of the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These approximations are independent of the distributions of income shocks and mimic similar derivations in BPP (who abstract from higher-order moments).
Permanent and transitory income shocks drive unexplained consumption growth in (4). The transmission parameters $\phi_{it}^{(1)}=\partial\Delta c_{it}/\partial\zeta_{it}$ and $\psi_{it}^{(1)}=\partial\Delta c_{it}/\partial v_{it}$ reflect the pass-through of permanent and transitory shocks, respectively, into consumption. Adopting BPP’s terminology, $1-\phi_{it}^{(1)}$ is then the extent of insurance to permanent shocks (i.e., the fraction of a permanent shock that does not translate into a consumption fluctuation), while $1-\psi_{it}^{(1)}$ is the extent of insurance to transitory shocks. Therefore, $\phi_{it}^{(1)}=\psi_{it}^{(1)}=0$ implies that consumption is fully insured to income shocks, while $\phi_{it}^{(1)}=\psi_{it}^{(1)}=1$ implies no insurance. By contrast, values $\phi_{it}^{(1)},\psi_{it}^{(1)}\in(0,1)$ reflect partial insurance: a given income shock translates into a consumption fluctuation, albeit less than one-to-one.
The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things. The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure. [^8] In addition, households in Blundell et al. (2016) have access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences; both are reasons why the transmission parameters may vary with $i$ or $t$. We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A. So while the pass-through of shocks depends on wealth, it does not depend on the magnitude or sign of the shock itself. Function (4) thus implies that a household can insure big or small, good or bad shocks similarly, a point to which we return subsequently.
In addition to income shocks, unexplained consumption growth is also driven by $\xi_{it}$. This reflects unobserved (but non-stochastic) consumption taste heterogeneity across households, independent of the income shocks. We let the distribution of $\xi_{it}$ across households have zero mean and variance $\sigma_{\xi_{t}}^{2}$, as in BPP. We also let its third and fourth central moments be $\gamma_{\xi_{t}}$ and $\kappa_{\xi_{t}}$ respectively; these will appear subsequently in expressions for the skewness and kurtosis of consumption growth.
2.3 Quadratic consumption function
The linear consumption function offers an approximation to the true consumption response to income shocks when such shocks do not move the household too far away from its deterministic consumption profile. Function (4) relies on a linearization of marginal utility (the problem’s first-order conditions) around that deterministic profile; but such a linearization will be inaccurate in the presence of tail income risk, e.g., when extreme shocks may induce large shifts in consumption as a consequence of concave utility. [^9] This echoes Carroll (2001)’s critique of the log-linearized Euler equation. Kaplan and Violante (2010) and Blundell et al. (2013) assess the approximations in BPP across alternative scenarios, though not in one with respect to the severity of income shocks.
To address this issue, we generalize the consumption function to a nonlinear (quadratic) specification, given by
$\Delta c_{it}=\xi_{it}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\zeta_{it}^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\zeta_{it}v_{it},$ (5)
where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously. The quadratic function (5) stems from a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A. These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter.
[…]
In this generalized framework, $\phi_{it}^{(1)}$ no longer represents the pass-through of permanent shocks to consumption. Instead, the pass-through is given by
$\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}=\phi_{it}^{(1)}+2\phi_{it}^{(2)}\zeta_{it}+\omega_{it}^{(22)}v_{it},$ (6)
which depends on multiple parameters and on the magnitude and sign of the income shocks. As such, there is an asymmetric pass-through of good versus bad shocks into consumption. Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest. At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a higher pass-through to consumption relative to a positive shock of the same magnitude. The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response. The size of this bias depends on the magnitude of $\phi^{(2)}$ (a more negative $\phi^{(2)}$ implies bigger bias) and on the size of the permanent shock itself (a more severe bad shock implies bigger bias). When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
A similar asymmetry exists in the transmission of good versus bad transitory income shocks. In this case, the pass-through into consumption is given by
$\frac{\partial\Delta c_{it}}{\partial v_{it}}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
which again depends on multiple parameters and on the magnitude and sign of the income shocks. Importantly, this generalized pass-through does not rule out the possibility of full insurance to transitory shocks, a recurrent empirical finding in the literature. Suppose, for example, that $\psi^{(1)}\approx\psi^{(2)}\approx 0$; then $\partial\Delta c_{it}/\partial v_{it}\approx 0$ at $\zeta_{it}=0$ (the average permanent shock), indicating full insurance to transitory shocks.
3 Identification
Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters. The set of parameters includes the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^11] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments. We will be clear whenever those terms refer to standardized moments. the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$). We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups.
For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth. In the case of the quadratic consumption function, we also assume that income shocks are not Bernoulli. For consistent estimation of such moments, particularly over demographic groups subsequently, we further assume non-vanishing shares in each group asymptotically and knowledge of group membership. To carry out statistical inference, we finally assume existence of up to eighth moments.
Two important aspects of the data must be addressed before proceeding to identification: the biennial nature of the modern PSID and measurement error.
As we describe in section 4, the PSID provides data only every two years within our time period. This necessitates to recast the income and consumption growth processes of section 2 to this lower frequency, a straightforward operation that we carry out in appendix B. Nevertheless, for consistency with the earlier literature, we provide the identifying statements subsequently as if the data come at a yearly frequency. The complete identifying statements at the lower frequency of the modern PSID appear in appendix B.
Measurement error in income and consumption is more challenging. In the permanent-transitory specification of income, the moments of classical income error are not separately identifiable from the moments of the transitory shock. This necessitates that we restrict the income error to be Gaussian and retrieve its variance from validation studies of the PSID, as in Meghir and Pistaferri (2004). We return to this in section 4. By contrast, the moments of classical consumption error are identifiable; this introduces three additional parameters for the variance, skewness, and kurtosis of consumption error ($\sigma_{u^{c}_{t}}^{2}$, $\gamma_{u^{c}_{t}}$, $\kappa_{u^{c}_{t}}$). For simplicity, however, we subsequently show identification as if income and consumption errors are not present; we provide our full identifying statements with measurement error in appendix B.
3.1 Income process parameters
Given the income process in (3) and the properties of shocks, the second and higher-order moments of the cross-sectional distribution of income growth are given by
${\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2})$ $=\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2},$
${\operatorname*{\mathbb{E}}}((\Delta y_{it})^{3})$ $=\gamma_{\zeta_{t}}+\gamma_{v_{t}}-\gamma_{v_{t-1}},$
${\operatorname*{\mathbb{E}}}((\Delta y_{it})^{4})$ $=\kappa_{\zeta_{t}}+\kappa_{v_{t}}+\kappa_{v_{t-1}}+6\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})+6\sigma_{v_{t}}^{2}\sigma_{v_{t-1}}^{2}.$
The second moment depends on the variance of permanent shocks at $t$ and the sum of variances of transitory shocks at $t$ and $t-1$; it thus reflects the dispersion of one permanent and two transitory shocks. The fourth moment reflects, similarly, the dispersion and kurtosis of both types of shocks. By contrast, the third moment is driven by skewness in the permanent shock. Barring strong non stationarities in the distribution of transitory shocks, the difference between $\gamma_{v_{t}}$ and $\gamma_{v_{t-1}}$ will be small (zero if the distribution is stationary), leaving $\gamma_{\zeta_{t}}$, the skewness in permanent shocks, to drive entirely the third moment of income.
The above expressions convey information on the parameters of the income process. Formally, the variance of shocks is identified as
$\sigma_{\zeta_{t}}^{2}$ $={\operatorname*{\mathbb{E}}}(\Delta y_{it}\times(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1})),$
$\sigma_{v_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta y_{it}\times\Delta y_{it+1}).$
The long sum in the first line strips income growth at $t$ of its contemporary transitory shock; its covariance with $\Delta y_{it}$ thus picks up the variance of the permanent shock. The covariance of two consecutive income growths in the second line identifies the variance of the transitory shocks due to mean reversion in the shock.
In a similar way, the skewness of shocks is identified as
$\gamma_{\zeta_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1})),$
$\gamma_{v_{t}}$ $=-{\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times\Delta y_{it+1}),$
while, conditional on identification of the second moments, kurtosis is jointly identified from
$\kappa_{\zeta_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{4})-\kappa_{v_{t}}-\kappa_{v_{t-1}}-6\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})-6\sigma_{v_{t}}^{2}\sigma_{v_{t-1}}^{2},$
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}$
$-\sigma_{v_{t-1}}^{2}\sigma_{v_{t}}^{2}-\sigma_{v_{t-1}}^{2}\sigma_{v_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
There are several over-identifying restrictions for both skewness and kurtosis.
3.2 Linear consumption function parameters
Given the linear consumption function in (4) and the properties of shocks, the second and higher-order moments of the distribution of consumption growth are given by
${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$ $=(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}+(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}+\sigma_{\xi_{t}}^{2},$
${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{3})$ $=(\phi_{t}^{(1)})^{3}\gamma_{\zeta_{t}}+(\psi_{t}^{(1)})^{3}\gamma_{v_{t}}+\gamma_{\xi_{t}},$
${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{4})$ $=(\phi_{t}^{(1)})^{4}\kappa_{\zeta_{t}}+(\psi_{t}^{(1)})^{4}\kappa_{v_{t}}+\kappa_{\xi_{t}}+6(\phi_{t}^{(1)})^{2}(\psi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}$
$+6(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}\sigma_{\xi_{t}}^{2}+6(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}\sigma_{\xi_{t}}^{2}.$
The moments of the distribution of consumption growth reflect the underlying distributions of income shocks, as well as the distribution of unobserved taste heterogeneity. In general, larger dispersion, skewness, or kurtosis in the distributions of income shocks implies larger dispersion, skewness, or kurtosis in the distribution of consumption growth. The link between the two depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B. Formally, the transmission parameter of permanent shocks is identified as
$\phi^{(1)}_{t}={\operatorname*{\mathbb{E}}}(\Delta c_{it}\times(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2},$
while the transmission parameter of transitory shocks is identified as
$\psi^{(1)}_{t}=-{\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2};$
in both cases, the variance of shocks is given by the earlier expressions. [^12] The variance of consumption taste heterogeneity is further given by $\sigma_{\xi_{t}}^{2}={\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})-(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}-(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}$. Its skewness and kurtosis are similarly identified from the higher moment of consumption growth. These statements reflect the intuition behind identification in BPP: the covariance between consumption and (appropriately defined) income growth captures the strength of the relation between the two series and, consequently, the pass-through of income shocks into consumption. [^13] An equivalent view is that the partial insurance parameters are identified from a regression of consumption growth on income growth, using permanent or future income as instruments for income growth.
This workhorse covariance, however, is a second moment of the joint distribution of consumption and income and, as such, neglects information in the tails. If most income shocks are either zero or rather bad, identifying partial insurance based on second moments alone may be misleading. However, the partial insurance parameters are heavily over-identified by higher-order moments of the marginal and joint distributions of consumption and income; inspect, for example, the expressions for ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{3})$ or ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{4})$. We will thus subsequently exploit these higher-order moments in the context of the linear consumption function in our first attempt to understand the implications of non-Gaussianity for partial insurance.
3.3 Quadratic consumption function parameters
As per the quadratic specification in (5) and the properties of shocks, the second moment of the distribution of consumption growth is given by
${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$ $=(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}+(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}+2\phi_{t}^{(1)}\phi_{t}^{(2)}\gamma_{\zeta_{t}}+2\psi_{t}^{(1)}\psi_{t}^{(2)}\gamma_{v_{t}}+(\phi_{t}^{(2)})^{2}\kappa_{\zeta_{t}}+(\psi_{t}^{(2)})^{2}\kappa_{v_{t}}$
$+(\omega_{t}^{(22)})^{2}\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\phi_{t}^{(2)}\psi_{t}^{(2)}\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{\xi_{t}}^{2}.$
This depends not only on the variance of permanent and transitory income shocks, as in the linear case, but also on their skewness and kurtosis. Contrary to the linear case, however, the higher-order moments of consumption growth depend on higher than fourth-order moments of the shocks, which we do not model.
Identification of the partial insurance parameters comes from the joint moments of consumption and income, conceptually similar to identification in the linear case. Multiple moments contribute to the identification of any single parameter. Formally, the parameters are given by
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)\\ {\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)\\ {\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+1}\right)\\ {\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it+1})^{2}\right)\\ {\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+1}\right)\\ \end{array}\right),$ (8)
where square matrix $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^14] The matrix of coefficients is given by $\mathbf{A}=\left(\begin{array}[]{lllll}\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ \gamma_{\zeta_{t}}&\gamma_{v_{t}}&\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t-1}}^{2})&2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{v_{t+1}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(\sigma_{\zeta_{t+1}}^{2}+\sigma_{v_{t+1}}^{2})&0\\ 0&-\gamma_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}&-\kappa_{v_{t}}&-\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ , [^15] The variance of consumption taste heterogeneity is identified from ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$. Its skewness and kurtosis enter the third and fourth moments of consumption growth. Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting. We establish in appendix B that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway). [^16] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank. Assuming away Bernoulli distributions is thus an identifying assumption. The matrix is invertible even if the distributions of shocks are normal and have the same variance between them. In the nonlinear specification, tail shocks are associated with a potentially different pass-through from that of average shocks; but identification of such pass-through does not require those shocks to be non-Gaussian.
As seen in (8) and matrix $\mathbf{A}$, the transmission parameters can no longer be identified with second-order income moments alone. Higher-order moments of income play an essential role in identification, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails. By contrast, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth. This is an advantage over methods that require observing the full consumption distribution (e.g., Arellano et al., 2017, need the entire consumption-income joint distribution) because, while in principle we may estimate higher moments of consumption growth in the PSID, we cannot be certain how reliable these estimates are. This is in contrast to higher income moments, for which Guvenen et al. (2021) is a natural benchmark given their use of administrative data. We return to this point in section 4.
How biased are the parameter estimates if one assumes a linear specification when the true consumption function is quadratic? The workhorse covariance that identifies the pass-through of permanent shocks in BPP is
${\operatorname*{\mathbb{E}}}(\Delta c_{it}\times(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}=\phi_{t}^{(1)}+\phi_{t}^{(2)}(\gamma_{\zeta_{t}}/\sigma_{\zeta_{t}}^{2}).$
This fails to identify the true pass-through of permanent shocks in the presence of nonlinear transmission, i.e., expression (6). Moreover, this workhorse covariance is biased for $\phi_{t}^{(1)}$ if permanent shocks are non-Gaussian. If $\gamma_{\zeta_{t}}<0$, i.e., permanent shocks are left skewed as we find subsequently, and $\phi_{t}^{(2)}<0$, as our empirical estimates suggest, linearity biases $\phi_{t}^{(1)}$ upwards. Similarly, the identifying condition for $\psi_{t}^{(1)}$ in BPP becomes
$-{\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$
also failing to identify the true pass-through of transitory shocks in the presence of nonlinear transmission, i.e., expression (7). If $\gamma_{v_{t}}<0$ and $\psi_{t}^{(2)}>0$, as our baseline estimates suggest, linearity biases $\psi_{t}^{(1)}$ downwards. This helps explain the discrepancy between BPP (who find little pass-through of transitory shocks) and studies based on natural experiments (who find large consumption responses to said shocks). [^17] The discrepancy is between studies that use covariance restrictions in the joint distribution of income and consumption (they find small pass-through of transitory shocks) and studies that employ natural experiments (e.g., tax refunds or lotteries; they find large pass-through). Commault (2022) explains the discrepancy on the basis of neglected effects from past transitory shocks on consumption, while Crawley and Kuchler (2023) explain it on the basis of neglected time aggregation in the data. Both are alternative types of misspecification in the linear consumption function/empirical implementation vis-à-vis our focus on nonlinearities here.
4 Empirical implementation
4.1 Data
As is evident from the previous section, panel data on income and consumption are needed to identify the moments of the permanent and transitory shocks and the degrees of insurance to them. The Panel Study of Income Dynamics (PSID) provides such data for a representative and continuously evolving sample of U.S. households since $1968$. For most of this period, the PSID collected consumption expenditures only on food items (food representing a declining fraction of the consumer’s basket over time). Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016). We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when writing this paper; the data are available only biennially. This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$. We focus on a sample that is representative of the U.S. population, eliminating the supplementary low-income subsample (SEO). Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors. Naturally, we require non-missing data on income, expenditure, and basic demographics. [^18] We also remove observations with extreme jumps in income or consumption in consecutive years (jump: $\geq$10-fold increase, or $\geq$90reduction, or the bottom 0.25% of the distribution of products of growth rates between $t$ and $t-1$), i.e., large increases followed by large drops, likely reflecting measurement error. Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
Table 1: Descriptive statistics
Baseline sample Low wealth High wealth No college Some college
Mean Median Mean Mean Mean Mean
(1) (2) (3) (4) (5) (6)
Consumption
Nondurable+services 26,766 22,895 23,106 29,852 22,137 28,908
Food at home 8,256 7,609 7,861 8,580 7,462 8,630
Utilities 3,515 3,238 3,270 3,721 3,394 3,571
Food out 3,318 2,511 2,683 3,840 2,649 3,634
Health 1,614 863 1,390 1,803 1,429 1,699
Transportation 5,866 4,900 5,398 6,261 5,436 6,065
Education 3,202 0 1,529 4,614 1,262 4,100
Childcare 977 0 937 1,009 543 1,181
Rent (or rent eq.) 18,105 13,821 10,949 23,991 11,647 21,149
Home insurance 894 747 654 1,092 702 985
Income
Disposable income 142,910 113,581 98,527 179,416 97,675 164,239
Taxable+transfer inc. 137,277 109,003 93,358 173,401 90,810 159,187
Earnings 118,423 97,154 87,005 144,266 78,565 137,217
Male 81,857 62,185 56,817 102,452 50,199 96,783
Female 36,567 29,183 30,187 41,814 28,366 40,433
Wealth 547,761 191,193 69,731 950,823 284,036 669,788
Demographics
% working (male) 0.89 1.00 0.92 0.88 0.86 0.91
% working (female) 0.79 1.00 0.81 0.78 0.79 0.79
Age (male) 46 45 43 49 46 46
% some college (male) 0.68 1.00 0.55 0.79 0.00 1.00
% some college (female) 0.71 1.00 0.62 0.78 0.44 0.84
Periods per household 5.63 5.00 4.89 6.45 5.21 5.86
Observations 20,866 9,417 11,449 6,686 14,180
Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 19991999199919992019201920192019. Monetary amounts are expressed in \$2018. Rent for owners is calculated at 6% of the self-reported value of the house.
Table 1 presents summary statistics over the sample period 19991999199919992019201920192019. About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education. Our measure of income is household disposable income, namely the sum of taxable income (e.g., earnings and asset income), transfers, and social security income of the couple and of other family members, minus taxes paid. Average disposable income is \$142,910 in 2018 prices. [^19] We express all monetary figures in 2018 prices because earnings are 1-year retrospective and our latest data are from 2019. We deflate using the Consumer Price Index from the Bureau of Labor Statistics. Our measure of consumption is real expenditure on nondurable goods and services, comprising food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses (gasoline, parking, insurance), education, and child care. Average real consumption is \$26,766, with food at home accounting for about $30\%$ of this. Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it. Including housing does not change any of the subsequent results meaningfully. Finally, table 1 reports summary statistics also for samples sorted by wealth and education; we postpone the discussion of those samples to section 5.4.
Aside from the PSID, we use a similarly selected sample drawn from the CEX to facilitate comparison with BPP. We defer the discussion of this secondary data to section 4.3.
4.2 Estimation
Estimation follows a three-step approach. In the first step, we compute income and consumption growth net of their predictable components, respectively, $\mathbf{X}_{it}$ and $\mathbf{Z}_{it}$ in model notation, by regressing real income and consumption growth on a rich set of time effects and household characteristics. [^20] The controls include year and year-of-birth dummies, dummies for education, race, region, family size and number of children (level and change from last period), employment status (current and past), presence of outside dependents, and presence of income recipients other than the husband and wife (both current and past). The dummies for education, race, and region are interacted with the year dummies. The residuals from these regressions constitute the empirical counterparts to unexplained income and consumption growth, $\Delta y_{it}$ and $\Delta c_{it}$, respectively. In the second step, we estimate the parameters of the income process – i.e., the second and higher moments of permanent and transitory shocks – using income data alone. In the third step, we estimate the transmission parameters of shocks and the moments of taste heterogeneity and consumption measurement error; we do so using income and consumption data conditional on the parameters of the income process from the second stage.
We estimate the second- and third-step parameters using GMM. Given the potentially poor small sample properties of optimal weighting, we use the identity matrix as our baseline weighting choice. In practice, this matters little for the parameter estimates of either stage: using the optimal weighting matrix gives almost identical point estimates, suggesting the parameters are strongly identified. We report results from equal weighting subsequently, but we also present the main estimates using optimal GMM in appendix C.5.212121Diagonal weighting also delivers similar results (those results are omitted for brevity).
We present estimates of the transmission parameters separately for the linear and quadratic specifications of the consumption function. For the linear function, we first present results when only second moments of income and consumption growth are targeted, exactly as in BPP. This purports to replicate BPP over the more recent sample period. We next report results for the linear function when, in addition to the second moments, third and fourth moments are also targeted. This helps assess whether and how, through the lenses of the linear specification, information in the tails of the income distribution affects the measurement of partial insurance. We then move on to the results from the quadratic specification. Our choice to estimate the income process in an earlier step enables the income parameters to remain unchanged regardless of the specification of the consumption function fitted. Appendix B.4 lists the moments targeted in each case.
Given the multi-step approach, we conduct inference using the block bootstrap jointly over all stages. We thus account for serial correlation and heteroskedasticity of arbitrary form, as well as for our use of pre-estimated parameters in the later stages of the multi-step approach. To avoid the standard errors being affected by extreme draws, we apply a normal approximation to the inter-quartile range of bootstrap replications. [^22] Instead of estimating standard errors as the standard deviation of the parameters’ point estimates across replications, we compute them as the normalized interquartile range of the replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf. This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws. See Blundell et al. (2016) and Theloudis (2021) for details. Standard errors without the normal approximation are very similar.
As we explained in section 3, we must use an external estimate for the variance of income measurement error. We cannot otherwise separate the moments of the transitory shock from those of the measurement error. Following Meghir and Pistaferri (2004) and Blundell et al. (2016), we retrieve the error variance from Bound et al. (1994), a popular validation study of the PSID, setting it equal to $4\%$ of the variance of log income in any given year. [^23] Bound et al. (1994) survey workers from a single manufacturing firm in 1983 and 1987. They obtain information on earnings in a way that mimics the PSID questionnaire and coding practices and compare them to administrative data from the firm. There are two caveats in this approach: first, the sample of workers in the validation study comes from two decades prior to the data in this paper; second, we use estimates of error in male earnings to correct for error in household income.
4.3 Empirical moments of income and consumption
Table 2 reports the empirical second, third, and fourth moments of the cross-sectional distributions of unexplained income and consumption growth in the sample.
Income moments. Panel A presents the moments of income growth. The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant. [^24] In contrast to the identifying statements of section 3, the data are biennial so we report moments for income/consumption growth at a biennial frequency. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that this corresponds to a difference over two years. See appendix B for details. The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data. [^25] The second-order autocovariance is an order of magnitude smaller and only marginally significant. This feature and the biennial nature of the data, whereby the second-order autocovariance reflects correlations 4 calendar years apart, motivates our white noise specification for the transitory shock.
Table 2: Empirical moments of income and consumption growth
Baseline sample
Income data: PSID PSID
Consumption data: PSID imputed from CEX
(1) (2)
Panel A. Moments of income growth
${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.153 (0.005) 0.153 (0.005)
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.049 (0.002) -0.049 (0.002)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.143 (0.098) -0.143 (0.098)
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.009 (0.003) 0.009 (0.003)
${\operatorname*{\mathrm{Kurt}}}(\Delta y_{it})$ 10.032 (0.509) 10.032 (0.509)
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},(\Delta y_{it+1})^{2})$ 0.080 (0.007) 0.080 (0.007)
Panel B. Moments of consumption growth
${\operatorname*{\mathrm{Var}}}(\Delta c_{it})$ 0.126 (0.002) 0.323 (0.015)
${\operatorname*{\mathrm{Cov}}}(\Delta c_{it},\Delta c_{it+1})$0.044 (0.001) -0.140 (0.009)
${\operatorname*{\mathrm{Skew}}}(\Delta c_{it})$0.039 (0.034) -0.177 (0.083)
${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},\Delta c_{it+1})$0.003 (0.001) 0.190 (0.036)
${\operatorname*{\mathrm{Kurt}}}(\Delta c_{it})$ 4.445 (0.100) 17.109 (2.111)
${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},(\Delta c_{it+1})^{2})$ 0.027 (0.001) 0.909 (0.198)
Panel C. Joint moments
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta c_{it})$ 0.008 (0.001) 0.012 (0.002)
Notes: The table presents the second, third, and fourth moments of income and consumption growth in the baseline sample. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of consumption internally in the PSID; column 2 reports moments of consumption imputed from the CEX. See appendix C.3 for details on the imputation. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the data, this corresponds to a difference over two calendar years. Block bootstrap standard errors are in parentheses.
The third standardized moment of income growth is $-0.143$. Despite the small size of the cross-section and the inherent difficulties in estimating higher moments, the point estimate misses statistical significance at the 10% level only by a small margin. ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$, which reflects (minus) the third central moment of the transitory shock, is significant at the 5% level. The point estimates suggest that income growth exhibits left skewness, i.e., a long left tail. In simple terms, the left tail of the distribution of income growth (i.e., of income shocks) is longer than the right tail, so a negative shock is on average more severe, more unsettling, than a positive one. This is in line with Guvenen et al. (2021), who document strong left skewness in the distribution of male earnings growth using administrative data from the U.S. Social Security Administration.
The fourth standardized moment of income growth is $10.032$ and, as expected, highly significant. ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},(\Delta y_{it+1})^{2})$ is also very significant; this is larger than three times the square of the variance of transitory shocks, indicating excess kurtosis in those shocks. These estimates suggest that income growth is strongly leptokurtic. In simple terms, the distribution of income growth (i.e., of income shocks) has thicker tails and less mass in the shoulders relative to a Gaussian density. Shocks are thus either close to the center of the distribution, i.e., zero, or out in the tails. This is again in line with Guvenen et al. (2021), who document strong excess kurtosis in male earnings growth.
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate the comparison, appendix C.1 reports moments of male earnings growth in our sample. The variance is very similar between the PSID and the U.S. Social Security Administration data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) confirm this. Kurtosis, finally, is identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 2), but they are also observed in alternative measures of income, such as male earnings (table C.1, panel A) and household earnings (table C.1, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). These results corroborate (and, in fact, are very close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
Large income reductions do not just reflect unemployment, on which there is an abundant existing literature, but a multitude of events. Issues of reverse causality aside, appendix C.2 documents that households with large income reductions also experience the loss of a second job, worsening health, large drops in hours worked (possibly reflecting both demand and supply factors), and precarious housing. It is unclear how persistent those events are, but our subsequent decomposition of income to moments of permanent and transitory shocks will shed light on this.
Consumption moments. Panel B of table 2 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID. The variance of consumption growth is very similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error. The first-order autocovariance is $-0.044$ and highly significant. In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
[…]
The fourth standardized moment of consumption growth is $4.445$ and highly significant. There is excess kurtosis and the distribution is thus leptokurtic. In contrast to the distribution of income, however, excess kurtosis is less pronounced here.
The covariance between income and consumption growth is $0.008$. As we established in section 3, this moment is crucial for the identification of the pass-through of income shocks into consumption. While we cannot immediately gauge its magnitude, the strong statistical significance reflects that income shocks do transmit into consumption.
Unlike income, there are neither administrative data on consumption, nor has there been a validation study for consumption in the PSID. [^27] Andreski et al. (2014) and Li et al. (2010) compare consumption between the PSID and the CEX; Meyer and Sullivan (2023) offer a critical discussion of the CEX data against the PSID. As a consequence, we cannot benchmark our estimates of the consumption higher moments against the literature. The CEX, however, provides high quality consumption data for a representative sample of the U.S. population, although it lacks a long panel dimension. We can thus use the CEX to select a sample of households as close to our baseline as possible, and estimate similar cross-sectional moments across the surveys. The CEX is useful also for another reason. Since the PSID collected little consumption data prior to $1999$, the early literature on consumption insurance, notably BPP, resorted to imputing expenditure from the CEX into the PSID. To facilitate the comparison with BPP, we can use the CEX to replicate the consumption imputation into the PSID. This is not strictly needed in our case (the PSID provides comprehensive data over our sample period) but the imputation is nevertheless insightful, as we illustrate later.
Consumption in the CEX. We use CEX interview survey data over 1999-2019 and select a sample of stable married households that mimics closely our baseline selection in the PSID. We report details on the sample selection in appendix C.3.
Appendix table C.3 offers a comparison between the PSID and CEX samples over time. The two samples are very close with respect to demographics and labor market participation. Household income is higher in the PSID; BPP noted this and argued that it is due to the more comprehensive definition of income in the PSID. Food expenditure in the PSID is also higher, in contrast to BPP who find that food expenditure over the earlier period 1980–1992 is similar across the two surveys. Consequently, overall consumption in the PSID is, on average, higher by about 25% than a similarly selected sample in the CEX.
Appendix figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID and CEX samples. The patterns for the variance and kurtosis are quite similar in the two surveys over time, though the variance in the PSID is lower and the kurtosis is slightly higher. Skewness, however, is markedly different. Skewness of log consumption is strongly positive in the PSID, while it is mostly negative in the CEX; the patterns also diverge over time. Overall, these statistics suggest that consumption across otherwise similar samples in the PSID and CEX exhibits some notable differences over our sample period.
This comparison is not informative about the moments of consumption growth, which is the relevant variable in our case. The CEX lacks the panel dimension needed for our analysis, so it cannot be used to obtain measures of consumption growth. This is why BPP imputed consumption from the CEX into the PSID: the CEX lacked the necessary panel dimension while the PSID lacked comprehensive consumption data before 1999. We repeat BPP’s imputation over the recent years (our sample period 1999–2019). The imputation consists of the estimation of food demand in the CEX (food expenditure expressed as a function of prices and nondurable consumption). The demand equation is then inverted (nondurable consumption expressed as a function of prices and food) and used to impute nondurables in the PSID based on food expenditure and prices that are available in both surveys. In this way, we obtain an alternative, external measure of consumption in the PSID, and of consumption growth. Appendix C.3 provides details on the imputation.
The moments of imputed consumption growth are in column 2 of table 2. The variance at $0.323$ is two-and-a-half times higher than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The higher-order moments also differ markedly between the imputed and true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
Since the imputation is not strictly needed over our sample period, we use consumption internally available in the PSID as our baseline measure of consumption. The present analysis also suggests that we may not want to rely on higher-order moments of consumption, given the notable discrepancies in skewness of log consumption between the PSID and the CEX (this is independent of the imputation). As we established in section 3.3, an advantage of our identification approach over other methods is precisely its non-reliance on such higher-order consumption moments.
5 Results
We first present baseline results imposing stationarity in the distributions of income shocks, taste heterogeneity, consumption error, and a constant pass-through of income shocks over the lifecycle. [^28] Stationarity is not an identifying assumption but helps with precision, especially of higher moments. We subsequently relax stationarity/lifecycle-invariance in section 5.4.
5.1 Income process
We present estimates of the income process in three specifications. First, we target only second moments of income growth, which enables us to estimate the second moments of income shocks. This is similar to what BPP do over the earlier period 1980–1992. Second, we target second and third moments of income growth, which enables us to estimate also the skewness of income shocks. Finally, we target up to the fourth moments of income growth, which allows us to estimate all the parameters of the income process as presented in section 2.1. The results from each specification appear in table 3, columns 1-3, respectively. The full set of moments targeted in each case is shown in appendix B.4.1.
When only the second moments of income growth are targeted, we estimate the variance of permanent and transitory shocks at $0.027$ and $0.033$, respectively. These are very close to BPP and Meghir and Pistaferri (2004), who estimate these parameters over an earlier period. As the variance of income growth is driven by the variance of the contemporaneous permanent shock and the sum of the variances of transitory shocks in the current and previous periods, the contribution of the transitory shock to the variance of income growth is almost $2.5$-fold that of the permanent shock (as in Gottschalk and Moffitt, 2009).
Table 3: Estimates of the income process, baseline sample
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
Income data: PSID PSID PSID
(1) (2) (3)
Panel A. Permanent shocks
Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.030
(0.002) (0.002) (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 -0.004
(0.003) (0.003)
standardized -0.947 -0.831
(0.649) (0.571)
Kurtosis $\kappa_{\zeta}$ central 0.039
standardized 44.393
Panel B. Transitory shocks
Variance $\sigma_{v}^{2}$ 0.033 0.033 0.031
(0.002) (0.002) (0.003)
Skewness $\gamma_{v}$ central -0.008 -0.008
(0.003) (0.003)
standardized -1.310 -1.431
(0.510) (0.566)
Kurtosis $\kappa_{v}$ central 0.048
standardized 49.226
Notes: The table presents the estimates of the parameters of the income process, imposing stationarity over time. Column 1 targets only the second moments of income growth; column 2 targets also third moments; column 3 targets all moments up to fourth-order. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. Results from optimal GMM are similar (appendix table C.5).
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.310$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a much longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
Targeting all moments of income growth up to fourth-order leaves the third central moments unchanged and modifies the variances only by little. We estimate the kurtosis (fourth standardized moment) at $44.393$ and $49.226$ respectively. In both cases, the distributions are highly leptokurtic: they are sharp-pointed around the mean, have relatively little mass in the shoulders and very fat tails. This is consistent with Busch and Ludwig (2021), who use the PSID until 2012 and estimate numerically similar kurtosis coefficients for both shocks. Together with left skewness, large excess kurtosis implies that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad. Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
5.2 Linear consumption function
We present estimates of the linear consumption function, first, targeting only second moments of the joint income-consumption distribution and, subsequently, targeting second as well as higher-order moments of the joint distribution. The full set of moments targeted in each case is shown in appendix B.4.2.
Targeting second-order moments only. The estimation of the linear consumption function targeting second moments alone is effectively a replication of BPP over the recent years (1999-2019) while using consumption data that are internally available in the PSID. [^29] We use the pre-estimated second moments of income shocks from column 1 of table 3. This replication is important in its own right. The recent years offer superior consumption data (in the sense that we do not need to rely on an imputation from the CEX), thus reducing the scope of measurement/imputation error. The data, however, come at a biennial rather than annual frequency. Moreover, there have been many changes in taxation and redistribution policies since 1980–1992, the original period studied in BPP, which may matter for the extent of partial insurance available to households more recently.
The results are in column 1 of table 4. We estimate the transmission parameter of permanent shocks at $\phi^{(1)}=0.152$. A $10\%$ permanent change in income thus results in only $1.52\%$ shift in consumption. This reflects a very low pass-through of permanent shocks and a substantial degree of consumption insurance. While highly significant, the estimate for $\phi^{(1)}$ is strikingly different from BPP, whose baseline estimate of $0.64$, four times bigger, indicates a much larger pass-through of permanent shocks, hence a much lower degree of insurance. By contrast, we find evidence of full insurance to transitory shocks, similarly to BPP. We estimate the pass-through of said shocks at $\psi^{(1)}=-0.006$, statistically indistinguishable from zero. Therefore, transitory shocks do not pass through into consumption. [^30] Commault (2022) generalizes BPP’s linear consumption function by allowing past transitory shocks to enter the specification in (4). She then finds a strong consumption response to transitory income, similar to the response estimated from quasi-experimental data. We abstract from this generalization, as our focus here is on higher-order income moments and on non-linearities in the consumption response.
What explains the discrepancy in $\phi^{(1)}$ with BPP? As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period. To help decipher their role, we impute consumption from the CEX into the PSID in the same way BPP did for the earlier period, and we repeat the estimation; see appendix C.3 for details on the imputation. The results are in column 2 of table 4. The transmission parameter of permanent shocks doubles to $\phi^{(1)}=0.288$, while the transmission parameter of transitory shocks remains statistically zero (though at a larger magnitude). Using imputed consumption data thus inflates the pass-through of permanent shocks, substantially lowering the degree of partial insurance. This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true series (column 1) and the imputed data from the CEX (column 2). By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and also very close to BPP. The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. We relegate the details of the imputation and a discussion of a mechanism through which the imputation in particular inflates $\phi^{(1)}$ to appendix C.3.
Table 4: Estimates of the consumption function, baseline sample
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID CEX PSID PSID PSID
(1) (2) (3) (4) (5)
$\phi^{(1)}$ 0.152 0.288 0.160 0.112 0.134
(0.028) (0.044) (0.029) (0.041) (0.026)
$\psi^{(1)}$0.006 -0.109 -0.012 0.007 -0.003
(0.045) (0.078) (0.038) (0.056) (0.052)
$\phi^{(2)}$0.040
$\psi^{(2)}$ 0.015
$\omega^{(22)}$ 0.320
$\sigma^{2}_{\xi}$ 0.019 0.019 0.019 0.020 0.019
(0.001) (0.004) (0.001) (0.001) (0.001)
$\gamma_{\xi}$0.001 -0.001
(0.001) (0.001)
$\kappa_{\xi}$ 0.005
$\sigma^{2}_{u_{c}}$ 0.044 0.140 0.044 0.044 0.044
(0.002) (0.009) (0.002) (0.001) (0.002)
$\gamma_{u_{c}}$ 0.002 0.002
(0.001) (0.001)
$\kappa_{u_{c}}$ 0.013
Notes: The table presents the estimates of the parameters of the consumption function, imposing homogeneous transmission parameters over the lifecycle/in the cross-section and stationarity of taste heterogeneity and consumption measurement error. Columns 1-4 present parameter estimates in the linear function, while column 5 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. Except for column 2 where we use imputed data from the CEX, all other columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. Results from optimal GMM are similar (appendix table C.6).
Although the imputation inflates $\phi^{(1)}$, the inflated value in column 2 is still substantially lower than BPP. The PSID switched in 1997 from annual to biennial frequency. It is thus possible that at the lower frequency certain higher frequency income shocks are smoothed out or undone by subsequent shocks. To assess this, we would need to re-estimate our model on annual growth rates, which is impossible after 1997. However, we can use the original BPP data over 1980–1992 to estimate the model annually versus biennially. We do this in appendix C.4. While we obtain $\phi^{(1)}=0.64$ using annual growth rates (exactly as BPP), this drops to $\phi^{(1)}=0.4$–0.5 using biennial rates. The drop is sharper when we target time-invariant moments, which was recently also shown in Chatterjee et al. (2021). The biennial nature of the modern data thus explains another 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. [^31] In both annual and biennial cases, $\phi^{(1)}$ measures the pass-through of the permanent shock at an annual rate. This becomes clear in appendix B that shows the estimating equations with biennial data. The difference in the estimates of $\phi^{(1)}$ between the two cases arises because with biennial data we only measure an ‘aggregate’ pass-through within a given biennial period. Consequently, some higher-frequency shocks will be inevitably smoothed out at the observed lower frequency and their transmission will be muted. We attribute the remaining gap to calendar time effects, e.g., changes in taxes/benefits or other insurance channels between 1980–1992 and 1999–2019 (see Borella et al., 2022, for a discussion of such changes). It is worth noting that we are not the first paper to find a low pass-through of permanent shocks in the recent years; Arellano et al. (2023) use PSID data after 2005 and find a pass-through of about $0.3$ (specification closest to ours). If we zoom into the period after 2005, we estimate $\phi^{(1)}=0.21$ (results not shown for brevity), statistically indistinguishable from Arellano et al. (2023)’s estimate.
Introducing higher-order moments. The estimation of the linear consumption function targeting higher-order moments is our first attempt to uncover how the tails of income risk affect the pass-through of shocks into consumption. In other words, we want to measure the degree of partial insurance going beyond the workhorse income-consumption covariance to account for the non-Gaussian features of modern income dynamics. We introduce the higher-order moments incrementally. [^32] We use the corresponding pre-estimated income parameters from columns 2 and 3 of table 3 respectively.
Column 3 of table 4 presents the results from targeting third- in addition to second-order moments of consumption and income. The partial insurance parameters remain largely unchanged upon inclusion of the third moments. We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with second moments only) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero). We thus again find large partial insurance to permanent shocks (a $10\%$ permanent change in income results in only $1.6\%$ shift in consumption) and full insurance to transitory shocks. Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with second moments only. Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or consumption measurement error.
Column 4 of table 4 presents the results from targeting all income and consumption moments up to fourth-order. The earlier findings hold up. We estimate $\phi^{(1)}=0.112$ (versus $0.152$ and $0.160$ in the lower-order cases) and $\psi^{(1)}=0.007$ (indistinguishable from zero). A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption. Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth-order moments. The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks. [^33] Appendix table C.7 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 4. We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong, significant, and negative response to transitory shocks, which is hard to explain. We do not think those results are very credible; the imputation brings in substantial measurement/imputation error, which becomes worse the higher the order of the target moments is.
Clearly, targeting higher-order moments in the context of the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone. This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary. For example, Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets, while Busch and Ludwig (2021) find sizable welfare implications of higher-order risk. [^34] De Nardi et al. (2019) and Busch and Ludwig (2021) find higher insurance when shocks are non-normal, due to stronger precautionary motives. The latter caution against BPP’s linear framework, showing evidence of bias in the presence of tail risk. Our quadratic specification addresses their concern. One would thus expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events.
Yet, the linear consumption function leaves no room to such events. BPP rely on a log-linearization of the consumption rule around a deterministic lifecycle profile (appendix A). But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response. A linearized marginal utility, however, fails to pick up such large response (echoing Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments. In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates and justifies our higher-order specification subsequently.
To summarize our findings from the linear consumption function: 1.) we replicate BPP over the recent years and find much lower pass-through of permanent shocks (larger degree of partial insurance); the imputation of consumption from the CEX and the biennial nature of the modern PSID explain the bulk of the difference from BPP; 2.) targeting higher-order moments does not alter the degree of consumption partial insurance, which is not entirely surprising as tail risk has, by design, no role in the linear framework.
5.3 Quadratic consumption function
We now turn to the estimation of the quadratic consumption function. The results appear in column 5 of table 4; the full set of targeted moments is shown in appendix B.4.3.353535Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 3., [^36] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case, i.e., in column 4, table 4. This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model. This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2). However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway (section 3). Our estimates of the linear function in that case do not differ from those in column 4, table 4, so estimation is robust to the choice of moments. These results are available upon request.
We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, highly significant and close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification. We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero. As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial in the shocks. We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks (statistically significant at the margin for now), while we estimate a positive $\psi^{(2)}=0.015$ for transitory shocks (statistically insignificant for now). The coefficient $\omega^{(22)}$ on the product of permanent and transitory shocks is large but imprecisely estimated. The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case. [^37] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
The values of $\phi^{(2)}$ and $\psi^{(2)}$ have important implications for the pass-through of shocks into consumption. To help interpret the results, recall that the pass-through of permanent shocks is given by (6) and depends on $\phi^{(1)}$, $\phi^{(2)}$, $\omega^{(22)}$, and the magnitudes and signs of both income shocks. If, for simplicity, we set the transitory shock to its mean value ($v_{it}=0$) and plug in the parameter estimates from table 4, we get
$\partial\Delta c_{it}/\partial\zeta_{it}=0.134-0.08\times\zeta_{it}.$
Two main results follow.
First, bad (negative) permanent shocks are more hurtful than good (positive) ones because they are associated with a much larger transmission parameter into consumption. Consider, for example, a large permanent shock to household income, $\zeta_{it}=\pm 0.5$, approximately three standard deviations away from the mean. If the shock is negative ($\zeta_{it}=-0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$; by contrast, if the shock is positive ($\zeta_{it}=0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$. We can generalize this to bad and good shocks of any magnitude and conclude that
$\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<0}~{}>~{}\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta>0}.$
Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 3, but they also have much larger pass-through rates into consumption. Bad permanent shocks are thus double hurtful.
Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption. Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean. In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$. We can generalize this to bad shocks of any magnitude and conclude that
$\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<<0}~{}>~{}\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<0}.$
This result is not hard to comprehend. When a bad shock hits, households can insure consumption by running down savings or by using other formal or informal insurance arrangements. But there is clearly a limit to this. As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, thus resulting in a larger fraction of the shock eventually transmitting into consumption. [^38] The parameter estimates also imply that the pass-through of a good shock decreases with its magnitude. While this seems counterintuitive, recall that extreme good shocks are rare in our data, which follows from left skewness and excess kurtosis in table 3. We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios. For example, Bunn et al. (2018) find that the marginal propensity to consume from negative income shocks is larger than from positive shocks, while Christelis et al. (2019) and Fuster et al. (2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the shock. This literature, however, elicits responses to transitory shocks. By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
Why would a positive or small shock be better insured? While we cannot assess candidate mechanisms without a structural model, our best explanation is that people save a good income shock (thus consumption does not respond much) but they dissave in the presence of a bad one. As there is clearly a limit to how much households can dissave or borrow, especially in the presence of borrowing constraints, a large negative shock would typically affect consumption by relatively more. [^39] Busch and Ludwig (2021) simulate a lifecycle model with concave utility and borrowing constraints; they find larger pass-through of negative versus positive permanent shocks. While we do not model borrowing constraints here, a strength of our approach is that it enables us to measure partial insurance in the data without taking a stance on the precise underlying structural model/economic mechanism at play.
Moving on to transitory shocks, their pass-through is given by (7); this depends on $\psi^{(1)}$, $\psi^{(2)}$, $\omega^{(22)}$, and on the magnitudes and signs of both income shocks. If we set the permanent shock to its mean value ($\zeta_{it}=0$) and plug in the parameter estimates from table 4, we get
$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.$
Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample. This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption. This is similar to our findings for the linear case as well as to BPP. By contrast, this is not the case for certain well-defined subsamples, but we postpone the discussion of the subsamples for section 5.4.
Despite the lack of statistical significance, the partial derivative with respect to transitory shocks has, at face value, an interesting implication: large good transitory shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero. While we should be cautious in drawing conclusions about extreme good transitory shocks (given left skewness and excess kurtosis, large good transitory shocks are not present in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and other semi-structural studies that use covariance restrictions on survey data. The former studies typically exploit large and infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021). In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions. Our nonlinear consumption function reconciles the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
To summarize our findings from the quadratic consumption function: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do. For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein. By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way. [^40] Appendix table C.7 shows that similar conclusions are obtained if we use consumption imputed from the CEX. In particular, $\phi^{(2)}<0$ (albeit not statistically significant), with a magnitude similar to that obtained directly from the PSID ($-0.036$ versus $-0.04$), and $\psi^{(2)}>0$.
5.4 Sensitivity to age, wealth, and education
We first consider how partial insurance changes with age, by repeating the estimation over three, almost equally sized, subsamples of households with the male spouse between 30–40, 41–50, and 51–65 years old respectively. [^41] We target the same moments as earlier, broken down by the age of the male spouse. The moments of income shocks are also age-specific but we do not report those results for brevity. Intuitively, the transmission parameters of shocks should exhibit some age dependence since wealth accumulates over the lifecycle, which in turn influences the household’s ability to insure against income fluctuations.
The results appear in table 5. Column 1 presents results from the linear specification when only second moments are targeted. Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65. A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are included in the estimation. This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results. [^42] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters. By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details. By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
Table 5: Estimates of the consumption function, by age
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID
(1) (2) (3) (4)
$\phi^{(1)}$ 30-40 0.154 (0.045) 0.164 (0.051) 0.089 (0.058) 0.149 (0.050)
41-50 0.126 (0.051) 0.146 (0.055) 0.079 (0.089) 0.115 (0.049)
51-65 0.180 (0.049) 0.175 (0.051) 0.136 (0.061) 0.113 (0.057)
$\psi^{(1)}$ 30-40 0.094 (0.079) 0.083 (0.073) 0.095 (0.107) 0.105 (0.090)
41-50 0.010 (0.079) -0.016 (0.072) -0.075 (0.061) 0.019 (0.090)
51-65 -0.046 (0.070) -0.044 (0.066) 0.037 (0.101) -0.063 (0.088)
$\phi^{(2)}$ 30-40 0.006 (0.040)
41-50 -0.008 (0.063)
51-65 -0.105 (0.054)
$\psi^{(2)}$ 30-40 0.036 (0.061)
41-50 0.048 (0.046)
51-65 -0.052 (0.079)
$\omega^{(22)}$ 30-40 -1.090 (1.465)
41-50 -0.290 (1.195)
51-65 3.003 (1.306)
$\sigma^{2}_{\xi}$ 30-40 0.014 (0.002) 0.014 (0.002) 0.016 (0.002) 0.013 (0.004)
41-50 0.022 (0.002) 0.022 (0.002) 0.022 (0.002) 0.022 (0.002)
51-65 0.021 (0.002) 0.021 (0.002) 0.023 (0.002) 0.010 (0.008)
$\gamma_{\xi}$ 30-40 -0.001 (0.001) -0.001 (0.001)
41-50 0.006 (0.001) 0.006 (0.002)
51-65 -0.006 (0.002) -0.006 (0.002)
$\kappa_{\xi}$ 30-40 0.004 (0.002)
41-50 0.007 (0.002)
51-65 0.006 (0.002)
$\sigma^{2}_{u_{c}}$ 30-40 0.046 (0.003) 0.046 (0.003) 0.046 (0.003) 0.046 (0.003)
41-50 0.040 (0.002) 0.040 (0.002) 0.040 (0.002) 0.040 (0.002)
51-65 0.043 (0.002) 0.043 (0.002) 0.043 (0.002) 0.043 (0.002)
$\gamma_{u_{c}}$ 30-40 0.000 (0.002) 0.000 (0.002)
41-50 0.002 (0.002) 0.002 (0.002)
51-65 0.004 (0.002) 0.004 (0.002)
$\kappa_{u_{c}}$ 30-40 0.014 (0.003)
41-50 0.012 (0.002)
51-65 0.012 (0.002)
Notes: The table presents the estimates of the parameters of the consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse (the underlying income moments also vary over these age brackets). Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
Column 4 presents the results from the quadratic specification. In this case, $\phi^{(1)}$ decreases monotonically with age, so the average permanent shock (the shock close to the mean of the distribution) impacts consumption gradually less as households age. In line with theory, this pattern reflects the age-increasing role of financial wealth for self-insurance. Perhaps more importantly, $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases monotonically with age. $\phi^{(2)}$ drops from $0.006$ to $-0.008$ as we move from ages 30–40 to ages 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.105$ in the 51–65 age bracket. As a result, large negative permanent shocks impact consumption more severely at older than at younger ages.
The last finding echoes Guvenen et al. (2021), who show that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the low-earners or the young. In fact, we also find larger left skewness in income growth as households age in the PSID and more negative skewness in permanent shocks among the old. [^43] The coefficient of skewness of permanent shocks more than triples in magnitude from $-0.65$ to $-2.39$ as we move from ages 30–40 to 51–65. The coefficient of skewness of transitory shocks, by contrast, does not change much with the age – its estimate is $-1.43$ and $-1.45$, respectively, over ages 30–40 and 51–65. So, not only do permanent shocks become more negative as households age, but they also transmit into consumption at increasingly larger rates. In other words, not only is there more room for income to fall among older earners, but this also causes even greater harm to consumption. By contrast, despite a similar pattern for transitory shocks, their transmission parameters remain statistically indistinguishable from zero over the entire lifecycle.
We next consider two subsamples formed on the basis of household wealth. We define a household as ‘low wealth’ (‘high wealth’) if its average wealth over its entire observation window is less (more) than median real wealth in the sample over 19991999199919992019201920192019. Wealth is the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt. Table 1 provides summary statistics for the wealth subsamples. Average wealth in the first group is \$69,731 in 2018 prices, while the second group has 14 times more wealth, an average of \$950,823. Income and earnings are twice as high in the second group; consumption is only 30% higher.
Table 6, columns 1-2, presents the results in the quadratic model. [^44] Appendix table C.8, columns 1-4, presents the parameter estimates in the linear model. Three points emerge. First, $\phi^{(1)}$ is more than twice as large among the low vis-à-vis high wealth households ($0.212$ versus $0.091$). As $\phi^{(1)}$ measures the transmission of the average permanent shock, low wealth households exhibit a larger pass-through of average shocks, i.e., they are less able to insure against them, compared to high wealth households. Second, $\phi^{(2)}<0$ for both groups, albeit statistically insignificant. With the caveat of insignificance, the magnitude of $\phi^{(2)}$ is bigger among the wealthy. Large negative permanent shocks thus impact consumption of the wealthy more severely than they do among the less wealthy. As in the case of older households previously, there is more room for consumption to fall sharply among the wealthier. Third, $\psi^{(2)}$, the transmission parameter of the typical transitory shock away from the mean, is negative, large, and statistically significant among the less wealthy. Transitory shocks thus have an asymmetric pass-through among lower wealth households, with large bad shocks reducing consumption substantially more than good shocks increase it. So while average transitory shocks are fully insured among low wealth households, larger negative transitory shocks are not: they reduce consumption quite substantially as low wealth households lack the self-insurance means to smooth out even transitory fluctuations.
Table 6: Estimates of the consumption function, by wealth and education
Consumption fn.: Quadratic
Subsample: Low wealth High wealth No college Some college
Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID
(1) (2) (3) (4)
$\phi^{(1)}$ 0.212 0.091 0.219 0.105
(0.041) (0.033) (0.057) (0.032)
$\psi^{(1)}$0.081 0.017 -0.071 0.030
(0.090) (0.059) (0.079) (0.063)
$\phi^{(2)}$0.018 -0.039 -0.095 -0.023
(0.052) (0.035) (0.048) (0.035)
$\psi^{(2)}$0.108 0.031 0.050 0.012
(0.065) (0.041) (0.063) (0.040)
$\omega^{(22)}$ 1.116 -0.049 0.668 0.095
(0.874) (0.897) (1.067) (0.971)
$\sigma^{2}_{\xi}$ 0.013 0.022 0.013 0.021
(0.002) (0.002) (0.003) (0.002)
$\sigma^{2}_{u_{c}}$ 0.048 0.042 0.043 0.044
(0.002) (0.002) (0.003) (0.002)
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups (linear specification estimates appear in appendix table C.8). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
We finally turn to the role of education. We define a household as ‘no college’ (‘some college’) if the male spouse has not gone to college (has some college education). Table 1 provides summary statistics. Those with some college education earn 70% more income, they have more than twice as much wealth, and they consume about 30% more.
Table 6, columns 3-4, presents the results in the quadratic model. [^45] Appendix table C.8, columns 5-8, presents the parameter estimates in the linear model. Similar to the patterns observed by wealth, we find that $\phi^{(1)}$ is higher (double) for households without college education ($0.219$ versus $0.105$). The pass-through of the average permanent shock is therefore larger among the lower educated, presumably because of the lower wealth that these households hold. $\phi^{(2)}$ is negative in both groups, suggesting that large bad permanent shocks transmit into consumption more than equal-sized good shocks. However, the magnitude of the estimate in the lower education group is over four-fold larger (and statistically significant) compared to the higher education group. Permanent shocks away from the mean are therefore way more hurtful among the lower educated. By contrast, the transmission parameters of all transitory shocks are statistically indistinguishable from zero in both groups.
6 Welfare costs of tail income risk
In the previous section, we established that income risk is negatively skewed and highly leptokurtic, and features a nonlinear pass-through to consumption. In this section, we calculate the welfare costs of such risk, particularly contrasting the implications of the linear and nonlinear consumption functions. To do so, we first draw a sequence of income shocks over the lifecycle, we then use the different consumption specifications to simulate responses to those shocks, and, finally, we use a parametric utility function to calculate household welfare. Our interest is in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
We draw permanent and transitory income shocks for 50,000 households from mixtures of normal distributions, namely
$\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob. }p_{\zeta}\\ {\cal N}(\mu_{\zeta,2},\sigma_{\zeta,2})\text{ with prob. }1-p_{\zeta}\\ \end{cases}\text{ and }~{}v_{it}\sim\begin{cases}{\cal N}(\mu_{v,1},\sigma_{v,1})\text{ with prob. }p_{v}\\ {\cal N}(\mu_{v,2},\sigma_{v,2})\text{ with prob. }1-p_{v}.\\ \end{cases}$
Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021). There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of the shocks’ distributions in the PSID (column 3, table 3). We assume $\mu_{1}<0$ for identification. Appendix table D.1 reports the parameter estimates and the (excellent) model fit. In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
We simulate alternative consumption profiles in response to the income shocks using the linear and nonlinear specifications
$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$
$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\zeta_{it}^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\zeta_{it}v_{it}.$
These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate directly in our baseline sample. We simulate consumption over a working life of 35 years, starting at age 30. We also simulate a retirement period of 20 more years (starting at age 65), in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. [^46] We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample. Our baseline simulations use the time-invariant partial insurance parameters from table 4. Appendix table D.3 relaxes time-invariance.
We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion (CRRA) utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, or logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$. We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk. This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation
${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}.$
${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles, respectively reflecting less (Gaussian) and more (non-Gaussian) income risk. We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of a linear consumption function. Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
In a second scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the quadratic specification in response to Gaussian and non-Gaussian income shocks, respectively. The estimate of $\alpha_{i}$ here reflects the welfare cost of tail risk in the presence of nonlinear transmission. Given the evidence that larger negative shocks have a stronger impact on consumption when nonlinear transmission is allowed for, we expect larger welfare costs of tail risk in this scenario.
In a third scenario, the less risky profile ${\cal C}_{i}^{*}$ reflects the linear specification in response to Gaussian risk, while the riskier profile ${\cal C}_{i}$ is simulated from the quadratic specification in response to non-Gaussian shocks. Here $\alpha_{i}$ reflects both the welfare cost of tail risk and the bias induced by neglecting the possibility of nonlinear transmission of such risk.
Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shock they experience, in the case of $\gamma=2$. The first row presents the average cost across all households, unconditional on a given shock. This is close to zero across all three scenarios, or even negative in some cases, suggesting that households may in fact prefer tail over Gaussian risk. Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance. As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the empirical variance of shocks (or the variance of Gaussian risk). For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
Table 7: Welfare costs of non-Gaussian relative to Gaussian income risk
Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -0.78 -0.86 0.65
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.74 6.31 7.71
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.28 8.53 9.90
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.43 11.74 13.06
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 4. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
The next rows zoom in on households who experience large negative permanent shocks at least once in their lifetime. These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them. Those who face a permanent shock at least three standard deviations below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.74% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly. This increases to 6.28% and 8.43% for shocks respectively four and five standard deviations below the mean – all shown in column 1. While these costs are already sizable, they increase further by 55% to 62% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.06%) for those households who experience the worst shocks. Those estimates would be even larger if we considered the welfare cost of tail risk relative to a zero risk scenario, as, e.g., in Guvenen et al. (2024).
The welfare estimates are robust to the choice of $\gamma$ (appendix table D.2) or the use of age-specific partial insurance parameters when simulating various consumption profiles (appendix table D.3). Tail income risk consistently implies large welfare costs. The econometrician will greatly underestimate such costs if she neglects the nonlinear transmission into consumption, for which we found ample evidence in the earlier sections.
7 Conclusion
We measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution. Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers. Our focus is on measuring partial insurance without taking a stance on the transmission mechanisms from income to consumption. We use recent income and consumption data from the PSID (1999–2019) and establish several new insights.
First, permanent and transitory shocks to disposable household income exhibit negative skewness and very large excess kurtosis.
Second, in estimating BPP over the recent years, we find a lower pass-through of permanent shocks (larger partial insurance) and full insurance to transitory ones. We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
Third, in introducing third- and fourth-order income and consumption moments into the linear consumption model does not alter the above results.
Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock. By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy.
Fifth, tail risk incurs substantial welfare costs, up to 1/8 of lifetime consumption.
In summary, the paper offers an analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and severity of the shock. To measure the transmission, we exploit information in higher-order income moments but we do not otherwise need to observe the entire consumption-income joint distribution (e.g., in contrast to Arellano et al., 2017). Our results reconcile the traditional partial insurance measurement literature, in which BPP is the default methodology, with both the quantitative literature (e.g., Guvenen et al., 2024) and the more empirical one based on surveys or experiments (e.g., Fuster et al., 2020), both of which document asymmetries in the pass-through of good and bad, or small and large shocks. Overall, our results suggest that tail income risk matters substantially for consumption.
Appendix A Derivation of the consumption function
In this appendix we derive the consumption function under a permanent-transitory specification for log income. We follow similar derivations in Blundell et al. (2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein. We show below how to derive the most general (quadratic) consumption function (5); the linear function (4) is a special case of this. We let a time unit in the model correspond to one calendar year. However, data in the PSID in recent years are available only every second year, so we adapt the derivations to reflect this.
Let the utility function take the form $U(C_{it};\mathbf{Z}_{it})=\widetilde{U}(C_{it}\exp(-\mathbf{Z}_{it}^{\prime}\boldsymbol{\alpha}))$, which simplifies the subsequent statements. The first-order conditions of the household problem (1) s.t. (2) in a generic period $t+s$ are
[…]
where $\widetilde{U}_{C}$ is the marginal utility of consumption and $\widetilde{C}_{it}=C_{it}\exp(-\mathbf{Z}_{it}^{\prime}\boldsymbol{\alpha})$. $\lambda_{it}$ is the Lagrange multiplier on the sequential budget constraint (the marginal utility of wealth).
A.1 Taylor approximation to optimality conditions
Approximation to static optimality condition. Applying logs to the static condition and taking a first difference over time yields $\Delta\ln\widetilde{U}_{C}(\widetilde{C}_{it+s})-\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})=\Delta\ln\lambda_{it+s}$. A first-order Taylor expansion of $\ln\widetilde{U}_{C}(\widetilde{C}_{it+s})$ around $\widetilde{C}_{it+s-1}$ yields
$\Delta\ln\widetilde{U}_{C}(\widetilde{C}_{it+s})$ $\approx\frac{\widetilde{U}_{CC}(\widetilde{C}_{it+s-1})}{\widetilde{U}_{C}(\widetilde{C}_{it+s-1})}\exp(-\mathbf{Z}_{it+s-1}^{\prime}\boldsymbol{\alpha})(C_{it+s}-C_{it+s-1})$
$\approx\frac{\widetilde{U}_{CC}(\widetilde{C}_{it+s-1})}{\widetilde{U}_{C}(\widetilde{C}_{it+s-1})}C_{it+s-1}\exp(-\mathbf{Z}_{it+s-1}^{\prime}\boldsymbol{\alpha})\Delta\ln C_{it+s}=\theta^{-1}_{it+s-1}\Delta\ln C_{it+s},$
where $\widetilde{U}_{CC}$ is the second derivative of the utility function and $\theta_{it}=\widetilde{U}_{CC}^{-1}(\widetilde{C}_{it})\widetilde{C}^{-1}_{it}\widetilde{U}_{C}(\widetilde{C}_{it})$ is the inverse of the consumption substitution elasticity. Plugging the last expression in the log-linearized first-order condition yields $\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})=\theta_{it+s-1}\Delta\ln\lambda_{it+s}$.
Approximation to Euler equation. The approximation to the intertemporal condition involves future expectations. Let $\exp(\Gamma)=\beta^{-1}(1+r)^{-1}$ for some $\Gamma$. A second-order Taylor expansion of $\exp(\ln\lambda_{it+s+1})$ around $\ln\lambda_{it+s}+\Gamma$ yields
$\lambda_{it+s+1}\approx\frac{\lambda_{it+s}}{\beta(1+r)}\left(1+(\Delta\ln\lambda_{it+s+1}-\Gamma)+\frac{1}{2}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}\right).$
Taking expectations at time $t+s$ and noting that ${\operatorname*{\mathbb{E}}}_{t+s}\lambda_{it+s+1}=\lambda_{it+s}\beta^{-1}(1+r)^{-1}$, we obtain
${\operatorname*{\mathbb{E}}}_{t+s}\Delta\ln\lambda_{it+s+1}\approx\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t+s}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}\Rightarrow\Delta\ln\lambda_{it+s+1}\approx\omega_{it+s+1}+\epsilon_{it+s+1}.$
The first term $\omega_{it+s+1}=\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}$ captures the impact of interest rates, impatience, and precautionary motives on the growth of the marginal utility of wealth. To maintain tractability, we assume that $\omega$ is non-stochastic but possibly heterogeneous across households. The second term is an expectation error with ${\operatorname*{\mathbb{E}}}_{t}(\epsilon_{it+1})=0$; it captures idiosyncratic revisions to the marginal utility of wealth upon arrival of income shocks.
Combining the approximations to the optimality conditions, we obtain
$\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})$ $=\theta_{it+s-1}(\omega_{it+s}+\epsilon_{it+s})$
$\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$ $=\xi_{it+s}+\theta_{it+s-1}\epsilon_{it+s}$
$\Delta c_{it+s}$ $=\xi_{it+s}+\theta_{it+s-1}\epsilon_{it+s},$ (A.1)
where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$. We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to interest rates and impatience (discounting), the second term reflects unobserved consumption taste heterogeneity. We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path. Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of log consumption on appropriate observables (demographics, time fixed effects).
A.2 Taylor approximation to lifetime budget constraint
Let $G(\boldsymbol{\xi})=\ln\sum_{s=0}^{N}\exp\xi_{s}$ for $\boldsymbol{\xi}=(\xi_{0},\xi_{1},\dots,\xi_{N})^{\prime}$. Applying a second-order Taylor expansion of $G(\boldsymbol{\xi})$ around a deterministic $\boldsymbol{\xi}^{0}$, and taking expectations given information ${\cal I}$, yields
$\begin{split}{\operatorname*{\mathbb{E}}}_{\cal I}G(\boldsymbol{\xi})\approx G(\boldsymbol{\xi}^{0})&+\sum_{s=0}^{N}\frac{\exp\xi^{0}_{s}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}{\operatorname*{\mathbb{E}}}_{\cal I}(\xi_{s}-\xi^{0}_{s})\\ &+\frac{1}{2}\sum_{s=0}^{N}\sum_{\ell=0}^{N}\frac{\exp\xi^{0}_{s}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}\left(\delta_{s\ell}-\frac{\exp\xi^{0}_{\ell}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}\right){\operatorname*{\mathbb{E}}}_{\cal I}(\xi_{s}-\xi^{0}_{s})(\xi_{\ell}-\xi^{0}_{\ell}),\end{split}$ (A.2)
where $\delta_{s\ell}$ is the Kronecker delta ($\delta_{s\ell}=1$ if $s=\ell$, $\delta_{s\ell}=0$ otherwise). We apply this approximation to both sides of the intertemporal budget constraint.
Assuming expectations away, the logarithm of the left hand side of the budget constraint (2) in a generic period $t$ is $\ln\left(A_{it}+\sum_{s=0}^{T-t}\exp\left(\ln Y_{it+s}-s\ln(1+r)\right)\right)$. Letting
$\xi_{s}$ =$=\begin{cases}\ln A_{it+s}&\text{ for }s=0\\ \ln Y_{it+s-1}-(s-1)\ln(1+r)\qquad\qquad&\text{ for }s=1,\dots,T-t+1\\ \end{cases}$
$\xi^{0}_{s}$ =$=\begin{cases}{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it+s}&\text{ for }s=0\\ {\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s-1}-(s-1)\ln(1+r)&\text{ for }s=1,\dots,T-t+1,\end{cases}$
and declaring the information set to contain time $t$ information, we follow (A.2) to write
${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)\approx\ln\left(\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})+\sum_{s=0}^{T-t}\exp\left({\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s}-s\ln(1+r)\right)\right)$
$+\pi_{it}{\operatorname*{\mathbb{E}}}_{t}(\ln A_{it}-{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})$
$+(1-\pi_{it})\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})$
$+\frac{1}{2}\pi_{it}(1-\pi_{it}){\operatorname*{\mathbb{E}}}_{t}(\ln A_{it}-{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})^{2}$
$+\frac{1}{2}\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
$-\pi_{it}(1-\pi_{it})\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln A_{it}-{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s}).$
The notation is: $\pi_{it}=\frac{Q_{1t}}{Q_{1t}+Q_{2t}}$ is equal to the expected share of financial wealth $Q_{1t}=\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})$ in the household’s total financial and human wealth at $t$, the latter defined as $Q_{2t}=\sum_{\kappa=0}^{T-t}\exp\left({\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\kappa}-\kappa\ln(1+r)\right)$, i.e., the household’s expected lifetime income for the remaining of life; $\vartheta_{it+s}^{Y}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln Y_{it+s}-s\ln(1+r)\right)/Q_{2t}$ is an annuitization factor equal to the expected share of time $t+s$ household income in total lifetime income. $\pi_{it}$ and $\vartheta_{it+s}^{Y}$ pertain to expectations at $t-2$ so they are both known at $t-2$ and later.
Using the income process and assuming the deterministic observables away (they cancel out in a first difference in expectations at a later stage), we reach $\ln Y_{it+s}=\ln Y_{it-2}+\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}-v_{it-2}$, therefore $\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it-2}=\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}$. It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})=\zeta_{it}+\zeta_{it-1}+\vartheta_{it}^{Y}v_{it}$ because ${\operatorname*{\mathbb{E}}}_{t}\zeta_{it+\kappa}={\operatorname*{\mathbb{E}}}_{t}v_{it+\kappa}=0$ for $\kappa>0$, and $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$ by construction. The second-order term in income is more involved. This is given by
$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}\left((\zeta_{it}+\zeta_{it-1})^{2}+(\zeta_{it}+\zeta_{it-1})(v_{it+s}+v_{it+\ell})+v_{it+s}v_{it+\ell}\right)$
$=(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+2\pi_{it}(1-\pi_{it})\vartheta_{it}^{Y}(\zeta_{it}+\zeta_{it-1})v_{it}+(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}.$
We derive this making use of $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$, $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}\vartheta_{it+\ell}^{Y}=(\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{Y})=1$, and the expectation of future shocks being zero by construction.
Plugging these results in the approximation of the previous page, noting that assets at $t$ are ‘beginning-of-period’ and thus independent of the time $t$ income shocks, and taking a first difference in expectations between $t-2$ and $t$ yields
${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)-{\operatorname*{\mathbb{E}}}_{t-2}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)$ $\approx(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}$
$+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}$
$+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}$
$+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.$
The linear terms in the first line are the first-order approximation of BPP (adapted for biennial data); the quadratic terms in the next lines are the refinement from the second-order approximation.
Assuming expectations away, the logarithm of the right hand side of the budget constraint (2) is $\ln\sum_{s=0}^{T-t}\exp(\ln C_{it+s}-s\ln(1+r))$. Adopting the notation of (A.2), we let $\xi_{s}=\ln C_{it+s}-s\ln(1+r)$ and $\xi_{s}^{0}={\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s}-s\ln(1+r)$. Letting the information set contain time $t$ information, it follows that
${\operatorname*{\mathbb{E}}}_{t}\ln\sum_{s=0}^{T-t}\frac{C_{it+s}}{(1+r)^{s}}$ $\approx\ln\sum_{s=0}^{T-t}\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s}-s\ln(1+r))$
$+\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}{\operatorname*{\mathbb{E}}}_{t}(\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s})$
$+\frac{1}{2}\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C}){\operatorname*{\mathbb{E}}}_{t}(\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s})(\ln C_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+\ell}),$
where $\vartheta_{it+s}^{C}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+s}-s\ln(1+r)\right)/\sum_{\kappa=0}^{T-t}\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+\kappa}-\kappa\ln(1+r)\right)$ is an annuitization factor equal to the expected share of time $t+s$ consumption in total lifetime consumption. $\vartheta_{it+s}^{C}$ pertains to expectations at $t-2$, so it is known at $t-2$ and later.
Using the linearized consumption function in (A.1) to replace $\ln C_{it+s}$ recursively, we reach $\ln C_{it+s}=\ln C_{it-2}+\sum_{\tau=-1}^{s}\Psi_{it+\tau}+\sum_{\tau=-1}^{s}\xi_{it+\tau}+\sum_{\tau=-1}^{s}\theta_{it+\tau-1}\epsilon_{it+\tau}$, where $\Psi_{it+\tau}=\theta_{it+\tau-1}\Delta(\mathbf{Z}_{it+\tau}^{\prime}\boldsymbol{\alpha})+\Xi_{it+\tau}$, therefore $\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s}=\sum_{\tau=-1}^{s}\theta_{it+\tau-1}\epsilon_{it+\tau}$ because $\Psi_{it}$ and $\xi_{it}$ are non-stochastic by assumption and they cancel out in the first difference. It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}{\operatorname*{\mathbb{E}}}_{t}(\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s})=\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1}$ because ${\operatorname*{\mathbb{E}}}_{t}\epsilon_{it+\kappa}=0$ for $\kappa>0$ and $\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}=1$ by construction. The second-order term is more involved; it is given by
$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C}){\operatorname*{\mathbb{E}}}_{t}(\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s})(\ln C_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+\ell})$
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})(\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1})^{2}=0$
We derive this by noting that future expectations errors are zero by assumption, and that $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})=\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}\delta_{ss}+\sum_{s=0}^{T-t}\sum_{\ell\neq s}\vartheta_{it+s}^{C}\delta_{s\ell}-(\sum_{s=0}^{T-t}\vartheta_{it+s}^{C})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{C})=0$. Plugging these results in the approximation above and taking a first difference in expectations between $t-2$ and $t$ yields
${\operatorname*{\mathbb{E}}}_{t}\ln\sum_{s=0}^{T-t}\frac{C_{it+s}}{(1+r)^{s}}-{\operatorname*{\mathbb{E}}}_{t-2}\ln\sum_{s=0}^{T-t}\frac{C_{it+s}}{(1+r)^{s}}\approx\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1}.$
A.3 Biennial consumption growth
Given the biennial nature of our data, it follows from (A.1) that observed consumption growth is $\Delta^{2}c_{it+s}=\xi_{it+s}+\xi_{it+s-1}+\theta_{it+s-1}\epsilon_{it+s}+\theta_{it+s-2}\epsilon_{it+s-1}$. The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a difference between $t$ and $t-2$. Dropping the index $s$ and replacing the expectations errors with the quadratic expression in income shocks (the budget must balance so we bring the two sides of the budget constraint together), we obtain the final expression for consumption growth with biennial survey data, given by
$\begin{split}\Delta^{2}c_{it}&\approx\xi_{it}+\xi_{it-1}+(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}\\ &+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}\\ &+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.\end{split}$ (A.3)
The first line of (A.3) corresponds to the linear consumption function (4), adapted here for the biennial nature of the data and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}.$
The transmission parameters of income shocks are given by $\phi_{it}^{(1)}=1-\pi_{it}$ and $\psi_{it}^{(1)}=(1-\pi_{it})\vartheta_{it}^{Y}$, where $\pi_{it}$ and $\vartheta_{it}^{Y}$ were defined previously.
The quadratic terms in the next two lines of (A.3) reflect the refinement from the higher-order approximation. This higher-order approximation leads to the quadratic consumption function (5), adapted here for the biennial nature of the data and given compactly by
$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\left(\zeta_{it}+\zeta_{it-1}\right)^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\left(\zeta_{it}+\zeta_{it-1}\right)v_{it}.$
The additional transmission parameters of income shocks are given by $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\pi_{it}$, $\psi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})$, and $\omega_{it}^{(22)}=(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}$.
Appendix B Identification details
This appendix provides detailed identification statements based on our use of the recent biennial PSID data. There are two points of departure from section 3 in the paper.
First, we recast the income and consumption processes to reflect the biennial frequency of the modern PSID. Unexplained income growth is thus given by $\Delta^{2}y_{it}=\Delta y_{it}+\Delta y_{it-1}=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}$, with $\Delta y_{it}$ given by (3). The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a first difference in variable $x$ between $t$ and $t-2$, i.e., the observed frequency.
Second, we allow for classical measurement error. We let observed income $y_{it}^{*}$ and consumption $c_{it}^{*}$ be the sum of their corresponding true value and measurement error, namely
$y_{it}^{*}=y_{it}+u_{it}^{y}\qquad\text{and}\qquad c_{it}^{*}=c_{it}+u_{it}^{c}.$
The moments of income measurement error are not separately identifiable from the moments of the transitory shock (Blundell et al., 2008). This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
${\operatorname*{\mathbb{E}}}((u^{y}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{y}_{t}}^{2}&\text{for }m=2\\ 0&\text{for }m=3\\ 3(\sigma_{u^{y}_{t}}^{2})^{2}&\text{for }m=4\\ \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}((u^{c}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{c}_{t}}^{2}&\text{for }m=2\\ \gamma_{u^{c}_{t}}&\text{for }m=3\\ \kappa_{u^{c}_{t}}&\text{for }m=4.\\ \end{cases}$
We retrieve the variance of income error from the validation study of the PSID in Bound et al. (1994), as we explain in section 4. We do not need similar assumptions for the consumption error, whose moments we identify below. We further assume that income and consumption errors are mutually independent (but note discussion of consumption imputation from the CEX), independent over time, and independent of the income shocks and taste heterogeneity.
These choices lead to our processes for observed income growth, consumption growth in the linear specification, and consumption growth in the quadratic specification, given by
$\Delta^{2}y_{it}$ $=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}+\Delta^{2}u^{y}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}+\Delta^{2}u^{c}_{it},$
$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}$
$+\phi_{it}^{(2)}(\zeta_{it}+\zeta_{it-1})^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}(\zeta_{it}+\zeta_{it-1})v_{it}+\Delta^{2}u^{c}_{it},$
respectively. These are the theoretical expressions we bring to the data.
Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks that comprise a two-year period. We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distribution is stationary. Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$. For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, and $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$. These last restrictions are not needed for identification with biennial data, but they simplify the illustration.
B.1 Income process parameters
The second and third moments of shocks are then given by
$\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
$\gamma_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\gamma_{v_{t}}$ $=-{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it+2}).$
The fourth moments are more involved and are given by
$\kappa_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})-\kappa_{v_{t}}-3(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
$-12\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-12\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$
$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it+2})^{2})-4(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
$-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-8\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}.$
B.2 Linear consumption function parameters
Given the income process, the linear consumption function, and the properties of shocks, the transmission parameters are given by
$\phi^{(1)}_{t}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})/\sigma_{\zeta_{t}}^{2}\qquad\text{and}\qquad\psi^{(1)}_{t}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})/\sigma_{v_{t}}^{2}.$
The variance of taste heterogeneity and consumption measurement error are identified as
$\sigma_{\xi_{t}}^{2}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}c_{it+\kappa})-(\phi^{(1)}_{t})^{2}\sigma_{\zeta_{t}}^{2}-\frac{1}{2}(\psi^{(1)}_{t})^{2}\sigma_{v_{t}}^{2}\quad\text{and}\quad\sigma_{u^{c}_{t}}^{2}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2}).$
The third and fourth moments of taste heterogeneity and consumption error are identified by higher-order consumption moments through expressions analogous to those above.
B.3 Quadratic consumption function parameters
Given the income process, the quadratic consumption function, and the properties of income shocks, the transmission parameters are given by
$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$
where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks. [^1] With biennial data, matrix $\mathbf{A}$ is given by (different from $\mathbf{A}$ with annual data in the paper) $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
$\text{det}(\mathbf{A})$ $=$ $-8\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\Big{[}\Big{(}k_{\zeta_{t}}$ $-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{]}.$
Assuming that the shocks are not drawn from a Bernoulli distribution, kurtosis is bounded below by $\gamma^{2}/\sigma^{2}+\sigma^{4}$. It follows that
$k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}-\sigma_{\zeta_{t}}^{4}$ $>0,$ (B.1)
$k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}-\sigma_{v_{t}}^{4}$ $>0.$ (B.2)
Note also that
$6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0,$ (B.3)
$2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0.$ (B.4)
Adding the inequalities (B.1)+(B.3) and (B.2)+(B.4) by parts, it follows that $k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$ and $k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$, so the bracketed terms in the first line of the determinant are both strictly positive. It then follows that
$\Big{(}k_{\zeta_{t}}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>$
$\Big{(}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+\sigma_{\zeta_{t}}^{4}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+\sigma_{v_{t}}^{4}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
$\Big{(}6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
$4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}+4\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}+4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}$ $>0,$
where the first inequality follows from the lower bound on the kurtosis. The term inside the squared brackets in the determinant is thus strictly positive, and $\text{det}(\mathbf{A})<0$. The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli. [^2] If the distributions of shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$, so the determinant would not always be strictly away from zero.
An interesting special case arises when the distributions of permanent and transitory shocks are normal and have the same variance, and there is no measurement error. In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}<0$ (where $\sigma_{t}$ is the dispersion of the generic distribution), $\phi_{t}^{(1)}=(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$. Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of a typical permanent shock away from the mean, is separately identified from $\phi_{t}^{(1)}$. Therefore, identification does not require the shocks to be non-Gaussian. Despite normality in this case, tail shocks are still associated with a different pass-through from that of average shocks. Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
Upon identification of the transmission parameters, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2})$ identifies $\sigma_{u^{c}_{t}}^{2}$, while ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ identifies $\sigma_{\xi_{t}}^{2}$. Higher moments of the consumption error or of $\xi$ involve higher than fourth-order moments of income, which we do not model.
B.4 Targeted moments
There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model. Table B.1 lists the moments targeted in the various specifications of the income process, while table B.2 lists the moments targeted in the different specifications of the consumption model. For simplicity, we assume stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters. These are not identifying assumptions but they make the illustration below more compact by allowing us to pool common terms together.
Table B.1: Targeted moments of income
Moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$
2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$
3 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$
4 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$
5 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$
6 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$
7 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$
9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the various specifications of the income process. $\Delta^{2}y_{t}=\Delta y_{t}+\Delta y_{t-1}=y_{t}-y_{t-2}$ indicates a first difference of $y$ between periods $t$ and $t-2$.
B.4.1 Moments of income
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ $=2\sigma_{\zeta}^{2}+2\sigma_{v}^{2}+2\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ $=-\sigma_{v}^{2}-\sigma_{{u}^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ $=2\gamma_{\zeta}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ $=\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ $=-\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$ $=2\kappa_{\zeta}+2\kappa_{v}+6(\sigma_{\zeta}^{2})^{2}+6(\sigma_{v}^{2})^{2}+12(\sigma_{u^{y}}^{2})^{2}+24\sigma_{\zeta}^{2}\sigma_{v}^{2}+24\sigma_{v}^{2}\sigma_{u^{y}}^{2}+24\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$ $=\kappa_{v}+4(\sigma_{\zeta}^{2})^{2}+3(\sigma_{v}^{2})^{2}+6(\sigma_{u^{y}}^{2})^{2}+8\sigma_{\zeta}^{2}\sigma_{v}^{2}+8\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
Table B.2: Targeted joint moments of consumption and income
Cons. fn.: Linear Quadratic
Moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th (income shocks)
1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$
2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$
3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$
4 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$
5 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$
6 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times\Delta^{2}c_{it-2})$
7 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$
8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$
9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$
10 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$
11 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$
12 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$
13 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$
14 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$
15 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$
16 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$
17 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
18 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$
19 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$
20 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$
21 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$
22 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$
23 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
24 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$
25 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$
26 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$
Notes: The table lists the moments targeted in the estimation of the different specifications of the consumption process. In all cases, the notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates the first difference of the generic variable $x$ between periods $t$ and $t-2$.
B.4.2 Moments of consumption – linear function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=-\sigma_{{u}^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ $=2\gamma_{\xi}+2(\phi^{(1)})^{3}\gamma_{\zeta}+(\psi^{(1)})^{3}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\gamma_{u^{c}}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-\gamma_{u^{c}}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ $=2(\phi^{(1)})^{2}\gamma_{\zeta}+(\psi^{(1)})^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-(\psi^{(1)})^{2}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=0$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=0$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$ $=2\kappa_{\xi}+2(\phi^{(1)})^{4}\kappa_{\zeta}+(\psi^{(1)})^{4}\kappa_{v}+2\kappa_{u^{c}}+6(\sigma_{\xi}^{2})^{2}+6(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+6(\sigma_{u^{c}}^{2})^{2}$
$+24(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+12(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+24\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}$
$+12(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+24(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}$
$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+\kappa_{u^{c}}+3(\sigma_{u^{c}}^{2})^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\kappa_{\zeta}+6(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}+8\phi^{(1)}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\kappa_{\zeta}+6\phi^{(1)}(\sigma_{\zeta}^{2})^{2}+\psi^{(1)}\kappa_{v}+3\psi^{(1)}(\sigma_{v}^{2})^{2}$
$+6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$ $=2(\phi^{(1)})^{3}\kappa_{\zeta}+6(\phi^{(1)})^{3}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{3}\kappa_{v}+12\phi^{(1)}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+6\psi^{(1)}\sigma_{\xi}^{2}\sigma_{v}^{2}$
$+6\phi^{(1)}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\kappa_{v}-3\psi^{(1)}(\sigma_{v}^{2})^{2}-6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-(\psi^{(1)})^{3}\kappa_{v}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{\xi}^{2}-6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{v}^{2}\sigma_{\zeta}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
B.4.3 Moments of consumption – quadratic function
${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2(\phi^{(2)})^{2}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2})+(\psi^{(2)})^{2}\kappa_{v}$
$+2(\omega^{(22)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}+4\phi^{(1)}\phi^{(2)}\gamma_{\zeta}+2\psi^{(1)}\psi^{(2)}\gamma_{v}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=4(\phi^{(2)})^{2}(\sigma_{\zeta}^{2})^{2}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma^{2}_{v}+(\psi^{(2)})^{2}(\sigma^{2}_{v})^{2}-\sigma_{u^{c}}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}+2\phi^{(2)}\gamma_{\zeta}+\psi^{(2)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}-\psi^{(2)}\gamma_{v}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}+2\phi^{(2)}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2})$
$+\psi^{(2)}(\kappa_{v}+(\sigma_{v}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{v}^{2}\sigma_{u^{y}}^{2})+4\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}+4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}(2\sigma_{\zeta}^{2}+\sigma_{v}^{2}+2\sigma_{u^{y}}^{2}))$
${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+2\psi^{(2)}\sigma_{v}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}\sigma_{u^{y}}^{2})-2\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
Appendix C Empirical details and additional results
C.1 Comparison with administrative moments
Table C.1 reports the empirical second, third, and fourth moments of the cross-sectional distribution of unexplained growth in male earnings (panel A) and household earnings (panel B), as opposed to the moments of household disposable income in table 2 in the paper. Column 1 presents moments from our baseline sample in the PSID, while column 2 presents the corresponding statistics, wherever available, from the U.S. Social Security Administration data reported in Guvenen et al. (2021).
Table C.1: Empirical moments, male earnings and household earnings
Baseline sample Guvenen et al. (2021)
Income data: PSID Social Security Admin.
(1) (2)
Panel A. Moments of male earnings growth
${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.271 (0.010) 0.323
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.065 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.569 (0.128) -1.039
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.087 (0.010)
${\operatorname*{\mathrm{Kurt}}}(\Delta y_{it})$ 13.711 (0.551) 13.494
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},(\Delta y_{it+1})^{2})$ 0.253 (0.022)
Panel B. Moments of household earnings growth
${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.233 (0.009)
${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.056 (0.004)
${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.728 (0.142)
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.045 (0.008)
${\operatorname*{\mathrm{Kurt}}}(\Delta y_{it})$ 14.728 (0.822)
${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},(\Delta y_{it+1})^{2})$ 0.176 (0.017)
Notes: The table presents the second, third, and fourth moments of earnings growth. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of biennial earnings growth in our baseline sample in the PSID (we maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the PSID, this corresponds to a difference over two calendar years in this case), while column 2 reports moments of male earnings growth from the U.S. Social Security Administration data. The latter moments concern the one-year male earnings growth, unconditional on past earnings, and correspond to averages over ages 30-55 of, respectively, dispersion squared, skewness, and kurtosis, as reported in online appendix tables C23-C25 in Guvenen et al. (2021). Block bootstrap standard errors are in parentheses.
C.2 Sources of tail income shocks
Table C.2 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis. Columns 1 & 2 report the empirical probability that unemployment, disability, bad health, etc, occur in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively. Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) also experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.
Table C.2: Sources of tail income shocks
Household income growth
bottom 10% top 90% difference
empirical probabilities $p$-value
(1) (2) (3)
Unemployed $t$ (male) 0.082 0.034 0.000
Unemployed $t-1$ (male) 0.084 0.036 0.000
Unemployed $t$ (female) 0.081 0.037 0.000
Unemployed $t-1$ (female) 0.081 0.041 0.000
Main job laid off (male) 0.054 0.039 0.003
Second job laid off (male) 0.153 0.117 0.000
Main job laid off (female) 0.073 0.059 0.023
Second job laid off (female) 0.096 0.095 0.921
Physical disability (male) 0.130 0.103 0.001
Physical disability (female) 0.121 0.120 0.840
Physical disability (both) 0.032 0.025 0.077
Overall health got worse (male) 0.129 0.110 0.036
Overall health got worse (female) 0.121 0.104 0.052
Eviction from home 0.049 0.031 0.000
Hours growth (male) -0.309 0.031 0.000
Hours growth (female) -0.228 0.023 0.000
Notes: The table reports the empirical probability of the various events listed in the rows of the table in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$. Given the biennial nature of the PSID, $\Delta y_{it}$ corresponds to a difference over two calendar years.
C.3 CEX data and imputation in the PSID
The Consumer Expenditure Survey accounts for about $95\%$ of all household expenditures from a highly disaggregated list of consumption goods and services (see Meyer and Sullivan, 2023, for a recent overview). The CEX is run by the Census Bureau and the Bureau of Labor Statistics, and it forms the basis for the calculation of the consumer’s price index.
Sample selection and variables. To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019.333To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID. We select a sample of households that mimics closely our baseline selection in the PSID. Specifically, we select continuously married couples with the male spouse aged 30 to 65. We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the consumption imputation subsequently. Similar to the consumption measure in the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare. As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters. We assign the final observation to the calendar year to which the underlying quarterly data mostly correspond to. Our final sample consists of 31,751 observations.
Table C.3 presents a comparison of means between the PSID and CEX samples for a host of household characteristics over the period 1999–2019. The two samples are very close with respect to age, family size, number of children, race, education, region of residence, or labor market participation. Household earnings are higher in the PSID than in the CEX; BPP noted this also for the earlier period (1980–1992) and argued that it is due to the more comprehensive definition of income in the PSID. Food expenditure in the PSID is higher, on average, by about 60% than food expenditure in the CEX. This is in contrast to BPP who find that food expenditure is similar across the two surveys over 1980–1992. This results in consumption being about 25% higher in the PSID than in the CEX.
Table C.3: Comparison of sample means, PSID and CEX
1999 2009 2019
PSID CEX PSID CEX PSID CEX
Earnings 72,829 59,011 109,047 84,383 130,488 106,316
Consumption 16,115 10,631 22,317 15,454 27,315 18,592
Food 7,434 4,461 9,306 6,090 12,575 7,636
Age 44.70 45.76 46.40 47.59 46.69 48.40
Family size 3.39 3.54 3.24 3.48 3.42 3.44
# children 1.13 1.21 1.01 1.11 1.19 1.07
% white 0.93 0.88 0.91 0.86 0.91 0.84
% high school dropout 0.09 0.12 0.06 0.12 0.07 0.10
% high school graduate 0.28 0.27 0.25 0.25 0.22 0.21
% college dropout 0.62 0.62 0.69 0.63 0.71 0.69
% Northeast 0.17 0.18 0.18 0.19 0.16 0.17
% Midwest 0.32 0.24 0.30 0.25 0.31 0.21
% South 0.33 0.32 0.32 0.34 0.33 0.33
% West 0.18 0.25 0.20 0.22 0.20 0.28
% working (male) 0.89 0.88 0.91 0.93 0.90 0.92
% working (female) 0.81 0.77 0.79 0.81 0.79 0.79
Notes: The table presents a comparison of means in the PSID and CEX samples in 1999, 2009, and 2019.
Figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID (blue solid line) and CEX samples (black dotted line). Although the patterns for the variance and kurtosis are quite similar in the two surveys over time, consumption in the PSID is less volatile relative to the CEX but slightly more leptokurtic. For skewness, there is disagreement both in the trends over time and, importantly, with respect to the sign, with the PSID and CEX samples featuring positive and weakly negative skewness, respectively. Overall, these statistics (in particular the skewness statistics) suggest that consumption across otherwise similarly selected samples in the PSID and CEX exhibits notable differences; this calls into question any conclusions on the pass-through drawn from methods that use higher-order moments of consumption to identify the response to income shocks.
Consumption imputation from CEX into PSID. The previous comparison is not directly informative about the moments of consumption growth, which is the relevant variable in our model. The CEX lacks the panel dimension needed for our analysis, so it cannot be used to obtain measures of annual (or biennial) consumption growth. [^4] The CEX consists of repeated cross sections, in which households are interviewed over 4 quarters. There is also an initial training quarter, which is rarely used in empirical analyses. We thus follow BPP and impute nondurable consumption from the CEX into the PSID.
Figure C.1: Moments of log consumption in the PSID and CEX
(a) Variance of log consumption
(b) Skewness of log consumption
(c) Kurtosis of log consumption
Notes: The figure plots the moments of log consumption in the PSID (blue solid line; left axis) and the CEX (black dotted line for the true series and black dashed-dotted line for the imputed series; right axis).
The imputation involves two distinct steps. In the first step, we estimate food demand in the CEX. Specifically, we let real food expenditure be a log-linear function of the log of nondurable expenditure $c$, relative prices $\mathbf{p}$, and demographics $\mathbf{W}$, namely
$f_{it}=\beta(D_{it})c_{it}+\mathbf{p}_{t}^{\prime}\boldsymbol{\gamma}+\mathbf{W}_{it}^{\prime}\boldsymbol{\mu}+e_{it},$ (C.1)
where $f_{it}$ is the log of real annual food expenditure for household $i$ in the CEX in year $t$, $\beta$ is the budget elasticity of food demand, which we allow to shift with time and household characteristics $D_{it}$, and $e_{it}$ captures unobserved heterogeneity in food demand. [^5] Prices $\mathbf{p}$ include the relative prices of food, public transportation, utilities, medical services, and childcare. Demographics $\mathbf{W}$ include a quadratic polynomial in age and dummy variables for the number of children, family size, education, region of residence, race, and cohort. $D_{it}$ includes year dummies and dummies for education and the number of children. To address possible endogeneity of and measurement error in total expenditure, we follow an instrumental variables approach using as instruments the mean and standard deviation of household disposable income and the standard deviation of spousal wages by cohort, education, year, the former interacted with time, education, and number of children dummies. [^6] These instruments are slightly different from those used in BPP, namely the cohort-education-year specific averages of male and female hourly wages, interacted with time, education, and kids dummies. The use of their instruments over our sample period often resulted in convergence problems.
In the second step, we invert (C.1) under the assumption that food demand is monotonic in total expenditure. We can then use the inverted equation to predict nondurable expenditure $c$ for a given level of food expenditure, given prices and demographics. Since the CEX and PSID samples are representative of the same underlying population, and since food expenditure, prices, and demographics are available in both surveys, we can use the inverted (C.1) to impute consumption into the PSID. This allows us to obtain an alternative, external measure of consumption in the PSID, and thus of consumption growth.
Figure C.1 plots the variance, skewness, and kurtosis of log consumption imputed from the CEX (black dashed-dotted line). The magnitudes of the various moments are quite different between the true PSID and the imputed series, as are also the trends. Imputed consumption is more volatile than the true PSID, skewness is negative (as opposed to positive in the PSID) and peaks down in 2007–2011, kurtosis is often more than twice as high as kurtosis in the PSID and peaks upwards in 2007–2011.
Table 2 in the text reports the second and higher-moments of consumption growth in the imputed series, residualized through similar first-stage regressions as in the true data. The main qualitative features of the distributions remain similar across the true PSID data and the imputed series, although the moments of imputed consumption are several-fold accentuated relative to the true data. The imputation likely imparts substantial measurement/imputation error to consumption, which casts doubt on the reliability of the partial insurance estimates derived from imputed consumption data from the CEX.
An explanation for why the imputation inflates $\boldsymbol{\phi_{t}^{(1)}}$. Our empirical findings suggest that estimation of the linear specification using imputed consumption data inflates $\phi_{t}^{(1)}$ relative to estimation using the true PSID series. One reason behind this (though likely not the only one) is neglected correlation between measurement error in income and expenditure.
Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$. The positive correlation may be the result of, e.g., recall bias that affects similarly all monetary amounts. This is something we have not allowed for in the paper, and neither have BPP. While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors). This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification. However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. Clearly, correlation between consumption and income errors causes positive bias even in the baseline, i.e., without the use of imputed consumption, as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$. We show below that this bias is multiplicatively larger when one uses imputed consumption.
In the simplest form without covariates, imputed log consumption in the PSID is given by $c_{it}^{\text{imp}}=\kappa_{t}(f_{it}-\pi_{t})$, where $f_{it}$ is log food expenditure in the PSID, $\pi_{t}=\mathbf{p}_{t}^{\prime}\boldsymbol{\gamma}$ is an intercept that depends on prices, and $\kappa_{t}=\beta(D_{it})^{-1}$ is the inverse budget elasticity of food consumption. Suppose that $\pi_{t}=0$ for simplicity, so $c_{it}^{\text{imp}}=\kappa_{t}f_{it}$. As $\beta(D_{it})$ is typically less than 1 (e.g., Blundell et al., 2004, BPP), it follows that $\kappa_{t}>1$. Blundell et al. (2004) discuss extensively the demand estimation in the CEX, and provide an in-depth characterization of the bias induced by endogenous expenditure or non-classical measurement error in the CEX. However, they do not discuss measurement error in food expenditure in the PSID, and in particular the case when such error may correlate with error in income.
Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{c})$. We assume for simplicity that measurement error in food expenditure is the same as measurement error in nondurable consumption. This is not crucial for what follows; what is crucial is that measurement error in income varies jointly with measurement error in food expenditure in a similar way as it does with measurement error in nondurable consumption. Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error. One can show that, with imputed consumption, the moment that is used to identify $\phi_{t}^{(1)}$ in fact identify $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$. As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation in this environment inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
C.4 Comparison using the original BPP data
Table C.4 uses the original BPP data over 1980–1992 (taken from the replication package of Blundell et al., 2008, available online) and does three things.
First, we replicate BPP over the original data/period of time, using BPP’s annual growth rates of income and consumption. This is column 1 in table C.4. We obtain numerically similar results to BPP, namely $\phi^{(1)}=0.64$ while $\psi^{(1)}$ is statistically zero.
Second, we calculate biennial growth rates of income and consumption and re-estimate the linear model. This is column 2 in table C.4; this estimation is analogous to the linear function estimation in column 1 of table 4, albeit using the original BPP sample. Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B). $\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains qualitatively unaffected. With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years. Some higher frequency shocks will be muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period. In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
Third, we re-estimate the linear model assuming that the target second-order income and consumption moments are time-invariant. This differs from the original BPP exercise where the moments vary with time, but it is otherwise similar to our main exercise in the paper given that we cannot precisely estimate higher-order moments on a year-by-year basis. The results are in columns 3-4 in table C.4. Time-invariance further reduces the transmission parameter of permanent shocks ($\phi^{(1)}$ drops by another 20% to $0.404$), bringing the pass-through closer to the value we estimate in the modern data. Taking averages of moments over time attenuates the largest income-consumption co-movements; consequently, the model can fit the data using, ceteris paribus, smaller transmission parameters. This is similar to Chatterjee et al. (2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in the BPP target moments.
There are some caveats comparing the exercise in table C.4 and our results with the most recent data in table 4. First, our main exercise uses consumption data internally available in the PSID, while table C.4 uses consumption imputed from the CEX. We showed earlier that the imputation imparts substantial imputation error. Second, while the estimation in table C.4 does not correct the income moments for measurement error (in line with BPP), our main exercise in table C.4 does so (in line with Blundell et al., 2016). Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
Table C.4: Replication of BPP, original BPP data 1980-1992
Consumption fn.: Linear
Income moments: 2nd 2nd 2nd 2nd
Consumption data: CEX CEX CEX CEX
Time variability: Time-varying moments Time-invariant moments
Growth rates: annual biennial annual biennial
(original BPP)
(1) (2) (3) (4)
$\phi^{(1)}$ 0.644 0.518 0.504 0.404
(0.079) (0.068) (0.076) (0.057)
$\psi^{(1)}$ 0.024 0.117 0.043 0.158
(0.044) (0.050) (0.044) (0.051)
$\theta^{\#}$ 0.112 0 0.115 0
(0.025) (not identified) (0.026) (not identified)
Notes: The table presents the estimates of the parameters of the linear consumption function, using the original BPP data over 1980–1992 (PSID income data, consumption data imputed from the CEX). The data are available through Blundell et al. (2008)’s online replication package. Columns 1-2 allow the target second-order income and consumption moments to vary with calendar time, while columns 3-4 target said moments imposing time-invariance. Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates. In line with BPP, estimation is done via diagonally weighted GMM; asymptotic standard errors are in parentheses.
# $\theta$ is the moving average parameter of the transitory shock. BPP assume that the transitory component of income is given by $v_{it}=\epsilon_{it}+\theta\epsilon_{it-1}$, where $\epsilon_{it}$ is the transitory shock. Given their use of annual growth rates, $\theta$ is identified by the first-order income autocovariance. Our use of biennial data in this paper rules out identification of $\theta$, which we thus set to 0.
C.5 Additional results
Tables C.5 and C.6 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM). As such, these tables accompany tables 3 and 4 in the paper, which report results from equally weighted GMM. The results are qualitatively (and most often numerically) unchanged from those reported in the paper.
Table C.7 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX. As such, this table accompanies table 4 in the main text, which reports results using consumption internally available in the PSID. In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data directly available in the PSID, regardless of whether we target only second- or also higher-order moments. This echoes our earlier discussion about the imputation inflating the pass-through of shocks. We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases. Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude ($-0.418$) and statistically significant. This is counterintuitive and raises further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved. We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function. However, $\phi^{(2)}<0$ in that case with a magnitude very similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant). This is, once again, indicative of the asymmetric pass-through of negative versus positive permanent shocks. Similar to our baseline, we also find $\psi^{(2)}>0$.
Table C.5: Estimates of the income process, baseline sample, optimal weighting
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
Income data: PSID PSID PSID
(1) (2) (3)
Panel A. Permanent shocks
Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.025
(0.002) (0.002) (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 -0.004
(0.003) (0.003)
standardized -0.895 -0.997
(0.633) (0.760)
Kurtosis $\kappa_{\zeta}$ central 0.037
standardized 61.636
Panel B. Transitory shocks
Variance $\sigma_{v}^{2}$ 0.033 0.033 0.034
(0.002) (0.002) (0.002)
Skewness $\gamma_{v}$ central -0.008 -0.008
(0.003) (0.003)
standardized -1.351 -1.324
(0.517) (0.485)
Kurtosis $\kappa_{v}$ central 0.049
standardized 41.579
Notes: The table presents the estimates of the parameters of the income process, imposing stationarity over time. Column 1 targets only the second moments of income growth; column 2 targets also third moments; column 3 targets all moments up to fourth-order. Estimation is via GMM using the optimal weighting matrix; block bootstrap standard errors are in parentheses. This table accompanies table 3 in the paper, which shows the baseline results from equally weighted GMM.
Table C.6: Estimates of the consumption function, baseline sample, optimal weighting
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: PSID CEX PSID PSID PSID
(1) (2) (3) (4) (5)
$\phi^{(1)}$ 0.152 0.288 0.149 0.146 0.169
(0.028) (0.044) (0.025) (0.029) (0.030)
$\psi^{(1)}$0.006 -0.109 -0.015 -0.004 -0.018
(0.045) (0.078) (0.037) (0.041) (0.047)
$\phi^{(2)}$0.037
$\psi^{(2)}$0.007
$\omega^{(22)}$ 0.740
$\sigma^{2}_{\xi}$ 0.019 0.019 0.019 0.018 0.018
(0.001) (0.004) (0.001) (0.001) (0.001)
$\gamma_{\xi}$0.001 -0.001
(0.001) (0.001)
$\kappa_{\xi}$ 0.005
$\sigma^{2}_{u_{c}}$ 0.044 0.140 0.043 0.043 0.044
(0.002) (0.009) (0.001) (0.001) (0.001)
$\gamma_{u_{c}}$ 0.002 0.002
(0.001) (0.001)
$\kappa_{u_{c}}$ 0.012
Notes: The table presents the estimates of the parameters of the consumption function, imposing homogeneous transmission parameters over the lifecycle/in the cross-section and stationarity of taste heterogeneity and consumption measurement error. Columns 1-4 present parameter estimates in the linear function, while column 5 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. Except for column 2 where we use imputed data from the CEX, all other columns use consumption data internally available in the PSID. Estimation is via GMM using the optimal weighting matrix; block bootstrap standard errors are in parentheses. This table accompanies table 4 in the paper, which shows the baseline results from equally weighted GMM.
Table C.7: Estimates of the consumption function, baseline sample, imputed consumption
Consumption fn.: Linear Quadratic
Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
Consumption data: CEX CEX CEX CEX
(1) (2) (3) (4)
$\phi^{(1)}$ 0.288 0.339 0.404 0.260
(0.044) (0.074) (0.118) (0.043)
$\psi^{(1)}$0.109 -0.155 -0.418 -0.108
(0.078) (0.098) (0.247) (0.086)
$\phi^{(2)}$0.036
$\psi^{(2)}$ 0.025
$\omega^{(22)}$0.157
$\sigma^{2}_{\xi}$ 0.019 0.018 0.000 0.019
(0.004) (0.004) (0.004) (0.004)
$\gamma_{\xi}$0.016 -0.016
(0.007) (0.007)
$\kappa_{\xi}$ 0.000
$\sigma^{2}_{u_{c}}$ 0.140 0.140 0.152 0.140
(0.009) (0.009) (0.012) (0.009)
$\gamma_{u_{c}}$0.185 -0.185
(0.037) (0.037)
$\kappa_{u_{c}}$ 0.816
Notes: The table presents the estimates of the parameters of the consumption function, imposing homogeneous transmission parameters over the lifecycle/in the cross-section and stationarity of taste heterogeneity and consumption measurement error. Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table. All columns use consumption imputed from the CEX (see appendix C.3). Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. This table accompanies table 4 in the paper, which shows the baseline results using consumption internally available in the PSID.
Table C.8 presents estimation results for the linear consumption function in subsamples formed by wealth and education. As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification. Columns 1-4 present parameter estimates among low and high wealth households. $\phi^{(1)}$ is consistently smaller among the wealthier, regardless of which set of moments are targeted, indicating the self-insurance role of assets in that group. $\psi^{(1)}$ is statistically indistinguishable from zero in both groups. Columns 5-8 present parameter estimates among households with and without college education. The results do not present a clear pattern across groups. When only second moments of income and consumption are targeted, $\phi^{(1)}$ is larger for households without college education – $0.225$ versus $0.123$. However, when all moments up to fourth-order are included, $\phi^{(1)}$ is almost identical in the two groups.
Table C.8: Estimates of the linear consumption function, by wealth and education
Consumption fn.: Linear
By wealth By education
Subsample: Low wealth High wealth No college Some college
Income moments: 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th
Consumption data: PSID PSID PSID PSID PSID PSID PSID PSID
(1) (2) (3) (4) (5) (6) (7) (8)
$\phi^{(1)}$ 0.223 0.148 0.111 0.099 0.225 0.107 0.123 0.113
(0.040) (0.045) (0.037) (0.055) (0.055) (0.086) (0.033) (0.042)
$\psi^{(1)}$0.021 0.047 0.011 0.018 -0.070 -0.041 0.024 0.031
(0.066) (0.062) (0.052) (0.076) (0.074) (0.086) (0.050) (0.067)
$\sigma^{2}_{\xi}$ 0.014 0.015 0.023 0.023 0.014 0.016 0.021 0.022
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002) (0.001) (0.001)
$\gamma_{\xi}$0.002 $0$0.002 0.000
(0.001) (0.001) (0.001) (0.001)
$\kappa_{\xi}$ 0.003 0.007 0.003 0.006
(0.001) (0.001) (0.002) (0.001)
$\sigma^{2}_{u_{c}}$ 0.048 0.048 0.042 0.042 0.043 0.043 0.044 0.044
(0.002) (0.002) (0.002) (0.002) (0.003) (0.003) (0.002) (0.002)
$\gamma_{u_{c}}$ 0.003 0.003 0.003 0.002
(0.002) (0.001) (0.002) (0.001)
$\kappa_{u_{c}}$ 0.015 0.012 0.015 0.012
(0.002) (0.002) (0.003) (0.001)
Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary across wealth and education groups. This table accompanies the quadratic specification estimates in table 6 in the paper. For reasons of brevity, we do not report results when second- and third-order moments of income and consumption are targeted in the estimation; these are available from the authors upon request. Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019. No college versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
Appendix D Welfare costs details
Table D.1 presents the parameter estimates and the model fit of the quantitative model for income shocks (mixture of normal distributions). Table D.2 reports average welfare costs for different values of the CRRA coefficient $\gamma$. Table D.3 calculates the welfare costs respecting the age-specific nature of the partial insurance parameters.
Table D.1: Normal mixtures, parameter estimates and model fit
Panel A. Parameter estimates
Permanent $p_{\zeta}$ $\mu_{\zeta,1}$ $\sigma_{\zeta,1}$ $\mu_{\zeta,2}$ $\sigma_{\zeta,2}$
0.058 -0.049 0.683 0.003 0.053
Transitory $p_{v}$ $\mu_{v,1}$ $\sigma_{v,1}$ $\mu_{v,2}$ $\sigma_{v,2}$
0.059 -0.093 0.713 0.006 0.020
Panel B. Model fit permanent shocks
Model Data
Variance $\sigma_{\zeta}^{2}$ 0.030 0.030
Skewness $\gamma_{\zeta}$ central -0.004 -0.004
standardized -0.757 -0.831
Kurtosis $\kappa_{\zeta}$ central 0.038 0.039
standardized 43.146 44.393
Panel C. Model fit transitory shocks
Model Data
Variance $\sigma_{v}^{2}$ 0.031 0.031
Skewness $\gamma_{v}$ central -0.008 -0.008
standardized -1.520 -1.431
Kurtosis $\kappa_{v}$ central 0.047 0.048
standardized 49.630 49.226
Notes: The table reports the parameter estimates (panel A) and the model fit (panels B & C) of the quantitative model for income shocks (mixture of normal distributions). Estimation is done via the method of simulated moments, based on 50,000 random draws for each type of income shock.
Table D.2: Welfare costs of non-Gaussian relative to Gaussian income risk, varying $\gamma$
Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
Panel A. $\boldsymbol{\gamma=1}$
All households 100% -0.86 -0.95 0.64
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.93 6.46 7.94
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.50 8.71 10.15
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.69 11.90 13.29
Panel B. $\boldsymbol{\gamma=3}$
All households 100% -0.72 -0.77 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.55 6.15 7.48
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.05 8.36 9.66
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.16 11.57 12.83
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B). We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 4. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific transmission parameters
Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
% households# (1) (2) (3)
All households 100% -1.13 -1.15 -0.24
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.27 5.66 6.51
$\min\{\zeta_{it}\}<-0.70$ 29.11% 5.72 7.71 8.54
$\min\{\zeta_{it}\}<-0.95$ 17.07% 7.85 10.78 11.59
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using the age-specific transmission parameters of table 5. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.

### v4

@@
 Abstract
-We measure the extent of consumption insurance to income shocks accounting for high-order moments of the income distribution.
-We derive a nonlinear consumption function, in which the extent of insurance varies with the sign and magnitude of income shocks.
-Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption compared to 9% of an equal-sized positive shock – and the pass-through increases as the shock worsens.
-We find similar patterns for transitory shocks among the least wealthy.
-Households are willing to sacrifice more than 1/8 of lifetime consumption to eliminate tail income risk.
-Our results are consistent with surveys of consumption responses to hypothetical events and suggest that tail risk matters substantially for consumption.
+We measure consumption insurance against income shocks accounting for high-order moments of the income distribution.
+We derive a nonlinear consumption function, where the extent of insurance varies with the shocks’ sign and magnitude.
+Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen.
+Households would sacrifice over 1/8 of lifetime consumption to eliminate tail income risk.
+Our results align with survey responses to hypothetical events and suggest that tail risk matters substantially for consumption.
 Keywords: Income risk, skewness, kurtosis, consumption, partial insurance, PSID
-JEL Classification: D12, D15, D31, E21
+JEL Classification: D15, D31, E21
 1 Introduction
@@
 In a seminal paper, Blundell et al.
-(2008) –henceforth, BPP– introduce an empirical methodology to measure the pass-through of income shocks of varying persistence into consumption.
+(2008) – henceforth BPP – introduce an empirical methodology to measure the pass-through of income shocks of varying persistence into consumption.
 Using Panel Study of Income Dynamics (PSID) and Consumer Expenditure Survey (CEX) data over 1980–1992, BPP find evidence of partial insurance to permanent shocks and full insurance to transitory ones.
@@
 (2021) use data from the U.S.
-Social Security Administration and establish two new important facts about the distribution of unexplained income growth, namely the distribution of income shocks.
-First, income growth is very negatively skewed, i.e., it has a long left tail, so there are far more people in the data who experience large negative than large positive shocks.
-Second, income growth exhibits excess kurtosis, i.e., fat tails, so there are far more people in the data who face either small or large income shocks, than people who face moderate ones, relative to a Gaussian density.
+Social Security Administration and establish important new facts about the distribution of unexplained income growth, namely the distribution of income shocks.
+They show that income growth is very negatively skewed, i.e., it has a long left tail, so far more people experience large negative than large positive shocks.
+Moreover, income growth exhibits excess kurtosis, i.e., fat tails, so far more people face either small or large shocks, than moderate ones, relative to a Gaussian density.
 These are now benchmark facts about income dynamics, observed in many countries (e.g., De Nardi et al., 2021, compare the U.S.
 and the Netherlands).
-If most shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks.
-In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution, respectively, may matter for the pass-through of shocks into consumption.
+If most income shocks away from the mean are large and bad shocks, their effect on consumption might, intuitively, differ from that of small/moderate shocks.
+In other words, left skewness and excess kurtosis, which are statements about the third and fourth moments of the income distribution respectively, may matter for the pass-through of shocks into consumption.
 This is precisely what we investigate in this paper.
 Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks).
-This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution.
-The second moment is a poor summary of the tails and, thus, abstracts from the modern higher-order facts that characterize income dynamics.
+This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution and, thus, abstracts from the modern higher-order facts that characterize income dynamics (the second moment is a poor summary of the tails).
 We offer four main contributions:
@@
 We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks.
-We characterize and empirically measure the consumption pass-through of income shocks; in this environment, the degree of partial insurance depends on the magnitude and sign of shocks.
+We characterize and empirically measure the consumption pass-through of income shocks, which, in this environment, depends explicitly on the magnitude and sign of shocks.
 While a nonlinear consumption function has been estimated before, most notably in Arellano et al.
-(2017), identification in our case is far less data demanding relative to them: we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution.
-This has certain important advantages, which we discuss below.
+(2017) – henceforth ABB, our framework is far less data demanding (we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution) and offers a more transparent link with the higher-order features of the distribution of income shocks.
+Both points have certain important advantages, which we discuss below.
 Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk.
-De Nardi et al.
+We use parametric utility to calculate these costs, in line with De Nardi et al.
 (2019), Busch and Ludwig (2024), and Guvenen et al.
-(2024) provide similar calculations through structural models that fully specify the insurance channels at the nexus of income and consumption.
-By contrast, we remain agnostic about specific insurance mechanisms; we estimate an ‘aggregate’ of all partial insurance available to households, and then use a parametric utility function to estimate the welfare costs of tail risk.
+(2024), but we base these calculations on the pass-through measure from the quadratic consumption function of the previous step.
+This measures an ‘aggregate’ of all partial insurance available to households, so it is not specific to a particular structural mechanism or a particular parameterization of preferences and expectations.
 Our baseline empirical implementation uses income and consumption panel data from the PSID over the years 1999–2019.
@@
 Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad.
+However, these features vary by age and by the household’s position in the income distribution.
 Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP.
 A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP.
-In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using the recent PSID data but a different method, Arellano et al.
-(2023) also find a low pass-through of permanent shocks.
-We discuss this point further when we present our empirical results.
-We attribute one third of the difference from BPP to the consumption imputation from the CEX that BPP use.
+In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al.
+(2023) find a similarly low pass-through of permanent shocks.
+We discuss this further when we present our empirical results.
+We attribute about a third of the difference from BPP to the consumption imputation from the CEX that BPP use.
 The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth.
-Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute higher frequency shocks and their transmission into consumption.
+Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute the transmission of higher frequency shocks.
 The last third reflects the different calendar times.
 Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance.
-This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare implications of tail risk (e.g., Barro, 2009; De Nardi et al., 2019).
+This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare costs of tail risk (e.g., Barro, 2009; De Nardi et al., 2019).
 One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find.
@@
 In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters.
-We draw three main conclusions.
-First, bad permanent shocks are more hurtful than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$.
-Second, the pass-through of bad shocks increases with the severity of the shock.
-Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also have increasingly larger transmission rates as their magnitude worsens.
-Third, transitory shocks provide mixed evidence.
-They are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021).
-Among low wealth households, however, transitory shocks shift consumption significantly, and negative shocks pass through at increasingly larger rates than positive ones.
+We draw three main conclusions: 1.) Bad permanent shocks are more unsettling than equal-sized good ones because they have larger transmission rates into consumption – a three standard deviations negative shock (a 50% permanent income cut) has a pass-through of $0.17$ (17% of the shock transmits into consumption), while a positive shock of similar magnitude has a pass-through of $0.09$.
+2.) The pass-through of bad shocks increases with the severity of the shock.
+Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also transmit increasingly more as their magnitude worsens.
+3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021).
+Among low wealth low educated households, however, transitory shocks, and in particular negative ones, shift consumption significantly.
 Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one.
-As there is clearly a limit to how much one can dissave or borrow, a large negative shock ends up affecting consumption by relatively more.
-The asymmetric transmission of bad versus good permanent shocks intensifies with the age of the household.
+There is clearly a limit to how much one can dissave or borrow, so a large negative shock ends up affecting consumption by relatively more.
+The asymmetric transmission of bad versus good permanent shocks varies with age and the household’s position in the income distribution.
 While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks.
-Severely bad permanent shocks thus cause more harm among older than younger households.
+Severely bad permanent shocks thus cause more harm among the old.
+Moreover, bad permanent shocks from a given position in the bottom tail of the distribution of income risk cause more harm among the rich (households in the top of the income distribution) than among the poor (households in the bottom); the converse is true for good shocks.
 Guvenen et al.
-(2021) show that income skewness becomes more negative with the age of the earner – there is more room for income to fall as older workers have higher incomes; we find a similar pattern in our data too.
-In turn, we document that there is also more room for consumption to fall, so bad shocks are double hurtful among the old.
-Overall, the results from our quadratic specification are indicative of limited insurance to left tail income risk.
-Moreover, if the econometrician assumed a linear consumption function, she would overestimate the true extent of insurance to it, i.e., she would infer excessive insurance.
+(2021) show that income skewness becomes more negative with age and income level – there is more room for income to fall among older or higher income workers.
+In turn, we document that there is also more room for consumption to fall.
+Among the old, this is because both the rate of transmission and the severity of downwards income risk increase with age; among the rich, however, the asymmetric transmission is driven mostly by the increasing severity of downwards risk along the income distribution.
 […]
@@
 The seminal paper in the first branch, namely BPP, sparked a large volume of research.
-[^3] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical distribution of consumption and income, so they are early precursors in this literature.
+[^4] Hall and Mishkin (1982), Deaton and Paxson (1994) and Blundell and Preston (1998) characterize the empirical joint distribution of consumption and income, so they are early precursors in this literature.
 Blundell et al.
-(2016) extend BPP to an environment with wage risk to measure the contribution of family labor supply to partial insurance.
-Theloudis (2021) extends this further to allow for household heterogeneity, which he estimates using information in the third moments.
+(2016) measure partial insurance against wage risk, allowing for family labor supply.
+Theloudis (2021) allows for preference heterogeneity in partial insurance, which he estimates using information in the third moments.
 Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find.
-Hryshko and Manovskii (2022) carry out an interesting subsample analysis, identifying sets of households in the PSID with different degrees of insurance based on certain features of their incomes.
-Crawley and Kuchler (2023) offer another empirical extension, addressing neglected time aggregation in income and consumption in the PSID.
+Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes.
+Crawley and Kuchler (2023) address neglected time aggregation in income and consumption in the PSID.
 The previous papers restrict income shocks to transmit linearly into consumption.
-Arellano et al.
-(2017); Arellano et al.
-(2023) generalize the income and consumption process by letting permanent shocks feature nonlinear persistence and nonlinear transmission into consumption.
-[^4] The latter paper advances the former by allowing for flexible heterogeneity and unbalanced panels.
-Using recent PSID data, they measure the consumption response to income at about 0.3 on average (parameter closest to our definition of pass-through; specification with heterogeneity in the second paper), and the response varies over the distribution of shocks with evidence of asymmetries similar to ours.
-Identification is quite demanding as they require observing the entire consumption-income joint distribution; our case, by contrast, is far simpler: we need second moments of consumption and up to fourth moments of income.
-This is appealing in itself (we use more accessible information) and because, as we show later, higher-order features of the consumption distribution exhibit notable discrepancies across survey data, calling into question the conclusions drawn from them.
-The price of our econometrically more accessible approach is that we have to maintain a linear income process (though one that is subject to tail shocks), while Arellano et al.
-(2017) do not.
-Yet, we learn that nonlinearity in income is not crucial for obtaining asymmetric responses to income shocks.
+ABB and Arellano et al.
+(2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission.
+Using the recent PSID and a quantile regression approach, they measure the marginal propensity to consume from persistent income, which varies along the income distribution.
+They simulate the transmission of shocks and document asymmetries between bad/good shocks or between rich/poor, similar to ours.
+Our papers, however, differ in important dimensions.
+First, identification in ABB is quite demanding as they require observing the entire consumption-income joint distribution, while we need only second moments of consumption and up to fourth moments of income.
+This is appealing because we use more accessible information and because, unlike income, we cannot assess the higher-order features of consumption against an established population benchmark.
+The price of our econometrically more accessible approach is that, unlike ABB, we have to maintain a linear income process (but subject to tail shocks).
+Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
+Second, ABB measure the pass-through over quantiles of the distribution of income shocks, without explicitly studying the higher-order features of their distribution.
+We show subsequently that seemingly different quantiles, e.g., shocks in the 10th or 90th percentile as in ABB, may in fact be very similar to one another when the underlying distribution is skewed and leptokurtic.
+Our model allows us to estimate pass-through rates that explicitly shift with the sign and size of income shocks and can thus describe more transparently how shocks of different signs/sizes shift consumption.
 The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models.
 [^5] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand.
-The early works focus on self-insurance, providing mixed evidence for how much insurance the model permits.
-Kaplan and Violante (2010) find less insurance compared to BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance.
-More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durable purchases (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021).
-Except Madera (2019), these papers abstract from higher income moments due to the computational difficulties involved.
+Early works focus on self-insurance.
+Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance.
+More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021).
+Except Madera (2019), these papers abstract from higher income moments due to computational difficulties.
 De Nardi et al.
-(2019) embed a non-Gaussian income process into a model of consumption/savings and show that tail income risk increases partial insurance due to stronger precautionary motives.
-In a similar model, Busch and Ludwig (2021) target income skewness and kurtosis, they distinguish between good and bad shocks, and they show that the latter are worse insured.
+(2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk increases partial insurance due to stronger precautionary motives.
+In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis; they distinguish between good and bad shocks and show that the latter are worse insured.
 Guvenen et al.
-(2024) find that tail shocks are worse insured than average shocks, but the benchmark methods of measuring this, effectively BPP, overstate the true extent of insurance.
-Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; this enables us to measure the overall degree of insurance from all mechanisms together keeping our model partially unspecified, while accounting for higher-order features of the income distribution.
+(2024) find that tail shocks are worse insured than average shocks but the benchmark method in BPP overstates the true extent of insurance.
+Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; we measure aggregate insurance from all mechanisms together, while accounting for higher-order features of the income distribution, keeping our model parametrically unspecified.
 The asymmetric response to bad and good shocks is recently also found in surveys that elicit responses to hypothetical events.
@@
 [^6] Given our finding of heterogeneous responses to shocks, our paper is also broadly related to the literature that studies MPC heterogeneity (e.g., Lewis et al., 2022; Crawley and Kuchler, 2023).
-The paper is also related to the newer income dynamics literature that examines the non-Gaussian features of the income distribution.
-Geweke and Keane (2000) and Bonhomme and Robin (2010) are early works that use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth.
+The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects.
+Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth.
 Guvenen et al.
-(2021) use data from the population of working males in the U.S.
-to provide further evidence of left skewness and excess kurtosis in earnings growth, both accentuated with age and the level of earnings.
+(2021) use administrative data in the U.S.
+to provide further evidence of left skewness and excess kurtosis in earnings growth, accentuated with age and the level of earnings.
 Guvenen et al.
@@
 While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption.
-We thus contribute to a growing literature on the various economics/finance implications of tail risk, e.g.
-for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
-In the remainder of the paper, section 2 presents the model while section 3 establishes identification of the parameters.
-Section 4 discusses the empirical implementation, section 5 presents the empirical results, and section 6 discusses welfare.
+We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
+In the remainder of the paper, section 2 presents the model and section 3 establishes identification of the parameters.
+Section 4 discusses the empirical implementation, section 5 presents the results, and section 6 discusses welfare.
 Section 7 concludes.
-The appendix provides theoretical derivations and additional results.
+The supplemental appendix provides additional derivations and results.
 2 Lifecycle model with tail income risk
@@
 $A_{i0}+{\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{Y_{it}}{(1+r)^{t}}={\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{C_{it}}{(1+r)^{t}},$ (2)
-where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y$, and $T$ is the end of the horizon.
+where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon.
 We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.
@@
 $\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction.
-$\gamma$ denotes the skewness/asymmetry of the distributions, i.e., the relative lengths of the upper and lower tails.
-Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distributions to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
+$\gamma$ denotes the skewness/asymmetry of the distribution, i.e., the relative length of the upper and lower tails.
+Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
 We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al.
 (2021) provide some evidence for that.
-Otherwise, shocks across households are identically distributed within $t$.
+Otherwise, shocks across households are identically distributed given $t$.
 Permanent and transitory shocks are mutually independent and independent over time.
@@
 The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things.
+[^8] In addition, households in Blundell et al.
+(2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences.
 The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure.
-[^8] In addition, households in Blundell et al.
-(2016) have access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences; both are reasons why the transmission parameters may vary with $i$ or $t$.
 We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A.
@@
 These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter.
+[^10] We do not go beyond a quadratic approximation for reasons of tractability and because higher-order terms in a regression of $\Delta c_{it}$ on a polynomial in $\Delta y_{it}$ are imprecisely estimated.
 […]
@@
 The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response.
-The size of this bias depends on the magnitude of $\phi^{(2)}$ (a more negative $\phi^{(2)}$ implies bigger bias) and on the size of the permanent shock itself (a more severe bad shock implies bigger bias).
+The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the permanent shock itself.
 When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
-A similar asymmetry exists in the transmission of good versus bad transitory income shocks. In this case, the pass-through into consumption is given by
+A similar asymmetry exists in the transmission of good versus bad transitory shocks. In this case, the pass-through into consumption is given by
 $\frac{\partial\Delta c_{it}}{\partial v_{it}}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
@@
 the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$).
-We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups.
+We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the distribution of income.
 For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth.
@@
 The covariance of two consecutive income growths in the second line identifies the variance of the transitory shocks due to mean reversion in the shock.
-In a similar way, the skewness of shocks is identified as
+In a similar way, skewness is identified from
 $\gamma_{\zeta_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1})),$
@@
 $\kappa_{\zeta_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{4})-\kappa_{v_{t}}-\kappa_{v_{t-1}}-6\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})-6\sigma_{v_{t}}^{2}\sigma_{v_{t-1}}^{2},$
-$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}$
-$-\sigma_{v_{t-1}}^{2}\sigma_{v_{t}}^{2}-\sigma_{v_{t-1}}^{2}\sigma_{v_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
+$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}$
+$-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
 There are several over-identifying restrictions for both skewness and kurtosis.
@@
 This depends not only on the variance of permanent and transitory income shocks, as in the linear case, but also on their skewness and kurtosis.
-Contrary to the linear case, however, the higher-order moments of consumption growth depend on higher than fourth-order moments of the shocks, which we do not model.
+Contrary to the linear case, however, the higher-order moments ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{3})$ and ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{4})$ depend on higher than fourth-order moments of the shocks, which we do not model.
 Identification of the partial insurance parameters comes from the joint moments of consumption and income, conceptually similar to identification in the linear case.
@@
 Assuming away Bernoulli distributions is thus an identifying assumption.
-The matrix is invertible even if the distributions of shocks are normal and have the same variance between them.
-In the nonlinear specification, tail shocks are associated with a potentially different pass-through from that of average shocks; but identification of such pass-through does not require those shocks to be non-Gaussian.
-As seen in (8) and matrix $\mathbf{A}$, the transmission parameters can no longer be identified with second-order income moments alone.
-Higher-order moments of income play an essential role in identification, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails.
-By contrast, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth.
-This is an advantage over methods that require observing the full consumption distribution (e.g., Arellano et al., 2017, need the entire consumption-income joint distribution) because, while in principle we may estimate higher moments of consumption growth in the PSID, we cannot be certain how reliable these estimates are.
+Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
+In the nonlinear specification, tail shocks are associated with a different pass-through from that of average shocks, regardless of the precise nature of the underlying distribution of shocks.
+The transmission parameters, however, can no longer be identified with second-order income moments alone – inspect matrix $\mathbf{A}$ in (8).
+Skewness and kurtosis, both defined also in the Gaussian case, are crucial for identifying the pass-through of such shocks, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails.
+In appendix B, we investigate what different levels of income skewness and kurtosis imply for the transmission parameters.
+Contrary to income, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth.
+This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is no established population benchmark to assess them against.
 This is in contrast to higher income moments, for which Guvenen et al.
-(2021) is a natural benchmark given their use of administrative data.
-We return to this point in section 4.
+(2021) is a natural benchmark given their use of administrative population data.
 How biased are the parameter estimates if one assumes a linear specification when the true consumption function is quadratic?
@@
 If $\gamma_{v_{t}}<0$ and $\psi_{t}^{(2)}>0$, as our baseline estimates suggest, linearity biases $\psi_{t}^{(1)}$ downwards.
-This helps explain the discrepancy between BPP (who find little pass-through of transitory shocks) and studies based on natural experiments (who find large consumption responses to said shocks).
-[^17] The discrepancy is between studies that use covariance restrictions in the joint distribution of income and consumption (they find small pass-through of transitory shocks) and studies that employ natural experiments (e.g., tax refunds or lotteries; they find large pass-through).
-Commault (2022) explains the discrepancy on the basis of neglected effects from past transitory shocks on consumption, while Crawley and Kuchler (2023) explain it on the basis of neglected time aggregation in the data.
-Both are alternative types of misspecification in the linear consumption function/empirical implementation vis-à-vis our focus on nonlinearities here.
+This may help bridge the discrepancy between covariance based methods (who find little pass-through of transitory shocks) and studies based on natural experiments (who usually find large consumption responses to said shocks).
+Our estimation procedure, which we describe in the next section, is linked transparently to our identification argument.
+Effectively, we bring (8) to the data, along with some over-identifying restrictions, using moments estimation.
 4 Empirical implementation
@@
 households since $1968$.
-For most of this period, the PSID collected consumption expenditures only on food items (food representing a declining fraction of the consumer’s basket over time).
+For most of this period, the PSID collected expenditures only on food items.
 Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016).
-We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when writing this paper; the data are available only biennially.
+We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when we started working on this paper; the data are available only biennially.
 This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
@@
 We focus on a sample that is representative of the U.S.
-population, eliminating the supplementary low-income subsample (SEO).
+population, eliminating the supplementary low-income subsample.
 Households are also removed in the event of a break-up or remarriage; our analysis thus focuses on income risk rather than on divorce or other household dissolution factors.
 Naturally, we require non-missing data on income, expenditure, and basic demographics.
-[^18] We also remove observations with extreme jumps in income or consumption in consecutive years (jump: $\geq$10-fold increase, or $\geq$90reduction, or the bottom 0.25% of the distribution of products of growth rates between $t$ and $t-1$), i.e., large increases followed by large drops, likely reflecting measurement error.
+[^18] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90reduction), which likely also reflect measurement error.
+Our results are not sensitive to this sample selection choice.
 Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
-Table 1: Descriptive statistics
-Baseline sample Low wealth High wealth No college Some college
-Mean Median Mean Mean Mean Mean
-(1) (2) (3) (4) (5) (6)
-Consumption
-Nondurable+services 26,766 22,895 23,106 29,852 22,137 28,908
-Food at home 8,256 7,609 7,861 8,580 7,462 8,630
-Utilities 3,515 3,238 3,270 3,721 3,394 3,571
-Food out 3,318 2,511 2,683 3,840 2,649 3,634
-Health 1,614 863 1,390 1,803 1,429 1,699
-Transportation 5,866 4,900 5,398 6,261 5,436 6,065
-Education 3,202 0 1,529 4,614 1,262 4,100
-Childcare 977 0 937 1,009 543 1,181
-Rent (or rent eq.) 18,105 13,821 10,949 23,991 11,647 21,149
-Home insurance 894 747 654 1,092 702 985
-Income
-Disposable income 142,910 113,581 98,527 179,416 97,675 164,239
-Taxable+transfer inc. 137,277 109,003 93,358 173,401 90,810 159,187
-Earnings 118,423 97,154 87,005 144,266 78,565 137,217
-Male 81,857 62,185 56,817 102,452 50,199 96,783
-Female 36,567 29,183 30,187 41,814 28,366 40,433
-Wealth 547,761 191,193 69,731 950,823 284,036 669,788
-Demographics
-% working (male) 0.89 1.00 0.92 0.88 0.86 0.91
-% working (female) 0.79 1.00 0.81 0.78 0.79 0.79
-Age (male) 46 45 43 49 46 46
-% some college (male) 0.68 1.00 0.55 0.79 0.00 1.00
-% some college (female) 0.71 1.00 0.62 0.78 0.44 0.84
-Periods per household 5.63 5.00 4.89 6.45 5.21 5.86
-Observations 20,866 9,417 11,449 6,686 14,180
-Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 19991999199919992019201920192019.
-Monetary amounts are expressed in \$2018.
-Rent for owners is calculated at 6% of the self-reported value of the house.
-Table 1 presents summary statistics over the sample period 19991999199919992019201920192019.
+Appendix table C.1 presents summary statistics over the sample period 19991999199919992019201920192019.
 About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education.
@@
 Housing, including rent and home insurance, is the main expenditure we exclude; we do so for comparability with BPP who also exclude it.
-Including housing does not change any of the subsequent results meaningfully.
-Finally, table 1 reports summary statistics also for samples sorted by wealth and education; we postpone the discussion of those samples to section 5.4.
+Including housing does not meaningfully change the subsequent results.
 Aside from the PSID, we use a similarly selected sample drawn from the CEX to facilitate comparison with BPP. We defer the discussion of this secondary data to section 4.3.
 4.2 Estimation
-Estimation follows a three-step approach.
-In the first step, we compute income and consumption growth net of their predictable components, respectively, $\mathbf{X}_{it}$ and $\mathbf{Z}_{it}$ in model notation, by regressing real income and consumption growth on a rich set of time effects and household characteristics.
+Estimation proceeds as follows.
+In a first step, we compute income and consumption growth net of their predictable components, $\mathbf{X}_{it}$ and $\mathbf{Z}_{it}$ in model notation, by regressing real income and consumption growth on a rich set of time effects and household characteristics.
 [^20] The controls include year and year-of-birth dummies, dummies for education, race, region, family size and number of children (level and change from last period), employment status (current and past), presence of outside dependents, and presence of income recipients other than the husband and wife (both current and past).
 The dummies for education, race, and region are interacted with the year dummies.
-The residuals from these regressions constitute the empirical counterparts to unexplained income and consumption growth, $\Delta y_{it}$ and $\Delta c_{it}$, respectively.
-In the second step, we estimate the parameters of the income process – i.e., the second and higher moments of permanent and transitory shocks – using income data alone.
-In the third step, we estimate the transmission parameters of shocks and the moments of taste heterogeneity and consumption measurement error; we do so using income and consumption data conditional on the parameters of the income process from the second stage.
+The residuals from these regressions are the empirical counterparts to unexplained income and consumption growth, $\Delta y_{it}$ and $\Delta c_{it}$, respectively.
+In a second step, we estimate the parameters of the income process – i.e., the second and higher moments of permanent and transitory shocks – using income data alone.
+In a third step, we estimate the transmission parameters of shocks and the moments of taste heterogeneity and consumption measurement error.
 We estimate the second- and third-step parameters using GMM.
@@
 Appendix B.4 lists the moments targeted in each case.
-Given the multi-step approach, we conduct inference using the block bootstrap jointly over all stages.
-We thus account for serial correlation and heteroskedasticity of arbitrary form, as well as for our use of pre-estimated parameters in the later stages of the multi-step approach.
-To avoid the standard errors being affected by extreme draws, we apply a normal approximation to the inter-quartile range of bootstrap replications.
-[^22] Instead of estimating standard errors as the standard deviation of the parameters’ point estimates across replications, we compute them as the normalized interquartile range of the replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf.
-This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws.
-See Blundell et al.
-(2016) and Theloudis (2021) for details.
-Standard errors without the normal approximation are very similar.
+Unless noted otherwise, we conduct inference for the main parameters using block bootstrap.
+We thus account for serial correlation and heteroskedasticity of arbitrary form and for parameter uncertainty from earlier steps of the estimation process.
+GMM (asymptotic) standard errors clustered at the household level are very similar, suggesting that earlier-step parameter uncertainty is negligible (the income parameters are tightly estimated).
+[^22] We compute bootstrap standard errors as the normalized interquartile range of bootstrap replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf.
+This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws (Blundell et al., 2016; Theloudis, 2021).
+Bootstrap standard errors without this approximation are similar.
+For the subsample analysis (tables 5-6 and corresponding tables in the appendix), we calculate GMM (asymptotic) standard errors clustered at the household level to avoid numerical instability in bootstrap, driven by vanishing shares of certain subsamples, e.g., the wealthy non-college graduates, in many bootstrap draws.
+For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors, as in BPP.
 As we explained in section 3, we must use an external estimate for the variance of income measurement error.
@@
 Table 1 reports the empirical second, third, and fourth moments of the cross-sectional distributions of unexplained income and consumption growth in the sample.
+[^24] In contrast to the identifying statements of section 3, the data are biennial so we report moments for income/consumption growth at a biennial frequency.
+We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that this corresponds to a difference over two years.
+See appendix B for details.
 Income moments.
@@
 The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant.
-[^24] In contrast to the identifying statements of section 3, the data are biennial so we report moments for income/consumption growth at a biennial frequency.
-We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that this corresponds to a difference over two years.
-See appendix B for details.
 The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data.
@@
 Table 1: Empirical moments of income and consumption growth
-Baseline sample
 Income data: PSID PSID
@@
 ${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta c_{it})$ 0.008 (0.001) 0.012 (0.002)
-Notes: The table presents the second, third, and fourth moments of income and consumption growth in the baseline sample.
+Notes: The table presents the second, third, and fourth moments of income and consumption growth.
 Skewness and kurtosis correspond to the third and fourth standardized moments respectively.
@@
 To facilitate the comparison, appendix C.2 reports moments of male earnings growth in our sample.
-The variance is very similar between the PSID and the U.S.
-Social Security Administration data, at $0.271$ and $0.323$ respectively.
+The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively.
 Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data.
@@
 (2021) confirm this.
-Kurtosis, finally, is identical between the two, at $13.711$ and $13.494$ respectively.
+Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
 Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B).
 Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions).
-These results corroborate (and, in fact, are very close numerically with) findings by De Nardi et al.
+These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al.
 (2019); De Nardi et al.
 (2021) over an earlier period of the PSID.
-Large income reductions do not just reflect unemployment, on which there is an abundant existing literature, but a multitude of events.
-Issues of reverse causality aside, appendix C.2 documents that households with large income reductions also experience the loss of a second job, worsening health, large drops in hours worked (possibly reflecting both demand and supply factors), and precarious housing.
+Large income reductions do not just reflect unemployment, but a multitude of events.
+Issues of reverse causality aside, appendix C.3 documents that households with large income reductions also experience the loss of a second job, worsening health, large drops in hours worked (reflecting both demand and supply factors), and precarious housing.
 It is unclear how persistent those events are, but our subsequent decomposition of income to moments of permanent and transitory shocks will shed light on this.
@@
 Panel B of table 1 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID.
-The variance of consumption growth is very similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error.
+The variance of consumption growth is similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error.
 The first-order autocovariance is $-0.044$ and highly significant.
@@
 Consequently, overall consumption in the PSID is, on average, higher by about 25% than a similarly selected sample in the CEX.
-Appendix figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID and CEX samples.
-The patterns for the variance and kurtosis are quite similar in the two surveys over time, though the variance in the PSID is lower and the kurtosis is slightly higher.
-Skewness, however, is markedly different.
-Skewness of log consumption is strongly positive in the PSID, while it is mostly negative in the CEX; the patterns also diverge over time.
-Overall, these statistics suggest that consumption across otherwise similar samples in the PSID and CEX exhibits some notable differences over our sample period.
+Appendix figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID and CEX.
+The patterns for the variance and kurtosis are similar in the two samples over time, though the variance in the PSID is lower and the kurtosis is slightly higher.
+Skewness of log consumption, however, is markedly different: it is strongly positive in the PSID but mostly negative in the CEX; the patterns also diverge over time.
 This comparison is not informative about the moments of consumption growth, which is the relevant variable in our case.
@@
 This is why BPP imputed consumption from the CEX into the PSID: the CEX lacked the necessary panel dimension while the PSID lacked comprehensive consumption data before 1999.
-We repeat BPP’s imputation over the recent years (our sample period 1999–2019).
-The imputation consists of the estimation of food demand in the CEX (food expenditure expressed as a function of prices and nondurable consumption).
+We repeat BPP’s imputation over the recent years, i.e., our sample period 1999–2019.
+The imputation involves the estimation of food demand in the CEX (food expenditure expressed as a function of prices and nondurable consumption).
 The demand equation is then inverted (nondurable consumption expressed as a function of prices and food) and used to impute nondurables in the PSID based on food expenditure and prices that are available in both surveys.
@@
 In fact, the variance of imputed consumption growth is more than double the variance of income growth.
-The higher-order moments also differ markedly between the imputed and true series.
+The higher-order moments differ markedly between the imputed and true series.
 Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series.
@@
 Since the imputation is not strictly needed over our sample period, we use consumption internally available in the PSID as our baseline measure of consumption.
-The present analysis also suggests that we may not want to rely on higher-order moments of consumption, given the notable discrepancies in skewness of log consumption between the PSID and the CEX (this is independent of the imputation).
-As we established in section 3.3, an advantage of our identification approach over other methods is precisely its non-reliance on such higher-order consumption moments.
+The present analysis also suggests that we may not want to rely on higher-order moments of consumption, given the discrepancies in skewness of log consumption between the PSID and the CEX, which are independent of the imputation.
+As we established in section 3.3, an advantage of our approach is precisely its non-reliance on higher-order consumption moments.
 5 Results
 We first present baseline results imposing stationarity in the distributions of income shocks, taste heterogeneity, consumption error, and a constant pass-through of income shocks over the lifecycle.
-[^28] Stationarity is not an identifying assumption but helps with precision, especially of higher moments.
-We subsequently relax stationarity/lifecycle-invariance in section 5.4.
+We subsequently relax stationarity/lifecycle-invariance in section 5.3.282828Stationarity is not an identifying assumption but helps with precision, especially of higher moments.
 5.1 Income process
 We present estimates of the income process in three specifications.
-First, we target only second moments of income growth, which enables us to estimate the second moments of income shocks.
+First, we target only the second moments of income growth, which enables us to estimate the variance of income shocks.
 This is similar to what BPP do over the earlier period 1980–1992.
-Second, we target second and third moments of income growth, which enables us to estimate also the skewness of income shocks.
-Finally, we target up to the fourth moments of income growth, which allows us to estimate all the parameters of the income process as presented in section 2.1.
-The results from each specification appear in table 3, columns 1-3, respectively.
-The full set of moments targeted in each case is shown in appendix B.4.1.
+Second, we target second and third moments of income growth, which enables us to estimate also the skewness of shocks.
+Finally, we target up to the fourth moments, which allows us to estimate all the parameters of the income process as presented in section 2.1.
+The results are in table 2, while the full set of moments targeted in each case is shown in appendix B.4.1.
 When only the second moments of income growth are targeted, we estimate the variance of permanent and transitory shocks at $0.027$ and $0.033$, respectively.
@@
 As the variance of income growth is driven by the variance of the contemporaneous permanent shock and the sum of the variances of transitory shocks in the current and previous periods, the contribution of the transitory shock to the variance of income growth is almost $2.5$-fold that of the permanent shock (as in Gottschalk and Moffitt, 2009).
-Table 3: Estimates of the income process, baseline sample
+Table 2: Estimates of the income process
 Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
@@
 Panel A. Permanent shocks
-Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.030
-(0.002) (0.002) (0.002)
-Skewness $\gamma_{\zeta}$ central -0.004 -0.004
-(0.003) (0.003)
-standardized -0.947 -0.831
-(0.649) (0.571)
-Kurtosis $\kappa_{\zeta}$ central 0.039
-standardized 44.393
+Variance $\sigma_{\zeta}^{2}$ 0.027 (0.002) 0.027 (0.002) 0.030 (0.002)
+Skewness $\gamma_{\zeta}$ central -0.004 (0.003) -0.004 (0.003)
+standardized -0.947 (0.649) -0.831 (0.571)
+Kurtosis $\kappa_{\zeta}$ central 0.039 (0.007)
+standardized 44.393 (11.080)
 Panel B. Transitory shocks
-Variance $\sigma_{v}^{2}$ 0.033 0.033 0.031
-(0.002) (0.002) (0.003)
-Skewness $\gamma_{v}$ central -0.008 -0.008
-(0.003) (0.003)
-standardized -1.310 -1.431
-(0.510) (0.566)
-Kurtosis $\kappa_{v}$ central 0.048
-standardized 49.226
+Variance $\sigma_{v}^{2}$ 0.033 (0.002) 0.033 (0.002) 0.031 (0.003)
+Skewness $\gamma_{v}$ central -0.008 (0.003) -0.008 (0.003)
+standardized -1.310 (0.510) -1.431 (0.566)
+Kurtosis $\kappa_{v}$ central 0.048 (0.005)
+standardized 49.226 (10.119)
 Notes: The table presents the estimates of the parameters of the income process, imposing stationarity over time.
@@
 Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged.
-We estimate the skewness (third standardized moment) at $-0.947$ and $-1.310$ for permanent and transitory shocks, respectively.
-Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a much longer left than right tail.
+We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively.
+Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail.
 This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period.
+These results vary with the level of income, but we postpone this discussion for later.
 Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme.
@@
 Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
-Targeting all moments of income growth up to fourth-order leaves the third central moments unchanged and modifies the variances only by little.
+Targeting all moments of income growth up to fourth-order leaves the third central moments unchanged and modifies the variances by little.
 We estimate the kurtosis (fourth standardized moment) at $44.393$ and $49.226$ respectively.
@@
 This is consistent with Busch and Ludwig (2024), who use the PSID until 2012 and estimate numerically similar kurtosis coefficients for both shocks.
-Together with left skewness, large excess kurtosis implies that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad.
+Seen together, large left skewness and excess kurtosis imply that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad.
 Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
 5.2 Linear consumption function
-We present estimates of the linear consumption function, first, targeting only second moments of the joint income-consumption distribution and, subsequently, targeting second as well as higher-order moments of the joint distribution.
-The full set of moments targeted in each case is shown in appendix B.4.2.
+We present estimates of the linear consumption function, first, targeting only second moments of the joint income-consumption distribution and, subsequently, targeting second as well as higher-order moments.
+The full set of targeted moments is shown in appendix B.4.2.
 Targeting second-order moments only.
@@
 As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period.
-To help decipher their role, we impute consumption from the CEX into the PSID in the same way BPP did for the earlier period, and we repeat the estimation; see appendix C.3 for details on the imputation.
+To help decipher their role, we impute consumption from the CEX into the modern PSID in the same way BPP did for the earlier period, and we repeat the estimation.
 The results are in column 2 of table 3.
-The transmission parameter of permanent shocks doubles to $\phi^{(1)}=0.288$, while the transmission parameter of transitory shocks remains statistically zero (though at a larger magnitude).
-Using imputed consumption data thus inflates the pass-through of permanent shocks, substantially lowering the degree of partial insurance.
-This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true series (column 1) and the imputed data from the CEX (column 2).
-By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and also very close to BPP.
+The pass-through of permanent shocks doubles to $\phi^{(1)}=0.288$, while the pass-through of transitory shocks remains statistically zero (but at a larger magnitude).
+Using imputed consumption data thus inflates the transmission of permanent shocks and lowers the degree of insurance.
+This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true and the imputed data.
+By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and very close to BPP.
 The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate.
-We relegate the details of the imputation and a discussion of a mechanism through which the imputation in particular inflates $\phi^{(1)}$ to appendix C.3.
-Table 4: Estimates of the consumption function, baseline sample
+Appendix C.4 discusses the details of the imputation and a mechanism through which it inflates $\phi^{(1)}$ in particular (the imputation involves the inversion of a demand system, so certain types of measurement error may inflate the imputed income-consumption covariance – evident also in table 1).
+Table 3: Estimates of the consumption function
 Consumption fn.: Linear Quadratic
@@
 Results from optimal GMM are similar (appendix table C.7).
-Although the imputation inflates $\phi^{(1)}$, the inflated value in column 2 is still substantially lower than BPP.
-The PSID switched in 1997 from annual to biennial frequency.
-It is thus possible that at the lower frequency certain higher frequency income shocks are smoothed out or undone by subsequent shocks.
-To assess this, we would need to re-estimate our model on annual growth rates, which is impossible after 1997.
-However, we can use the original BPP data over 1980–1992 to estimate the model annually versus biennially.
-We do this in appendix C.4.
-While we obtain $\phi^{(1)}=0.64$ using annual growth rates (exactly as BPP), this drops to $\phi^{(1)}=0.4$–0.5 using biennial rates.
-The drop is sharper when we target time-invariant moments, which was recently also shown in Chatterjee et al.
-(2021).
-The biennial nature of the modern data thus explains another 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate.
-[^31] In both annual and biennial cases, $\phi^{(1)}$ measures the pass-through of the permanent shock at an annual rate.
-This becomes clear in appendix B that shows the estimating equations with biennial data.
-The difference in the estimates of $\phi^{(1)}$ between the two cases arises because with biennial data we only measure an ‘aggregate’ pass-through within a given biennial period.
-Consequently, some higher-frequency shocks will be inevitably smoothed out at the observed lower frequency and their transmission will be muted.
-We attribute the remaining gap to calendar time effects, e.g., changes in taxes/benefits or other insurance channels between 1980–1992 and 1999–2019 (see Borella et al., 2022, for a discussion of such changes).
-It is worth noting that we are not the first paper to find a low pass-through of permanent shocks in the recent years; Arellano et al.
-(2023) use PSID data after 2005 and find a pass-through of about $0.3$ (specification closest to ours).
-If we zoom into the period after 2005, we estimate $\phi^{(1)}=0.21$ (results not shown for brevity), statistically indistinguishable from Arellano et al.
-(2023)’s estimate.
+[…]
 Introducing higher-order moments.
 The estimation of the linear consumption function targeting higher-order moments is our first attempt to uncover how the tails of income risk affect the pass-through of shocks into consumption.
-In other words, we want to measure the degree of partial insurance going beyond the workhorse income-consumption covariance to account for the non-Gaussian features of modern income dynamics.
+In other words, we want to measure the degree of partial insurance targeting moments beyond the workhorse income-consumption covariance in order to account for the non-Gaussian features of modern income dynamics.
 We introduce the higher-order moments incrementally.
@@
 Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income.
-The partial insurance parameters remain largely unchanged upon inclusion of the third moments.
-We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with second moments only) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero).
-We thus again find large partial insurance to permanent shocks (a $10\%$ permanent change in income results in only $1.6\%$ shift in consumption) and full insurance to transitory shocks.
-Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with second moments only.
-Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or consumption measurement error.
+The transmission parameters remain largely unchanged upon inclusion of the third moments.
+We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with only second moments) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero).
+We thus again find large partial insurance to permanent shocks and full insurance to transitory shocks.
+Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with only second moments.
+Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or measurement error.
 Column 4 of table 3 presents the results from targeting all income and consumption moments up to fourth-order.
@@
 A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption.
-Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth-order moments.
+Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth moments.
 The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks.
@@
 This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary.
-For example, Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets, while Busch and Ludwig (2021) find sizable welfare implications of higher-order risk.
-[^34] De Nardi et al.
+[^34] Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets.
+However, De Nardi et al.
 (2019) and Busch and Ludwig (2024) find higher insurance when shocks are non-normal, due to stronger precautionary motives.
@@
 Our quadratic specification addresses their concern.
-One would thus expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events.
+One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events.
 Yet, the linear consumption function leaves no room to such events.
@@
 But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response.
-A linearized marginal utility, however, fails to pick up such large response (echoing Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments.
-In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates and justifies our higher-order specification subsequently.
-To summarize our findings from the linear consumption function: 1.) we replicate BPP over the recent years and find much lower pass-through of permanent shocks (larger degree of partial insurance); the imputation of consumption from the CEX and the biennial nature of the modern PSID explain the bulk of the difference from BPP; 2.) targeting higher-order moments does not alter the degree of consumption partial insurance, which is not entirely surprising as tail risk has, by design, no role in the linear framework.
+A linearized marginal utility fails to pick up such large response (Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments.
+In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates our higher-order specification subsequently.
+To summarize our findings from the linear consumption function: 1.) we replicate BPP over the recent years and find much lower pass-through of permanent shocks (larger degree of partial insurance); the imputation of consumption from the CEX and the biennial nature of the modern PSID explain the bulk of the difference from BPP; 2.) targeting higher-order moments does not alter the overall degree of consumption partial insurance, which is not entirely surprising as tail risk has, by design, no role in the linear framework.
 5.3 Quadratic consumption function
@@
 This raises an issue of comparability between the linear and quadratic specifications, as the target moments differ (appendix table B.2).
-However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway (section 3).
+However, we can estimate the linear function targeting the exact same moments that we use in the quadratic case, as the linear function is identified by a smaller set of lower-order moments anyway.
 Our estimates of the linear function in that case do not differ from those in column 4, table 3, so estimation is robust to the choice of moments.
 These results are available upon request.
-We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, highly significant and close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification.
+5.3.1 Main results
+We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification.
 We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero.
-As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial in the shocks.
-We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks (statistically significant at the margin for now), while we estimate a positive $\psi^{(2)}=0.015$ for transitory shocks (statistically insignificant for now).
-The coefficient $\omega^{(22)}$ on the product of permanent and transitory shocks is large but imprecisely estimated.
+As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
+We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks and a positive (but statistically insignificant for now) $\psi^{(2)}=0.015$ for transitory shocks.
+The coefficient $\omega^{(22)}$ on the product of the two shocks is large but very imprecisely estimated.
 The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case.
@@
 Two main results follow.
-First, bad (negative) permanent shocks are more hurtful than good (positive) ones because they are associated with a much larger transmission parameter into consumption.
-Consider, for example, a large permanent shock to household income, $\zeta_{it}=\pm 0.5$, approximately three standard deviations away from the mean.
-If the shock is negative ($\zeta_{it}=-0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$; by contrast, if the shock is positive ($\zeta_{it}=0.5$), then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$.
+First, bad (negative) permanent shocks are more unsettling than good (positive) ones because they are associated with a larger transmission parameter into consumption.
+Consider, for example, a large permanent shock to household income, $|\zeta_{it}|=0.5$, about three standard deviations away from the mean.
+If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.$=0.036$, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.$=0.036$).
 We can generalize this to bad and good shocks of any magnitude and conclude that
 $\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<0}~{}>~{}\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta>0}.$
-Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 3, but they also have much larger pass-through rates into consumption.
+Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption.
 Bad permanent shocks are thus double hurtful.
@@
 Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean.
-In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$.
+In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$).
 We can generalize this to bad shocks of any magnitude and conclude that
 $\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<<0}~{}>~{}\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<0}.$
-This result is not hard to comprehend.
-When a bad shock hits, households can insure consumption by running down savings or by using other formal or informal insurance arrangements.
-But there is clearly a limit to this.
-As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, thus resulting in a larger fraction of the shock eventually transmitting into consumption.
-[^38] The parameter estimates also imply that the pass-through of a good shock decreases with its magnitude.
-While this seems counterintuitive, recall that extreme good shocks are rare in our data, which follows from left skewness and excess kurtosis in table 3.
+Why would a positive or small shock be better insured?
+When a good shock hits, households save part of it and consumption does not respond too much.
+When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements.
+But there is clearly a limit to how much they can dissave or borrow.
+As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption.
+A large negative shock thus affects consumption by relatively more than a small or positive shock does.
+[^38] The results also imply that the pass-through of a good shock decreases with its magnitude.
+The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks.
+However, large left skewness and excess kurtosis imply that extreme good shocks are rare.
 We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
@@
 (2019) and Fuster et al.
-(2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the shock.
+(2020) find that responses to losses are much larger than responses to gains, and the magnitude of the response increases with the magnitude of the bad shock.
 This literature, however, elicits responses to transitory shocks.
 By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
-Why would a positive or small shock be better insured?
-While we cannot assess candidate mechanisms without a structural model, our best explanation is that people save a good income shock (thus consumption does not respond much) but they dissave in the presence of a bad one.
-As there is clearly a limit to how much households can dissave or borrow, especially in the presence of borrowing constraints, a large negative shock would typically affect consumption by relatively more.
-[^39] Busch and Ludwig (2021) simulate a lifecycle model with concave utility and borrowing constraints; they find larger pass-through of negative versus positive permanent shocks.
-While we do not model borrowing constraints here, a strength of our approach is that it enables us to measure partial insurance in the data without taking a stance on the precise underlying structural model/economic mechanism at play.
 Moving on to transitory shocks, their pass-through is given by (7); this depends on $\psi^{(1)}$, $\psi^{(2)}$, $\omega^{(22)}$, and on the magnitudes and signs of both income shocks.
 If we set the permanent shock to its mean value ($\zeta_{it}=0$) and plug in the parameter estimates from table 3, we get
-$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.$
+$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\$
 Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample.
@@
 This is similar to our findings for the linear case as well as to BPP.
-By contrast, this is not the case for certain well-defined subsamples, but we postpone the discussion of the subsamples for section 5.4.
-Despite the lack of statistical significance, the partial derivative with respect to transitory shocks has, at face value, an interesting implication: large good transitory shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero.
-While we should be cautious in drawing conclusions about extreme good transitory shocks (given left skewness and excess kurtosis, large good transitory shocks are not present in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and other semi-structural studies that use covariance restrictions on survey data.
-The former studies typically exploit large and infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021).
+By contrast, this is not the case for certain well-defined subsamples, which we analyze in the next subsections.
+Despite the lack of statistical significance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero.
+While we should be cautious in drawing conclusions about large good transitory shocks (given left skewness and excess kurtosis, such shocks are rare in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and similar studies that employ covariance restrictions on survey data.
+The former typically exploit large infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021).
 In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions.
@@
 By contrast, transitory shocks, be they good or bad, small or large, do not affect consumption, at least not in a statistically significant way.
-[^40] Appendix table C.7 shows that similar conclusions are obtained if we use consumption imputed from the CEX.
-In particular, $\phi^{(2)}<0$ (albeit not statistically significant), with a magnitude similar to that obtained directly from the PSID ($-0.036$ versus $-0.04$), and $\psi^{(2)}>0$.
-5.4 Sensitivity to age, wealth, and education
-We first consider how partial insurance changes with age, by repeating the estimation over three, almost equally sized, subsamples of households with the male spouse between 30–40, 41–50, and 51–65 years old respectively.
+[^39] Similar conclusions are obtained if we use imputed consumption from the CEX.
+In particular, $\phi^{(2)}<0$, with a magnitude similar to that obtained from the PSID, and $\psi^{(2)}\approx>0$ (appendix table C.8).
+5.3.2 Results by age and position in the income distribution
+We now consider how partial insurance changes with age, by repeating the estimation over three, almost equally sized, subsamples of households with the male spouse between 30–40, 41–50, and 51–65 years old respectively.
 [^40] We target the same moments as earlier, broken down by the age of the male spouse.
-The moments of income shocks are also age-specific but we do not report those results for brevity.
+The moments of income shocks are also age-specific (appendix table C.9).
 Intuitively, the transmission parameters of shocks should exhibit some age dependence since wealth accumulates over the lifecycle, which in turn influences the household’s ability to insure against income fluctuations.
-The results appear in table 5.
-Column 1 presents results from the linear specification when only second moments are targeted.
-Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65.
-A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are included in the estimation.
-This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results.
-[^42] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters.
-By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details.
-By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
+Table 4 reports results from the quadratic specification (see table C.10 for the linear specification).
+$\phi^{(1)}$ decreases with age, so the average permanent shock impacts consumption less as households age.
+In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle.
+$\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age.
+It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket.
+As a result, large negative permanent shocks impact consumption more severely at older ages.
+Consider, again, $\zeta_{it}=-0.5$.
+At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
 Table 4: Estimates of the consumption function, by age
-Consumption fn.: Linear Quadratic
-Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
-Consumption data: PSID PSID PSID PSID
-(1) (2) (3) (4)
-$\phi^{(1)}$ 30-40 0.154 (0.045) 0.164 (0.051) 0.089 (0.058) 0.149 (0.050)
-41-50 0.126 (0.051) 0.146 (0.055) 0.079 (0.089) 0.115 (0.049)
-51-65 0.180 (0.049) 0.175 (0.051) 0.136 (0.061) 0.113 (0.057)
-$\psi^{(1)}$ 30-40 0.094 (0.079) 0.083 (0.073) 0.095 (0.107) 0.105 (0.090)
-41-50 0.010 (0.079) -0.016 (0.072) -0.075 (0.061) 0.019 (0.090)
-51-65 -0.046 (0.070) -0.044 (0.066) 0.037 (0.101) -0.063 (0.088)
-$\phi^{(2)}$ 30-40 0.006 (0.040)
-41-50 -0.008 (0.063)
-51-65 -0.105 (0.054)
-$\psi^{(2)}$ 30-40 0.036 (0.061)
-41-50 0.048 (0.046)
-51-65 -0.052 (0.079)
-$\omega^{(22)}$ 30-40 -1.090 (1.465)
-41-50 -0.290 (1.195)
-51-65 3.003 (1.306)
-$\sigma^{2}_{\xi}$ 30-40 0.014 (0.002) 0.014 (0.002) 0.016 (0.002) 0.013 (0.004)
-41-50 0.022 (0.002) 0.022 (0.002) 0.022 (0.002) 0.022 (0.002)
-51-65 0.021 (0.002) 0.021 (0.002) 0.023 (0.002) 0.010 (0.008)
-$\gamma_{\xi}$ 30-40 -0.001 (0.001) -0.001 (0.001)
-41-50 0.006 (0.001) 0.006 (0.002)
-51-65 -0.006 (0.002) -0.006 (0.002)
-$\kappa_{\xi}$ 30-40 0.004 (0.002)
-41-50 0.007 (0.002)
-51-65 0.006 (0.002)
-$\sigma^{2}_{u_{c}}$ 30-40 0.046 (0.003) 0.046 (0.003) 0.046 (0.003) 0.046 (0.003)
-41-50 0.040 (0.002) 0.040 (0.002) 0.040 (0.002) 0.040 (0.002)
-51-65 0.043 (0.002) 0.043 (0.002) 0.043 (0.002) 0.043 (0.002)
-$\gamma_{u_{c}}$ 30-40 0.000 (0.002) 0.000 (0.002)
-41-50 0.002 (0.002) 0.002 (0.002)
-51-65 0.004 (0.002) 0.004 (0.002)
-$\kappa_{u_{c}}$ 30-40 0.014 (0.003)
-41-50 0.012 (0.002)
-51-65 0.012 (0.002)
-Notes: The table presents the estimates of the parameters of the consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse (the underlying income moments also vary over these age brackets).
-Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table.
+Consumption fn.: Quadratic
+Age household head: 30–40 41–50 51–65
+Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
+Consumption data: PSID PSID PSID
+(1) (2) (3)
+$\phi^{(1)}$ 0.149 (0.050) 0.115 (0.049) 0.113 (0.057)
+$\psi^{(1)}$ 0.105 (0.090) 0.019 (0.090) -0.063 (0.085)
+$\phi^{(2)}$ 0.006 (0.038) -0.008 (0.063) -0.104 (0.052)
+$\psi^{(2)}$ 0.036 (0.061) 0.048 (0.046) -0.052 (0.073)
+$\omega^{(22)}$1.089 (1.465) -0.294 (1.195) 2.984 (0.791)
+$\sigma^{2}_{\xi}$ 0.013 (0.004) 0.022 (0.002) 0.010 (0.005)
+$\sigma^{2}_{u_{c}}$ 0.046 (0.003) 0.040 (0.002) 0.043 (0.002)
+Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the household head (male spouse).
+The underlying income moments also vary over these age brackets (appendix table C.9).
 All columns use consumption data internally available in the PSID.
 Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
-Column 4 presents the results from the quadratic specification.
-In this case, $\phi^{(1)}$ decreases monotonically with age, so the average permanent shock (the shock close to the mean of the distribution) impacts consumption gradually less as households age.
-In line with theory, this pattern reflects the age-increasing role of financial wealth for self-insurance.
-Perhaps more importantly, $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases monotonically with age.
-$\phi^{(2)}$ drops from $0.006$ to $-0.008$ as we move from ages 30–40 to ages 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.105$ in the 51–65 age bracket.
-As a result, large negative permanent shocks impact consumption more severely at older than at younger ages.
-The last finding echoes Guvenen et al.
-(2021), who show that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age.
-There is more room for income to fall sharply in this case than among the low-earners or the young.
-In fact, we also find larger left skewness in income growth as households age in the PSID and more negative skewness in permanent shocks among the old.
-[^43] The coefficient of skewness of permanent shocks more than triples in magnitude from $-0.65$ to $-2.39$ as we move from ages 30–40 to 51–65.
-The coefficient of skewness of transitory shocks, by contrast, does not change much with the age – its estimate is $-1.43$ and $-1.45$, respectively, over ages 30–40 and 51–65.
-So, not only do permanent shocks become more negative as households age, but they also transmit into consumption at increasingly larger rates.
-In other words, not only is there more room for income to fall among older earners, but this also causes even greater harm to consumption.
-By contrast, despite a similar pattern for transitory shocks, their transmission parameters remain statistically indistinguishable from zero over the entire lifecycle.
-We next consider two subsamples formed on the basis of household wealth.
-We define a household as ‘low wealth’ (‘high wealth’) if its average wealth over its entire observation window is less (more) than median real wealth in the sample over 19991999199919992019201920192019.
-Wealth is the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt.
-Table 1 provides summary statistics for the wealth subsamples.
-Average wealth in the first group is \$69,731 in 2018 prices, while the second group has 14 times more wealth, an average of \$950,823.
-Income and earnings are twice as high in the second group; consumption is only 30% higher.
-Table 6, columns 1-2, presents the results in the quadratic model.
-[^44] Appendix table C.8, columns 1-4, presents the parameter estimates in the linear model.
+Linear specification estimates appear in appendix table C.10.
+Guvenen et al.
+(2021) document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age.
+There is more room for income to fall sharply in this case than among the young.
+In our data, skewness of permanent shocks becomes more negative as households age; it more than triples in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9).
+The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption technology itself changes ($\phi^{(2)}$ more negative).
+The type of events that underlie large income downturns at older ages – such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption technology that differs between the old and the young.
+This does not mean that moderate bad shocks are worse insured among the old; for small shocks, the decreasing pattern in $\phi^{(1)}$ more than counteracts the increasing magnitude of $|\phi^{(2)}|$.
+Finally, the transmission parameters of transitory shocks remain statistically indistinguishable from zero over the entire lifecycle.
+We next consider how partial insurance changes with the household’s position in the income distribution before the shocks hit at $t$, i.e., in the distribution of $y_{it-1}$.
+Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, unlike ABB whose nonlinear income process naturally allows this.
+We thus take an empirical approach and estimate the model over different cuts of the distribution of $y_{it-1}$.
+Sample size considerations prevent us from being overly granular; we tried estimation over deciles, quartiles, and tertiles, and the latter was the most robust.
+For brevity we report results in the bottom (‘poor’) and top (‘rich’) tertiles.
+[^41] We target the same moments as earlier, broken down by tertiles of the distribution of past income $y_{it-1}$.
+The moments of shocks are also income level-specific (appendix table C.9).
+Table 5 reports results from the quadratic specification (see table C.11 for the linear specification).
 Three points emerge.
-First, $\phi^{(1)}$ is more than twice as large among the low vis-à-vis high wealth households ($0.212$ versus $0.091$).
-As $\phi^{(1)}$ measures the transmission of the average permanent shock, low wealth households exhibit a larger pass-through of average shocks, i.e., they are less able to insure against them, compared to high wealth households.
-Second, $\phi^{(2)}<0$ for both groups, albeit statistically insignificant.
-With the caveat of insignificance, the magnitude of $\phi^{(2)}$ is bigger among the wealthy.
-Large negative permanent shocks thus impact consumption of the wealthy more severely than they do among the less wealthy.
-As in the case of older households previously, there is more room for consumption to fall sharply among the wealthier.
-Third, $\psi^{(2)}$, the transmission parameter of the typical transitory shock away from the mean, is negative, large, and statistically significant among the less wealthy.
-Transitory shocks thus have an asymmetric pass-through among lower wealth households, with large bad shocks reducing consumption substantially more than good shocks increase it.
-So while average transitory shocks are fully insured among low wealth households, larger negative transitory shocks are not: they reduce consumption quite substantially as low wealth households lack the self-insurance means to smooth out even transitory fluctuations.
+First, $\phi^{(1)}$ among the rich is half the value of the poor, so the average permanent shock affects consumption less strongly among the rich.
+This reflects again the self-insurance role of wealth given that income-richer households are also wealthier.
+Second, $\phi^{(2)}$ is negative, confirming the baseline finding that negative permanent shocks have larger pass-through than equal-sized positive shocks.
+Third, the transmission of transitory shocks shows mixed patterns but it is never statistically significant.
+At face value, $\phi^{(2)}$ is more negative among the poor, suggesting that a bad shock of a given size transmits more strongly among them.
+While factually correct given the estimates, this provides an incomplete view of the underlying dynamics because the distribution of income risk is vastly different between poor and rich.
+Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but very negative among the rich, with a coefficient of $-4.86$ (multiple times larger than baseline skewness, and highly significant).
+One may thus study how large bad/good shocks from a given position in the underlying distribution, for example shocks in the 3rd and 97th percentiles, affect consumption in each case.
+Simulations are needed to infer the specific shocks.
+[^42] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9.
+We provide details on similar simulations in section 6.
+Adopting this approach, the bad shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) but it shifts $c_{it}$ by $-0.035$, twice as much, among the rich (pass-through $0.084$).
+By contrast, the good shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but shifts $c_{it}$ by only $0.006$ among the rich (pass-through $0.057$).
+These findings agree with ABB, who simulate the transmission of 10- and 90-percentile shocks and also find a stronger response to bad news among households in the top end of the income distribution and a stronger response to good news among households in the bottom end.
+The reported $\Delta c_{it}$ are also numerically in the same ballpark (impulse responses with heterogeneity in their appendix).
+Our results thus suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
+Table 5: Estimates of the consumption function, over position in income distribution
+Consumption fn.: Quadratic
+Lag income $y_{it-1}$: Bottom tertile Top tertile
+Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th
+Consumption data: PSID PSID
+(1) (2)
+$\phi^{(1)}$ 0.113 (0.029) 0.063 (0.100)
+$\psi^{(1)}$ 0.186 (0.199) 0.275 (0.200)
+$\phi^{(2)}$0.064 (0.029) -0.025 (0.101)
+$\psi^{(2)}$ 0.210 (0.220) -0.268 (0.197)
+$\omega^{(22)}$ 1.965 (1.513) 1.072 (2.144)
+$\sigma^{2}_{\xi}$ 0.011 (0.005) 0.025 (0.004)
+$\sigma^{2}_{u_{c}}$ 0.044 (0.002) 0.041 (0.002)
+Notes: The table presents the estimates of the parameters of the quadratic consumption function in the bottom and top tertiles of the distribution of lag income $y_{it-1}$.
+The underlying income moments also vary over the income distribution (appendix table C.9).
+All columns use consumption data internally available in the PSID.
+Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses.
+Linear specification estimates appear in appendix table C.11.
+Our approach highlights two distinct forces at play in the response of consumption to income shocks: the underlying nature of income risk (i.e., the income parameters) and the consumption technology (i.e., the consumption parameters).
+The larger response of the rich to shocks in the bottom of the distribution is not because their consumption technology differs vastly from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates the response to said shocks), but mostly due to stronger downwards risk that the rich face.
+To infer that, one needs to evaluate the marginal distribution of income risk.
+ABB’s quantile method can flexibly estimate the consumption response over income quantiles and compare responses to, say, 10- and 90-percentile shocks, but it does not measure the higher-order features of the distribution these shocks come from.
+This may lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be very similar to one another since the underlying distribution is skewed and leptokurtic.
+For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked further out at the 3rd and 97th percentiles.
+Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
+$c_{it}=g_{t}(\underbrace{y_{it-1}+F_{t}^{-1}(u_{it})}_{y_{it}},~{}\dots),\$
+where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ is a uniform shock denoting the quantile of the shocks’ distribution $F_{t}$.
+They then simulate the consumption pass-through over different values for $y_{it-1}$ and $u_{it}$, without measuring $F_{t}$.
+Our model allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the sign and size of the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$.
+It can thus describe more transparently how income shocks of different signs/sizes shift consumption.
+How does a less negative $\phi^{(2)}$ among the rich reconcile with a more negative $\phi^{(2)}$ among the old, who are presumably also rich?
+The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between the nature of income risk and the consumption technology.
+Skewness of permanent shocks is twice as large among the rich ($-4.86$) as it is among the old ($-2.39$).
+The consumption distribution is not too different between the two, however, so the rich must have more insurance to a bad shock of given size.
+This explains why $\phi^{(2)}$ is four times smaller among them.
+In both cases (old, rich), bad permanent shocks from a given position in the distribution decrease consumption more strongly than among their counterparts (young, poor).
+5.3.3 Results by wealth and education
+We next consider four subsamples formed on the basis of household wealth and education.
+We use conventional definitions for ‘low’ versus ‘high’ wealth (the cutoff is median wealth in the sample) and ‘no’ versus ‘some’ college (the cutoff is 12 years of education).
+[^43] Wealth is defined as the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt.
+Appendix table C.1 provides summary statistics for these subsamples.
 Table 6: Estimates of the consumption function, by wealth and education
 Consumption fn.: Quadratic
-Subsample: Low wealth High wealth No college Some college
+Low wealth High wealth
+No college Some college No college Some college
 Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
@@
 (1) (2) (3) (4)
-$\phi^{(1)}$ 0.212 0.091 0.219 0.105
+$\phi^{(1)}$ 0.242 0.192 0.164 0.079
 (0.051) (0.052) (0.067) (0.034)
-$\psi^{(1)}$0.081 0.017 -0.071 0.030
+$\psi^{(1)}$0.251 0.015 0.006 0.004
 (0.131) (0.108) (0.085) (0.073)
-$\phi^{(2)}$0.018 -0.039 -0.095 -0.023
+$\phi^{(2)}$0.094 0.005 -0.076 -0.032
 (0.062) (0.052) (0.057) (0.041)
-$\psi^{(2)}$0.108 0.031 0.050 0.012
+$\psi^{(2)}$0.162 -0.001 0.177 0.001
 (0.087) (0.074) (0.089) (0.048)
-$\omega^{(22)}$ 1.116 -0.049 0.668 0.095
+$\omega^{(22)}$ 2.998 -0.660 -2.228 0.441
 (1.423) (1.530) (1.568) (0.995)
-$\sigma^{2}_{\xi}$ 0.013 0.022 0.013 0.021
+$\sigma^{2}_{\xi}$ 0.006 0.013 0.006 0.025
 (0.008) (0.002) (0.010) (0.002)
-$\sigma^{2}_{u_{c}}$ 0.048 0.042 0.043 0.044
+$\sigma^{2}_{u_{c}}$ 0.046 0.045 0.038 0.043
 (0.003) (0.003) (0.004) (0.002)
-Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups (linear specification estimates appear in appendix table C.8).
-Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019.
+Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups.
+The underlying income moments also vary by wealth and education (results not shown for brevity).
+Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 19991999199919992019201920192019.
 No versus some college is defined on the basis of the highest level of education attained by the male spouse.
 All columns use consumption data internally available in the PSID.
-Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
-We finally turn to the role of education.
-We define a household as ‘no college’ (‘some college’) if the male spouse has not gone to college (has some college education).
-Table 1 provides summary statistics.
-Those with some college education earn 70% more income, they have more than twice as much wealth, and they consume about 30% more.
-Table 6, columns 3-4, presents the results in the quadratic model.
-[^45] Appendix table C.8, columns 5-8, presents the parameter estimates in the linear model.
-Similar to the patterns observed by wealth, we find that $\phi^{(1)}$ is higher (double) for households without college education ($0.219$ versus $0.105$).
-The pass-through of the average permanent shock is therefore larger among the lower educated, presumably because of the lower wealth that these households hold.
-$\phi^{(2)}$ is negative in both groups, suggesting that large bad permanent shocks transmit into consumption more than equal-sized good shocks.
-However, the magnitude of the estimate in the lower education group is over four-fold larger (and statistically significant) compared to the higher education group.
-Permanent shocks away from the mean are therefore way more hurtful among the lower educated.
-By contrast, the transmission parameters of all transitory shocks are statistically indistinguishable from zero in both groups.
+Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses.
+Linear specification estimates appear in appendix table C.12.
+Table 6 reports results from the quadratic model (see table C.12 for the linear model).
+Three points stand out.
+First, $\phi^{(1)}$ decreases monotonically as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college).
+As $\phi^{(1)}$ measures the transmission of the average permanent shock, these groups possess gradually more insurance to average shocks; this reflects the insurance role of wealth, which increases monotonically across the groups.
+Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move across groups, with the exception of group 2 (low wealth and some college) where it equals zero.
+Negative shocks thus mostly transmit more strongly than positive shocks, but wealthier and more educated people have more insurance to bad shocks of given size.
+These people face, however, much stronger downwards income risk.
+This pattern is consistent with the gradient between poor and rich in table 5.
+Third, the transmission parameters of transitory shocks are mostly statistically insignificant, with two exceptions.
+In group 1 (low wealth, no college), $\psi^{(1)}$ is significant but negative, which is counterintuitive at first, yet $\psi^{(2)}$ is also negative, large, and significant.
+Large negative transitory shocks thus reduce consumption substantially among low wealth, low education households.
+In group 3 (high wealth, no college), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones.
+We argued earlier that this helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in studies that employ covariance restrictions.
+Note that the size of group 3 is by far the smallest in our sample.
+5.3.4 Link with structural model
+Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, and income distributions.
+Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
+Madera (2019) finds that earnings and nondurable consumption are correlated in the middle of the earnings distribution, but less so in the tails compared to earnings and durable consumption.
+Negative shocks are associated with larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
+De Nardi et al.
+(2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings (thus also enables more insurance to a given shock).
+They show that the BPP method that relies on second-order moments implies a substantially higher pass-through rate than the true rate in the model.
+Busch and Ludwig (2024) simulate a model with borrowing constraints.
+We do not model such constraints, yet a strength of our approach is that it enables us to measure partial insurance without taking a stance on the precise economic mechanisms at play.
+They find that the marginal propensity to consume increases with the severity of a negative permanent shock as households move towards the constraint, while it decreases with the magnitude of a positive shock.
+Their pass-through rates are slightly higher than ours, though their model does not feature other insurance over self-insurance.
+Guvenen et al.
+(2024) simulate a model with retirement, stochastic lifetimes, and altruism.
+The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, higher than ours but lower than BPP.
+They show that measuring pass-through as in BPP overestimates the true rate by 1.5 times.
+Without non-employment shocks, agents are willing to give up almost about 22.3% of consumption to live in a world without risk.
+We read these results as not too different from ours.
+Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow, so those shocks end up affecting consumption by relatively more than other shocks.
+The magnitude of our pass-through rates seems uniformly lower than theirs (to some extent this is because these models focus on self-insurance), though other quasi reduced-form studies using the recent PSID, most notably ABB, find similar rates to ours.
 6 Welfare costs of tail income risk
-In the previous section, we established that income risk is negatively skewed and highly leptokurtic, and features a nonlinear pass-through to consumption.
-In this section, we calculate the welfare costs of such risk, particularly contrasting the implications of the linear and nonlinear consumption functions.
-To do so, we first draw a sequence of income shocks over the lifecycle, we then use the different consumption specifications to simulate responses to those shocks, and, finally, we use a parametric utility function to calculate household welfare.
-Our interest is in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
-We draw permanent and transitory income shocks for 50,000 households from mixtures of normal distributions, namely
+In this section, we calculate the welfare costs of income risk, which, as we found previously, is negatively skewed, highly leptokurtic, and features a nonlinear pass-through to consumption.
+We are particularly interested in contrasting the implications of linear and nonlinear transmission.
+To do so, we draw a sequence of income shocks over the lifecycle, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare.
+Our interest lies in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
+Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021, 2024).
+We draw permanent and transitory income shocks for 50,000 households from a mixture of two normal distributions, namely
 $\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob.
@@
 }1-p_{v}.\\ \end{cases}$
-Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021).
-There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of the shocks’ distributions in the PSID (column 3, table 3).
+There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2).
 We assume $\mu_{1}<0$ for identification.
@@
 These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock).
-In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate directly in our baseline sample.
+In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data.
 We simulate consumption over a working life of 35 years, starting at age 30.
-We also simulate a retirement period of 20 more years (starting at age 65), in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
-[^46] We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample.
+In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
+[^44] We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
+We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample.
 In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample.
-Our baseline simulations use the time-invariant partial insurance parameters from table 4.
-Appendix table D.3 relaxes time-invariance.
-We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion (CRRA) utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, or logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$.
+We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, and logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$.
 We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk.
-This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation
-${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}.$
-${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles, respectively reflecting less (Gaussian) and more (non-Gaussian) income risk.
+This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation ${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}$.
+Here, ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles that reflect less (Gaussian) and more (non-Gaussian) income risk, respectively.
 We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
 In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively.
-The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of a linear consumption function.
+The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission.
 Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
@@
 Here $\alpha_{i}$ reflects both the welfare cost of tail risk and the bias induced by neglecting the possibility of nonlinear transmission of such risk.
-Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shock they experience, in the case of $\gamma=2$.
+Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shocks they experience, in the case of $\gamma=2$.
 The first row presents the average cost across all households, unconditional on a given shock.
@@
 Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance.
-As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the empirical variance of shocks (or the variance of Gaussian risk).
+As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the variance of Gaussian risk.
 For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
@@
 % households# (1) (2) (3)
-All households 100% -0.78 -0.86 0.65
-$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.74 6.31 7.71
-$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.28 8.53 9.90
-$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.43 11.74 13.06
+All households 100% -0.81 -0.88 0.66
+$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
+$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
+$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
 Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$.
@@
 #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
-The next rows zoom in on households who experience large negative permanent shocks at least once in their lifetime.
+The next rows zoom in on households who experience large negative permanent shocks at least once in their life.
 These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them.
-Those who face a permanent shock at least three standard deviations below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.74% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly.
-This increases to 6.28% and 8.43% for shocks respectively four and five standard deviations below the mean – all shown in column 1.
-While these costs are already sizable, they increase further by 55% to 62% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.06%) for those households who experience the worst shocks.
-Those estimates would be even larger if we considered the welfare cost of tail risk relative to a zero risk scenario, as, e.g., in Guvenen et al.
+Those who face a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.8% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly.
+This increases to 6.36% and 8.54% for shocks respectively $4\sigma$ and $5\sigma$ below the mean – all shown in column 1.
+These costs are already sizable, and they increase further by about 60% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.23%) for those who experience the worst shocks.
+Those costs would be even larger if we considered a counterfactual zero risk scenario as in Guvenen et al.
 (2024).
-The welfare estimates are robust to the choice of $\gamma$ (appendix table D.2) or the use of age-specific partial insurance parameters when simulating various consumption profiles (appendix table D.3).
-Tail income risk consistently implies large welfare costs.
-The econometrician will greatly underestimate such costs if she neglects the nonlinear transmission into consumption, for which we found ample evidence in the earlier sections.
+These welfare calculations are ‘back-of-envelope’, as they neglect heterogeneity in income risk and consumption transmission, they treat wealth as exogenous, and they do not account for borrowing constraints.
+Appendix table D.2 repeats the calculations allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline.
+Appendix table D.3 calculates the welfare costs by wealth; the least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs, as expected.
+As the likelihood that a borrowing constraint binds depends on wealth or age, the age- and wealth-specific results are indirectly reflective also of such constraints.
+Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs lie in the middle of the wealth-specific costs, so exogenous wealth is unlikely to have caused a serious bias in our calculations.
+The welfare estimates are also robust to small variations in the choice of $\gamma$ (appendix table D.4).
 7 Conclusion
@@
 Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers.
-Our focus is on measuring partial insurance without taking a stance on the transmission mechanisms from income to consumption.
-We use recent income and consumption data from the PSID (1999–2019) and establish several new insights.
+We use recent data from the PSID and establish several new findings.
 First, permanent and transitory shocks to disposable household income exhibit negative skewness and very large excess kurtosis.
-Second, in estimating BPP over the recent years, we find a lower pass-through of permanent shocks (larger partial insurance) and full insurance to transitory ones.
+Second, in estimating a linear consumption model over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones.
 We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
 By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy.
+The nature of income risk and consumption transmission vary by age and the household’s position in the income distribution.
 Fifth, tail risk incurs substantial welfare costs, up to 1/8 of lifetime consumption.
-In summary, the paper offers an analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and severity of the shock.
-To measure the transmission, we exploit information in higher-order income moments but we do not otherwise need to observe the entire consumption-income joint distribution (e.g., in contrast to Arellano et al., 2017).
-Our results reconcile the traditional partial insurance measurement literature, in which BPP is the default methodology, with both the quantitative literature (e.g., Guvenen et al., 2024) and the more empirical one based on surveys or experiments (e.g., Fuster et al., 2020), both of which document asymmetries in the pass-through of good and bad, or small and large shocks.
-Overall, our results suggest that tail income risk matters substantially for consumption.
+In summary, the paper offers a straightforward analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and size of the shock.
+To do so, we exploit higher-order income moments but we do not otherwise need to observe the entire consumption-income joint distribution.
+Our results consistently point to asymmetries in the pass-through of good and bad shocks, or by age and income level, suggesting that tail income risk matters substantially for consumption.
 Appendix A Derivation of the consumption function
@@
 We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path.
-Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of log consumption on appropriate observables (demographics, time fixed effects).
+Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of consumption growth on appropriate observables (demographics, time fixed effects).
 A.2 Taylor approximation to lifetime budget constraint
@@
 $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C}){\operatorname*{\mathbb{E}}}_{t}(\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s})(\ln C_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+\ell})$
-$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})(\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1})^{2}=0$
+$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})(\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1})^{2}=0.$
 We derive this by noting that future expectations errors are zero by assumption, and that $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})=\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}\delta_{ss}+\sum_{s=0}^{T-t}\sum_{\ell\neq s}\vartheta_{it+s}^{C}\delta_{s\ell}-(\sum_{s=0}^{T-t}\vartheta_{it+s}^{C})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{C})=0$.
@@
 $y_{it}^{*}=y_{it}+u_{it}^{y}\qquad\text{and}\qquad c_{it}^{*}=c_{it}+u_{it}^{c}.$
-The moments of income measurement error are not separately identifiable from the moments of the transitory shock (Blundell et al., 2008).
+The moments of income measurement error are not separately identifiable from the moments of the transitory shock.
 This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
@@
 respectively. These are the theoretical expressions we bring to the data.
-Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks that comprise a two-year period.
-We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distribution is stationary.
+Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks within a two-year period.
+We thus assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t-1}}^{2}$, $\gamma_{\zeta_{t}}\approx\gamma_{\zeta_{t-1}}$, and $\kappa_{\zeta_{t}}\approx\kappa_{\zeta_{t-1}}$, which are exactly true if the underlying distributions are stationary.
 Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$.
@@
 B.1 Income process parameters
-The second and third moments of shocks are then given by
+The second and third moments of shocks are given by
 $\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
@@
 Given the income process, the quadratic consumption function, and the properties of income shocks, the transmission parameters are given by
-$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$
+$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$ (B.1)
 where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks.
@@
 where the first inequality follows from the lower bound on the kurtosis.
-The term inside the squared brackets in the determinant is thus strictly positive, and $\text{det}(\mathbf{A})<0$.
+The term inside the squared brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$.
 The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli.
-[^2] If the distributions of shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$, so the determinant would not always be strictly away from zero.
-An interesting special case arises when the distributions of permanent and transitory shocks are normal and have the same variance, and there is no measurement error.
-In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}<0$ (where $\sigma_{t}$ is the dispersion of the generic distribution), $\phi_{t}^{(1)}=(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$.
-Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of a typical permanent shock away from the mean, is separately identified from $\phi_{t}^{(1)}$.
-Therefore, identification does not require the shocks to be non-Gaussian.
-Despite normality in this case, tail shocks are still associated with a different pass-through from that of average shocks.
+[^2] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant would not always be strictly away from zero.
+How do different levels of skewness and kurtosis connect with the transmission parameters?
+Consider the transmission of permanent shocks, abstracting from transitory shocks (i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$).
+$\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$, which corresponds to the first equation in (B.1).
+Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks.
+Suppose $\phi^{(1)}_{t}$ is fixed.
+We highlight three cases:
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$.
+Letting $\gamma_{\zeta_{t}}<0$ as in the data, this equation restricts $\phi^{(2)}_{t}=0$.
+Tail shocks thus transmit similarly to non-tail ones, i.e., this configuration identifies a linear consumption process.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$.
+Consumption growth and permanent shocks co-vary more strongly than in the linear consumption process, i.e., there is excess co-movement.
+For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks.
+Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement.
+Ceteris paribus, a less skewed distribution of shocks ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement.
+If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$.
+Consumption growth and permanent shocks co-move less strongly than in the linear process, i.e., there is excess smoothness.
+For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks.
+These are the largest shocks to income, so muting their pass-through suppresses the covariance between consumption and income growth.
+The opposite is true if $\gamma_{\zeta_{t}}>0$.
+Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1).
+Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and squared shocks so it captures the pass-through of shocks of large magnitude.
+We highlight again three cases (for $\phi^{(1)}_{t}$ fixed):
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$.
+This identifies $\phi^{(2)}_{t}=0$, so tail shocks transmit similarly to non-tail ones.
+In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is directly implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\zeta_{it-1})$, so this configuration identifies the linear consumption process.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$ is negative and large.
+This suggests that large shocks (of which negative shocks are the most likely) transmit strongly with a negative coefficient.
+It thus identifies $\phi^{(2)}_{t}<0$.
+Ceteris paribus, a less leptokurtic income distribution (lower $\kappa_{\zeta_{t}}$) turns $\phi^{(2)}_{t}$ more negative; a bell-shaped distribution is associated with fewer large shocks, thus necessitating a stronger transmission of said shocks to maintain a given excess co-movement.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$.
+This identifies $\phi^{(2)}_{t}>0$ for analogous reasons.
+For given excess co-movement, the smaller the kurtosis of permanent shocks, the larger the value of $\phi^{(2)}_{t}$ must be to achieve said excess co-movement.
+In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by this and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity.
+The sensitivity analysis is thus more nuanced so simulations can prove useful.
+Starting from the baseline estimates of the income process (table 2), reducing the left skewness of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by almost 10% (more negative).
+Independently reducing the kurtosis of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by 60% (more negative).
+The response of $\phi^{(1)}_{t}$ is much smaller, confirming that the values of skewness and kurtosis are crucial for identifying the pass-through of tail shocks.
+An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error.
+In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the generic distribution.
+Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$.
+Therefore, identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks.
+Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks.
 Analogous expressions separately identify $\psi_{t}^{(1)}$ and $\psi_{t}^{(2)}$.
 Upon identification of the transmission parameters, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2})$ identifies $\sigma_{u^{c}_{t}}^{2}$, while ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ identifies $\sigma_{\xi_{t}}^{2}$.
-Higher moments of the consumption error or of $\xi$ involve higher than fourth-order moments of income, which we do not model.
+Higher moments of the consumption error or of the taste shock involve higher than fourth-order moments of income, which we do not model.
 B.4 Targeted moments
 There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model.
-Table B.1 lists the moments targeted in the various specifications of the income process, while table B.2 lists the moments targeted in the different specifications of the consumption model.
-For simplicity, we assume stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters.
-These are not identifying assumptions but they make the illustration below more compact by allowing us to pool common terms together.
+Table B.1 lists the income moments targeted in each case, while table B.2 lists the moments targeted in the different specifications of the consumption model.
+For simplicity, we illustrate the next statements assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters.
+These are not identifying assumptions but they make the illustration more compact by allowing us to pool common terms together.
 Table B.1: Targeted moments of income
@@
 $+12(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+24(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}$
-$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-$+\kappa_{u^{c}}+3(\sigma_{u^{c}}^{2})^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}$
+$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
@@
 Appendix C Empirical details and additional results
-C.1 Comparison with administrative moments
-Table C.1 reports the empirical second, third, and fourth moments of the cross-sectional distribution of unexplained growth in male earnings (panel A) and household earnings (panel B), as opposed to the moments of household disposable income in table 2 in the paper.
-Column 1 presents moments from our baseline sample in the PSID, while column 2 presents the corresponding statistics, wherever available, from the U.S.
-Social Security Administration data reported in Guvenen et al.
+C.1 Summary statistics in the PSID sample
+Table C.1 presents summary statistics in the PSID sample over the period 1999–2019 (columns 1–2), as well as in subsamples sorted by wealth and education (columns 3–6).
+Table C.1: Descriptive statistics, PSID
+Low wealth High wealth
+Baseline sample No college Some college No college Some college
+mean median mean mean mean mean
+(1) (2) (3) (4) (5) (6)
+Consumption
+Nondurable+services 26,766 22,895 21,280 24,566 23,675 31,479
+Food at home 8,256 7,609 7,438 8,210 7,505 8,870
+Utilities 3,515 3,238 3,351 3,205 3,471 3,787
+Food out 3,318 2,511 2,387 2,928 3,109 4,037
+Health 1,614 863 1,321 1,445 1,623 1,850
+Transportation 5,866 4,900 5,283 5,489 5,711 6,406
+Education 3,202 0 957 1,986 1,810 5,352
+Childcare 977 0 546 1,259 537 1,136
+Rent (or rent eq.) 18,105 13,821 8,910 12,631 16,449 26,020
+Home insurance 894 747 585 712 908 1,142
+Income
+Disposable income 142,910 113,581 84,414 110,176 120,931 195,158
+Taxable+transfer inc. 137,277 109,003 77,707 106,276 113,789 189,446
+Earnings 118,423 97,154 71,504 99,798 90,950 158,616
+Male 81,857 62,185 46,025 65,725 57,519 114,545
+Female 36,567 29,183 25,479 34,074 33,430 44,070
+Wealth 547,761 191,193 68,561 70,665 670,948 1,024,577
+Demographics
+% working (male) 0.89 1.00 0.89 0.94 0.82 0.89
+% working (female) 0.79 1.00 0.79 0.82 0.78 0.78
+Age (male) 46 45 44 42 50 48
+% some college (male) 0.68 1.00 0.00 1.00 0.00 1.00
+% some college (female) 0.71 1.00 0.42 0.79 0.48 0.87
+Periods per household 5.63 5.00 4.90 4.88 5.86 6.62
+Observations 20,866 4,258 5,159 2,428 9,021
+Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 19991999199919992019201920192019.
+Monetary amounts are expressed in \$2018.
+Rent for owners is calculated at 6% of the self-reported value of the house.
+C.2 Comparison with administrative moments
+Table C.2 reports the empirical second, third, and fourth moments of the cross-sectional distribution of unexplained growth in male earnings (panel A) and household earnings (panel B), as opposed to the moments of household disposable income in table 1.
+Column 1 presents moments in the PSID sample, while column 2 presents the corresponding statistics, wherever available, from the U.S.
+Social Security Administration data of Guvenen et al.
 (2021).
@@
 Skewness and kurtosis correspond to the third and fourth standardized moments respectively.
-Column 1 reports moments of biennial earnings growth in our baseline sample in the PSID (we maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the PSID, this corresponds to a difference over two calendar years in this case), while column 2 reports moments of male earnings growth from the U.S.
+Column 1 reports moments of biennial earnings growth in the PSID (we maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the PSID, this corresponds to a difference over two calendar years), while column 2 reports moments of male earnings growth from the U.S.
 Social Security Administration data.
-The latter moments concern the one-year male earnings growth, unconditional on past earnings, and correspond to averages over ages 30-55 of, respectively, dispersion squared, skewness, and kurtosis, as reported in online appendix tables C23-C25 in Guvenen et al.
+The latter concern the one-year male earnings growth, unconditional on past earnings, and correspond to averages over ages 30-55 of, respectively, dispersion squared, skewness, and kurtosis, as reported in appendix tables C23-C25 in Guvenen et al.
 (2021).
 Block bootstrap standard errors are in parentheses.
-C.2 Sources of tail income shocks
+C.3 Sources of tail income shocks
 Table C.3 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis.
-Columns 1 & 2 report the empirical probability that unemployment, disability, bad health, etc, occur in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively.
-Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) also experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.
+Columns 1 & 2 report average unemployment, disability, bad health, etc, in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively.
+Column 3 reports the $p$-value of the difference between the two; in most cases, the bottom 10% (effectively households facing large income reductions) experience unemployment, worsening health, large drops in hours, and precarious housing at much higher rates than the rest.
 Table C.3: Sources of tail income shocks
-Household income growth
-bottom 10% top 90% difference
-empirical probabilities $p$-value
+Disposable income growth
+Bottom 10% Top 90%
+mean mean $p$-value $\Delta$mean
 (1) (2) (3)
-Unemployed $t$ (male) 0.082 0.034 0.000
-Unemployed $t-1$ (male) 0.084 0.036 0.000
-Unemployed $t$ (female) 0.081 0.037 0.000
-Unemployed $t-1$ (female) 0.081 0.041 0.000
-Main job laid off (male) 0.054 0.039 0.003
-Second job laid off (male) 0.153 0.117 0.000
-Main job laid off (female) 0.073 0.059 0.023
-Second job laid off (female) 0.096 0.095 0.921
-Physical disability (male) 0.130 0.103 0.001
-Physical disability (female) 0.121 0.120 0.840
-Physical disability (both) 0.032 0.025 0.077
-Overall health got worse (male) 0.129 0.110 0.036
-Overall health got worse (female) 0.121 0.104 0.052
-Eviction from home 0.049 0.031 0.000
+% unemployed $t$ (male) 0.082 0.034 0.000
+% unemployed $t-1$ (male) 0.084 0.036 0.000
+% unemployed $t$ (female) 0.081 0.037 0.000
+% unemployed $t-1$ (female) 0.081 0.041 0.000
+% main job laid off (male) 0.054 0.039 0.003
+% second job laid off (male) 0.153 0.117 0.000
+% main job laid off (female) 0.073 0.059 0.023
+% second job laid off (female) 0.096 0.095 0.921
+% physical disability (male) 0.130 0.103 0.001
+% physical disability (female) 0.121 0.120 0.840
+% physical disability (both) 0.032 0.025 0.077
+% overall health got worse (male) 0.129 0.110 0.036
+% overall health got worse (female) 0.121 0.104 0.052
+% eviction from home 0.049 0.031 0.000
 Hours growth (male) -0.309 0.031 0.000
 Hours growth (female) -0.228 0.023 0.000
-Notes: The table reports the empirical probability of the various events listed in the rows of the table in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$.
+Notes: The table reports the sample mean (mostly empirical probabilities) of the various events listed in the rows of the table in the bottom 10% and top 90% of the distribution of residual income growth $\Delta y_{it}$.
 Given the biennial nature of the PSID, $\Delta y_{it}$ corresponds to a difference over two calendar years.
-C.3 CEX data and imputation in the PSID
-The Consumer Expenditure Survey accounts for about $95\%$ of all household expenditures from a highly disaggregated list of consumption goods and services (see Meyer and Sullivan, 2023, for a recent overview).
-The CEX is run by the Census Bureau and the Bureau of Labor Statistics, and it forms the basis for the calculation of the consumer’s price index.
+C.4 CEX data and imputation in the PSID
+The Consumer Expenditure Survey (CEX) accounts for about $95\%$ of all household expenditures from a highly disaggregated list of consumption goods and services (see Meyer and Sullivan, 2023, for a recent overview).
+The CEX is run by the Census Bureau and the Bureau of Labor Statistics and forms the basis for the calculation of the consumer’s price index.
 Sample selection and variables.
-To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019.333To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID.
+To parallel the design of the PSID sample, we use CEX interview data between 1999 and 2019.
 We select a sample of households that mimics closely our baseline selection in the PSID.
+[^3] To improve the quality of the food demand estimation subsequently, we keep all calendar years between 1999 and 2019, i.e., not just the odd ones as in the PSID.
 Specifically, we select continuously married couples with the male spouse aged 30 to 65.
-We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the consumption imputation subsequently.
-Similar to the consumption measure in the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare.
+We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the imputation subsequently.
+Similar to the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare.
 As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters.
@@
 Figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID (blue solid line) and CEX samples (black dotted line).
-Although the patterns for the variance and kurtosis are quite similar in the two surveys over time, consumption in the PSID is less volatile relative to the CEX but slightly more leptokurtic.
-For skewness, there is disagreement both in the trends over time and, importantly, with respect to the sign, with the PSID and CEX samples featuring positive and weakly negative skewness, respectively.
-Overall, these statistics (in particular the skewness statistics) suggest that consumption across otherwise similarly selected samples in the PSID and CEX exhibits notable differences; this calls into question any conclusions on the pass-through drawn from methods that use higher-order moments of consumption to identify the response to income shocks.
+The patterns for the variance and kurtosis are quite similar in the two surveys over time, although consumption in the PSID is less volatile but slightly more leptokurtic.
+For skewness, there is disagreement both in the trends over time and the sign.
+The PSID features positive while the CEX weakly negative skewness.
+Overall, these statistics (in particular skewness) suggest that consumption across otherwise similarly selected samples in the PSID and CEX exhibits some notable differences.
+Figure C.1: Moments of log consumption in the PSID and CEX
+(a) Variance of log consumption
+(b) Skewness of log consumption
+(c) Kurtosis of log consumption
+Notes: The figure plots the moments of log consumption in the PSID (blue solid line; left), the CEX (black dotted line; right), and imputed from the CEX into the PSID (black dashed-dotted line; right axis).
 Consumption imputation from CEX into PSID.
@@
 We thus follow BPP and impute nondurable consumption from the CEX into the PSID.
-Figure C.1: Moments of log consumption in the PSID and CEX
-(a) Variance of log consumption
-(b) Skewness of log consumption
-(c) Kurtosis of log consumption
-Notes: The figure plots the moments of log consumption in the PSID (blue solid line; left axis) and the CEX (black dotted line for the true series and black dashed-dotted line for the imputed series; right axis).
 The imputation involves two distinct steps.
 In the first step, we estimate food demand in the CEX.
-Specifically, we let real food expenditure be a log-linear function of the log of nondurable expenditure $c$, relative prices $\mathbf{p}$, and demographics $\mathbf{W}$, namely
+Specifically, we let real food expenditure be a log-linear function of the log of nondurable expenditure $c_{it}$, relative prices $\mathbf{p}_{t}$, and demographics $\mathbf{W}_{it}$, namely
 $f_{it}=\beta(D_{it})c_{it}+\mathbf{p}_{t}^{\prime}\boldsymbol{\gamma}+\mathbf{W}_{it}^{\prime}\boldsymbol{\mu}+e_{it},$ (C.1)
 where $f_{it}$ is the log of real annual food expenditure for household $i$ in the CEX in year $t$, $\beta$ is the budget elasticity of food demand, which we allow to shift with time and household characteristics $D_{it}$, and $e_{it}$ captures unobserved heterogeneity in food demand.
-[^5] Prices $\mathbf{p}$ include the relative prices of food, public transportation, utilities, medical services, and childcare.
-Demographics $\mathbf{W}$ include a quadratic polynomial in age and dummy variables for the number of children, family size, education, region of residence, race, and cohort.
+[^5] Prices $\mathbf{p}_{t}$ include the relative prices of food, public transportation, utilities, medical services, and childcare.
+Demographics $\mathbf{W}_{it}$ include a quadratic polynomial in age and dummy variables for the number of children, family size, education, region of residence, race, and cohort.
 $D_{it}$ includes year dummies and dummies for education and the number of children.
-To address possible endogeneity of and measurement error in total expenditure, we follow an instrumental variables approach using as instruments the mean and standard deviation of household disposable income and the standard deviation of spousal wages by cohort, education, year, the former interacted with time, education, and number of children dummies.
-[^6] These instruments are slightly different from those used in BPP, namely the cohort-education-year specific averages of male and female hourly wages, interacted with time, education, and kids dummies.
+To address possible endogeneity of nondurable expenditure, we follow an instrumental variables approach using as instruments the mean and standard deviation of household disposable income and the standard deviation of spousal wages by cohort, education, year, the former interacted with time, education, and number of children dummies.
+[^6] These instruments are different from those used in BPP, namely the cohort-education-year specific averages of male and female hourly wages, interacted with time, education, and kids dummies.
 The use of their instruments over our sample period often resulted in convergence problems.
-In the second step, we invert (C.1) under the assumption that food demand is monotonic in total expenditure.
-We can then use the inverted equation to predict nondurable expenditure $c$ for a given level of food expenditure, given prices and demographics.
+In the second step, we invert (C.1) under the assumption that food demand is monotonic in nondurable expenditure.
+We can then use the inverted equation to predict $c_{it}$ for a given level of food expenditure, given prices and demographics.
 Since the CEX and PSID samples are representative of the same underlying population, and since food expenditure, prices, and demographics are available in both surveys, we can use the inverted (C.1) to impute consumption into the PSID.
@@
 Our empirical findings suggest that estimation of the linear specification using imputed consumption data inflates $\phi_{t}^{(1)}$ relative to estimation using the true PSID series.
-One reason behind this (though likely not the only one) is neglected correlation between measurement error in income and expenditure.
+One reason behind this is neglected correlation between measurement error in income and expenditure.
 Suppose that consumption and income measurement errors are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{c},u_{it}^{y})=\sigma_{u_{t}}>0$.
@@
 However, if the errors are correlated, then this target moment in fact identifies $\phi_{t}^{(1)\text{psid*}}=\phi_{t}^{(1)}+\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$.
-Clearly, correlation between consumption and income errors causes positive bias even in the baseline, i.e., without the use of imputed consumption, as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$.
-We show below that this bias is multiplicatively larger when one uses imputed consumption.
+Clearly, correlation between consumption and income errors causes positive bias as $\phi_{t}^{(1)\text{psid*}}>\phi_{t}^{(1)}$.
+This is true in the baseline, i.e., without the use of imputed consumption, but the bias becomes multiplicatively larger with imputed consumption, as we show below.
 In the simplest form without covariates, imputed log consumption in the PSID is given by $c_{it}^{\text{imp}}=\kappa_{t}(f_{it}-\pi_{t})$, where $f_{it}$ is log food expenditure in the PSID, $\pi_{t}=\mathbf{p}_{t}^{\prime}\boldsymbol{\gamma}$ is an intercept that depends on prices, and $\kappa_{t}=\beta(D_{it})^{-1}$ is the inverse budget elasticity of food consumption.
@@
 However, they do not discuss measurement error in food expenditure in the PSID, and in particular the case when such error may correlate with error in income.
-Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{c})$.
-We assume for simplicity that measurement error in food expenditure is the same as measurement error in nondurable consumption.
-This is not crucial for what follows; what is crucial is that measurement error in income varies jointly with measurement error in food expenditure in a similar way as it does with measurement error in nondurable consumption.
+Suppose that food expenditure in the PSID is measured with error, so imputed consumption is $c_{it}^{\text{imp*}}=\kappa_{t}(f_{it}+u_{it}^{f})$.
+Suppose that food consumption and income error are positively correlated, i.e., ${\operatorname*{\mathrm{Cov}}}(u_{it}^{f},u_{it}^{y})=\sigma_{u_{t}}>0$.
 Because $\kappa_{t}>1$, the consumption imputation magnifies the role of the consumption measurement error.
-One can show that, with imputed consumption, the moment that is used to identify $\phi_{t}^{(1)}$ in fact identify $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$.
-As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation in this environment inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
-C.4 Comparison using the original BPP data
+One can show that, with imputed consumption, the moment used to identify $\phi_{t}^{(1)}$ in fact identifies $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$.
+As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
+C.5 Comparison using the original BPP data
 Table C.5 uses the original BPP data over 1980–1992 (taken from the replication package of Blundell et al., 2008, available online) and does three things.
@@
 With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years.
-Some higher frequency shocks will be muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period.
+Some higher frequency shocks are muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period.
 In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
@@
 The results are in columns 3-4 in table C.5.
-Time-invariance further reduces the transmission parameter of permanent shocks ($\phi^{(1)}$ drops by another 20% to $0.404$), bringing the pass-through closer to the value we estimate in the modern data.
+Time-invariance further reduces the transmission parameter of permanent shocks ($\phi^{(1)}$ drops to $0.404$), bringing the pass-through closer to the value we estimate in the modern data.
 Taking averages of moments over time attenuates the largest income-consumption co-movements; consequently, the model can fit the data using, ceteris paribus, smaller transmission parameters.
 This is similar to Chatterjee et al.
-(2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in the BPP target moments.
+(2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in BPP’s target moments.
 There are some caveats comparing the exercise in table C.5 and our results with the most recent data in table 3.
@@
 We showed earlier that the imputation imparts substantial imputation error.
-Second, while the estimation in table C.4 does not correct the income moments for measurement error (in line with BPP), our main exercise in table C.4 does so (in line with Blundell et al., 2016).
+Second, while the estimation in table C.5 does not correct the income moments for measurement error (in line with BPP), our main exercise in table 3 does so (in line with Blundell et al., 2016).
 Finally, the use of biennial growth rates does not allow the identification of $\theta$, the moving average parameter of the transitory shock, so the transitory shock is $MA(0)$ in our baseline ($\theta=0$), while it is $MA(1)$ in the original BPP exercise.
@@
 Columns 1 and 3 estimate the model using annual growth rates of income and consumption, while columns 2 and 4 use biennial rates.
-In line with BPP, estimation is done via diagonally weighted GMM; asymptotic standard errors are in parentheses.
+In line with BPP, estimation is done via diagonally weighted GMM; GMM standard errors are in parentheses.
+The very small unimportant numerical differences between column 1 herein and the original BPP are likely due to different software and optimization algorithms (we use Matlab and fmincon).
 #$\theta$ is the moving average parameter of the transitory shock.
@@
 Our use of biennial data in this paper rules out identification of $\theta$, which we thus set to 0.
-C.5 Additional results
-Tables C.5 and C.6 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM).
-As such, these tables accompany tables 3 and 4 in the paper, which report results from equally weighted GMM.
-The results are qualitatively (and most often numerically) unchanged from those reported in the paper.
+C.6 Additional results
+Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM).
+As such, these tables accompany tables 2 and 3 in the paper, which report results from equally weighted GMM.
+The results are qualitatively (and most often numerically) similar to those reported in the paper.
 Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX.
@@
 Similar to our baseline, we also find $\psi^{(2)}>0$.
-Table C.5: Estimates of the income process, baseline sample, optimal weighting
+Table C.9 presents the estimates of the parameters of the income process by age and the household’s position in the income distribution.
+Skewness of all shocks becomes more negative, both economically and statistically, with age.
+Skewness of permanent shocks is weakly positive in the bottom income tertile but turns very negative in the top tertile.
+The opposite is true for transitory shocks; skewness is negative in the bottom income tertile but strongly positive in the top.
+Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution.
+Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65.
+A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted.
+This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results.
+[^7] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters.
+By contrast, when they generalize this to a quadratic trend, they report a worsening of their results without including further details.
+By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
+Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, possess more insurance than the poor.
+Table C.6: Estimates of the income process, optimal weighting
 Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
@@
 Panel A. Permanent shocks
-Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.025
-(0.002) (0.002) (0.002)
-Skewness $\gamma_{\zeta}$ central -0.004 -0.004
-(0.003) (0.003)
-standardized -0.895 -0.997
-(0.633) (0.760)
-Kurtosis $\kappa_{\zeta}$ central 0.037
-standardized 61.636
+Variance $\sigma_{\zeta}^{2}$ 0.027 (0.002) 0.027 (0.002) 0.025 (0.002)
+Skewness $\gamma_{\zeta}$ central -0.004 (0.003) -0.004 (0.003)
+standardized -0.895 (0.633) -0.997 (0.760)
+Kurtosis $\kappa_{\zeta}$ central 0.037 (0.007)
+standardized 61.636 (14.018)
 Panel B. Transitory shocks
-Variance $\sigma_{v}^{2}$ 0.033 0.033 0.034
-(0.002) (0.002) (0.002)
-Skewness $\gamma_{v}$ central -0.008 -0.008
-(0.003) (0.003)
-standardized -1.351 -1.324
-(0.517) (0.485)
-Kurtosis $\kappa_{v}$ central 0.049
-standardized 41.579
+Variance $\sigma_{v}^{2}$ 0.033 (0.002) 0.033 (0.002) 0.034 (0.002)
+Skewness $\gamma_{v}$ central -0.008 (0.003) -0.008 (0.003)
+standardized -1.351 (0.517) -1.324 (0.485)
+Kurtosis $\kappa_{v}$ central 0.049 (0.005)
+standardized 41.579 (7.052)
 Notes: The table presents the estimates of the parameters of the income process, imposing stationarity over time.
@@
 This table accompanies table 2 in the paper, which shows the baseline results from equally weighted GMM.
-Table C.6: Estimates of the consumption function, baseline sample, optimal weighting
+Table C.7: Estimates of the consumption function, optimal weighting
 Consumption fn.: Linear Quadratic
@@
 This table accompanies table 3 in the paper, which shows the baseline results from equally weighted GMM.
-Table C.7: Estimates of the consumption function, baseline sample, imputed consumption
+Table C.8: Estimates of the consumption function, imputed consumption
 Consumption fn.: Linear Quadratic
@@
 Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
-This table accompanies table 4 in the paper, which shows the baseline results using consumption internally available in the PSID.
-Table C.8 presents estimation results for the linear consumption function in subsamples formed by wealth and education.
+This table accompanies table 3 in the paper, which shows the baseline results using consumption internally available in the PSID (column 2 in table 3 is the same as column 1 in the present table).
+Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education.
 As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification.
-Columns 1-4 present parameter estimates among low and high wealth households.
-$\phi^{(1)}$ is consistently smaller among the wealthier, regardless of which set of moments are targeted, indicating the self-insurance role of assets in that group.
-$\psi^{(1)}$ is statistically indistinguishable from zero in both groups.
-Columns 5-8 present parameter estimates among households with and without college education.
-The results do not present a clear pattern across groups.
-When only second moments of income and consumption are targeted, $\phi^{(1)}$ is larger for households without college education – $0.225$ versus $0.123$.
-However, when all moments up to fourth-order are included, $\phi^{(1)}$ is almost identical in the two groups.
+The odd numbered columns present parameter estimates when only second-order moments are targeted.
+$\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across the groups.
+This is similar to what we found in the quadratic specification.
+However, as soon as we target up to fourth-order moments in the linear specification, shown in the even numbered columns, this monotonic pattern is lost.
+Transitory shocks transmit strongly in group 1, and with a negative coefficient.
+Table C.9: Estimates of the income process, by age and position in income distribution
+Age Position in income distribution $y_{it-1}$
+30–40 41–50 51–65 Bottom tertile Top tertile
+Income moments: 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th 2nd, 3rd, 4th
+Income data: PSID PSID PSID PSID PSID
+(1) (2) (3) (4) (5)
+Panel A. Permanent shocks
+Variance $\sigma_{\zeta}^{2}$ 0.028 0.026 0.034 0.037 0.032
+(0.004) (0.004) (0.005) (0.004) (0.003)
+Skewness $\gamma_{\zeta}$ central -0.003 0.004 -0.015 0.010 -0.028
+(0.006) (0.005) (0.006) (0.005) (0.004)
+standardized -0.653 0.926 -2.386 1.344 -4.851
+(1.291) (1.114) (1.167) (0.783) (0.987)
+Kurtosis $\kappa_{\zeta}$ central 0.043 0.030 0.058 0.052 0.039
+(0.014) (0.013) (0.016) (0.013) (0.008)
+standardized 56.275 45.537 50.945 37.627 36.974
+(23.787) (22.709) (21.605) (13.058) (10.586)
+Panel B. Transitory shocks
+Variance $\sigma_{v}^{2}$ 0.033 0.027 0.036 0.020 0.020
+(0.005) (0.004) (0.005) (0.004) (0.003)
+Skewness $\gamma_{v}$ central -0.008 -0.005 -0.010 -0.024 0.021
+(0.006) (0.005) (0.005) (0.005) (0.004)
+standardized -1.427 -1.178 -1.453 -8.246 7.333
+(1.057) (1.088) (0.782) (3.074) (2.235)
+Kurtosis $\kappa_{v}$ central 0.056 0.040 0.050 0.014 0.011
+(0.013) (0.007) (0.008) (0.008) (0.006)
+standardized 52.863 53.730 38.832 33.974 27.286
+(19.670) (16.619) (12.566) (24.531) (16.935)
+Notes: The table presents the estimates of the parameters of the income process, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse, and, separately, by the household’s position in the income distribution (bottom and top tertiles of the distribution of lag income $y_{it-1}$).
+We target all income moments up to fourth-order; results when targeting only lower order moments are available upon request.
+Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses.
+This table accompanies tables 4 and 5 in the paper, as well as the following tables C.10 and C.11.
+Table C.10: Estimates of the linear consumption function, by age
+Consumption fn.: Linear
+Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
+Consumption data: PSID PSID PSID
+(1) (2) (3)
+$\phi^{(1)}$ 30-40 0.154 (0.045) 0.164 (0.051) 0.089 (0.058)
+41-50 0.126 (0.051) 0.146 (0.055) 0.079 (0.089)
+51-65 0.180 (0.049) 0.175 (0.051) 0.136 (0.061)
+$\psi^{(1)}$ 30-40 0.094 (0.079) 0.083 (0.073) 0.095 (0.107)
+41-50 0.010 (0.079) -0.016 (0.072) -0.075 (0.061)
+51-65 -0.046 (0.070) -0.044 (0.066) 0.037 (0.101)
+$\sigma^{2}_{\xi}$ 30-40 0.014 (0.002) 0.014 (0.002) 0.016 (0.002)
+41-50 0.022 (0.002) 0.022 (0.002) 0.022 (0.002)
+51-65 0.021 (0.002) 0.021 (0.002) 0.023 (0.002)
+$\gamma_{\xi}$ 30-40 -0.001 (0.001) -0.001 (0.001)
+41-50 0.006 (0.001) 0.006 (0.002)
+51-65 -0.006 (0.002) -0.006 (0.002)
+$\kappa_{\xi}$ 30-40 0.004 (0.002)
+41-50 0.007 (0.002)
+51-65 0.006 (0.002)
+$\sigma^{2}_{u_{c}}$ 30-40 0.046 (0.003) 0.046 (0.003) 0.046 (0.003)
+41-50 0.040 (0.002) 0.040 (0.002) 0.040 (0.002)
+51-65 0.043 (0.002) 0.043 (0.002) 0.043 (0.002)
+$\gamma_{u_{c}}$ 30-40 0.000 (0.002) 0.000 (0.002)
+41-50 0.002 (0.002) 0.002 (0.002)
+51-65 0.004 (0.002) 0.004 (0.002)
+$\kappa_{u_{c}}$ 30-40 0.014 (0.003)
+41-50 0.012 (0.002)
+51-65 0.012 (0.002)
+Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary over ages 30–40, 41–50, and 51–65 of the male spouse.
+The underlying income moments also vary over these age brackets (table C.9).
+All columns use consumption data internally available in the PSID.
+Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
+This table accompanies table 4 in the paper, which shows results from the quadratic specification.
+Table C.11: Estimates of the linear consumption function, over position in income distribution
+Consumption fn.: Linear
+Lag income $y_{it-1}$: Bottom tertile Top tertile
+Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd 2nd, 3rd 2nd, 3rd, 4th
+Consumption data: PSID PSID PSID PSID PSID PSID
+(1) (2) (3) (4) (5) (6)
+$\phi^{(1)}$ 0.103 0.083 0.063 0.102 0.060 0.036
+(0.031) (0.032) (0.035) (0.036) (0.034) (0.041)
+$\psi^{(1)}$0.045 0.013 0.025 -0.004 0.101 0.071
+(0.086) (0.038) (0.054) (0.093) (0.048) (0.073)
+$\sigma^{2}_{\xi}$ 0.014 0.014 0.014 0.025 0.025 0.026
+(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
+$\gamma_{\xi}$0.001 -0.001 0.002 0.002
+(0.001) (0.001) (0.001) (0.001)
+$\kappa_{\xi}$ 0.003 0.009
+(0.002) (0.002)
+$\sigma^{2}_{u_{c}}$ 0.044 0.044 0.044 0.041 0.041 0.041
+(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
+$\gamma_{u_{c}}$ 0.003 0.003 0.003 0.003
+(0.002) (0.002) (0.002) (0.002)
+$\kappa_{u_{c}}$ 0.011 0.010
+(0.002) (0.002)
+Notes: The table presents the estimates of the parameters of the linear consumption function in the bottom and top tertiles of the distribution of lag income $y_{it-1}$.
+The underlying income moments also vary over the income distribution (table C.9).
+All columns use consumption data internally available in the PSID.
+Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses.
+This table accompanies table 5 in the paper, which shows results from the quadratic specification.
 Table C.12: Estimates of the linear consumption function, by wealth and education
 Consumption fn.: Linear
-By wealth By education
-Subsample: Low wealth High wealth No college Some college
+Low wealth High wealth
+Subsample: No college Some college No college Some college
 Income moments: 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th 2nd 2nd, 3rd, 4th
@@
 (1) (2) (3) (4) (5) (6) (7) (8)
-$\phi^{(1)}$ 0.223 0.148 0.111 0.099 0.225 0.107 0.123 0.113
+$\phi^{(1)}$ 0.242 0.143 0.198 0.153 0.171 0.028 0.099 0.106
 (0.053) (0.042) (0.053) (0.056) (0.071) (0.089) (0.041) (0.055)
-$\psi^{(1)}$0.021 0.047 0.011 0.018 -0.070 -0.041 0.024 0.031
+$\psi^{(1)}$0.178 -0.164 0.015 0.073 0.056 0.142 0.003 0.004
 (0.107) (0.077) (0.089) (0.091) (0.089) (0.131) (0.059) (0.071)
-$\sigma^{2}_{\xi}$ 0.014 0.015 0.023 0.023 0.014 0.016 0.021 0.022
+$\sigma^{2}_{\xi}$ 0.014 0.016 0.013 0.014 0.014 0.015 0.025 0.026
 (0.003) (0.003) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
-$\gamma_{\xi}$0.002 $0$0.002 0.000
+$\gamma_{\xi}$0.003 -0.001 -0.003 0.000
 (0.002) (0.001) (0.002) (0.001)
-$\kappa_{\xi}$ 0.003 0.007 0.003 0.006
+$\kappa_{\xi}$ 0.005 0.000 0.000 0.009
 (0.002) (0.002) (0.002) (0.002)
-$\sigma^{2}_{u_{c}}$ 0.048 0.048 0.042 0.042 0.043 0.043 0.044 0.044
+$\sigma^{2}_{u_{c}}$ 0.045 0.045 0.045 0.045 0.038 0.038 0.043 0.043
 (0.003) (0.003) (0.003) (0.003) (0.004) (0.004) (0.002) (0.002)
-$\gamma_{u_{c}}$ 0.003 0.003 0.003 0.002
+$\gamma_{u_{c}}$0.001 0.004 0.010 0.002
 (0.003) (0.002) (0.005) (0.001)
-$\kappa_{u_{c}}$ 0.015 0.012 0.015 0.012
+$\kappa_{u_{c}}$ 0.013 0.013 0.018 0.011
 (0.003) (0.002) (0.006) (0.002)
-Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary across wealth and education groups.
+Notes: The table presents the estimates of the parameters of the linear consumption function, allowing them to vary by wealth and education.
+The underlying income moments also vary by wealth and education (results not shown for brevity).
+We do not report results when up to third-order moments of income and consumption are targeted in the estimation; these are available upon request.
+Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019.
+No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling).
+All columns use consumption data internally available in the PSID.
+Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses.
 This table accompanies the quadratic specification estimates in table 6 in the paper.
-For reasons of brevity, we do not report results when second- and third-order moments of income and consumption are targeted in the estimation; these are available from the authors upon request.
-Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019.
-No college versus some college is defined on the basis of the highest level of education attained by the male spouse.
-All columns use consumption data internally available in the PSID.
-Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
 Appendix D Welfare costs details
-Table D.1 presents the parameter estimates and the model fit of the quantitative model for income shocks (mixture of normal distributions).
-Table D.2 reports average welfare costs for different values of the CRRA coefficient $\gamma$.
-Table D.3 calculates the welfare costs respecting the age-specific nature of the partial insurance parameters.
+Table D.1 presents the parameter estimates and model fit of the quantitative model for income (mixture of normal distributions).
+Table D.2 calculates the welfare costs allowing for heterogeneity in income risk and consumption transmission by age, while table D.3 allows for heterogeneity by wealth.
+Table D.4 reports average welfare costs for different values of the utility coefficient $\gamma$.
 Table D.1: Normal mixtures, parameter estimates and model fit
@@
 Estimation is done via the method of simulated moments, based on 50,000 random draws for each type of income shock.
+Table D.2: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific income risk and transmission parameters
+Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
+${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
+Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
+% households# (1) (2) (3)
+All households 100% -1.02 -1.06 0.31
+$\min\{\zeta_{it}\}<-0.50$ 39.18% 3.87 6.07 7.35
+$\min\{\zeta_{it}\}<-0.70$ 29.23% 5.14 8.29 9.54
+$\min\{\zeta_{it}\}<-0.95$ 18.28% 6.79 11.55 12.75
+Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$.
+We simulate consumption for 50,000 households using age-specific income parameters (table C.9) and consumption transmission parameters (table 4).
+The first row presents the average costs across all households.
+The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
+#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, by wealth
+Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$
+${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
+Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
+% households# (1) (2) (3)
+Panel A. Low wealth
+All households 100% -1.58 -1.80 6.14
+$\min\{\zeta_{it}\}<-0.50$ 40.96% 6.78 7.43 14.64
+$\min\{\zeta_{it}\}<-0.70$ 26.87% 9.61 10.63 17.60
+$\min\{\zeta_{it}\}<-0.95$ 12.99% 13.12 14.72 21.36
+Panel B. High wealth
+All households 100% -0.62 -0.69 -0.16
+$\min\{\zeta_{it}\}<-0.50$ 47.18% 3.06 4.59 5.09
+$\min\{\zeta_{it}\}<-0.70$ 35.12% 4.18 6.42 6.91
+$\min\{\zeta_{it}\}<-0.95$ 21.91% 5.73 9.06 9.54
+Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B).
+Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019.
+We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text.
+However, the consumption parameters can be inferred from table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample.
+The first row presents the average costs across all households in the wealth group.
+The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
+#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
 Table D.4: Welfare costs of non-Gaussian relative to Gaussian income risk, varying $\gamma$
@@
 Panel B. $\boldsymbol{\gamma=3}$
-All households 100% -0.72 -0.77 0.66
-$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.55 6.15 7.48
-$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.05 8.36 9.66
-$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.16 11.57 12.83
+All households 100% -0.81 -0.88 0.66
+$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
+$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
+$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
 Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B).
-We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 4.
+We simulate consumption for 50,000 households using the time-invariant consumption transmission parameters of table 3.
 The first row presents the average costs across all households.
@@
 #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
-Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific transmission parameters
-Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
-${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
-Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
-% households# (1) (2) (3)
-All households 100% -1.13 -1.15 -0.24
-$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.27 5.66 6.51
-$\min\{\zeta_{it}\}<-0.70$ 29.11% 5.72 7.71 8.54
-$\min\{\zeta_{it}\}<-0.95$ 17.07% 7.85 10.78 11.59
-Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$.
-We simulate consumption for 50,000 households using the age-specific transmission parameters of table 5.
-The first row presents the average costs across all households.
-The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.

### v5

@@
 Abstract
-We measure consumption insurance against income shocks accounting for high-order moments of the income distribution.
-We derive a nonlinear consumption function, where the extent of insurance varies with the shocks’ sign and magnitude.
-Using PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen.
-Households would sacrifice over 1/8 of lifetime consumption to eliminate tail income risk.
+We propose a simple framework to measure consumption insurance against income shocks that accounts for higher-order moments of the income distribution.
+We derive a nonlinear consumption function in which the extent of insurance varies with both the sign and magnitude of shocks.
+Using recent PSID data, we estimate an asymmetric pass-through of bad versus good permanent shocks – 17% of a $3\sigma$ negative shock transmits to consumption versus 9% of an equal-sized positive shock – with greater pass-through as shocks worsen.
+These consumption transmission rates further vary by age, wealth, and along the income distribution.
 Our results align with survey responses to hypothetical events and suggest that tail risk matters substantially for consumption.
@@
 How does consumption respond to income shocks?
-The answer to this question is important for determining the welfare effects of shifts in the income distribution, i.e., the link between income and consumption inequality, for understanding how families cope after adverse events, for designing government insurance, and for the dynamics of business cycles.
+This question is central to understanding the welfare effects of shifts in the income distribution, i.e., the link between income and consumption inequality, how families cope after adverse events, government insurance design, and the dynamics of business cycles.
 A consistent empirical finding is that consumption is partially insured against income shocks, so income fluctuations translate into consumption fluctuations less than one-to-one (e.g., Blundell et al., 2008; Heathcote et al., 2014).
@@
 With a few exceptions discussed subsequently, most measures of consumption insurance do not distinguish between good and bad shocks, or between small and large ones.
-However, recent evidence shows that the distribution of income risk has a long and fat left tail; that is, there is an asymmetrically large probability that households face some really bad income shocks.
-This paper measures the consumption response to income shocks precisely when the distribution of said shocks may exhibit long and fat tails.
+However, recent evidence shows that income risk has a long and fat left tail; that is, there is an asymmetrically large probability that households face some really bad income shocks.
+This paper proposes a simple framework to measure the consumption response to income shocks, precisely when the distribution of said shocks may exhibit long and fat tails.
 In a seminal paper, Blundell et al.
@@
 This is precisely what we investigate in this paper.
-Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting for higher-order moments of their distribution (i.e., tail shocks).
-This is in contrast to the methodology in BPP, which revolves around the second moments of the joint income-consumption distribution and, thus, abstracts from the modern higher-order facts that characterize income dynamics (the second moment is a poor summary of the tails).
-We offer four main contributions:
-First, we re-estimate BPP using recent PSID data over 1999–2019, targeting the original income and consumption second moments.
-However, while BPP imputed consumption from the CEX (the PSID previously lacked comprehensive consumption data), such data are now internally available in the PSID.
-We subsequently use this data to study tail income risk, so it is natural to start with a ‘replication’ of BPP using the newer data.
-This ‘replication’ is insightful in its own right not only because of the different time period, but also because we no longer need to rely on imputed consumption from the CEX.
-Second, we introduce into BPP the third and fourth moments of the joint distribution of income and consumption, in addition to the second moments that BPP originally use.
+Specifically, we measure the degree of consumption insurance to income shocks of varying persistence, accounting not only for second- but also higher-order moments of their distribution (i.e., tail shocks).
+We offer three main contributions:
+First, we re-estimate BPP using recent PSID data over 1999–2019, targeting the original second moments of income and consumption, and compare the results with those based on imputed consumption from the contemporaneous CEX.
+The PSID previously lacked detailed consumption data, so BPP relied on an imputation from the CEX.
+However, such data are now internally available in the PSID.
+We subsequently use this data to study tail income risk, so it is natural to start with a ‘replication’ of BPP using the updated PSID.
+This exercise is insightful because of the different time period and the comparison with the CEX.
+Second, we introduce into BPP the third and fourth moments of the joint distribution of income and consumption, in addition to the second moments above.
 This is our first attempt to go beyond the workhorse income-consumption covariance and measure partial insurance targeting additional information in the tails of the distribution.
@@
 BPP log-linearize the policy rule of a lifecycle consumption/savings problem to obtain a consumption function that depends linearly on income shocks.
-Our first and second contributions above revolve firmly around such a linear framework.
-We argue (as also does Carroll, 2001) that linearity is a poor approximation to behavior when income is subject to tail shocks.
-We then derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks.
-We characterize and empirically measure the consumption pass-through of income shocks, which, in this environment, depends explicitly on the magnitude and sign of shocks.
-While a nonlinear consumption function has been estimated before, most notably in Arellano et al.
-(2017) – henceforth ABB, our framework is far less data demanding (we need second moments of consumption and up to fourth moments of income, rather than observing the entire consumption-income joint distribution) and offers a more transparent link with the higher-order features of the distribution of income shocks.
-Both points have certain important advantages, which we discuss below.
-Fourth, we estimate the welfare costs of tail income risk by calculating what fraction of lifetime consumption a household is willing to give up to avoid tail risk.
-We use parametric utility to calculate these costs, in line with De Nardi et al.
-(2019), Busch and Ludwig (2024), and Guvenen et al.
-(2024), but we base these calculations on the pass-through measure from the quadratic consumption function of the previous step.
-This measures an ‘aggregate’ of all partial insurance available to households, so it is not specific to a particular structural mechanism or a particular parameterization of preferences and expectations.
-Our baseline empirical implementation uses income and consumption panel data from the PSID over the years 1999–2019.
+Our first two contributions above revolve firmly around this linear framework.
+However, linearity may be a poor approximation to behavior when income is subject to tail shocks.
+We thus derive a higher-order (quadratic) consumption function that allows for nonlinear transmission of shocks.
+We characterize and empirically measure the consumption pass-through of income shocks, which now depends explicitly on the shocks’ magnitude and sign.
+While a nonlinear consumption function has been estimated before – most notably in Arellano et al.
+(2017), henceforth ABB – our simple framework offers key advantages: it separates the nature of income risk from the nature of the consumption response as distinct drivers of consumption growth along the income, age, and wealth distributions; it is far less data demanding (second moments of consumption and up to fourth moments of income suffice, with no need to observe the full consumption-income joint distribution or wealth); and it links identification and estimation transparently, aiding interpretation.
+We elaborate on these points below.
+Our core empirical exercises use income and consumption panel data from the PSID from 1999 to 2019.
 We focus on a representative sample of stable married couples that is similar to the sample in BPP.
@@
 (2021).
-We then obtain five main findings:
+For the re-estimation of BPP in the recent years, we also use a comparable sample from the CEX.
+We obtain four main findings:
 First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively.
-Left skewness implies that a negative shock is typically more unsettling, more hurtful than a positive shock, because it is further away from the zero mean.
+Left skewness implies that a negative shock is typically more unsettling than a positive shock, because it is farther away from the zero mean.
 Excess kurtosis implies that more shocks are concentrated either in the middle of the distribution or far out in the tails, relative to the shoulders.
-Taken together, large left skewness and excess kurtosis suggest that the realizations of income shocks tend to be either close to their zero mean or far out in the left tail; that is, income shocks for most households are either small/average or quite bad.
+Taken together, large left skewness and kurtosis suggest that for most households the realizations of income shocks tend to be either small (i.e., close to the zero mean) or far out in the left tail.
 However, these features vary by age and by the household’s position in the income distribution.
-Second, in estimating BPP in the recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP.
+Second, in estimating BPP using the updated PSID in recent years, we find full insurance to transitory shocks and a small pass-through of permanent shocks at 0.15 versus 0.64 in BPP.
 A 10% permanent income cut thus reduces consumption by only 1.5%, vis-à-vis 6.4% in BPP.
-In other words, we find much larger partial insurance to permanent shocks in the recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al.
+In other words, we find much larger partial insurance to permanent shocks in recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al.
 (2023) find a similarly low pass-through of permanent shocks.
 We discuss this further when we present our empirical results.
-We attribute about a third of the difference from BPP to the consumption imputation from the CEX that BPP use.
+We attribute about a third of the gap from BPP to the consumption imputation from the CEX that BPP use.
 The imputation brings in large amounts of measurement error, inflating the covariance between income and consumption growth.
-Another third of the difference arises because the modern PSID provides data biennially (versus annually previously), which tends to mute the transmission of higher frequency shocks.
+Another third of the gap arises because the modern PSID provides data biennially, versus annually previously, which tends to mute the transmission of higher frequency shocks.
 The last third reflects the different calendar times.
 Third, introducing the third and fourth moments of income and consumption growth into BPP’s linear consumption setting does not change our estimates of partial insurance.
-This is counterintuitive in light of the literature that highlights, in multiple settings, the large welfare costs of tail risk (e.g., Barro, 2009; De Nardi et al., 2019).
-One would thus expect that targeting higher-order moments would reduce the degree of partial insurance, which is not what we find.
+This is counterintuitive in light of the literature that highlights, in various settings, the large welfare costs of tail risk (e.g., Barro, 2009; Busch and Ludwig, 2024).
+One would thus expect that targeting higher-order moments reduces the degree of partial insurance, which is not what we find.
 This result is suggestive of a potential misspecification in the linear consumption function, which motivates our generalization of the methodology through the introduction of a higher-order consumption specification subsequently.
-Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of bad versus good shocks, and of small versus large ones.
+Fourth, in estimating our quadratic consumption function, we find an asymmetry in the pass-through of large negative shocks versus positive or smaller ones.
 In this environment, the pass-through depends on the magnitude and sign of shocks; and targeting higher-order income moments enables the identification of the transmission parameters.
@@
 2.) The pass-through of bad shocks increases with the severity of the shock.
-Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie further out in the left tail), but they also transmit increasingly more as their magnitude worsens.
+Bad permanent shocks are thus double hurtful: not only are they more extreme than good ones (they lie farther out in the left tail), but they also transmit more strongly as their magnitude worsens.
 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021).
-Among low wealth low educated households, however, transitory shocks, and in particular negative ones, shift consumption significantly.
 Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one.
@@
 The asymmetric transmission of bad versus good permanent shocks varies with age and the household’s position in the income distribution.
-While older households insure average permanent shocks better than younger ones, older households have larger transmission rates of extreme shocks.
-Severely bad permanent shocks thus cause more harm among the old.
-Moreover, bad permanent shocks from a given position in the bottom tail of the distribution of income risk cause more harm among the rich (households in the top of the income distribution) than among the poor (households in the bottom); the converse is true for good shocks.
-Guvenen et al.
-(2021) show that income skewness becomes more negative with age and income level – there is more room for income to fall among older or higher income workers.
-In turn, we document that there is also more room for consumption to fall.
-Among the old, this is because both the rate of transmission and the severity of downwards income risk increase with age; among the rich, however, the asymmetric transmission is driven mostly by the increasing severity of downwards risk along the income distribution.
-[…]
+While older households insure average permanent shocks better than younger ones, they respond more strongly to larger negative shocks.
+Moreover, bad permanent shocks in the lower tail of income risk cause more harm to the rich (households in the top of the income distribution) than to the poor (households in the bottom); the converse is true for good shocks.
+How a shock from a given position in the distribution affects consumption depends on both the sign/size of the shock and the extent of its transmission.
+A key advantage of our moments-based approach is that it separates income risk and consumption transmission, allowing us to analyze each independently.
+Among the rich, the asymmetry is driven by an increasing severity of downside risk along the income distribution, rather than an increasing consumption transmission.
+Among the old, however, the asymmetry is because both the rate of transmission and the severity of downwards income risk increase with age.
+The asymmetric transmission of bad versus good permanent shocks also varies with wealth and education.
+Transmission of a given average or tail shock decreases as we move from the least to the most wealthy/educated, so wealthier and highly educated households are better insured to all types of permanent shocks.
+However, the nature of income risk differs vastly across groups: the wealthy/educated face more severe downwards income risk, ultimately leading to larger consumption declines when bad shocks hit.
+With left skewness and excess kurtosis, large negative shocks are more likely and more severe than positive ones, so households would rather remove such tail risk from their portfolio.
+A back-of-the-envelop calculation suggests that households facing large downwards risk would give up 4%-8% of lifetime consumption when income risk transmits linearly and rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission.
+There are several estimates of the welfare costs of income risk in the literature.
+Two settings in which the baseline environment concerns idiosyncratic tail income risk of similar nature and severity to ours are Busch and Ludwig (2024) and Guvenen et al.
+(2024).
+Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk (a more distant counterfactual from ours).
+[^3] These papers calculate welfare costs using full structural models in which the consumption responses depend on endogenously accumulated wealth.
+By contrast, we rely on approximate consumption rules, which take wealth as given.
+We show in appendix D that this is unlikely to be quantitatively important., [^4] Welfare cost estimates of aggregate risk include Barro (2009), who find that agents can give up about 20% of GDP each year to eliminate rare disasters, and Georgarakos et al.
+(2025), who find that households can sacrifice 5%-6% of lifetime consumption to eliminate business cycle fluctuations.
 There is a large literature in macro and labor economics that studies the transmission of income shocks into consumption, i.e., the link between income and consumption inequality.
 There are two distinct strands in this literature: a quasi-reduced form branch that measures the degree of partial insurance in the data without fully specifying the underlying insurance mechanisms, and a structural branch that fully models specific insurance channels.
+The present paper falls in the first branch.
 The seminal paper in the first branch, namely BPP, sparked a large volume of research.
@@
 Blundell et al.
-(2016) measure partial insurance against wage risk, allowing for family labor supply.
-Theloudis (2021) allows for preference heterogeneity in partial insurance, which he estimates using information in the third moments.
-Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response, similar to what experimental studies find.
-Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes.
-Crawley and Kuchler (2023) address neglected time aggregation in income and consumption in the PSID.
-The previous papers restrict income shocks to transmit linearly into consumption.
+(2016) measure partial insurance against wage risk, Theloudis (2021) allows for preference heterogeneity in partial insurance, and Commault (2022) allows consumption to respond to past transitory shocks and finds a strong response similar to what experimental studies find.
+Hryshko and Manovskii (2022) identify sets of households in the PSID with different degrees of insurance based on certain features of their incomes, and Crawley and Kuchler (2023) address neglected time aggregation in the PSID data.
+These papers restrict income shocks to transmit linearly into consumption.
 ABB and Arellano et al.
-(2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission.
-Using the recent PSID and a quantile regression approach, they measure the marginal propensity to consume from persistent income, which varies along the income distribution.
-They simulate the transmission of shocks and document asymmetries between bad/good shocks or between rich/poor, similar to ours.
-Our papers, however, differ in important dimensions.
-First, identification in ABB is quite demanding as they require observing the entire consumption-income joint distribution, while we need only second moments of consumption and up to fourth moments of income.
-This is appealing because we use more accessible information and because, unlike income, we cannot assess the higher-order features of consumption against an established population benchmark.
-The price of our econometrically more accessible approach is that, unlike ABB, we have to maintain a linear income process (but subject to tail shocks).
-Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
-Second, ABB measure the pass-through over quantiles of the distribution of income shocks, without explicitly studying the higher-order features of their distribution.
-We show subsequently that seemingly different quantiles, e.g., shocks in the 10th or 90th percentile as in ABB, may in fact be very similar to one another when the underlying distribution is skewed and leptokurtic.
-Our model allows us to estimate pass-through rates that explicitly shift with the sign and size of income shocks and can thus describe more transparently how shocks of different signs/sizes shift consumption.
+(2023) generalize the income and consumption processes by letting income feature nonlinear persistence, which varies with the level of income and its shock, and nonlinear transmission that is more flexible than ours.
+Using recent PSID data and quantile methods, they measure the marginal propensity to consume from income along the income distribution.
+To quantify the pass-through of income shocks, they simulate the model and find asymmetries in the transmission of bad and good shocks, or between rich and poor households, similar to ours.
+Our third contribution shares ABB’s goal of measuring nonlinear transmission, but differs in three key dimensions.
+First, ABB compare consumption growth across quantiles of income and income shocks, without distinguishing between the nature of income risk and that of the consumption response.
+For example, the rich appear to update consumption more strongly than the poor after a lower-tail shock, but this can reflect more severe downside risk or/and a stronger transmission independent of the shock.
+Our moments-based approach separates these channels, allowing us to analyze each independently.
+In this example, we find richer households respond less to a given shock (more insurance), but face worse downside risk, leading to larger overall consumption changes.
+Second, the quantile methods in ABB require observing the full consumption-income joint distribution, and wealth, while our approach requires only second moments of consumption and up to fourth moments of income.
+This makes our framework less data demanding, appealing both on its own and because, unlike income, we cannot benchmark higher-order consumption moments against population data.
+The price of our econometrically more accessible approach is that, unlike ABB, we maintain a linear income process (but subject to tail shocks).
+Yet, we find that nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetry across different shocks and income levels.
+Third, our consumption insurance parameters are directly estimable in the data without simulations, and tightly linked to their identification.
+They are comparable to the existing literature on partial insurance and ‘portable’ as they describe transparently how shocks of different signs/magnitudes shift consumption.
 The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models.
@@
 Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; we measure aggregate insurance from all mechanisms together, while accounting for higher-order features of the income distribution, keeping our model parametrically unspecified.
-The asymmetric response to bad and good shocks is recently also found in surveys that elicit responses to hypothetical events.
+The asymmetric response to bad and good shocks is recently found in surveys that elicit responses to hypothetical events.
 Bunn et al.
@@
 (2020) find that responses to losses are larger than to gains, and the response increases with the severity of the loss.
-We document that these patterns exist also in survey data, where, unlike elicitation surveys, income shocks are realized but unobserved.
+These papers consider observed hypothetical transitory shocks, while we document similar patterns in response to unobserved permanent shocks.
 [^7] Given our finding of heterogeneous responses to shocks, our paper is also broadly related to the literature that studies MPC heterogeneity (e.g., Lewis et al., 2022; Crawley and Kuchler, 2023).
 The paper is finally related to the newer income dynamics literature that focuses on non-Gaussian aspects.
-Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth.
-Guvenen et al.
-(2021) use administrative data in the U.S.
-to provide further evidence of left skewness and excess kurtosis in earnings growth, accentuated with age and the level of earnings.
+Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth, while Guvenen et al.
+(2021) provide even richer evidence from administrative data.
 Guvenen et al.
@@
 While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption.
-We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Constantinides and Ghosh, 2017), business cycles (McKay, 2017), consumer welfare (De Nardi et al., 2019), and labor market dynamics (Ai and Bhandari, 2021).
-In the remainder of the paper, section 2 presents the model and section 3 establishes identification of the parameters.
-Section 4 discusses the empirical implementation, section 5 presents the results, and section 6 discusses welfare.
-Section 7 concludes.
-The supplemental appendix provides additional derivations and results.
-2 Lifecycle model with tail income risk
-The framework through which we study the extent of partial insurance to income shocks is a consumption lifecycle model with idiosyncratic income risk.
-Household $i$ chooses lifetime consumption $\{C_{it}\}_{t=0}^{T}$ and savings $\{A_{it+1}\}_{t=0}^{T}$ to maximize its expected discounted utility over its lifecycle. The problem is formulated as
+We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Backus et al., 2011; Ghosh and Julliard, 2012; Constantinides and Ghosh, 2017), business cycles (McKay, 2017), and labor market dynamics (Ai and Bhandari, 2021).
+In the rest of the paper, section 2 presents the income and consumption processes and section 3 discusses identification.
+Section 4 discusses the implementation, section 5 presents the results, and section 6 concludes.
+The appendix has additional derivations and results.
+2 An organizing framework: a lifecycle model with tail income risk
+The framework through which we organize the discussion of partial insurance to income shocks is a consumption lifecycle model with idiosyncratic income risk.
+Household $i$ chooses a sequence of consumption $\{C_{it}\}_{t=0}^{T}$ and savings $\{A_{it+1}\}_{t=0}^{T}$ to maximize its expected discounted utility over its lifecycle. The problem is formulated as
 $\max_{\{C_{it},A_{it+1}\}_{t=0}^{T}}{\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\beta^{t}U(C_{it};\mathbf{Z}_{it}),$ (1)
@@
 $A_{i0}+{\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{Y_{it}}{(1+r)^{t}}={\operatorname*{\mathbb{E}}}_{0}\sum_{t=0}^{T}\frac{C_{it}}{(1+r)^{t}},$ (2)
-where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is beginning-of-time financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon.
+where $U(\cdot;\mathbf{Z}_{it})$ is utility over consumption, $\mathbf{Z}_{it}$ is a vector of observed taste shifters (e.g., age, education), $\beta$ is the discount factor, $r$ is the deterministic interest rate, $A_{i0}$ is initial financial wealth, ${\operatorname*{\mathbb{E}}}_{0}$ denotes expectations over uncertain future income $Y_{it}$, and $T$ is the end of the horizon.
 We assume that utility satisfies standard regularity conditions, in particular it is continuously twice differentiable, but we do not otherwise parameterize it.
@@
 $\ln Y_{it}=\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}+P_{it}+v_{it},$
-where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race); $\boldsymbol{\delta}$ is the loading factor of those variables on the logarithm of income.
+where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race) and $\boldsymbol{\delta}$ is its loading factor.
 The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$.
@@
 These modeling choices give rise to the familiar permanent-transitory formulation
-$\Delta y_{it}=\zeta_{it}+\Delta v_{it},$ (3)
+$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vspace{-1ex}$ (3)
 where $\Delta y_{it}=\Delta\ln Y_{it}-\Delta\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}$ is income growth net of the deterministic profile, i.e., unexplained income growth, and $\Delta$ is the first difference operator.
@@
 Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
-We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may exhibit different amounts of dispersion, skewness, or thickness of the tails of income risk; Guvenen et al.
+We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may differ in the amount and nature of income risk; Guvenen et al.
 (2021) provide some evidence for that.
@@
 where $\Delta c_{it}$ is the residual from a regression of real consumption growth, $\Delta\ln C_{it}$, on the taste observables $\mathbf{Z}_{it}$ (i.e., consumption growth net of its predictable profile).
-The linear function (4) stems from a log-linearization (i.e., a first-order Taylor series expansion) of the problem’s first-order conditions and the household budget constraint, as we show in appendix A.
+The linear function (4) is consistent with a log-linearization (i.e., a first-order Taylor series expansion) of the problem’s first-order conditions and the household budget constraint, as we show in appendix A.
 These approximations are independent of the distributions of income shocks and mimic similar derivations in BPP (who abstract from higher-order moments).
@@
 (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences.
-The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds the better they can self-insure.
+The idea is that a household can use wealth to smooth consumption when income shocks hit; and the more wealth one holds, the better they can self-insure.
 We establish this link from wealth to $\phi_{it}^{(1)}$, $\psi_{it}^{(1)}$ in appendix A.
@@
 2.3 Quadratic consumption function
-The linear consumption function offers an approximation to the true consumption response to income shocks when such shocks do not move the household too far away from its deterministic consumption profile.
-Function (4) relies on a linearization of marginal utility (the problem’s first-order conditions) around that deterministic profile; but such a linearization will be inaccurate in the presence of tail income risk, e.g., when extreme shocks may induce large shifts in consumption as a consequence of concave utility.
-[^9] This echoes Carroll (2001)’s critique of the log-linearized Euler equation.
-Kaplan and Violante (2010) and Blundell et al.
-(2013) assess the approximations in BPP across alternative scenarios, though not in one with respect to the severity of income shocks.
+The linear function offers an approximation to the true consumption rule when income shocks do not move the household too far away from its deterministic consumption profile.
+Function (4) is consistent with a linearization of marginal utility around that deterministic profile; but such a linearization will be inaccurate in the presence of tail income risk, e.g., when large shocks may induce large shifts in consumption as a consequence of concave utility.
+[^10] This echoes Carroll (2001)’s critique of approximate Euler equations.
+He argues that the -then- popular log-linear Euler equation, and to a lesser extent its quadratic counterpart, imply consumption behavior that deviates from that in the true underlying model.
+However, Kaplan and Violante (2010) and Blundell et al.
+(2013) assess the linear approximations in BPP across alternative scenarios, though not in one with respect to the severity of income shocks, and they find that they perform reasonably well.
+The second-order approximations herein aim at improving the inference drawn from these previously established approximations.
 To address this issue, we generalize the consumption function to a nonlinear (quadratic) specification, given by
@@
 where $\Delta c_{it}$ and $\xi_{it}$ are defined as previously.
-The quadratic function (5) stems from a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A.
+The quadratic function (5) is consistent with a second-order Taylor approximation to the problem’s first-order conditions and the household budget constraint, as we show in appendix A.
 These operations and resulting expressions generalize the original approximations in BPP and in several papers thereafter.
@@
 In this generalized framework, $\phi_{it}^{(1)}$ no longer represents the pass-through of permanent shocks to consumption. Instead, the pass-through is given by
-$\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}=\phi_{it}^{(1)}+2\phi_{it}^{(2)}\zeta_{it}+\omega_{it}^{(22)}v_{it},$ (6)
+$\partial\Delta c_{it}/\partial\zeta_{it}=\phi_{it}^{(1)}+2\phi_{it}^{(2)}\zeta_{it}+\omega_{it}^{(22)}v_{it},$ (6)
 which depends on multiple parameters and on the magnitude and sign of the income shocks.
@@
 Suppose, for example, that $\phi^{(1)}>0$ and $\phi^{(2)}<0$, as our results subsequently suggest.
-At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a higher pass-through to consumption relative to a positive shock of the same magnitude.
-The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of partial insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response.
-The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the permanent shock itself.
-When income risk is non-Gaussian (e.g., it is left tailed), the realization of shocks can be far from their mean, thereby causing large errors in the inference of the degree of partial insurance based on $\phi^{(1)}$ alone.
+At $v_{it}=0$ (the average transitory shock), a negative permanent shock has a larger pass-through to consumption relative to a positive shock of the same magnitude.
+The pass-through of negative shocks would thus be underestimated (i.e., excessive insurance concluded) if the econometrician assessed the extent of insurance on the basis of $\phi^{(1)}$ alone, ignoring the non-linearity in the consumption response.
+The size of this bias increases with the magnitude of $\phi^{(2)}$ and the size of the negative permanent shock (e.g., the extent of left skewness).
 A similar asymmetry exists in the transmission of good versus bad transitory shocks. In this case, the pass-through into consumption is given by
-$\frac{\partial\Delta c_{it}}{\partial v_{it}}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
+$\partial\Delta c_{it}/\partial v_{it}=\psi_{it}^{(1)}+2\psi_{it}^{(2)}v_{it}+\omega_{it}^{(22)}\zeta_{it},$ (7)
 which again depends on multiple parameters and on the magnitude and sign of the income shocks.
@@
 Suppose, for example, that $\psi^{(1)}\approx\psi^{(2)}\approx 0$; then $\partial\Delta c_{it}/\partial v_{it}\approx 0$ at $\zeta_{it}=0$ (the average permanent shock), indicating full insurance to transitory shocks.
+A final remark is needed here.
+The specific lifecycle model (1)-(2) is not the only structure that gives rise to the quadratic specification (5) or the subsequent econometric framework.
+Several models, e.g., models with liquidity constraints (Campbell and Hercowitz, 2019) or models for the wealthy hand-to-mouth (Kaplan and Violante, 2014), to offer just two examples, imply consumption policy rules of the form $C_{it}=g_{t}(Y_{it},\dots)$, motivating our empirical specification through a quadratic approximation to $g_{t}$.
+This light dependence on specific theory is why we used the term ‘quasi-reduced form’ to name our framework.
 3 Identification
 Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters.
-The set of parameters includes the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^12] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments.
+These include the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^13] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments.
 We will be clear whenever those terms refer to standardized moments.
-the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of the unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$).
-We initially restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the distribution of income.
+the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$).
+We first restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the income distribution.
 For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth.
 In the case of the quadratic consumption function, we also assume that income shocks are not Bernoulli.
-For consistent estimation of such moments, particularly over demographic groups subsequently, we further assume non-vanishing shares in each group asymptotically and knowledge of group membership.
-To carry out statistical inference, we finally assume existence of up to eighth moments.
+For consistent estimation of such moments, particularly over demographic groups subsequently, we further assume non-vanishing shares in each group asymptotically.
+For inference, we assume existence of up to eighth moments.
 Two important aspects of the data must be addressed before proceeding to identification: the biennial nature of the modern PSID and measurement error.
@@
 $\kappa_{\zeta_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{4})-\kappa_{v_{t}}-\kappa_{v_{t-1}}-6\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})-6\sigma_{v_{t}}^{2}\sigma_{v_{t-1}}^{2},$
-$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}$
-$-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
+$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta y_{it})^{2}\times(\Delta y_{it+1})^{2})-\sigma_{\zeta_{t}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t+1}}^{2}\sum_{\tau=-1}^{0}\sigma_{v_{t+\tau}}^{2}-\sigma_{v_{t-1}}^{2}\sum_{\tau=0}^{1}\sigma_{v_{t+\tau}}^{2}-\sigma_{\zeta_{t}}^{2}\sigma_{\zeta_{t+1}}^{2}-\sigma_{v_{t}}^{2}\sigma_{v_{t+1}}^{2}.$
 There are several over-identifying restrictions for both skewness and kurtosis.
@@
 The moments of the distribution of consumption growth reflect the underlying distributions of income shocks, as well as the distribution of unobserved taste heterogeneity.
-In general, larger dispersion, skewness, or kurtosis in the distributions of income shocks implies larger dispersion, skewness, or kurtosis in the distribution of consumption growth.
-The link between the two depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
-BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B.
+In general, larger dispersion, skewness, or kurtosis in the distributions of income shocks imply larger dispersion, skewness, or kurtosis in the distribution of consumption growth.
+The link between them depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
+BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B.2.
 Formally, the transmission parameter of permanent shocks is identified as
@@
 [^14] The variance of consumption taste heterogeneity is further given by $\sigma_{\xi_{t}}^{2}={\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})-(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}-(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}$.
-Its skewness and kurtosis are similarly identified from the higher moment of consumption growth.
+Its skewness and kurtosis are similarly identified from higher moments of consumption growth.
 These statements reflect the intuition behind identification in BPP: the covariance between consumption and (appropriately defined) income growth captures the strength of the relation between the two series and, consequently, the pass-through of income shocks into consumption.
@@
 3.3 Quadratic consumption function parameters
-As per the quadratic specification in (5) and the properties of shocks, the second moment of the distribution of consumption growth is given by
+As per the quadratic specification in (5) and the properties of shocks, the second moment of consumption growth is given by
 ${\operatorname*{\mathbb{E}}}((\Delta c_{it})^{2})$ $=(\phi_{t}^{(1)})^{2}\sigma_{\zeta_{t}}^{2}+(\psi_{t}^{(1)})^{2}\sigma_{v_{t}}^{2}+2\phi_{t}^{(1)}\phi_{t}^{(2)}\gamma_{\zeta_{t}}+2\psi_{t}^{(1)}\psi_{t}^{(2)}\gamma_{v_{t}}+(\phi_{t}^{(2)})^{2}\kappa_{\zeta_{t}}+(\psi_{t}^{(2)})^{2}\kappa_{v_{t}}$
@@
 Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting.
-We establish in appendix B that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway).
+We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway).
 [^18] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank.
@@
 Identification of the parameters does not require those shocks to be non-Gaussian; the matrix is invertible even if the distributions of shocks are normal and have the same variance between them (i.e., $\sigma_{\zeta_{t}}^{2}=\sigma_{v_{t}}^{2}$).
-In the nonlinear specification, tail shocks are associated with a different pass-through from that of average shocks, regardless of the precise nature of the underlying distribution of shocks.
-The transmission parameters, however, can no longer be identified with second-order income moments alone – inspect matrix $\mathbf{A}$ in (8).
-Skewness and kurtosis, both defined also in the Gaussian case, are crucial for identifying the pass-through of such shocks, which is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails.
-In appendix B, we investigate what different levels of income skewness and kurtosis imply for the transmission parameters.
+In the nonlinear specification, tail shocks exhibit a different pass-through compared to average shocks, regardless of the specific form of underlying income risk.
+As a result, their transmission parameters cannot be identified with second-order income moments alone.
+Skewness and kurtosis now become essential for identification, even if the income distribution is Gaussian.
+This is intuitive: we want to identify the impact of tail shocks, so we must use information on said tails.
+In appendix B.3, we explore how varying levels of income skewness and kurtosis influence the transmission parameters.
 Contrary to income, no higher-order moments of consumption are needed; the identifying moments in (8) are covariances between consumption growth and various powers of income growth.
-This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is no established population benchmark to assess them against.
+This is an advantage over quantile methods that require observing the entire consumption distribution, e.g., ABB, because, while we may estimate higher moments of consumption growth in the PSID, there is yet no established population benchmark to assess them against.
 This is in contrast to higher income moments, for which Guvenen et al.
@@
 This fails to identify the true pass-through of permanent shocks in the presence of nonlinear transmission, i.e., expression (6).
-Moreover, this workhorse covariance is biased for $\phi_{t}^{(1)}$ if permanent shocks are non-Gaussian.
-If $\gamma_{\zeta_{t}}<0$, i.e., permanent shocks are left skewed as we find subsequently, and $\phi_{t}^{(2)}<0$, as our empirical estimates suggest, linearity biases $\phi_{t}^{(1)}$ upwards.
-Similarly, the identifying condition for $\psi_{t}^{(1)}$ in BPP becomes
+Moreover, this workhorse covariance is biased for $\phi_{t}^{(1)}$ if permanent shocks exhibit skewness.
+If $\gamma_{\zeta_{t}}<0$ and $\phi_{t}^{(2)}<0$, as our empirical estimates suggest, linearity biases $\phi_{t}^{(1)}$ upwards.
+Similarly, BPP’s identifying condition for $\psi_{t}^{(1)}$ becomes
 $-{\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$
@@
 This may help bridge the discrepancy between covariance based methods (who find little pass-through of transitory shocks) and studies based on natural experiments (who usually find large consumption responses to said shocks).
-Our estimation procedure, which we describe in the next section, is linked transparently to our identification argument.
+Our estimation procedure, which we describe in the next section, is linked transparently to the above identification argument.
 Effectively, we bring (8) to the data, along with some over-identifying restrictions, using moments estimation.
@@
 We then move on to the results from the quadratic specification.
-Our choice to estimate the income process in an earlier step enables the income parameters to remain unchanged regardless of the specification of the consumption function fitted.
+Our choice to estimate the income process separately in an earlier step enables the income parameters to remain unchanged regardless of the specification of the consumption function fitted.
 Appendix B.4 lists the moments targeted in each case.
@@
 For the subsample analysis (tables 5-6 and corresponding tables in the appendix), we calculate GMM (asymptotic) standard errors clustered at the household level to avoid numerical instability in bootstrap, driven by vanishing shares of certain subsamples, e.g., the wealthy non-college graduates, in many bootstrap draws.
-For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors, as in BPP.
+For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors as in BPP.
 As we explained in section 3, we must use an external estimate for the variance of income measurement error.
@@
 (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income.
-To facilitate the comparison, appendix C.2 reports moments of male earnings growth in our sample.
+To facilitate a comparison, appendix C.2 reports moments of male earnings growth in our sample.
 The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively.
@@
 While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al.
-(2021) confirm this.
+(2021) also confirmed this.
 Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
@@
 Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions).
+[^27] Using earnings as the relevant income variable in the model does not materially change our subsequent results, in particular the asymmetry in the transmission of bad versus good (earnings) shocks.
 These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al.
@@
 The moments of imputed consumption growth are in column 2 of table 1.
-The variance at $0.323$ is two-and-a-half times higher than the variance in the true series (column 1).
+The variance at $0.323$ is two-and-a-half times larger than the variance in the true series (column 1).
 In fact, the variance of imputed consumption growth is more than double the variance of income growth.
-The higher-order moments differ markedly between the imputed and true series.
+The magnitude of the higher-order moments is also multiple times larger in the imputed versus the true series.
 Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series.
@@
 This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period.
-These results vary with the level of income, but we postpone this discussion for later.
-Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme.
+These results vary with the level of income, but we postpone that discussion for later.
+Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are farther away from the zero mean – they are more extreme.
 This is true for both permanent and transitory shocks.
@@
 The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate.
-Appendix C.4 discusses the details of the imputation and a mechanism through which it inflates $\phi^{(1)}$ in particular (the imputation involves the inversion of a demand system, so certain types of measurement error may inflate the imputed income-consumption covariance – evident also in table 1).
+Appendix C.4 discusses the details of the imputation and a mechanism through which in particular it inflates $\phi^{(1)}$: the imputation involves the inversion of a demand system, so certain types of measurement error can inflate the imputed income-consumption covariance – evident also in table 1.
 Table 3: Estimates of the consumption function
@@
 [^34] We use the corresponding pre-estimated income parameters from columns 2 and 3 of table 2 respectively.
-Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income.
-The transmission parameters remain largely unchanged upon inclusion of the third moments.
+Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income in the PSID.
+The transmission parameters remain largely unchanged.
 We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with only second moments) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero).
@@
 [^35] Appendix table C.8 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 3.
-We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong, significant, and negative response to transitory shocks, which is hard to explain.
-We do not think those results are very credible; the imputation brings in substantial measurement/imputation error, which becomes worse the higher the order of the target moments is.
-Clearly, targeting higher-order moments in the context of the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone.
-This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary.
+We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong negative response to transitory shocks, which is hard to explain.
+We do not think those results are very credible; the imputation brings in substantial measurement and imputation error, which likely becomes worse the higher the order of the target moments is.
+Clearly, targeting higher-order moments within the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone.
+This is counterintuitive given the extent of tail income risk in the data and ample evidence in the literature to the contrary.
 [^36] Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets.
@@
 Our quadratic specification addresses their concern.
-One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events.
-Yet, the linear consumption function leaves no room to such events.
-BPP rely on a log-linearization of the consumption rule around a deterministic lifecycle profile (appendix A).
-But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response.
-A linearized marginal utility fails to pick up such large response (Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments.
+One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events, even if such risk triggers a stronger precautionary motive.
+Yet, the linear consumption function leaves no room to such events, because it reflects a linearization of the consumption rule around a deterministic lifecycle profile.
+Tail events, however, move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response.
+A linearized marginal utility fails to pick up such large response, which in turn explains why we do not see much action upon including higher-order moments.
 In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates our higher-order specification subsequently.
@@
 5.3.1 Main results
-We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of a polynomial in the permanent shock) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification.
-We estimate the transmission parameter of the average transitory shock (the coefficient on the linear part of a polynomial in the transitory shock) at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero.
+We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification.
+We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero.
 As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
-We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks and a positive (but statistically insignificant for now) $\psi^{(2)}=0.015$ for transitory shocks.
-The coefficient $\omega^{(22)}$ on the product of the two shocks is large but very imprecisely estimated.
+We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones.
+$\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently.
+By contrast, $\psi^{(2)}$ is almost always insignificant, barring the few exceptions we highlight below.
+The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise.
 The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case.
 [^39] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
-The values of $\phi^{(2)}$ and $\psi^{(2)}$ have important implications for the pass-through of shocks into consumption.
-To help interpret the results, recall that the pass-through of permanent shocks is given by (6) and depends on $\phi^{(1)}$, $\phi^{(2)}$, $\omega^{(22)}$, and the magnitudes and signs of both income shocks.
+At face value, $\phi^{(2)}$ has important implications for the pass-through of permanent shocks into consumption.
+To help interpret the results, recall that the pass-through is given by (6) and depends on $\phi^{(1)}$, $\phi^{(2)}$, $\omega^{(22)}$, and the magnitudes and signs of both income shocks.
 If, for simplicity, we set the transitory shock to its mean value ($v_{it}=0$) and plug in the parameter estimates from table 3, we get
@@
 $\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta<0}~{}>~{}\left.\frac{\partial\Delta c_{it}}{\partial\zeta_{it}}\right|_{\zeta>0}.$
-Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., further away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption.
+Therefore, bad permanent shocks are not just more extreme than good shocks, i.e., farther away from the zero mean – see left skewness in table 2, but they also have larger pass-through rates into consumption.
 Bad permanent shocks are thus double hurtful.
 Second, severe bad shocks are more hurtful than moderate ones because they too are associated with larger transmission parameters into consumption.
-Contrast, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean.
+Compare, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean.
 In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$).
@@
 When a bad shock hits, however, households insure themselves by running down savings or by using other formal or informal insurance arrangements.
-But there is clearly a limit to how much they can dissave or borrow.
+Yet there is clearly a limit to how much they can dissave or borrow.
 As the magnitude of the bad shock worsens, these insurance arrangements will be eventually exhausted, resulting in a larger fraction of the shock transmitting into consumption.
@@
 This literature, however, elicits responses to transitory shocks.
-By contrast, we establish these asymmetries also in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
+By contrast, we establish these asymmetries in response to permanent shocks in a context in which the shocks are not directly observed (namely in survey data).
 Moving on to transitory shocks, their pass-through is given by (7); this depends on $\psi^{(1)}$, $\psi^{(2)}$, $\omega^{(22)}$, and on the magnitudes and signs of both income shocks.
@@
 This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption.
-This is similar to our findings for the linear case as well as to BPP.
-By contrast, this is not the case for certain well-defined subsamples, which we analyze in the next subsections.
-Despite the lack of statistical significance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero.
+Despite their statistical insignificance, the transmission parameters of transitory shocks have, at face value, an interesting implication: good and large shocks have a numerically larger pass-through than smaller ones, whose pass-through is economically zero.
 While we should be cautious in drawing conclusions about large good transitory shocks (given left skewness and excess kurtosis, such shocks are rare in our sample), this implication helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in BPP and similar studies that employ covariance restrictions on survey data.
-The former typically exploit large infrequent transitory episodes (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021).
+The former typically exploit large infrequent transitory gains (e.g., tax rebates or lotteries as in Parker et al., 2013; Misra and Surico, 2014; Fagereng et al., 2021), which are unlikely to be present in the pooled survey data that the latter studies exploit (e.g., BPP, Blundell et al., 2016; Theloudis, 2021).
 In an environment where the transmission of shocks depends on their magnitude, these two types of settings will lead to opposite conclusions.
-Our nonlinear consumption function reconciles the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
-To summarize our findings from the quadratic consumption function: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do.
+Our framework can potentially reconcile the two by permitting transitory shocks to pass through to consumption at a varying rate that depends on the shocks’ magnitude and sign.
+To summarize our findings from the quadratic function up to this point: 1.) bad permanent shocks reduce consumption by more than equal-sized good permanent shocks increase it; 2.) severe bad permanent shocks reduce consumption by more than moderate ones do.
 For a given value of $\phi^{(1)}$, the econometrician will overestimate the true degree of partial insurance against bad permanent shocks if she assumes a linear consumption function instead of the nonlinear specification herein.
@@
 Table 4 reports results from the quadratic specification (see table C.10 for the linear specification).
-$\phi^{(1)}$ decreases with age, so the average permanent shock impacts consumption less as households age.
+$\phi^{(1)}$ decreases with age, so the average permanent shock affects consumption less as households age.
 In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle.
@@
 Linear specification estimates appear in appendix table C.10.
-Guvenen et al.
-(2021) document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age.
+In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9).
+This is consistent with Guvenen et al.
+(2021), who document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age.
 There is more room for income to fall sharply in this case than among the young.
-In our data, skewness of permanent shocks becomes more negative as households age; it more than triples in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9).
-The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption technology itself changes ($\phi^{(2)}$ more negative).
-The type of events that underlie large income downturns at older ages – such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption technology that differs between the old and the young.
-This does not mean that moderate bad shocks are worse insured among the old; for small shocks, the decreasing pattern in $\phi^{(1)}$ more than counteracts the increasing magnitude of $|\phi^{(2)}|$.
+The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption response itself changes ($\phi^{(2)}$ becomes more negative).
+The type of events that underlie large income downturns at older ages, such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption response that differs between the old and the young.
+This does not mean that moderate bad shocks are worse insured among the old; for small or moderate shocks, the decreasing pattern in $\phi^{(1)}$ offsets the increasing magnitude of $|\phi^{(2)}|$.
 Finally, the transmission parameters of transitory shocks remain statistically indistinguishable from zero over the entire lifecycle.
-We next consider how partial insurance changes with the household’s position in the income distribution before the shocks hit at $t$, i.e., in the distribution of $y_{it-1}$.
-Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, unlike ABB whose nonlinear income process naturally allows this.
-We thus take an empirical approach and estimate the model over different cuts of the distribution of $y_{it-1}$.
+Next, we consider how partial insurance varies over the household’s position in the income distribution prior to the realization of shocks, i.e., the distribution of $y_{it-1}$.
+Our linear income process and first-difference approach does not directly link the transmission parameters to the levels of income, therefore we take an empirical approach and estimate the model over different quantiles of $y_{it-1}$.
 Sample size considerations prevent us from being overly granular; we tried estimation over deciles, quartiles, and tertiles, and the latter was the most robust.
@@
 [^43] We target the same moments as earlier, broken down by tertiles of the distribution of past income $y_{it-1}$.
-The moments of shocks are also income level-specific (appendix table C.9).
+The moments of shocks are also income level-specific; we report them in appendix table C.9.
 Table 5 reports results from the quadratic specification (see table C.11 for the linear specification).
@@
 Third, the transmission of transitory shocks shows mixed patterns but it is never statistically significant.
-At face value, $\phi^{(2)}$ is more negative among the poor, suggesting that a bad shock of a given size transmits more strongly among them.
-While factually correct given the estimates, this provides an incomplete view of the underlying dynamics because the distribution of income risk is vastly different between poor and rich.
-Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but very negative among the rich, with a coefficient of $-4.86$ (multiple times larger than baseline skewness, and highly significant).
-One may thus study how large bad/good shocks from a given position in the underlying distribution, for example shocks in the 3rd and 97th percentiles, affect consumption in each case.
-Simulations are needed to infer the specific shocks.
+At face value, $\phi^{(2)}$ is more negative among the poor, and highly significant, suggesting that a given bad shock transmits more strongly among them.
+However, this provides only a partial picture of the underlying dynamics because the nature of income risk is vastly different between poor and rich.
+Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of $-4.86$ (multiple times larger than the baseline, and highly significant).
+One may thus study how large bad/good shocks from a given position in the underlying distributions, e.g., shocks in the 3rd or 97th percentiles, affect consumption in each case.
+Simulations are needed to infer the value of the specific shocks.
 [^44] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9.
-We provide details on similar simulations in section 6.
-Adopting this approach, the bad shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) but it shifts $c_{it}$ by $-0.035$, twice as much, among the rich (pass-through $0.084$).
-By contrast, the good shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but shifts $c_{it}$ by only $0.006$ among the rich (pass-through $0.057$).
-These findings agree with ABB, who simulate the transmission of 10- and 90-percentile shocks and also find a stronger response to bad news among households in the top end of the income distribution and a stronger response to good news among households in the bottom end.
-The reported $\Delta c_{it}$ are also numerically in the same ballpark (impulse responses with heterogeneity in their appendix).
-Our results thus suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
+We provide details on similar simulations in appendix D.
+Adopting this approach, the lower-tail shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by $-0.035$, twice as much, among the rich (pass-through $0.084$).
+By contrast, the upper-tail shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but by only $0.006$ among the rich (pass-through $0.057$).
+In other words, households in the top of the income distribution (the rich) exhibit a stronger consumption change in response to bad news, while households in the bottom (the poor) exhibit a stronger consumption change in response to good news.
+These findings agree with ABB, who simulate the transmission of shocks in the 10th and 90th percentiles.
+The reported $\Delta c_{it}$ – impulse responses with heterogeneity in their appendix figure S30 – are in the same ballpark, especially for bad shocks.
+While this is reassuring, our findings further suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
+Our approach also highlights two distinct forces at play: the underlying nature of income risk (i.e., income parameters) and the nature of the consumption response (i.e., consumption parameters).
+The larger consumption change among the rich after a lower-tail shock is not because their consumption response differs materially from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates said shocks), but mostly due to stronger downside risk that they face (the magnitude of the bad shock is simply larger in this case).
 Table 5: Estimates of the consumption function, over position in income distribution
@@
 Linear specification estimates appear in appendix table C.11.
-Our approach highlights two distinct forces at play in the response of consumption to income shocks: the underlying nature of income risk (i.e., the income parameters) and the consumption technology (i.e., the consumption parameters).
-The larger response of the rich to shocks in the bottom of the distribution is not because their consumption technology differs vastly from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates the response to said shocks), but mostly due to stronger downwards risk that the rich face.
-To infer that, one needs to evaluate the marginal distribution of income risk.
-ABB’s quantile method can flexibly estimate the consumption response over income quantiles and compare responses to, say, 10- and 90-percentile shocks, but it does not measure the higher-order features of the distribution these shocks come from.
-This may lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be very similar to one another since the underlying distribution is skewed and leptokurtic.
-For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked further out at the 3rd and 97th percentiles.
+To distinguish between these two forces, one must evaluate the marginal distribution of income risk.
+While ABB’s quantile method can flexibly estimate consumption changes over different income quantiles – and compare changes after, say, 10- and 90-percentile shocks – it does so without explicitly accounting for the higher-order features of the distribution these shocks come from.
 Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
 $c_{it}=g_{t}(\underbrace{y_{it-1}+F_{t}^{-1}(u_{it})}_{y_{it}},~{}\dots),\$
-where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ is a uniform shock denoting the quantile of the shocks’ distribution $F_{t}$.
-They then simulate the consumption pass-through over different values for $y_{it-1}$ and $u_{it}$, without measuring $F_{t}$.
-Our model allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the sign and size of the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$.
+where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ denotes the quantile of the shocks’ distribution $F_{t}$.
+They then simulate the consumption pass-through over different values of $y_{it-1}$ and $u_{it}$, without accounting for $F_{t}$.
+This may lead inadvertently to comparisons between consumption responses to very different income shocks, even if those shocks come from the same position in their respective distributions (the nature of income risk varies over the distribution of income).
+[^45] Apparently, this may also lead inadvertently to comparisons between consumption responses to seemingly different shocks, when these shocks may in fact be similar to one another.
+For instance, our estimates for skewness and kurtosis reveal that the 10- and 90-percentile shocks are both negative and have almost the same size among the poor; this is why we looked farther out at the 3rd and 97th percentiles.
+By contrast, our approach allows us to evaluate the second and higher moments of $F_{t}$ and estimate pass-through rates that explicitly shift with the economically relevant shock, i.e., with $\zeta_{it}=F_{t}^{-1}(u_{it})$.
 It can thus describe more transparently how income shocks of different signs/sizes shift consumption.
-How does a less negative $\phi^{(2)}$ among the rich reconcile with a more negative $\phi^{(2)}$ among the old, who are presumably also rich?
-The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between the nature of income risk and the consumption technology.
-Skewness of permanent shocks is twice as large among the rich ($-4.86$) as it is among the old ($-2.39$).
-The consumption distribution is not too different between the two, however, so the rich must have more insurance to a bad shock of given size.
-This explains why $\phi^{(2)}$ is four times smaller among them.
-In both cases (old, rich), bad permanent shocks from a given position in the distribution decrease consumption more strongly than among their counterparts (young, poor).
+How does a less negative $\phi^{(2)}$ among the rich (smaller response to larger bad shocks) reconcile with a more negative $\phi^{(2)}$ among the old (stronger response), who are presumably also rich?
+The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between income risk and consumption response.
+Skewness of permanent shocks is twice as large among the rich ($-4.86$) as among the old ($-2.39$).
+The consumption distribution is similar for both, so the rich must have overall more insurance to a given bad shock.
+This explains why $\phi^{(2)}$’s magnitude is four times smaller among them.
+In both cases (old, rich), bad permanent shocks from a given position in the distribution reduce consumption more than among their counterparts (young, poor).
 5.3.3 Results by wealth and education
@@
 We use conventional definitions for ‘low’ versus ‘high’ wealth (the cutoff is median wealth in the sample) and ‘no’ versus ‘some’ college (the cutoff is 12 years of education).
+Appendix table C.1 provides summary statistics for these subsamples.
 [^46] Wealth is defined as the sum of the value of housing, vehicles, business, investment, pension, savings, and other assets, net of mortgage and medical, student, and credit card debt.
-Appendix table C.1 provides summary statistics for these subsamples.
 Table 6: Estimates of the consumption function, by wealth and education
@@
 Three points stand out.
-First, $\phi^{(1)}$ decreases monotonically as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college).
-As $\phi^{(1)}$ measures the transmission of the average permanent shock, these groups possess gradually more insurance to average shocks; this reflects the insurance role of wealth, which increases monotonically across the groups.
-Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move across groups, with the exception of group 2 (low wealth and some college) where it equals zero.
-Negative shocks thus mostly transmit more strongly than positive shocks, but wealthier and more educated people have more insurance to bad shocks of given size.
-These people face, however, much stronger downwards income risk.
-This pattern is consistent with the gradient between poor and rich in table 5.
-Third, the transmission parameters of transitory shocks are mostly statistically insignificant, with two exceptions.
-In group 1 (low wealth, no college), $\psi^{(1)}$ is significant but negative, which is counterintuitive at first, yet $\psi^{(2)}$ is also negative, large, and significant.
-Large negative transitory shocks thus reduce consumption substantially among low wealth, low education households.
-In group 3 (high wealth, no college), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones.
-We argued earlier that this helps reconcile the large pass-through found in quasi-experimental studies with the negligible pass-through found in studies that employ covariance restrictions.
-Note that the size of group 3 is by far the smallest in our sample.
-5.3.4 Link with structural model
-Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, and income distributions.
+First, $\phi^{(1)}$ decreases monotonically from group 1 (low wealth, no college) to group 4 (high wealth, some college).
+$\phi^{(1)}$ measures the transmission of the average permanent shock, so these groups possess gradually more insurance to average shocks.
+This reflects the insurance role of wealth, which increases monotonically across groups.
+Second, $\phi^{(2)}$ is negative but increases (i.e., less negative) as we move rightwards across groups, with the exception of group 2 (low wealth, some college) where it equals zero.
+Negative shocks thus transmit weakly more strongly than positive shocks among everyone, but wealthier and more educated agents have more insurance to bad shocks of a given size.
+They also face, however, much stronger downside income risk, a pattern consistent with the gradient in income risk between poor and rich in the earlier discussion.
+Third, the transmission parameters of transitory shocks are mostly insignificant, with two exceptions.
+In group 1 (low wealth, no college), $\psi^{(1)}$ and $\psi^{(2)}$ are both significant and negative, leading to the odd prediction that larger bad shocks may increase consumption slightly while smaller good shocks reduce it.
+In group 3 (high wealth, no college; the smallest group in our sample), $\psi^{(2)}$ is positive and significant, suggesting that large good shocks have a larger pass-through than smaller ones (thus opposite from group 1).
+Both cases are at odds with the bulk of estimates for the transmission of transitory shocks throughout the paper that suggest that such shocks are fully insured, so these results remain a puzzle.
+5.3.4 Link with structural model and some estimates of welfare costs
+Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or income distribution.
 Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
-Madera (2019) finds that earnings and nondurable consumption are correlated in the middle of the earnings distribution, but less so in the tails compared to earnings and durable consumption.
-Negative shocks are associated with larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
+Madera (2019) finds that earnings and nondurables are correlated mainly in the middle of the earnings distribution, with weaker correlations in the tails – especially when compared to earnings and durables.
+This suggests that tail shocks pass through to durables (which we do not model) more than to nondurables.
+Negative shocks exhibit larger pass-through rates than equal-sized positive shocks, albeit at twice the magnitude that we estimate here.
 De Nardi et al.
-(2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings (thus also enables more insurance to a given shock).
-They show that the BPP method that relies on second-order moments implies a substantially higher pass-through rate than the true rate in the model.
+(2019) find that tail risk has large welfare costs because, among other things, it increases precautionary savings enabling more insurance to a given shock.
+They show that the BPP method relying on the workhorse consumption-income covariance implies a substantially higher pass-through rate than the true rate in the model.
 Busch and Ludwig (2024) simulate a model with borrowing constraints.
-We do not model such constraints, yet a strength of our approach is that it enables us to measure partial insurance without taking a stance on the precise economic mechanisms at play.
-They find that the marginal propensity to consume increases with the severity of a negative permanent shock as households move towards the constraint, while it decreases with the magnitude of a positive shock.
-Their pass-through rates are slightly higher than ours, though their model does not feature other insurance over self-insurance.
+While we do not explicitly model such constraints, a strength of our approach is that it measures insurance without relying on specific economic mechanisms.
+They find that the marginal propensity to consume rises with the size of a negative permanent shock as agents approach the constraint, but falls with the size of a positive shock.
+Their pass-through rates are slightly higher than ours, though their model excludes other form of insurance beyond self-insurance.
 Guvenen et al.
 (2024) simulate a model with retirement, stochastic lifetimes, and altruism.
-The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, higher than ours but lower than BPP.
-They show that measuring pass-through as in BPP overestimates the true rate by 1.5 times.
-Without non-employment shocks, agents are willing to give up almost about 22.3% of consumption to live in a world without risk.
+The consumption response varies with the size and sign of shocks; the pass-through of the average negative permanent shock is between 0.29 and 0.46, depending on age, which is higher than ours but lower than BPP.
+They show that measuring pass-through as in BPP overestimates the true transmission rate by 1.5 times.
 We read these results as not too different from ours.
-Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow, so those shocks end up affecting consumption by relatively more than other shocks.
-The magnitude of our pass-through rates seems uniformly lower than theirs (to some extent this is because these models focus on self-insurance), though other quasi reduced-form studies using the recent PSID, most notably ABB, find similar rates to ours.
-6 Welfare costs of tail income risk
-In this section, we calculate the welfare costs of income risk, which, as we found previously, is negatively skewed, highly leptokurtic, and features a nonlinear pass-through to consumption.
-We are particularly interested in contrasting the implications of linear and nonlinear transmission.
-To do so, we draw a sequence of income shocks over the lifecycle, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare.
-Our interest lies in the costs of non-Gaussian relative to (counterfactual) Gaussian income risk.
-Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021, 2024).
-We draw permanent and transitory income shocks for 50,000 households from a mixture of two normal distributions, namely
-$\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob.
-}p_{\zeta}\\ {\cal N}(\mu_{\zeta,2},\sigma_{\zeta,2})\text{ with prob.
-}1-p_{\zeta}\\ \end{cases}\text{ and }~{}v_{it}\sim\begin{cases}{\cal N}(\mu_{v,1},\sigma_{v,1})\text{ with prob.
-}p_{v}\\ {\cal N}(\mu_{v,2},\sigma_{v,2})\text{ with prob.
-}1-p_{v}.\\ \end{cases}$
-There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2).
-We assume $\mu_{1}<0$ for identification.
-Appendix table D.1 reports the parameter estimates and the (excellent) model fit.
-In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
-We simulate alternative consumption profiles in response to the income shocks using the linear and nonlinear specifications
-$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$
-$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\zeta_{it}^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\zeta_{it}v_{it}.$
-These reflect the model-derived linear and quadratic consumption functions (4) and (5), with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock).
-In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data.
-We simulate consumption over a working life of 35 years, starting at age 30.
-In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
-[^44] We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
-We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample.
-In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample.
-We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, and logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$.
-We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk.
-This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation ${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}$.
-Here, ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles that reflect less (Gaussian) and more (non-Gaussian) income risk, respectively.
-We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
-In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively.
-The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission.
-Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is much higher relative to a Gaussian density.
-In a second scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the quadratic specification in response to Gaussian and non-Gaussian income shocks, respectively.
-The estimate of $\alpha_{i}$ here reflects the welfare cost of tail risk in the presence of nonlinear transmission.
-Given the evidence that larger negative shocks have a stronger impact on consumption when nonlinear transmission is allowed for, we expect larger welfare costs of tail risk in this scenario.
-In a third scenario, the less risky profile ${\cal C}_{i}^{*}$ reflects the linear specification in response to Gaussian risk, while the riskier profile ${\cal C}_{i}$ is simulated from the quadratic specification in response to non-Gaussian shocks.
-Here $\alpha_{i}$ reflects both the welfare cost of tail risk and the bias induced by neglecting the possibility of nonlinear transmission of such risk.
-Table 7 averages $\alpha_{i}$ over cross-sections of households defined by the magnitude of the permanent income shocks they experience, in the case of $\gamma=2$.
-The first row presents the average cost across all households, unconditional on a given shock.
-This is close to zero across all three scenarios, or even negative in some cases, suggesting that households may in fact prefer tail over Gaussian risk.
-Inspecting the parameter estimates in table D.1, tail risk is associated with a very small probability (about 5%) that shocks come from a distribution with negative mean and large variance, and a very high probability that they come from a distribution with small positive mean but tiny variance.
-As such, most households will never actually experience shocks from the first distribution; their shocks will come from the latter, whose variance is an order of magnitude smaller than the variance of Gaussian risk.
-For many, Gaussian risk in fact implies larger variability, so households are on average willing to pay to move to the non-Gaussian environment.
-Table 7: Welfare costs of non-Gaussian relative to Gaussian income risk
-Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
-${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
-Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
-% households# (1) (2) (3)
-All households 100% -0.81 -0.88 0.66
-$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
-$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
-$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
-Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$.
-We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 3.
-The first row presents the average costs across all households.
-The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
-The next rows zoom in on households who experience large negative permanent shocks at least once in their life.
-These are households who happen to draw from the first distribution in the mixture and the consequences of tail risk are clearly felt among them.
-Those who face a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once in their lifetime) will give up 4.8% of their lifetime consumption to live in a world with Gaussian risk when such risk is transmitted linearly.
-This increases to 6.36% and 8.54% for shocks respectively $4\sigma$ and $5\sigma$ below the mean – all shown in column 1.
-These costs are already sizable, and they increase further by about 60% in the presence of nonlinear transmission (column 3), exceeding an eighth of lifetime consumption (13.23%) for those who experience the worst shocks.
-Those costs would be even larger if we considered a counterfactual zero risk scenario as in Guvenen et al.
-(2024).
-These welfare calculations are ‘back-of-envelope’, as they neglect heterogeneity in income risk and consumption transmission, they treat wealth as exogenous, and they do not account for borrowing constraints.
-Appendix table D.2 repeats the calculations allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline.
-Appendix table D.3 calculates the welfare costs by wealth; the least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs, as expected.
-As the likelihood that a borrowing constraint binds depends on wealth or age, the age- and wealth-specific results are indirectly reflective also of such constraints.
-Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs lie in the middle of the wealth-specific costs, so exogenous wealth is unlikely to have caused a serious bias in our calculations.
-The welfare estimates are also robust to small variations in the choice of $\gamma$ (appendix table D.4).
+Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow.
+Therefore, such shocks end up affecting consumption more than other shocks.
+The magnitude of our pass-through rate of permanent shocks is uniformly lower than its structural counterpart (consumption transmission in the data likely reflects multiple insurance sources, beyond self-insurance and taxes examined in these models), yet other empirical studies using the updated PSID, e.g., ABB and Arellano et al.
+(2023), find similar rates to ours.
+Finally, we obtain back-of-the-envelope estimates of the welfare costs of tail income risk.
+Specifically, we simulate permanent income shocks from mixtures of normal distributions, we use those shocks to simulate consumption responses based on the quasi-reduced form consumption function in this paper, and we use power utility to calculate the discounted present value of a lifetime consumption stream associated with a given profile of income risk.
+Our baseline choice is tail risk similar to what we estimate in the data, while our counterfactual is Gaussian risk with similar variance.
+We then estimate the welfare costs of tail risk as the fraction of lifetime consumption a household facing Gaussian risk (counterfactual) would give up to avoid tail risk (baseline).
+We provide full details in appendix D.
+Households facing large downside risk (ranging from a permanent shock 3 standard deviations below the mean at least once in their lifetime – 40% of households experience this – to a shock 5 standard deviations below the mean – a year’s long unemployment, 17% of households experience this at least once) would give up 4%-8% of lifetime consumption when income risk transmits linearly.
+This rises to 6%-12%, almost up to 1/8 of lifetime consumption, in the presence of nonlinear transmission.
+These estimates vary across the income and wealth distributions (appendix D), but the results remain in the same ballpark.
+Although we calculate the welfare costs outside of a full structural model (among other things, taking wealth as given; hence ‘back-of-the-envelope’), our estimates seem to fall in the same ballpark as estimates from the structural papers described above.
+As a comparison, households in Busch and Ludwig (2024) sacrifice 0.5%-14% of lifetime consumption to live in a world with Gaussian risk, while households in Guvenen et al.
+(2024) sacrifice almost 22% of consumption to live in a world without any risk.
 6 Conclusion
-We measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution.
-Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented in many recent papers.
-We use recent data from the PSID and establish several new findings.
+We introduce a simple analytical framework to measure the degree of consumption partial insurance to income shocks, accounting for higher-order moments of the income distribution.
+Such moments, in particular left skewness and excess kurtosis, are important features of unexplained income growth, documented recently in the U.S.
+and other countries.
+We use recent income and consumption data from the PSID and establish several findings.
 First, permanent and transitory shocks to disposable household income exhibit negative skewness and very large excess kurtosis.
-Second, in estimating a linear consumption model over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones.
+Second, in estimating a linear consumption function over the recent years, we find a low pass-through of permanent shocks (large insurance) and full insurance to transitory ones.
 We attribute one third of the difference from BPP to the consumption imputation from the CEX, another third to the biennial nature of our data, and the remainder to time effects.
@@
 Fourth, in estimating a nonlinear consumption function, we find that bad permanent shocks transmit more strongly than good ones, and the pass-through increases with the severity of the shock.
-By contrast, transitory shocks are fully insured in the full sample, but they are not insured among the less wealthy.
-The nature of income risk and consumption transmission vary by age and the household’s position in the income distribution.
-Fifth, tail risk incurs substantial welfare costs, up to 1/8 of lifetime consumption.
+By contrast, transitory shocks are mostly fully insured.
+The nature of income risk and consumption transmission vary substantially by age and the household’s position in the income distribution.
 In summary, the paper offers a straightforward analytical framework to measure the transmission of income shocks into consumption, allowing such transmission to vary with the sign and size of the shock.
@@
 Appendix A Derivation of the consumption function
-In this appendix we derive the consumption function under a permanent-transitory specification for log income.
+This appendix shows how to derive the consumption function under a permanent-transitory specification for log income.
 We follow similar derivations in Blundell et al.
@@
 where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks.
-[^1] With biennial data, matrix $\mathbf{A}$ is given by (different from $\mathbf{A}$ with annual data in the paper) $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
+[^1] With biennial data, matrix $\mathbf{A}$ is different from $\mathbf{A}$ with annual data in the main paper and is given by $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
 $\text{det}(\mathbf{A})$ $=$ $-8\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\Big{[}\Big{(}k_{\zeta_{t}}$ $-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
@@
 [^2] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant would not always be strictly away from zero.
-How do different levels of skewness and kurtosis connect with the transmission parameters?
-Consider the transmission of permanent shocks, abstracting from transitory shocks (i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$).
-$\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$, which corresponds to the first equation in (B.1).
+How do different levels of skewness and kurtosis influence the transmission parameters?
+Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$.
+$\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$.
 Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks.
@@
   • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$.
-Consumption growth and permanent shocks co-vary more strongly than in the linear consumption process, i.e., there is excess co-movement.
+Consumption growth and permanent shocks co-vary more strongly than the linear consumption process suggests, i.e., there is excess co-movement.
 For $\gamma_{\zeta_{t}}<0$, this configuration identifies $\phi^{(2)}_{t}<0$ and implies that negative shocks transmit more strongly than other shocks.
 Negative shocks are often the largest shocks (due to left skewness), which explains the excess co-movement.
-Ceteris paribus, a less skewed distribution of shocks ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement.
+Ceteris paribus, a less skewed distribution ($\gamma_{\zeta_{t}}$ drops in magnitude) implies a more negative $\phi^{(2)}_{t}$, because a more symmetric distribution associated with fewer large shocks requires a stronger transmission of said shocks to maintain a given excess co-movement.
 If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
   • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$.
-Consumption growth and permanent shocks co-move less strongly than in the linear process, i.e., there is excess smoothness.
+Consumption growth and permanent shocks co-move less strongly than the linear process suggests, i.e., there is excess smoothness.
 For $\gamma_{\zeta_{t}}<0$, this identifies $\phi^{(2)}_{t}>0$ and suggests a muted pass-through of negative shocks.
@@
 For given excess co-movement, the smaller the kurtosis of permanent shocks, the larger the value of $\phi^{(2)}_{t}$ must be to achieve said excess co-movement.
-In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by this and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity.
-The sensitivity analysis is thus more nuanced so simulations can prove useful.
+In practice, $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are jointly pinned down by these and other equations of system (B.1), and they are also determined jointly with the transmission parameters of transitory shocks, to which analogous statements apply – skipped here for brevity.
+The sensitivity analysis is thus more nuanced, so simulations can prove useful.
 Starting from the baseline estimates of the income process (table 2), reducing the left skewness of permanent shocks by half increases the magnitude of $\phi^{(2)}_{t}$ by almost 10% (more negative).
@@
 An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error.
-In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the generic distribution.
+In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution.
 Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$.
-Therefore, identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks.
+Identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks.
 Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks.
@@
 C.3 Sources of tail income shocks
-Table C.3 provides evidence on the sources of large income reductions that contribute to income left skewness and excess kurtosis.
+Table C.3 presents evidence on the underlying sources of large income reductions that drive the left skewness and excess kurtosis of income growth.
 Columns 1 & 2 report average unemployment, disability, bad health, etc, in the bottom 10% and top 90% of the distribution of residual income changes $\Delta y_{it}$, respectively.
@@
 We require non-missing data on expenditure and basic demographics, and we drop those with zero food expenditure given the nature of the imputation subsequently.
-Similar to the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare.
+As in the PSID, we define consumption as the sum of real expenditure on nondurable goods and services, namely food (at home and outside), utilities, out-of-pocket health expenses, public transport, vehicle expenses, education, and daycare.
 As the CEX collects quarterly data over four rolling quarters, we generate annual consumption as the sum of consumption over four quarters.
@@
 Table C.4 presents a comparison of means between the PSID and CEX samples for a host of household characteristics over the period 1999–2019.
-The two samples are very close with respect to age, family size, number of children, race, education, region of residence, or labor market participation.
+The two samples are very close with respect to age, family size, number of children, race, education, region of residence, and labor market participation.
 Household earnings are higher in the PSID than in the CEX; BPP noted this also for the earlier period (1980–1992) and argued that it is due to the more comprehensive definition of income in the PSID.
@@
 As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
-C.5 Comparison using the original BPP data
+C.5 Estimation using the original BPP data
 Table C.5 uses the original BPP data over 1980–1992 (taken from the replication package of Blundell et al., 2008, available online) and does three things.
-First, we replicate BPP over the original data/period of time, using BPP’s annual growth rates of income and consumption.
+First, we replicate BPP over the original data/period of time, using our code but BPP’s annual growth rates of income and consumption.
 This is column 1 in table C.5.
@@
 Note that despite the use of biennial data, $\phi^{(1)}$ and $\psi^{(1)}$ still express the pass-through at annual rates (see appendix B).
-$\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains qualitatively unaffected.
+$\phi^{(1)}$ drops by 20% to $0.518$, while $\psi^{(1)}$ remains small but turns statistically significant.
 With biennial data, we only identify the pass-though of an ‘aggregate’ permanent shock over two years.
 Some higher frequency shocks are muted at the observed lower frequency, given that we can only observe an average of two one-year permanent shocks over the two-year period.
-In other words, the use of biennial data allows us to only measure an ‘aggregate’ or average pass-through within a given biennial period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
+In other words, the biennial data allow us to measure only an ‘aggregate’ or average pass-through within a given period, and some higher-frequency shocks are inevitably smoothed out at this observed lower frequency.
 Third, we re-estimate the linear model assuming that the target second-order income and consumption moments are time-invariant.
@@
 (2021) who also find smaller pass-through rates (higher degree of partial insurance) when they impose time-invariance in BPP’s target moments.
-There are some caveats comparing the exercise in table C.5 and our results with the most recent data in table 3.
+There are some caveats comparing the exercise in table C.5 and our results from the recent PSID in table 3.
 First, our main exercise uses consumption data internally available in the PSID, while table C.5 uses consumption imputed from the CEX.
@@
 Growth rates: annual biennial annual biennial
-(original BPP)
 (1) (2) (3) (4)
@@
 In line with BPP, estimation is done via diagonally weighted GMM; GMM standard errors are in parentheses.
-The very small unimportant numerical differences between column 1 herein and the original BPP are likely due to different software and optimization algorithms (we use Matlab and fmincon).
+The small unimportant numerical differences between column 1 herein and the original BPP estimates are likely due to different software and optimization algorithms (our code uses Matlab and fmincon while BPP use Gauss).
 #$\theta$ is the moving average parameter of the transitory shock.
@@
 As such, this table accompanies table 3 in the main text, which reports results using consumption internally available in the PSID.
-In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data directly available in the PSID, regardless of whether we target only second- or also higher-order moments.
+In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments.
 This echoes our earlier discussion about the imputation inflating the pass-through of shocks.
 We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases.
-Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude ($-0.418$) and statistically significant.
-This is counterintuitive and raises further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved.
+Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude and negative ($-0.418$).
+This is counterintuitive, raising further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved.
 We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function.
-However, $\phi^{(2)}<0$ in that case with a magnitude very similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant).
-This is, once again, indicative of the asymmetric pass-through of negative versus positive permanent shocks.
+$\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant).
+Once again, this is indicative of the asymmetric pass-through of negative versus positive permanent shocks.
 Similar to our baseline, we also find $\psi^{(2)}>0$.
@@
 The opposite is true for transitory shocks; skewness is negative in the bottom income tertile but strongly positive in the top.
-Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution.
+Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution, respectively.
 Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65.
@@
 By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
-Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, possess more insurance than the poor.
+Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
 Table C.6: Estimates of the income process, optimal weighting
@@
 As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification.
-The odd numbered columns present parameter estimates when only second-order moments are targeted.
-$\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across the groups.
+The odd-numbered columns present parameter estimates when only second-order moments are targeted.
+$\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across groups.
 This is similar to what we found in the quadratic specification.
-However, as soon as we target up to fourth-order moments in the linear specification, shown in the even numbered columns, this monotonic pattern is lost.
+However, as soon as we target up to fourth-order moments in the linear specification, shown in the even-numbered columns, this monotonic pattern is lost.
 Transitory shocks transmit strongly in group 1, and with a negative coefficient.
@@
 The underlying income moments also vary by wealth and education (results not shown for brevity).
-We do not report results when up to third-order moments of income and consumption are targeted in the estimation; these are available upon request.
+We do not report results when up to third-order moments of income and consumption are targeted, but these are available upon request.
 Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019.
@@
 This table accompanies the quadratic specification estimates in table 6 in the paper.
-Appendix D Welfare costs details
-Table D.1 presents the parameter estimates and model fit of the quantitative model for income (mixture of normal distributions).
-Table D.2 calculates the welfare costs allowing for heterogeneity in income risk and consumption transmission by age, while table D.3 allows for heterogeneity by wealth.
-Table D.4 reports average welfare costs for different values of the utility coefficient $\gamma$.
+Appendix D Some welfare costs of tail income risk
+Given our empirical, quasi reduced-form approach to measuring partial insurance, we calculate the welfare costs of income risk outside of a fully specified structural model.
+Specifically, we draw a sequence of income shocks over the lifetime, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare based on those consumption responses.
+Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear and nonlinear transmission.
+Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021, 2024).
+We draw permanent and transitory income shocks for 50,000 households from a mixture of two normal distributions, namely
+$\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob.
+}p_{\zeta}\\ {\cal N}(\mu_{\zeta,2},\sigma_{\zeta,2})\text{ with prob.
+}1-p_{\zeta}\\ \end{cases}\text{ and }~{}v_{it}\sim\begin{cases}{\cal N}(\mu_{v,1},\sigma_{v,1})\text{ with prob.
+}p_{v}\\ {\cal N}(\mu_{v,2},\sigma_{v,2})\text{ with prob.
+}1-p_{v}.\\ \end{cases}$
+There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2).
+We assume $\mu_{1}<0$ for identification.
+Table D.1 reports the parameter estimates and the (excellent) model fit.
+In the counterfactual scenario with Gaussian risk, we fix $p=0$, $\mu_{2}=0$, and set $\sigma_{2}^{2}$ equal to the empirical variance of each shock.
+We simulate alternative consumption profiles in response to the income shocks using the linear and nonlinear specifications
+$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it},$
+$\Delta\ln C_{it}$ $=\mu^{c}_{t}+\phi_{it}^{(1)}\zeta_{it}+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\zeta_{it}^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\zeta_{it}v_{it}.$
+These reflect the linear and quadratic consumption functions (4) and (5) in the paper, with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock).
+In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data.
+We simulate consumption over a working life of 35 years, starting at age 30.888We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
+We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample.
+In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample.
+In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
 Table D.1: Normal mixtures, parameter estimates and model fit
@@
 Estimation is done via the method of simulated moments, based on 50,000 random draws for each type of income shock.
+We calculate the (ex-ante) welfare implied by a given consumption profile ${\cal C}_{i}=\{C_{it}\}_{t=0}^{T}$ using constant coefficient of relative risk aversion utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\frac{C_{it}^{1-\gamma}-1}{1-\gamma}$ if $\gamma\neq 1$, and logarithmic utility ${\cal U}({\cal C}_{i})=\sum_{t=0}^{T}\beta^{t}\log(C_{it})$ if $\gamma=1$.
+We then calculate the welfare cost of tail risk as the percentage reduction in lifetime consumption that a household who faces Gaussian income risk is willing to accept in order to avoid tail (namely, non-Gaussian) risk.
+This is given by $\alpha_{i}$ for a given household $i$, which, similar to the welfare calculations in Wu and Krueger (2021), satisfies the indifference equation ${\cal U}\big{(}(1-\alpha_{i}){\cal C}_{i}^{*}\big{)}={\cal U}\big{(}{\cal C}_{i}\big{)}$.
+Here, ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are alternative consumption profiles that reflect less (Gaussian) and more (non-Gaussian) income risk, respectively.
+We consider three alternative scenarios over which we calculate the welfare costs of non-Gaussian income risk.
+In a first scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the linear specification in response to Gaussian and non-Gaussian income shocks, respectively.
+The estimate of $\alpha_{i}$ in this case reflects the welfare cost of tail risk in the context of linear transmission.
+Tail risk is still costly in this context because the probability of very negative shocks (reflecting the long left fat tail) is higher relative to a Gaussian density.
+In a second scenario, both ${\cal C}_{i}^{*}$ and ${\cal C}_{i}$ are simulated from the quadratic specification in response to Gaussian and non-Gaussian income shocks, respectively.
+The estimate of $\alpha_{i}$ here reflects the welfare cost of tail risk in the presence of nonlinear transmission.
+Given the evidence that larger negative shocks have a stronger impact on consumption when nonlinear transmission is allowed for, we expect larger welfare costs of tail risk in this scenario.
+In a third scenario, the less risky profile ${\cal C}_{i}^{*}$ reflects the linear specification in response to Gaussian risk, while the riskier profile ${\cal C}_{i}$ is simulated from the quadratic specification in response to non-Gaussian shocks.
+Here $\alpha_{i}$ reflects both the welfare cost of tail risk and the bias induced by neglecting the possibility of nonlinear transmission of such risk.
+Table D.2 averages $\alpha_{i}$ across households, for $\gamma=2$.
+The first row presents the unconditional average cost across all households.
+The following rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
+We also count the fraction of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+Two main points stand out.
+Table D.2: Welfare costs of non-Gaussian relative to Gaussian income risk
+Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
+${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
+Specification: ${\cal C}_{i}$ linear ${\cal C}_{i}$ quadratic ${\cal C}_{i}$ quadratic
+% households# (1) (2) (3)
+All households 100% -0.81 -0.88 0.66
+$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
+$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
+$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
+Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$.
+We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 3.
+The first row presents the average costs across all households.
+The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
+#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+First, the unconditional cost is small but negative in both the first (linear) and second (nonlinear) scenarios, suggesting households may in fact prefer tail over Gaussian risk.
+While this may seem counterintuitive, table D.1 shows that tail risk entails a small probability (about 5%) of shocks from a distribution with negative mean and large variance, and a high probability of shocks from one with a small positive mean and very low variance.
+Most households never actually experience shocks from the former; instead, they face the latter, whose variance is much smaller than that of Gaussian risk.
+Counterfactually, Gaussian risk implies larger variability for many, so households are on average willing to pay to move to the non-Gaussian setting.
+However, the average cost turns positive in scenario 3, which evaluates the consumption loss of tail risk and the bias from a misspecified consumption function.
+That is, neglecting non-Gaussian features and nonlinear transmission leads to an overestimation of consumer welfare by 0.66% of lifetime consumption.
+Second, the welfare costs become substantial for households who actually experience large negative shocks at least once in life – those drawing from the first distribution in the mixture.
+Households facing a permanent shock at least $3\sigma$ below the mean ($\min\{\zeta_{it}\}<-0.50$; 40.31% of households experience this at least once) give up 4.8% of lifetime consumption to avoid tail risk when risk is transmitted linearly (column 1, second row).
+This rises to 6.39% in the presence of nonlinear transmission (column 2, second row), as bad shocks transmit more strongly in this case.
+For worse events, costs increase further to almost 1/8 or 11.89% of lifetime consumption among households with at least one $5\sigma$ shock below the mean (column 2, fourth row; $\min\{\zeta_{it}\}<-0.95$; 17.07% of households experience this at least once), and are higher still in scenario 3 where we evaluate the welfare cost of tail risk and the bias from a misspecified consumption function.
+They are also (much) higher under a counterfactual zero risk scenario, like that explored in Guvenen et al.
+(2024).
+These findings recur when we account for heterogeneity in income risk, consumption transmission, or wealth.
+Table D.3 calculates the welfare costs allowing for age-specific risk and consumption transmission; the costs are in the same ballpark as the baseline described above.
+Table D.4 further calculates the welfare costs by wealth.
+The least wealthy have costs almost twice as large as the baseline, while the wealthier have smaller costs.
+This gradient is explained by the different degrees of consumption insurance between the two groups.
+As the likelihood that a borrowing constraint binds depends on wealth and age, the wealth- and age-specific results are indirectly reflective also of such constraints.
+Without a structural model, we cannot allow households to switch wealth groups endogenously (e.g., in response to an income shock), yet the baseline costs (table D.2) lie in the middle of the wealth-specific costs (table D.4), so exogenous wealth is unlikely to have caused a serious bias in our baseline calculations.
+The welfare estimates are also robust to small variations in $\gamma$ (table D.5).
 Table D.3: Welfare costs of non-Gaussian relative to Gaussian income risk, age-specific income risk and transmission parameters
@@
 Table D.4: Welfare costs of non-Gaussian relative to Gaussian income risk, by wealth
-Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$
+Average welfare costs ${\operatorname*{\mathbb{E}}}(\alpha_{i})$, $\gamma=2$
 ${\cal C}_{i}^{*}$ linear ${\cal C}_{i}^{*}$ quadratic ${\cal C}_{i}^{*}$ linear
@@
 We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text.
-However, the consumption parameters can be inferred from table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample.
+The consumption parameters can be proxied by table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample.
 The first row presents the average costs across all households in the wealth group.

### v6

@@
 We obtain four main findings:
-First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at $-0.95$ and $-1.31$, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively.
+First, permanent and transitory shocks to disposable household income are negatively skewed, with skewness coefficients at 0.83-0.83 and 1.43-1.43, respectively, and highly leptokurtic, with kurtosis coefficients at $44.39$ and $49.23$, respectively.
 Left skewness implies that a negative shock is typically more unsettling than a positive shock, because it is farther away from the zero mean.
@@
 In other words, we find much larger partial insurance to permanent shocks in recent years than originally found in 1980–1992.222Using recent PSID data but a different method, ABB and Arellano et al.
-(2023) find a similarly low pass-through of permanent shocks.
+(2023) find a similarly low pass-through of persistent shocks.
 We discuss this further when we present our empirical results.
@@
 3.) Transitory shocks are fully insured in the full sample, at least from a statistical point of view, although the point estimates of their transmission parameters suggest that large transitory windfalls shift consumption considerably, consistent with quasi-experimental evidence of large responses to transitory gains (e.g., Fagereng et al., 2021).
+These and previous findings are robust to allowing for shocks that have linearly persistent, though not necessarily permanent, effects on income.
 Overall, the results are consistent with an environment in which households save a good income shock but dissave in the presence of a bad one.
@@
 (2024).
-Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk (a more distant counterfactual from ours).
+Agents in the former paper sacrifice between 0.5% and 14% of lifetime consumption to live in a world with Gaussian risk (a similar counterfactual to ours), while households in the latter sacrifice almost 22% of consumption to live in a world without risk.
 [^3] These papers calculate welfare costs using full structural models in which the consumption responses depend on endogenously accumulated wealth.
@@
 They are comparable to the existing literature on partial insurance and ‘portable’ as they describe transparently how shocks of different signs/magnitudes shift consumption.
-The second branch of the literature studies specific insurance mechanisms through calibrated consumption lifecycle models.
+The second branch of the literature studies specific insurance mechanisms using calibrated consumption lifecycle models.
 [^6] Gourinchas and Parker (2002) and Krueger and Perri (2006) are early precursors in this strand.
-Early works focus on self-insurance.
-Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more insurance.
-More recent works study the insurance consequences of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), or added workers (Wu and Krueger, 2021).
-Except Madera (2019), these papers abstract from higher income moments due to computational difficulties.
+Early works focus on self-insurance: Kaplan and Violante (2010) find less insurance than BPP, while Guvenen and Smith (2014), who jointly estimate income risk and self-insurance, find more.
+Recent studies examine the insurance implications of labor supply (Heathcote et al., 2014), durables (Madera, 2019), health (Blundell et al., 2020), and added workers (Wu and Krueger, 2021).
+Except for Madera (2019), these papers abstract from higher income moments due to computational reasons.
 De Nardi et al.
-(2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk increases partial insurance due to stronger precautionary motives.
-In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis; they distinguish between good and bad shocks and show that the latter are worse insured.
+(2019) embed a non-Gaussian income process into a consumption/savings model and show that tail risk strengthens precautionary motives and increases partial insurance.
+In a similar model, Busch and Ludwig (2024) target income skewness and kurtosis, distinguishing good from bad shocks and showing the latter are worse insured.
 Guvenen et al.
-(2024) find that tail shocks are worse insured than average shocks but the benchmark method in BPP overstates the true extent of insurance.
-Unlike this quantitative literature, we do not take a stance on the insurance mechanisms available to households; we measure aggregate insurance from all mechanisms together, while accounting for higher-order features of the income distribution, keeping our model parametrically unspecified.
+(2024) find that tail shocks are worse insured than average ones but BPP’s benchmark method overstates insurance against them.
+Unlike this quantitative literature, we take no stance on specific insurance mechanisms; we measure aggregate insurance across all mechanisms while accounting for higher-order moments and keeping our model parametrically unspecified.
 The asymmetric response to bad and good shocks is recently found in surveys that elicit responses to hypothetical events.
@@
 Geweke and Keane (2000) and Bonhomme and Robin (2010) use the PSID to document, respectively, left skewness and excess kurtosis in earnings growth, while Guvenen et al.
-(2021) provide even richer evidence from administrative data.
+(2021) provide richer evidence from administrative data.
 Guvenen et al.
@@
 (2018) study the cyclical nature of skewness in several countries.
-While these papers are mostly descriptive of the higher-order features of income, we further document the implications of these features for consumption.
+While these papers are mostly descriptive of the higher-order features of income, we further document the implications for consumption.
 We thus contribute to a growing literature on the implications of tail risk, e.g., for growth (Barro, 2009), asset prices (Backus et al., 2011; Ghosh and Julliard, 2012; Constantinides and Ghosh, 2017), business cycles (McKay, 2017), and labor market dynamics (Ai and Bhandari, 2021).
-In the rest of the paper, section 2 presents the income and consumption processes and section 3 discusses identification.
-Section 4 discusses the implementation, section 5 presents the results, and section 6 concludes.
+In the rest of the paper, section 2 introduces the income and consumption processes and section 3 discusses their identification.
+Section 4 presents the empirical implementation, section 5 discusses the results, section 6 extends the main analysis to a more general income process, and section 7 concludes.
 The appendix has additional derivations and results.
@@
 where $\mathbf{X}_{it}$ is a vector of characteristics in the $t=0$ information set of the household that drive the deterministic profile of income over the lifecycle (e.g., age, education, race) and $\boldsymbol{\delta}$ is its loading factor.
-The remaining terms make up the stochastic component of income, consisting of permanent income $P_{it}$ and transitory shock $v_{it}$.
-We assume that permanent income has a unit root, $P_{it}=P_{it-1}+\zeta_{it}$, where $\zeta_{it}$ is the permanent income shock.
-These modeling choices give rise to the familiar permanent-transitory formulation
-$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vspace{-1ex}$ (3)
+The remaining terms make up the stochastic component of income, consisting of persistent income $P_{it}$ and transitory shock $v_{it}$.
+We assume that $P_{it}$ has a unit root, namely $P_{it}=P_{it-1}+\zeta_{it}$, so in fact $P_{it}$ is permanent income and $\zeta_{it}$ is the permanent income shock.
+This modeling choice gives rise to the familiar permanent-transitory formulation
+$\Delta y_{it}=\zeta_{it}+\Delta v_{it},\vskip-4.30554pt$ (3)
 where $\Delta y_{it}=\Delta\ln Y_{it}-\Delta\mathbf{X}_{it}^{\prime}\boldsymbol{\delta}$ is income growth net of the deterministic profile, i.e., unexplained income growth, and $\Delta$ is the first difference operator.
 [^8] See Meghir and Pistaferri (2011) for a comprehensive review of modern income processes.
+Section 6 verifies our main results using a more general income process that relaxes the unit root assumption.
 We let the permanent and transitory income shocks, $\zeta_{it}$ and $v_{it}$, be non-Gaussian, with first four central moments given by
-${\operatorname*{\mathbb{E}}}(\zeta_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{\zeta_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{\zeta_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{\zeta_{t}}&\text{for }m=4\\[-2.0pt] \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}(v_{it}^{m})=\begin{cases}0&\text{for }m=1\\[-2.0pt] \sigma_{v_{t}}^{2}&\text{for }m=2\\[-2.0pt] \gamma_{v_{t}}&\text{for }m=3\\[-2.0pt] \kappa_{v_{t}}&\text{for }m=4.\\[-2.0pt] \end{cases}\vspace{-1ex}$
+[…]
 $\sigma^{2}$ describes the dispersion of the distribution about its mean; the mean is zero by construction.
 $\gamma$ denotes the skewness/asymmetry of the distribution, i.e., the relative length of the upper and lower tails.
-Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $[-\sigma,\sigma]$, and thus characterizes the thickness of the tails.
+Finally, $\kappa$ denotes the kurtosis, i.e., the tendency of the distribution to amass away from the middle $\sigma,\sigma]$, and thus characterizes the thickness of the tails.
 We allow the central moments of shocks to depend on $t$ to reflect that different stages of the lifecycle may differ in the amount and nature of income risk; Guvenen et al.
@@
 As discussed in the introduction, several recent studies document that the cross-sectional distribution of income growth exhibits left skewness and excess kurtosis, i.e., a long and fat left tail.
-Our specification of the moments of income shocks enables our income process to be consistent with these facts, allowing us to subsequently decompose its second- and higher-order moments into the relative contributions of permanent and transitory shocks.
+Our specification of the moments of income shocks enables our income process to be consistent with these facts, allowing us to subsequently decompose the various income moments into the relative contributions of permanent and transitory shocks.
 2.2 Linear consumption function
@@
 The transmission parameters vary, in principle, with $i$ (household) and $t$ (age), reflecting heterogeneity in financial wealth, among other things.
-[^9] In addition, households in Blundell et al.
+[^9] Households in Blundell et al.
 (2016) differ in their access to external insurance and insurance through family labor supply, while households in Theloudis (2021) differ in their preferences.
@@
 3 Identification
-Our specification of the income and consumption processes imposes restrictions on the second and higher-order moments of income and consumption that can be used to identify the various parameters.
-These include the variance ($\sigma_{\zeta_{t}}^{2}$, $\sigma_{v_{t}}^{2}$), skewness ($\gamma_{\zeta_{t}}$, $\gamma_{v_{t}}$), and kurtosis ($\kappa_{\zeta_{t}}$, $\kappa_{v_{t}}$) of the distributions of income shocks, [^13] With a slight abuse of terminology, we use the terms skewness and kurtosis to refer to the third and fourth central moments.
-We will be clear whenever those terms refer to standardized moments.
-the partial insurance parameters in the linear ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$) and quadratic consumption functions ($\phi^{(1)}_{t}$, $\psi^{(1)}_{t}$, $\phi^{(2)}_{t}$, $\psi^{(2)}_{t}$, $\omega^{(22)}_{t}$), and the moments of unobserved taste heterogeneity ($\sigma_{\xi_{t}}^{2}$, $\gamma_{\xi_{t}}$, $\kappa_{\xi_{t}}$).
-We first restrict the partial insurance parameters to not vary across households but we subsequently explore their variation across certain demographic groups and by the household’s position in the income distribution.
+[…]
 For identification, we assume existence of up to fourth moments of the joint distribution of income and consumption growth.
@@
 As we describe in section 4, the PSID provides data only every two years within our time period.
-This necessitates to recast the income and consumption growth processes of section 2 to this lower frequency, a straightforward operation that we carry out in appendix B.
+This necessitates to recast the income and consumption growth processes of section 2 to this lower frequency, a straightforward operation that we carry out in appendices A and B.
 Nevertheless, for consistency with the earlier literature, we provide the identifying statements subsequently as if the data come at a yearly frequency.
@@
 ${\operatorname*{\mathbb{E}}}((\Delta y_{it})^{4})$ $=\kappa_{\zeta_{t}}+\kappa_{v_{t}}+\kappa_{v_{t-1}}+6\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{v_{t-1}}^{2})+6\sigma_{v_{t}}^{2}\sigma_{v_{t-1}}^{2}.$
-The second moment depends on the variance of permanent shocks at $t$ and the sum of variances of transitory shocks at $t$ and $t-1$; it thus reflects the dispersion of one permanent and two transitory shocks.
+The second moment depends on the variance of permanent shocks at $t$ and the sum of variances of transitory shocks at $t$ and t1t-1; it thus reflects the dispersion of one permanent and two transitory shocks.
 The fourth moment reflects, similarly, the dispersion and kurtosis of both types of shocks.
 By contrast, the third moment is driven by skewness in the permanent shock.
-Barring strong non stationarities in the distribution of transitory shocks, the difference between $\gamma_{v_{t}}$ and $\gamma_{v_{t-1}}$ will be small (zero if the distribution is stationary), leaving $\gamma_{\zeta_{t}}$, the skewness in permanent shocks, to drive entirely the third moment of income.
+Barring strong non stationarities in the distribution of transitory shocks, the difference between $\gamma_{v_{t}}$ and $\gamma_{v_{t-1}}$ will be small (zero if the distribution is stationary), leaving $\gamma_{\zeta_{t}}$, the skewness in permanent shocks, to drive the third moment of income.
 The above expressions convey information on the parameters of the income process. Formally, the variance of shocks is identified as
@@
 The link between them depends on the degree of insurance with respect to these shocks, so these moments provide restrictions that help identify the parameters of partial insurance.
-BPP discuss identification in the linear function at length; we provide a summary here while we relegate some details to appendix B.2.
+BPP discuss identification in the linear function at length; we provide a summary here while we relegate the details to appendix B.2.
 Formally, the transmission parameter of permanent shocks is identified as
@@
 Those depend on higher than fourth-order moments of income, which we do not model, so they are generally not identified in the current setting.
-We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank and invertible for any distribution of income shocks except Bernoulli (for which there is no empirical evidence anyway).
+We establish in appendix B.3 that matrix $\mathbf{A}$ is full rank –and hence invertible– for any distribution of income shocks other than Bernoulli (for which there is no empirical evidence anyway).
 [^18] If the distributions of both shocks are Bernoulli, then there exist some parameter configurations for which matrix $\mathbf{A}$ is not full rank.
@@
 Similarly, BPP’s identifying condition for $\psi_{t}^{(1)}$ becomes
-$-{\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$
+${\operatorname*{\mathbb{E}}}(\Delta c_{it}\times\Delta y_{it+1})/\sigma_{v_{t}}^{2}=\psi_{t}^{(1)}+\psi_{t}^{(2)}(\gamma_{v_{t}}/\sigma_{v_{t}}^{2}),$
 also failing to identify the true pass-through of transitory shocks in the presence of nonlinear transmission, i.e., expression (7).
@@
 For most of this period, the PSID collected expenditures only on food items.
-Since $1999$, however, the PSID was drastically redesigned in that it started collecting information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016).
+Since $1999$, however, the PSID was drastically redesigned to collect information over a broader set of consumption categories, currently comprising over 70% of consumption in the National Income and Product Accounts (Blundell et al., 2016).
 We, therefore, use the PSID over the period between $1999$ and $2019$, the last available wave when we started working on this paper; the data are available only biennially.
-This is in contrast to BPP, who use PSID data over 19801980198019801992199219921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
+This is in contrast to BPP, who use PSID data over 1980198019921992 and rely on consumption imputations from the Consumer Expenditure Survey (CEX).
 Our baseline sample consists of married couples in the PSID with a male spouse aged $30$ to $65$.
@@
 Naturally, we require non-missing data on income, expenditure, and basic demographics.
-[^19] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90reduction), which likely also reflect measurement error.
+[^19] We also remove observations in the bottom 0.25% of the distributions of $\Delta y_{it}\Delta y_{it+1}$ and $\Delta c_{it}\Delta c_{it+1}$, i.e., large increases followed by large drops, likely reflecting measurement error, and observations with extreme jumps in income or consumption (jump: $\geq$10-fold increase or $\geq$90% reduction), which likely also reflect measurement error.
 Our results are not sensitive to this sample selection choice.
 Our final sample consists of 20,866 household-year observations; on average, a household is observed for $5.6$ periods, which corresponds to about $10$ calendar years given the biennial nature of the data.
-Appendix table C.1 presents summary statistics over the sample period 19991999199919992019201920192019.
+Appendix table C.1 presents summary statistics over the sample period 1999199920192019.
 About $89\%$ of men and $79\%$ of women work for pay (but some households have intermittent employment); $68\%$ of men and $71\%$ of women have had some college education.
@@
 Panel A presents the moments of income growth.
-The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant.
+The variance is $0.153$, while the first-order autocovariance is 0.049-0.049 and highly significant.
 The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data.
@@
 ${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.153 (0.005) 0.153 (0.005)
-${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.049 (0.002) -0.049 (0.002)
-${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.143 (0.098) -0.143 (0.098)
+${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.049 (0.002) -0.049 (0.002)
+${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.143 (0.098) -0.143 (0.098)
 ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.009 (0.003) 0.009 (0.003)
@@
 ${\operatorname*{\mathrm{Var}}}(\Delta c_{it})$ 0.126 (0.002) 0.323 (0.015)
-${\operatorname*{\mathrm{Cov}}}(\Delta c_{it},\Delta c_{it+1})$0.044 (0.001) -0.140 (0.009)
-${\operatorname*{\mathrm{Skew}}}(\Delta c_{it})$0.039 (0.034) -0.177 (0.083)
-${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},\Delta c_{it+1})$0.003 (0.001) 0.190 (0.036)
+${\operatorname*{\mathrm{Cov}}}(\Delta c_{it},\Delta c_{it+1})$ -0.044 (0.001) -0.140 (0.009)
+${\operatorname*{\mathrm{Skew}}}(\Delta c_{it})$ -0.039 (0.034) -0.177 (0.083)
+${\operatorname*{\mathrm{Cov}}}((\Delta c_{it})^{2},\Delta c_{it+1})$ -0.003 (0.001) 0.190 (0.036)
 ${\operatorname*{\mathrm{Kurt}}}(\Delta c_{it})$ 4.445 (0.100) 17.109 (2.111)
@@
 Block bootstrap standard errors are in parentheses.
-The third standardized moment of income growth is $-0.143$.
+The third standardized moment of income growth is 0.143-0.143.
 Despite the small size of the cross-section and the inherent difficulties in estimating higher moments, the point estimate misses statistical significance at the 10% level only by a small margin.
-${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$, which reflects (minus) the third central moment of the transitory shock, is significant at the 5% level.
+${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$, which reflects (minus) the third central moment of the transitory shock, is significant at the 1% level.
 The point estimates suggest that income growth exhibits left skewness, i.e., a long left tail.
@@
 The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively.
-Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data.
-While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al.
-(2021) also confirmed this.
+Skewness is 0.569-0.569 (and highly statistically significant) versus 1.039-1.039 in the administrative data.
+While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID.
 Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
-Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B).
+Evidently, left skewness and excess kurtosis are not only features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B).
 Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions).
@@
 The variance of consumption growth is similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error.
-The first-order autocovariance is $-0.044$ and highly significant.
+The first-order autocovariance is 0.044-0.044 and highly significant.
 In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
@@
 The magnitude of the higher-order moments is also multiple times larger in the imputed versus the true series.
-Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series.
+Skewness in imputed consumption growth is 0.177-0.177 (versus 0.039-0.039 in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series.
 The imputation thus seems to impart substantial measurement/imputation error to consumption.
@@
 Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged.
-We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively.
+We estimate the skewness (third standardized moment) at 0.947-0.947 and 1.31-1.31 for permanent and transitory shocks, respectively.
 Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail.
@@
 This is consistent with Busch and Ludwig (2024), who use the PSID until 2012 and estimate numerically similar kurtosis coefficients for both shocks.
-Seen together, large left skewness and excess kurtosis imply that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad.
-Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
+Seen together, left skewness and excess kurtosis imply that most households face either negligible shocks (close to zero) or shocks that are quite bad.
+Ignoring these stark features may lead to erroneous inference of the degree of partial insurance, to which we now turn.
 5.2 Linear consumption function
@@
 (0.028) (0.044) (0.029) (0.041) (0.026)
-$\psi^{(1)}$0.006 -0.109 -0.012 0.007 -0.003
+$\psi^{(1)}$ -0.006 -0.109 -0.012 0.007 -0.003
 (0.045) (0.078) (0.038) (0.056) (0.052)
-$\phi^{(2)}$0.040
+$\phi^{(2)}$ -0.040
 $\psi^{(2)}$ 0.015
@@
 (0.001) (0.004) (0.001) (0.001) (0.001)
-$\gamma_{\xi}$0.001 -0.001
+$\gamma_{\xi}$ -0.001 -0.001
 (0.001) (0.001)
@@
 We now turn to the estimation of the quadratic consumption function.
-The results appear in column 5 of table 3; the full set of targeted moments is shown in appendix B.4.3.373737Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 2., [^38] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case, i.e., in column 4, table 3.
+The results appear in column 5 of table 3; the full set of targeted moments is shown in appendix B.4.3.373737Estimation of the quadratic consumption function requires up to 4th-order moments of the income shocks, so we use the pre-estimated income parameters from column 3, table 2.
+5.3.1 Main results
+We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.160.16 obtained from the linear specification.
+We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero.
+As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
+We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones.
+$\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently.
+By contrast, $\psi^{(2)}$ is almost always insignificant.
+The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise.
+The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case.
+[^38] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income., [^39] The consumption-income moments targeted in the quadratic case differ from the moments targeted in the most general linear case in column 4, table 3.
 This is because in the quadratic case certain 3rd- and 4th-order consumption moments require up to 8th moments of income shocks, which we do not model.
@@
 These results are available upon request.
-5.3.1 Main results
-We estimate the transmission parameter of the average permanent shock (the coefficient on the linear part of the consumption polynomial) at $\phi^{(1)}=0.134$, close to the values 0.110.110.110.110.160.160.160.16 obtained from the linear specification.
-We estimate the transmission parameter of the average transitory shock at $\psi^{(1)}=-0.003$, statistically indistinguishable from zero.
-As $\phi^{(1)}$ and $\psi^{(1)}$ remain largely unaltered relative to the linear case, interest lies in the transmission parameters of the typical shocks away from the middle of the distribution, i.e., the coefficients on the quadratic part of the polynomial.
-We estimate a negative $\phi^{(2)}=-0.04$ for permanent shocks but a positive $\psi^{(2)}=0.015$ for transitory ones.
-$\phi^{(2)}$ is statistically insignificant at the margin for now but turns highly significant subsequently.
-By contrast, $\psi^{(2)}$ is almost always insignificant, barring the few exceptions we highlight below.
-The coefficient $\omega^{(22)}$ on the product of the two shocks is large but always imprecise.
-The variance of taste heterogeneity $\sigma_{\xi}^{2}$ and consumption measurement error $\sigma_{u_{c}}^{2}$ remain effectively unchanged from the linear case.
-[^39] Skewness and kurtosis of taste heterogeneity and consumption error are not identified because they require higher-order moments of consumption, which depend on higher than fourth-order moments of income.
 At face value, $\phi^{(2)}$ has important implications for the pass-through of permanent shocks into consumption.
@@
 Consider, for example, a large permanent shock to household income, $|\zeta_{it}|=0.5$, about three standard deviations away from the mean.
-If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.$=0.036$, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.$=0.036$).
+If the shock is negative, $\partial\Delta c_{it}/\partial\zeta_{it}=0.174$ (s.e.=0.036=0.036, estimated by the delta method); by contrast, if the shock is positive, then $\partial\Delta c_{it}/\partial\zeta_{it}=0.094$ (s.e.=0.036=0.036).
 We can generalize this to bad and good shocks of any magnitude and conclude that
@@
 Compare, for example, a very large bad permanent shock, $\zeta_{it}=-0.7$, about four standard deviations away from the mean, with a moderate one, $\zeta_{it}=-0.07$, half standard deviation from the mean.
-In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.$=0.044$) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.$=0.026$).
+In the first case, $\partial\Delta c_{it}/\partial\zeta_{it}=0.19$ (s.e.=0.044=0.044) while in the latter case $\partial\Delta c_{it}/\partial\zeta_{it}=0.139$ (s.e.=0.026=0.026).
 We can generalize this to bad shocks of any magnitude and conclude that
@@
 The reason behind this counterintuitive result is that we lack sufficient degrees of freedom to separate among all combinations of large/small, positive/negative shocks.
-However, large left skewness and excess kurtosis imply that extreme good shocks are rare.
-We should thus not extrapolate the implications of the parameter estimates for the case of extreme good shocks, as such shocks are effectively not present in our sample.
+However, large left skewness and excess kurtosis imply that large good shocks are rare.
+We should thus not extrapolate the implications of the parameter estimates for the case of large good shocks, as such shocks are effectively not present in our sample.
 The asymmetric response to bad versus good income news has recently been documented in survey questions in which subjects directly report how they would adjust consumption in different hypothetical scenarios.
@@
 If we set the permanent shock to its mean value ($\zeta_{it}=0$) and plug in the parameter estimates from table 3, we get
-$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\$
-Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant in the full sample.
+$\partial\Delta c_{it}/\partial v_{it}=-0.003+0.03\times v_{it}.\\$
+Unlike the transmission parameters of permanent shocks, neither $\psi^{(1)}$ nor $\psi^{(2)}$ are statistically significant.
 This implies that transitory shocks, be they good or bad, moderate or extreme, do not pass through into consumption.
@@
 $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age.
-It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket.
+It drops from $0.006$ to 0.008-0.008 as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket.
 As a result, large negative permanent shocks impact consumption more severely at older ages.
 Consider, again, $\zeta_{it}=-0.5$.
-At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
+At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.=0.080=0.080), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.=0.077=0.077).
 Table 4: Estimates of the consumption function, by age
@@
 $\psi^{(2)}$ 0.036 (0.061) 0.048 (0.046) -0.052 (0.073)
-$\omega^{(22)}$1.089 (1.465) -0.294 (1.195) 2.984 (0.791)
+$\omega^{(22)}$ -1.089 (1.465) -0.294 (1.195) 2.984 (0.791)
 $\sigma^{2}_{\xi}$ 0.013 (0.004) 0.022 (0.002) 0.010 (0.005)
@@
 Linear specification estimates appear in appendix table C.10.
-In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9).
+In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from 0.65-0.65 to 2.39-2.39 between ages 30–40 and 51–65 (table C.9).
 This is consistent with Guvenen et al.
@@
 However, this provides only a partial picture of the underlying dynamics because the nature of income risk is vastly different between poor and rich.
-Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of $-4.86$ (multiple times larger than the baseline, and highly significant).
+Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of 4.86-4.86 (multiple times larger than the baseline, and highly significant).
 One may thus study how large bad/good shocks from a given position in the underlying distributions, e.g., shocks in the 3rd or 97th percentiles, affect consumption in each case.
@@
 We provide details on similar simulations in appendix D.
-Adopting this approach, the lower-tail shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by $-0.035$, twice as much, among the rich (pass-through $0.084$).
+Adopting this approach, the lower-tail shock shifts $c_{it}$ by 0.017-0.017 among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by 0.035-0.035, twice as much, among the rich (pass-through $0.084$).
 By contrast, the upper-tail shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but by only $0.006$ among the rich (pass-through $0.057$).
 In other words, households in the top of the income distribution (the rich) exhibit a stronger consumption change in response to bad news, while households in the bottom (the poor) exhibit a stronger consumption change in response to good news.
-These findings agree with ABB, who simulate the transmission of shocks in the 10th and 90th percentiles.
-The reported $\Delta c_{it}$ – impulse responses with heterogeneity in their appendix figure S30 – are in the same ballpark, especially for bad shocks.
-While this is reassuring, our findings further suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
-Our approach also highlights two distinct forces at play: the underlying nature of income risk (i.e., income parameters) and the nature of the consumption response (i.e., consumption parameters).
-The larger consumption change among the rich after a lower-tail shock is not because their consumption response differs materially from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates said shocks), but mostly due to stronger downside risk that they face (the magnitude of the bad shock is simply larger in this case).
 Table 5: Estimates of the consumption function, over position in income distribution
@@
 $\psi^{(1)}$ 0.186 (0.199) 0.275 (0.200)
-$\phi^{(2)}$0.064 (0.029) -0.025 (0.101)
+$\phi^{(2)}$ -0.064 (0.029) -0.025 (0.101)
 $\psi^{(2)}$ 0.210 (0.220) -0.268 (0.197)
@@
 Linear specification estimates appear in appendix table C.11.
+These findings agree with ABB, who simulate the transmission of shocks in the 10th and 90th percentiles.
+The reported $\Delta c_{it}$ – impulse responses with heterogeneity in their appendix figure S30 – are in the same ballpark, especially for bad shocks.
+While this is reassuring, our findings further suggest that ABB’s nonlinear persistence is not crucial for the main consumption results, i.e., the asymmetric impact of different shocks or at different income levels.
+Our approach also highlights two distinct forces at play: the underlying nature of income risk (i.e., income parameters) and the nature of the consumption response (i.e., consumption parameters).
+The larger consumption change among the rich after a lower-tail shock is not because their consumption response differs materially from the poor ($\phi^{(2)}$’s smaller magnitude among the rich in fact mitigates said shocks), but mostly due to stronger downside risk that they face (the magnitude of the bad shock is simply larger in this case).
 To distinguish between these two forces, one must evaluate the marginal distribution of income risk.
@@
 Technically, in the context of a linear income process, and assuming transitory shocks away and no interaction between the income level and its shock, ABB estimate
-$c_{it}=g_{t}(\underbrace{y_{it-1}+F_{t}^{-1}(u_{it})}_{y_{it}},~{}\dots),\$
+$c_{it}=g_{t}(\underbrace{y_{it-1}+F_{t}^{-1}(u_{it})}_{y_{it}},~\dots),\\$
 where $\zeta_{it}=F_{t}^{-1}(u_{it})$ is the persistent shock and $u_{it}$ denotes the quantile of the shocks’ distribution $F_{t}$.
@@
 The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between income risk and consumption response.
-Skewness of permanent shocks is twice as large among the rich ($-4.86$) as among the old ($-2.39$).
+Skewness of permanent shocks is twice as large among the rich (4.86-4.86) as among the old (2.39-2.39).
 The consumption distribution is similar for both, so the rich must have overall more insurance to a given bad shock.
@@
 (0.051) (0.052) (0.067) (0.034)
-$\psi^{(1)}$0.251 0.015 0.006 0.004
+$\psi^{(1)}$ -0.251 0.015 0.006 0.004
 (0.131) (0.108) (0.085) (0.073)
-$\phi^{(2)}$0.094 0.005 -0.076 -0.032
+$\phi^{(2)}$ -0.094 0.005 -0.076 -0.032
 (0.062) (0.052) (0.057) (0.041)
-$\psi^{(2)}$0.162 -0.001 0.177 0.001
+$\psi^{(2)}$ -0.162 -0.001 0.177 0.001
 (0.087) (0.074) (0.089) (0.048)
@@
 The underlying income moments also vary by wealth and education (results not shown for brevity).
-Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 19991999199919992019201920192019.
+Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999199920192019.
 No versus some college is defined on the basis of the highest level of education attained by the male spouse.
@@
 5.3.4 Link with structural model and some estimates of welfare costs
-Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or income distribution.
+Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or the income distribution.
 Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
@@
 Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow.
-Therefore, such shocks end up affecting consumption more than other shocks.
 The magnitude of our pass-through rate of permanent shocks is uniformly lower than its structural counterpart (consumption transmission in the data likely reflects multiple insurance sources, beyond self-insurance and taxes examined in these models), yet other empirical studies using the updated PSID, e.g., ABB and Arellano et al.
@@
 (2024) sacrifice almost 22% of consumption to live in a world without any risk.
+6 An alternative income process: persistent shocks
+Our results so far rely on a permanent-transitory income process that leaves no room for shocks to have persistent, yet not fully permanent, effects on income.
+This unit root restriction in income may be perceived as unappealing in the context of biennial data and non-Gaussian shocks, especially if the various moments of shocks, on which the partial insurance estimates depend, are sensitive to it.
+Moreover, Kaplan and Violante (2010) show that BPP’s method of measuring insurance to permanent shocks is biased in the presence of borrowing limits, but the bias lessens if the framework is adapted to measure insurance to persistent shocks.
+This is unlikely to explain the low magnitude of $\phi^{(1)}$ that we find here (the bias tends to increase the linear pass-through rate), but it is unclear how it might affect $\phi^{(2)}$.
+For both these reasons, we relax the unit root restriction in income in this section.
+Let persistent income $P_{it}$ follow an AR(1) process rather than a unit root.
+We may write this as $P_{it}=\rho P_{it-1}+\zeta_{it}$, where $0<\rho\leq 1$ measures income persistence.
+All other assumptions (including the distributions of income shocks) remain similar to the baseline.
+This modeling choice gives rise to the persistent-transitory formulation $\Delta y_{it}=(\rho-1)P_{it-1}+\zeta_{it}+\Delta v_{it}$ and the quasi-difference
+$\Delta_{\rho}y_{it}=\zeta_{it}+\Delta_{\rho}v_{it},$ (9)
+where $\Delta_{\rho}x_{t}=x_{t}-\rho x_{t-1}$ and $\zeta$ reflects a persistent rather than a permanent shock. Clearly, the permanent-transitory process is a special case of this for $\rho=1$.
+Despite the modification in the income process, the specification of the linear and quadratic consumption functions remains unchanged from the baseline processes (4) and (5), respectively.
+However, the structure that underlies the transmission parameters $\phi^{(1)}$, $\phi^{(2)}$, etc, explicitly involves $\rho$ in this case.
+Appendix A details the derivation of the consumption function with income persistence and some further adjustments when the data are observed biennially.
+The identification strategy for the income process must now be adjusted to account for the AR(1) component and the presence of an additional parameter, $\rho$.
+Formally, the variance of shocks and $\rho$ are jointly identified through
+$\sigma_{\zeta_{t}}^{2}$ $={\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times(\rho\Delta_{\rho}y_{it-1}+\Delta_{\rho}y_{it}+\rho^{-1}\Delta_{\rho}y_{it+1})),$
+$\sigma_{v_{t}}^{2}$ $=-\rho^{-1}{\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times\Delta_{\rho}y_{it+1}),$
+${\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times\Delta y_{it-2})$ $={\operatorname*{\mathbb{E}}}((y_{it}-\rho y_{it-1})\times(y_{it-2}-y_{it-3}))=0.$
+The first two moments appear slightly modified also in the baseline; the third moment is new and pins down $\rho$.
+An additional time period of data is needed (t3t-3, for a minimum of 5 periods in total) to separately identify income persistence from the variance of persistent shocks.
+The skewness and kurtosis of shocks are then identified from similarly modified higher-order income moments that mimic the baseline.
+Identification of the consumption transmission parameters is analogous: the quasi-difference $\Delta_{\rho}y_{it}$ replaces $\Delta y_{it}$, and $\sum_{\tau=-1}^{1}\rho^{-\tau}\Delta_{\rho}y_{it+\tau}$ replaces the long sum $\sum_{\tau=-1}^{1}\Delta y_{it+\tau}$ in all previous identifying statements, but otherwise the strategy is similar to the baseline.
+Appendix B shows the identifying statements in detail, including the adjustments needed for biennial data and measurement error; identification in the quadratic case again obtains for any distribution of income shocks except Bernoulli.
+Table 7: Estimates of the consumption function, income persistence
+Consumption fn.: Quadratic
+Income moments: 2nd, 3rd, 4th
+Income persistence: $\rho=0.931$ (s.e.=0.028=0.028)
+Consumption data: PSID
+$\phi^{(1)}$ 0.112 (0.024)
+$\psi^{(1)}$ -0.010 (0.066)
+$\phi^{(2)}$ -0.049 (0.026)
+$\psi^{(2)}$ 0.019 (0.037)
+$\omega^{(22)}$ 0.430 (0.847)
+$\sigma^{2}_{\xi}$ 0.019 (0.001)
+$\sigma^{2}_{u_{c}}$ 0.044 (0.002)
+Notes: The table presents the estimates of the parameters of the quadratic consumption function, when income features linear persistence.
+Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
+The estimates of the income process appear in appendix table C.13; linear consumption specification estimates appear in appendix table C.14.
+Appendix table C.13 reports the estimates of the income process from this specification.
+Three points emerge.
+First, when targeting only the second moments of income, we estimate $\rho=0.846$ (s.e.=0.038=0.038), which lies at the lower end of persistence estimates in the literature but within ABB’s nonlinear persistence range of 0.5–1 over a similar period.
+The biennial nature of the data likely influences this, even though $\rho$ reflects annual persistence.
+As $\rho$ is below its baseline value, the variance of persistent shocks is higher than the baseline to help match the variance of income growth in the data (yet the transitory variance is smaller).
+Second, adding third-order moments yields negative skewness for both shocks but leaves the estimates of $\rho$ and the variances unchanged.
+Third, targeting all moments up to fourth-order, we estimate $\rho=0.931$ (s.e.=0.028=0.028).
+The estimates of the variances align with the baseline values of table 2, as also do the skewness coefficients at -0.91 and -1.178 for the two types of shocks, respectively.
+We estimate kurtosis at 28.891 and 52.091; for persistent shocks, this is lower than the baseline, yet both distributions remain highly leptokurtic and left-skewed despite the AR(1) component.
+Table 7 reports the estimates of the quadratic consumption function (see table C.14 for the linear specification).
+We estimate the transmission parameter of the average persistent shock at $\phi^{(1)}=0.112$, somewhat lower than the value 0.134 when shocks are permanent.
+The average persistent shock is therefore slightly better insured than the average permanent shock.
+Kaplan and Violante (2010) similarly observe a drop in the pass-through rate (an increase in partial insurance) when $\rho$ falls and shocks become less permanent.
+We further estimate the transmission parameter of persistent shocks away from the middle of the distribution at $\phi^{(2)}=-0.049$, negative and larger in magnitude than the baseline value of -0.04.
+Negative persistent shocks remain more unsettling than positive ones: they are associated with a larger pass-through rate into consumption, one that increases with the severity of the shock.
+The asymmetric response to bad versus good income news is therefore qualitatively and numerically robust to the earlier unit root assumption.
+Finally, transitory shocks have economically small and statistically zero pass-through rates.
+[^47] In the context of the linear consumption specification (appendix table table C.14), we estimate $\phi^{(1)}=0.089$ when we target only second moments, which gradually increases to match its quadratic specification counterpart as we introduce higher-order moments.
+This is driven by the increase in the value of $\rho$ as we add higher moments.
+Kaplan and Violante (2010) also find lower pass-through rates (more insurance) at lower values of $\rho$.
+Transitory shocks remain fully insured, statistically., [^48] It is straightforward to further condition the results on age, position in the income distribution, wealth and education.
+We do not report these results for brevity.
 7 Conclusion
@@
 Appendix A Derivation of the consumption function
-This appendix shows how to derive the consumption function under a permanent-transitory specification for log income.
+This appendix shows how to derive the consumption function under the general AR(1) specification for log income in section 6; the consumption function based on the baseline permanent-transitory income process is a special case of this.
 We follow similar derivations in Blundell et al.
-(2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein.
-We show below how to derive the most general (quadratic) consumption function (5); the linear function (4) is a special case of this.
+(2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein and the relaxation of the unit root assumption for income.
+We show below how to derive the most general (quadratic) consumption function; the linear function is a special case of this.
 We let a time unit in the model correspond to one calendar year.
@@
 (2) in a generic period $t+s$ are
-[…]
-[…]
+$[C_{it+s}]$ $:{}$ $\quad\widetilde{U}_{C}(\widetilde{C}_{it+s})\exp(-\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})$ $=\lambda_{it+s}$
+$[A_{it+s+1}]$ $:{}$ $\quad\beta(1+r){\operatorname*{\mathbb{E}}}_{t+s}\lambda_{it+s+1}$ $=\lambda_{it+s},$
 where $\widetilde{U}_{C}$ is the marginal utility of consumption and $\widetilde{C}_{it}=C_{it}\exp(-\mathbf{Z}_{it}^{\prime}\boldsymbol{\alpha})$.
@@
 ${\operatorname*{\mathbb{E}}}_{t+s}\Delta\ln\lambda_{it+s+1}\approx\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t+s}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}\Rightarrow\Delta\ln\lambda_{it+s+1}\approx\omega_{it+s+1}+\epsilon_{it+s+1}.$
-The first term $\omega_{it+s+1}=\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}$ captures the impact of interest rates, impatience, and precautionary motives on the growth of the marginal utility of wealth.
+The first term $\omega_{it+s+1}=\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}$ captures the impact of the interest rate, impatience, and precautionary motives on the growth of the marginal utility of wealth.
 To maintain tractability, we assume that $\omega$ is non-stochastic but possibly heterogeneous across households.
@@
 where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$.
-We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to interest rates and impatience (discounting), the second term reflects unobserved consumption taste heterogeneity.
+We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to impatience (discounting) and the interest rate, the second term reflects unobserved consumption taste heterogeneity.
 We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path.
@@
 A.2 Taylor approximation to lifetime budget constraint
-Let $G(\boldsymbol{\xi})=\ln\sum_{s=0}^{N}\exp\xi_{s}$ for $\boldsymbol{\xi}=(\xi_{0},\xi_{1},\dots,\xi_{N})^{\prime}$.
-Applying a second-order Taylor expansion of $G(\boldsymbol{\xi})$ around a deterministic $\boldsymbol{\xi}^{0}$, and taking expectations given information ${\cal I}$, yields
+[…]
 $\begin{split}{\operatorname*{\mathbb{E}}}_{\cal I}G(\boldsymbol{\xi})\approx G(\boldsymbol{\xi}^{0})&+\sum_{s=0}^{N}\frac{\exp\xi^{0}_{s}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}{\operatorname*{\mathbb{E}}}_{\cal I}(\xi_{s}-\xi^{0}_{s})\\ &+\frac{1}{2}\sum_{s=0}^{N}\sum_{\ell=0}^{N}\frac{\exp\xi^{0}_{s}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}\left(\delta_{s\ell}-\frac{\exp\xi^{0}_{\ell}}{\sum_{\kappa=0}^{N}\exp\xi^{0}_{\kappa}}\right){\operatorname*{\mathbb{E}}}_{\cal I}(\xi_{s}-\xi^{0}_{s})(\xi_{\ell}-\xi^{0}_{\ell}),\end{split}$ (A.2)
@@
 Letting
-$\xi_{s}$ =$=\begin{cases}\ln A_{it+s}&\text{ for }s=0\\ \ln Y_{it+s-1}-(s-1)\ln(1+r)\qquad\qquad&\text{ for }s=1,\dots,T-t+1\\ \end{cases}$
-$\xi^{0}_{s}$ =$=\begin{cases}{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it+s}&\text{ for }s=0\\ {\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s-1}-(s-1)\ln(1+r)&\text{ for }s=1,\dots,T-t+1,\end{cases}$
+[…]
+$\xi^{0}_{s}$ $=\begin{cases}{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it+s}&\text{ for }s=0\\ {\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s-1}-(s-1)\ln(1+r)&\text{ for }s=1,\dots,T-t+1,\end{cases}$
 and declaring the information set to contain time $t$ information, we follow (A.2) to write
@@
 $+\frac{1}{2}\pi_{it}(1-\pi_{it}){\operatorname*{\mathbb{E}}}_{t}(\ln A_{it}-{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})^{2}$
-$+\frac{1}{2}\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
+$+\frac{1}{2}(1-\pi_{it})\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
 $-\pi_{it}(1-\pi_{it})\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln A_{it}-{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s}).$
 The notation is: $\pi_{it}=\frac{Q_{1t}}{Q_{1t}+Q_{2t}}$ is equal to the expected share of financial wealth $Q_{1t}=\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})$ in the household’s total financial and human wealth at $t$, the latter defined as $Q_{2t}=\sum_{\kappa=0}^{T-t}\exp\left({\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\kappa}-\kappa\ln(1+r)\right)$, i.e., the household’s expected lifetime income for the remaining of life; $\vartheta_{it+s}^{Y}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln Y_{it+s}-s\ln(1+r)\right)/Q_{2t}$ is an annuitization factor equal to the expected share of time $t+s$ household income in total lifetime income.
-$\pi_{it}$ and $\vartheta_{it+s}^{Y}$ pertain to expectations at $t-2$ so they are both known at $t-2$ and later.
-Using the income process and assuming the deterministic observables away (they cancel out in a first difference in expectations at a later stage), we reach $\ln Y_{it+s}=\ln Y_{it-2}+\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}-v_{it-2}$, therefore $\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it-2}=\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}$.
-It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})=\zeta_{it}+\zeta_{it-1}+\vartheta_{it}^{Y}v_{it}$ because ${\operatorname*{\mathbb{E}}}_{t}\zeta_{it+\kappa}={\operatorname*{\mathbb{E}}}_{t}v_{it+\kappa}=0$ for $\kappa>0$, and $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$ by construction.
+$\pi_{it}$ and $\vartheta_{it+s}^{Y}$ pertain to expectations at t2t-2 so they are both known at t2t-2 and later.
+Using the income process and assuming the deterministic observables away (they cancel out in a first difference in expectations at a later stage), we reach $\ln Y_{it+s}=\rho^{s+2}\ln Y_{it-2}+\sum_{\tau=-1}^{s}\rho^{s-\tau}\zeta_{it+\tau}+v_{it+s}-\rho^{s+2}v_{it-2}$, therefore $\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s}=\sum_{\tau=-1}^{s}\rho^{s-\tau}\zeta_{it+\tau}+v_{it+s}$, given that $s\geq-1$.
+It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})=\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})+\vartheta_{it}^{Y}v_{it}$, where $\Theta_{it}^{Y}=\sum_{\kappa=0}^{T-t}\vartheta_{it+\kappa}^{Y}\rho^{\kappa}$, because ${\operatorname*{\mathbb{E}}}_{t}\zeta_{it+\kappa}={\operatorname*{\mathbb{E}}}_{t}v_{it+\kappa}=0$ for $\kappa>0$.
 The second-order term in income is more involved.
 This is given by
-$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
-$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}\left((\zeta_{it}+\zeta_{it-1})^{2}+(\zeta_{it}+\zeta_{it-1})(v_{it+s}+v_{it+\ell})+v_{it+s}v_{it+\ell}\right)$
-$=(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+2\pi_{it}(1-\pi_{it})\vartheta_{it}^{Y}(\zeta_{it}+\zeta_{it-1})v_{it}+(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}.$
-We derive this making use of $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$, $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}\vartheta_{it+\ell}^{Y}=(\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{Y})=1$, and the expectation of future shocks being zero by construction.
-Plugging these results in the approximation of the previous page, noting that assets at $t$ are ‘beginning-of-period’ and thus independent of the time $t$ income shocks, and taking a first difference in expectations between $t-2$ and $t$ yields
-${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)-{\operatorname*{\mathbb{E}}}_{t-2}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)$ $\approx(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}$
-$+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}$
+$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
+$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}\left(\rho^{s}\rho^{\ell}(\zeta_{it}+\rho\zeta_{it-1})^{2}+(\zeta_{it}+\rho\zeta_{it-1})(\rho^{s}v_{it+\ell}+\rho^{\ell}v_{it+s})+v_{it+s}v_{it+\ell}\right)$
+$=\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}+2\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}+\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2},$
+where $\widetilde{\pi}_{it}=\widetilde{\Theta}_{it}^{Y}-(1-\pi_{it})(\Theta_{it}^{Y})^{2}$, $\widetilde{\Theta}_{it}^{Y}=\sum_{\kappa=0}^{T-t}\vartheta_{it+\kappa}^{Y}\rho^{2\kappa}$, and $\widetilde{\widetilde{\pi}}_{it}=1-(1-\pi_{it})\Theta_{it}^{Y}$.
+We derive this noting that the expectation of future shocks is zero by construction.
+Plugging these results in the approximation of the previous page, noting that assets at $t$ are ‘beginning-of-period’ and thus independent of the time $t$ income shocks, and taking a first difference in expectations between t2t-2 and $t$ yields
+${\operatorname*{\mathbb{E}}}_{t}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)-{\operatorname*{\mathbb{E}}}_{t-2}\ln\left(A_{it}+\sum_{s=0}^{T-t}\frac{Y_{it+s}}{(1+r)^{s}}\right)$ $\approx(1-\pi_{it})\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})$
+$+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}$
+$+\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}$
 $+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}$
-$+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.$
-The linear terms in the first line are the first-order approximation of BPP (adapted for biennial data); the quadratic terms in the next lines are the refinement from the second-order approximation.
+$+(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}.$
+The linear terms in the first two lines are the first-order approximation of BPP (adapted here for biennial data and income persistence); the quadratic terms in the next lines are the refinement from the second-order approximation.
 Assuming expectations away, the logarithm of the right hand side of the budget constraint (2) is $\ln\sum_{s=0}^{T-t}\exp(\ln C_{it+s}-s\ln(1+r))$.
@@
 where $\vartheta_{it+s}^{C}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+s}-s\ln(1+r)\right)/\sum_{\kappa=0}^{T-t}\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln C_{it+\kappa}-\kappa\ln(1+r)\right)$ is an annuitization factor equal to the expected share of time $t+s$ consumption in total lifetime consumption.
-$\vartheta_{it+s}^{C}$ pertains to expectations at $t-2$, so it is known at $t-2$ and later.
+$\vartheta_{it+s}^{C}$ pertains to expectations at t2t-2, so it is known at t2t-2 and later.
 Using the linearized consumption function in (A.1) to replace $\ln C_{it+s}$ recursively, we reach $\ln C_{it+s}=\ln C_{it-2}+\sum_{\tau=-1}^{s}\Psi_{it+\tau}+\sum_{\tau=-1}^{s}\xi_{it+\tau}+\sum_{\tau=-1}^{s}\theta_{it+\tau-1}\epsilon_{it+\tau}$, where $\Psi_{it+\tau}=\theta_{it+\tau-1}\Delta(\mathbf{Z}_{it+\tau}^{\prime}\boldsymbol{\alpha})+\Xi_{it+\tau}$, therefore $\ln C_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln C_{it+s}=\sum_{\tau=-1}^{s}\theta_{it+\tau-1}\epsilon_{it+\tau}$ because $\Psi_{it}$ and $\xi_{it}$ are non-stochastic by assumption and they cancel out in the first difference.
@@
 We derive this by noting that future expectations errors are zero by assumption, and that $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{C}(\delta_{s\ell}-\vartheta_{it+\ell}^{C})=\sum_{s=0}^{T-t}\vartheta_{it+s}^{C}\delta_{ss}+\sum_{s=0}^{T-t}\sum_{\ell\neq s}\vartheta_{it+s}^{C}\delta_{s\ell}-(\sum_{s=0}^{T-t}\vartheta_{it+s}^{C})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{C})=0$.
-Plugging these results in the approximation above and taking a first difference in expectations between $t-2$ and $t$ yields
+Plugging these results in the approximation above and taking a first difference in expectations between t2t-2 and $t$ yields
 ${\operatorname*{\mathbb{E}}}_{t}\ln\sum_{s=0}^{T-t}\frac{C_{it+s}}{(1+r)^{s}}-{\operatorname*{\mathbb{E}}}_{t-2}\ln\sum_{s=0}^{T-t}\frac{C_{it+s}}{(1+r)^{s}}\approx\theta_{it-1}\epsilon_{it}+\theta_{it-2}\epsilon_{it-1}.$
@@
 Given the biennial nature of our data, it follows from (A.1) that observed consumption growth is $\Delta^{2}c_{it+s}=\xi_{it+s}+\xi_{it+s-1}+\theta_{it+s-1}\epsilon_{it+s}+\theta_{it+s-2}\epsilon_{it+s-1}$.
-The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a difference between $t$ and $t-2$.
-Dropping the index $s$ and replacing the expectations errors with the quadratic expression in income shocks (the budget must balance so we bring the two sides of the budget constraint together), we obtain the final expression for consumption growth with biennial survey data, given by
-$\begin{split}\Delta^{2}c_{it}&\approx\xi_{it}+\xi_{it-1}+(1-\pi_{it})(\zeta_{it}+\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}\\ &+\frac{1}{2}(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}\\ &+(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}(\zeta_{it}+\zeta_{it-1})v_{it}.\end{split}$ (A.3)
-The first line of (A.3) corresponds to the linear consumption function (4), adapted here for the biennial nature of the data and given compactly by
-$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}.$
-The transmission parameters of income shocks are given by $\phi_{it}^{(1)}=1-\pi_{it}$ and $\psi_{it}^{(1)}=(1-\pi_{it})\vartheta_{it}^{Y}$, where $\pi_{it}$ and $\vartheta_{it}^{Y}$ were defined previously.
+The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a difference between $t$ and t2t-2.
+Dropping the index $s$ and replacing the expectations errors with the quadratic expression in income shocks (the budget must balance so we bring the two sides of the budget constraint together), we obtain the final expression for consumption growth with biennial survey data and income persistence, given by
+$\begin{split}\Delta^{2}c_{it}&\approx\xi_{it}+\xi_{it-1}+(1-\pi_{it})\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})+(1-\pi_{it})\vartheta_{it}^{Y}v_{it}\\ &+\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}+\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}\\ &+(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}.\end{split}$ (A.3)
+The first line of (A.3) corresponds to the linear consumption function, adapted here for income persistence and the biennial nature of the data, and given compactly by
+$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}.$
+The transmission parameters of income shocks are given by $\phi_{it}^{(1)}=(1-\pi_{it})\Theta_{it}^{Y}$ and $\psi_{it}^{(1)}=(1-\pi_{it})\vartheta_{it}^{Y}$.
+In the baseline, the permanent-transitory income process has $\rho=1$ and further restricts $\phi_{it}^{(1)}=1-\pi_{it}$ because $\Theta_{it}^{Y}=\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$.
 The quadratic terms in the next two lines of (A.3) reflect the refinement from the higher-order approximation.
-This higher-order approximation leads to the quadratic consumption function (5), adapted here for the biennial nature of the data and given compactly by
-$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}+\phi_{it}^{(2)}\left(\zeta_{it}+\zeta_{it-1}\right)^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\left(\zeta_{it}+\zeta_{it-1}\right)v_{it}.$
-The additional transmission parameters of income shocks are given by $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\pi_{it}$, $\psi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})$, and $\omega_{it}^{(22)}=(1-\pi_{it})\vartheta_{it}^{Y}\pi_{it}$.
+This leads to the quadratic consumption function, adapted here for income persistence and the biennial nature of the data, and given compactly by
+$\Delta^{2}c_{it}=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)+\psi_{it}^{(1)}v_{it}\\ +\phi_{it}^{(2)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}\left(\zeta_{it}+\rho\zeta_{it-1}\right)v_{it}.$
+The additional transmission parameters of income shocks are given by $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\widetilde{\pi}_{it}$, $\psi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})$, and $\omega_{it}^{(22)}=(1-\pi_{it})\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}$.
+The baseline permanent-transitory income process has $\rho=1$ and further restricts $\phi_{it}^{(2)}=\frac{1}{2}(1-\pi_{it})\pi_{it}$ and $\omega_{it}^{(22)}=(1-\pi_{it})\pi_{it}\vartheta_{it}^{Y}$, because $\widetilde{\Theta}_{it}^{Y}=\Theta_{it}^{Y}=\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$.
 Appendix B Identification details
-This appendix provides detailed identification statements based on our use of the recent biennial PSID data. There are two points of departure from section 3 in the paper.
-First, we recast the income and consumption processes to reflect the biennial frequency of the modern PSID.
-Unexplained income growth is thus given by $\Delta^{2}y_{it}=\Delta y_{it}+\Delta y_{it-1}=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}$, with $\Delta y_{it}$ given by (3).
-The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a first difference in variable $x$ between $t$ and $t-2$, i.e., the observed frequency.
-Second, we allow for classical measurement error. We let observed income $y_{it}^{*}$ and consumption $c_{it}^{*}$ be the sum of their corresponding true value and measurement error, namely
+This appendix provides detailed identification statements based on our use of the recent biennial PSID data. There are three points of departure from section 3 in the paper.
+First, we use the most general AR(1) specification for income, shown in section 6.
+The baseline permanent-transitory specification is a special case with $\rho=1$.
+Second, we allow for classical measurement error.
+We let observed income $y_{it}^{*}$ and consumption $c_{it}^{*}$ be the sum of their corresponding true value and measurement error, namely
 $y_{it}^{*}=y_{it}+u_{it}^{y}\qquad\text{and}\qquad c_{it}^{*}=c_{it}+u_{it}^{c}.$
 The moments of income measurement error are not separately identifiable from the moments of the transitory shock.
-This is why we have assumed the income error to be Gaussian (but we let the consumption error unrestricted), so we specify
-${\operatorname*{\mathbb{E}}}((u^{y}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{y}_{t}}^{2}&\text{for }m=2\\ 0&\text{for }m=3\\ 3(\sigma_{u^{y}_{t}}^{2})^{2}&\text{for }m=4\\ \end{cases}\qquad\text{and}\qquad{\operatorname*{\mathbb{E}}}((u^{c}_{it})^{m})=\begin{cases}0&\text{for }m=1\\ \sigma_{u^{c}_{t}}^{2}&\text{for }m=2\\ \gamma_{u^{c}_{t}}&\text{for }m=3\\ \kappa_{u^{c}_{t}}&\text{for }m=4.\\ \end{cases}$
-We retrieve the variance of income error from the validation study of the PSID in Bound et al.
+This is why we assume the income error to be Gaussian (but let the consumption error unrestricted), so we specify
+[…]
+We retrieve the variance of income measurement error from the validation study of the PSID in Bound et al.
 (1994), as we explain in section 4.
-We do not need similar assumptions for the consumption error, whose moments we identify below.
-We further assume that income and consumption errors are mutually independent (but note discussion of consumption imputation from the CEX), independent over time, and independent of the income shocks and taste heterogeneity.
-These choices lead to our processes for observed income growth, consumption growth in the linear specification, and consumption growth in the quadratic specification, given by
-$\Delta^{2}y_{it}$ $=\zeta_{it}+\zeta_{it-1}+\Delta^{2}v_{it}+\Delta^{2}u^{y}_{it},$
-$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}+\Delta^{2}u^{c}_{it},$
-$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}+\phi_{it}^{(1)}(\zeta_{it}+\zeta_{it-1})+\psi_{it}^{(1)}v_{it}$
-$+\phi_{it}^{(2)}(\zeta_{it}+\zeta_{it-1})^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}(\zeta_{it}+\zeta_{it-1})v_{it}+\Delta^{2}u^{c}_{it},$
-respectively. These are the theoretical expressions we bring to the data.
+We do not need similar assumptions for the consumption measurement error, whose moments we identify below.
+We further assume that income and consumption errors are mutually independent (but note the discussion of consumption imputation from the CEX in appendix C.4), independent over time, and independent of the income shocks and taste heterogeneity.
+Third, we recast the income and consumption processes to reflect the biennial frequency of the modern PSID.
+These choices lead to the theoretical expressions that we bring to the data, namely the process for income growth, the quasi-difference in income, consumption growth in the linear specification, and consumption growth in the quadratic specification, given respectively by
+$\Delta^{2}y_{it}$ $=(\rho^{2}-1)P_{it-2}$~ $+$ $~\zeta_{it}+\rho\zeta_{it-1}+\Delta^{2}v_{it}+\Delta^{2}u^{y}_{it},$
+$\Delta^{2}_{\rho}y_{it}$ $=$ $~\zeta_{it}+\rho\zeta_{it-1}+\Delta^{2}_{\rho}v_{it}+\Delta^{2}_{\rho}u^{y}_{it},$
+$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}$ $+$ $~\phi_{it}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})+\psi_{it}^{(1)}v_{it}+\Delta^{2}u^{c}_{it},$
+$\Delta^{2}c_{it}$ $=\xi_{it}+\xi_{it-1}$ $+$ $~\phi_{it}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})+\psi_{it}^{(1)}v_{it}$
+$+$ $~\phi_{it}^{(2)}(\zeta_{it}+\rho\zeta_{it-1})^{2}+\psi_{it}^{(2)}v_{it}^{2}+\omega_{it}^{(22)}(\zeta_{it}+\rho\zeta_{it-1})v_{it}+\Delta^{2}u^{c}_{it}.$
+The notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates a first difference in $x$ between $t$ and t2t-2, i.e., the observed frequency.
+The quasi-difference is defined as $\Delta^{2}_{\rho}x_{t}=\Delta_{\rho}x_{t}+\rho\Delta_{\rho}x_{t-1}=x_{t}-\rho^{2}x_{t-2}$, where $\Delta_{\rho}x_{t}=x_{t}-\rho x_{t-1}$ is the quasi-difference between $t$ and t1t-1.
 Due to the biennial nature of the data, we cannot separately identify the moments of the yearly permanent shocks within a two-year period.
@@
 Similarly, we assume that $\sigma_{\xi_{t}}^{2}\approx\sigma_{\xi_{t-1}}^{2}$.
-For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, and $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$.
-These last restrictions are not needed for identification with biennial data, but they simplify the illustration.
+For simplicity of the next statements, we also assume $\sigma_{\zeta_{t}}^{2}\approx\sigma_{\zeta_{t+1}}^{2}\approx\sigma_{\zeta_{t+2}}^{2}$, $\sigma_{v_{t}}^{2}\approx\sigma_{v_{t-2}}^{2}\approx\sigma_{v_{t+2}}^{2}$, $\sigma_{u^{y}_{t}}^{2}\approx\sigma_{u^{y}_{t-2}}^{2}\approx\sigma_{u^{y}_{t+2}}^{2}$, and $\kappa_{v_{t}}\approx\kappa_{v_{t-2}}$.
+These last restrictions are not needed for identification, but they simplify the illustration.
 B.1 Income process parameters
 The second and third moments of shocks are given by
-$\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
-$\gamma_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})\qquad$ $\text{and}\qquad\gamma_{v_{t}}$ $=-{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it+2}).$
+$\sigma_{\zeta_{t}}^{2}$ $=\frac{1}{1+\rho^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa})\qquad$ $\text{and}\qquad\sigma_{v_{t}}^{2}$ $=-\frac{1}{\rho^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2})-\sigma_{u^{y}_{t}}^{2},$
+$\gamma_{\zeta_{t}}$ $=\frac{1}{1+\rho^{3}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa})\qquad$ $\text{and}\qquad\gamma_{v_{t}}$ $=-\frac{1}{\rho^{2}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it+2}).$
 The fourth moments are more involved and are given by
-$\kappa_{\zeta_{t}}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})-\kappa_{v_{t}}-3(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
-$-12\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-12\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$
-$\kappa_{v_{t}}$ $={\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it+2})^{2})-4(\sigma_{\zeta_{t}}^{2})^{2}-3(\sigma_{v_{t}}^{2})^{2}-6(\sigma_{u^{y}_{t}}^{2})^{2}$
-$-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-8\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-12\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}.$
+$\kappa_{\zeta_{t}}$ $=\frac{1}{1+\rho^{4}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})-\frac{1+\rho^{8}}{1+\rho^{4}}\kappa_{v_{t}}-6\frac{\rho^{2}}{1+\rho^{4}}(\sigma_{\zeta_{t}}^{2})^{2}-6\frac{\rho^{4}}{1+\rho^{4}}(\sigma_{v_{t}}^{2})^{2}-3(1+\rho^{4})(\sigma_{u^{y}_{t}}^{2})^{2}$
+$-6(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-6(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-6(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$
+$\kappa_{v_{t}}$ $=\frac{1}{\rho^{4}}{\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it+2})^{2})-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(\sigma_{\zeta_{t}}^{2})^{2}-\frac{1+\rho^{4}+\rho^{8}}{\rho^{4}}(\sigma_{v_{t}}^{2})^{2}-\frac{1+4\rho^{4}+\rho^{8}}{\rho^{4}}(\sigma_{u^{y}_{t}}^{2})^{2}$
+$-2(1+\rho^{2})\frac{1+\rho^{4}}{\rho^{4}}\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}-2(1+\rho^{2})\frac{1+\rho^{4}}{\rho^{4}}\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}-\left(2\frac{(1+\rho^{4})^{2}}{\rho^{4}}+4\right)\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}.$
+In all cases, either we need an external estimate for $\rho$ (e.g., $\rho=1$ as in the permanent-transitory process) or treat $\rho$ as a free parameter to be jointly estimated with the moments of income shocks.
+In the latter case, $\rho$ is pinned down by
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})={\operatorname*{\mathbb{E}}}((y_{it}-\rho^{2}y_{it-2})\times(y_{it-4}-y_{it-6}))=0,$
+namely a “long” intertemporal covariance.
+Note that when $\rho$ is a free parameter, five time period of data are needed at a minimum (i.e., t6t-6 through $t+2$ biennially) to identify both the variance of persistent shocks and $\rho$ itself, versus four periods when $\rho$ is set externally.
 B.2 Linear consumption function parameters
-Given the income process, the linear consumption function, and the properties of shocks, the transmission parameters are given by
-$\phi^{(1)}_{t}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}y_{it+\kappa})/\sigma_{\zeta_{t}}^{2}\qquad\text{and}\qquad\psi^{(1)}_{t}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})/\sigma_{v_{t}}^{2}.$
+Given the general income process, the linear consumption function, and the properties of shocks, the transmission parameters are given by
+$\phi^{(1)}_{t}$ $=\frac{1}{(1+\rho^{2})\sigma_{\zeta_{t}}^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\rho^{-\kappa}\Delta^{2}_{\rho}y_{it+\kappa}),$
+$\psi^{(1)}_{t}$ $=-\frac{1}{\rho^{2}\sigma_{v_{t}}^{2}}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2}).$
 The variance of taste heterogeneity and consumption measurement error are identified as
-$\sigma_{\xi_{t}}^{2}=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}c_{it+\kappa})-(\phi^{(1)}_{t})^{2}\sigma_{\zeta_{t}}^{2}-\frac{1}{2}(\psi^{(1)}_{t})^{2}\sigma_{v_{t}}^{2}\quad\text{and}\quad\sigma_{u^{c}_{t}}^{2}=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2}).$
+$\sigma_{\xi_{t}}^{2}$ $=\frac{1}{2}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\sum_{\mathclap{\kappa=\{-2,0,2\}}}\Delta^{2}c_{it+\kappa})-(\phi^{(1)}_{t})^{2}\frac{1+\rho^{2}}{2}\sigma_{\zeta_{t}}^{2}-\frac{1}{2}(\psi^{(1)}_{t})^{2}\sigma_{v_{t}}^{2},$
+$\sigma_{u^{c}_{t}}^{2}$ $=-{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it+2}).$
 The third and fourth moments of taste heterogeneity and consumption error are identified by higher-order consumption moments through expressions analogous to those above.
 B.3 Quadratic consumption function parameters
-Given the income process, the quadratic consumption function, and the properties of income shocks, the transmission parameters are given by
-$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it}\times\Delta^{2}y_{it+2})\\ \end{array}\right),$ (B.1)
-where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks.
-[^1] With biennial data, matrix $\mathbf{A}$ is different from $\mathbf{A}$ with annual data in the main paper and is given by $\mathbf{A}$ $=\left(\begin{array}[]{lllll}2\sigma_{\zeta_{t}}^{2}&\sigma_{v_{t}}^{2}&2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&0\\ 2\gamma_{\zeta_{t}}&\gamma_{v_{t}}&2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2}))&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ 0&-\sigma_{v_{t}}^{2}&0&-\gamma_{v_{t}}&0\\ 0&\gamma_{v_{t}}&4\sigma_{\zeta_{t}}^{2}(\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&\kappa_{v_{t}}+\sigma_{v_{t}}^{2}(2\sigma_{\zeta_{t}}^{2}+\sigma_{v_{t}}^{2}+2\sigma_{u^{y}_{t}}^{2})&0\\ 0&-\gamma_{v_{t}}&-2\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})&-\kappa_{v_{t}}-\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}&-2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\\ \end{array}\right).$ The determinant of this matrix is given by
-$\text{det}(\mathbf{A})$ $=$ $-8\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\Big{[}\Big{(}k_{\zeta_{t}}$ $-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
-$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{]}.$
-Assuming that the shocks are not drawn from a Bernoulli distribution, kurtosis is bounded below by $\gamma^{2}/\sigma^{2}+\sigma^{4}$. It follows that
-$k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}-\sigma_{\zeta_{t}}^{4}$ $>0,$ (B.2)
-$k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}-\sigma_{v_{t}}^{4}$ $>0.$ (B.3)
+Given the general income process, the quadratic consumption function, and the properties of shocks, the transmission parameters are given by
+$\left(\begin{array}[]{l}\phi^{(1)}_{t}\\ \psi^{(1)}_{t}\\ \phi^{(2)}_{t}\\ \psi^{(2)}_{t}\\ \omega^{(22)}_{t}\end{array}\right)=\mathbf{A}^{-1}\left(\begin{array}[]{l}{\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it+2})^{2})\\ {\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2})\\ \end{array}\right),$ (B.1)
+where the matrix of coefficients $\mathbf{A}$ depends exclusively on the second, third, and fourth moments of income shocks, the AR(1) persistence parameter $\rho$, and the variance of income measurement error.
+[^1] While the paper shows matrix $\mathbf{A}$ for the case of annual data, unit root in income, and no measurement error, the general form of $\mathbf{A}=[a_{ij}]_{5\times 5}$ with biennial data, income persistence, and measurement error has $a_{15}=a_{31}=a_{33}=a_{35}=a_{41}=a_{45}=a_{51}=0$, and $a_{11}$ $=(1+\rho^{2})\sigma_{\zeta_{t}}^{2},$ $a_{12}$ $=\sigma_{v_{t}}^{2},$ $a_{13}$ $=(1+\rho^{3})\gamma_{\zeta_{t}},$ $a_{14}$ $=\gamma_{v_{t}},$ $a_{21}$ $=(1+\rho^{3})\gamma_{\zeta_{t}},$ $a_{22}$ $=\gamma_{v_{t}},$ $a_{23}$ $=(1+\rho^{4})\kappa_{\zeta_{t}}+6\rho^{2}(\sigma_{\zeta_{t}}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{24}$ $=\kappa_{v_{t}}+\rho^{4}(\sigma_{v_{t}}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$ $a_{25}$ $=2(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2},$ $a_{32}$ $=-\rho^{2}\sigma_{v_{t}}^{2},$ $a_{34}$ $=-\rho^{2}\gamma_{v_{t}},$ $a_{42}$ $=\rho^{4}\gamma_{v_{t}},$ $a_{43}$ $=(1+\rho^{2})^{2}(\sigma_{\zeta_{t}}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{44}$ $=\rho^{4}\kappa_{v_{t}}+(\sigma_{v_{t}}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2},$ $a_{52}$ $=-\rho^{2}\gamma_{v_{t}},$ $a_{53}$ $=-\rho^{2}(1+\rho^{2})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2}),$ $a_{54}$ $=-\rho^{2}(\kappa_{v_{t}}+\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}),$ $a_{55}$ $=-\rho^{2}(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}.$ Its determinant [^2] The simpler derivation for $\rho=1$ appears in older versions of the paper, see arxiv.org/abs/2306.13208.
+is given by
+$\text{det}(\mathbf{A})$ $=-\rho^{4}(1+\rho^{2})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\times$
+$\bigg\{$ $\bigg((1+\rho^{2})(1+\rho^{4})(k_{\zeta_{t}}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}-\sigma_{\zeta_{t}}^{4})+(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}$
+$+\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})\bigg)\times$
+$\bigg(\rho^{4}(k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}-\sigma_{v_{t}}^{4})+(1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
+$-\frac{(1+\rho^{2})^{4}}{\rho^{4}}(1+\rho^{4})\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}$
+$-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}$
+$-2\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg\}.$
+Assuming that the shocks are not drawn from a Bernoulli distribution, kurtosis is bounded below by $\gamma^{2}/\sigma^{2}+\sigma^{4}$. Given that $0<\rho\leq 1$, it follows that
+$(1+\rho^{2})(1+\rho^{4})(k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}-\sigma_{\zeta_{t}}^{4})$ $>0,$ (B.2)
+$\rho^{4}(k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}-\sigma_{v_{t}}^{4})$ $>0.$ (B.3)
 Note also that
-$6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0,$ (B.4)
-$2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0.$ (B.5)
-Adding the inequalities (B.2)+(B.4) and (B.3)+(B.5) by parts, it follows that $k_{\zeta_{t}}-\gamma_{\zeta_{t}}^{2}/\sigma_{\zeta_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$ and $k_{v_{t}}-\gamma_{v_{t}}^{2}/\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}>0$, so the bracketed terms in the first line of the determinant are both strictly positive.
-It then follows that
-$\Big{(}k_{\zeta_{t}}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}k_{v_{t}}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
-$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>$
-$\Big{(}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+\sigma_{\zeta_{t}}^{4}-\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+5\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+\sigma_{v_{t}}^{4}-\frac{\gamma_{v_{t}}^{2}}{\sigma_{v_{t}}^{2}}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+\sigma_{v_{t}}^{4}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
-$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
-$\Big{(}6\sigma_{\zeta_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{\zeta_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}\Big{(}2\sigma_{v_{t}}^{4}+2\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+2\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\Big{)}$
-$-8\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-4\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}-8\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-12\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
-$4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}+4\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}+4\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}$ $>0,$
-where the first inequality follows from the lower bound on the kurtosis.
-The term inside the squared brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$.
-The matrix of coefficients is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli.
-[^2] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant would not always be strictly away from zero.
+$(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}$ $\geq 0,$ (B.4)
+$\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})$ $>0,$ (B.5)
+$(1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0,$ (B.6)
+as the sum of mostly strictly positive terms.
+Adding the inequalities (B.2)+(B.4)+(B.5) and (B.3)+(B.6) by parts, it follows that the two long bracketed terms in the first lines of the determinant are strictly positive, so their product is also strictly positive.
+Given the lower bound on the kurtosis, the term inside the curly brackets is then strictly greater than
+$\bigg((\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}+\frac{1+\rho^{2}}{\rho^{4}}(1+2\rho^{2}+2\rho^{4}+6\rho^{6}+\rho^{8})\sigma_{\zeta_{t}}^{4}+\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{8})\sigma_{\zeta_{t}}^{2}(\sigma_{v_{t}}^{2}+\sigma_{u^{y}_{t}}^{2})\bigg)$ $\times$
+$\bigg((1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
+$-\frac{(1+\rho^{2})^{4}}{\rho^{4}}(1+\rho^{4})\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{6}$
+$-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}-\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{4}$
+$-2\frac{(1+\rho^{2})^{2}}{\rho^{4}}(1+\rho^{4})(1+\rho^{8})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{4}\sigma_{u^{y}_{t}}^{2}-2\frac{(1+\rho^{2})^{3}}{\rho^{4}}(1+\rho^{4}+\rho^{8})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $=$
+$(\rho^{2}-\rho)^{2}\frac{\gamma_{\zeta_{t}}^{2}}{\sigma_{\zeta_{t}}^{2}}\bigg((1+\rho^{4})\sigma_{v_{t}}^{4}+(1+\rho^{2})\sigma_{\zeta_{t}}^{2}\sigma_{v_{t}}^{2}+(1+\rho^{4})\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}\bigg)$
+$+4\rho^{2}(1+\rho^{2})^{2}\sigma_{\zeta_{t}}^{6}\sigma_{v_{t}}^{2}+4\rho^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{4}+4\rho^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta_{t}}^{4}\sigma_{v_{t}}^{2}\sigma_{u^{y}_{t}}^{2}$ $>0.$
+The last inequality follows from $0<\rho\leq 1$ and the fact that we are adding mostly strictly positive terms (and one weakly positive, the term involving $\gamma_{\zeta_{t}}^{2}$).
+The expression inside the curly brackets in the determinant is thus strictly positive, so $\text{det}(\mathbf{A})<0$.
+$\mathbf{A}$ is thus full rank and invertible for any distribution of permanent & transitory shocks except Bernoulli.
+[^3] If the distributions of both permanent and transitory shocks were Bernoulli, kurtosis would equal its lower bound $\gamma^{2}/\sigma^{2}+\sigma^{4}$ and the determinant may not always be strictly away from zero.
 How do different levels of skewness and kurtosis influence the transmission parameters?
-Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$.
-$\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})$.
-Absent any transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks.
+Consider the transmission of permanent shocks, abstracting from transitory shocks, i.e., $\psi^{(1)}_{t}=\psi^{(2)}_{t}=\omega^{(22)}_{t}=0$, and let $\rho=1$.
+$\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ are pinned down by the first equation in (B.1), namely $2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}+2\gamma_{\zeta_{t}}\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})$.
+Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and permanent shocks.
 Suppose $\phi^{(1)}_{t}$ is fixed.
 We highlight three cases:
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}=0$.
 Letting $\gamma_{\zeta_{t}}<0$ as in the data, this equation restricts $\phi^{(2)}_{t}=0$.
 Tail shocks thus transmit similarly to non-tail ones, i.e., this configuration identifies a linear consumption process.
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}>0$.
 Consumption growth and permanent shocks co-vary more strongly than the linear consumption process suggests, i.e., there is excess co-movement.
@@
 If $\gamma_{\zeta_{t}}>0$, this configuration identifies $\phi^{(2)}_{t}>0$ and analogous arguments apply to positive shocks.
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it})-2\sigma_{\zeta_{t}}^{2}\phi^{(1)}_{t}<0$.
 Consumption growth and permanent shocks co-move less strongly than the linear process suggests, i.e., there is excess smoothness.
@@
 The opposite is true if $\gamma_{\zeta_{t}}>0$.
-Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1).
+Another equation that pins down $\phi^{(1)}_{t}$ and $\phi^{(2)}_{t}$ is $2\gamma_{\zeta_{t}}\phi^{(1)}_{t}+2(\kappa_{\zeta_{t}}+\sigma_{\zeta_{t}}^{2}(3\sigma_{\zeta_{t}}^{2}+2\sigma_{v_{t}}^{2}))\phi^{(2)}_{t}={\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})$, which, abstracting from income measurement error, corresponds to the second line in (B.1).
 Absent transmission of transitory shocks, the right hand side reflects the strength of the covariance between consumption growth and squared shocks so it captures the pass-through of shocks of large magnitude.
 We highlight again three cases (for $\phi^{(1)}_{t}$ fixed):
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}=0$.
 This identifies $\phi^{(2)}_{t}=0$, so tail shocks transmit similarly to non-tail ones.
-In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is directly implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\zeta_{it-1})$, so this configuration identifies the linear consumption process.
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})$ is negative and large.
+In fact, ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})=2\gamma_{\zeta_{t}}\phi^{(1)}_{t}$ is implied by $\Delta^{2}c_{it}=\phi_{t}^{(1)}(\zeta_{it}+\rho\zeta_{it-1})$, so this configuration identifies the linear consumption process.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}<0$, which, for $\gamma_{\zeta_{t}}<0$, is true when ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})$ is negative and large.
 This suggests that large shocks (of which negative shocks are the most likely) transmit strongly with a negative coefficient.
@@
 Ceteris paribus, a less leptokurtic income distribution (lower $\kappa_{\zeta_{t}}$) turns $\phi^{(2)}_{t}$ more negative; a bell-shaped distribution is associated with fewer large shocks, thus necessitating a stronger transmission of said shocks to maintain a given excess co-movement.
-  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$.
+  • ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2})-2\gamma_{\zeta_{t}}\phi^{(1)}_{t}>0$.
 This identifies $\phi^{(2)}_{t}>0$ for analogous reasons.
@@
 The response of $\phi^{(1)}_{t}$ is much smaller, confirming that the values of skewness and kurtosis are crucial for identifying the pass-through of tail shocks.
-An interesting special case arises when the distributions of permanent and transitory shocks are normal, have the same variance between them, and there is no measurement error.
-In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times(\Delta y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta c_{it}\times\Delta y_{it}\times\Delta y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution.
+An interesting special case arises when $\rho=1$, the distributions of permanent and transitory shocks are normal and have the same variance, and there is no measurement error.
+In this case, $\text{det}(\mathbf{A})=-288\sigma_{t}^{16}$, $\phi_{t}^{(1)}=(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\right)+(1/(2\sigma_{t}^{2})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it+2}\right)$ and $\phi_{t}^{(2)}=(1/(12\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times(\Delta^{2}_{\rho}y_{it})^{2}\right)+(1/(6\sigma_{t}^{4})){\operatorname*{\mathbb{E}}}\left(\Delta^{2}c_{it}\times\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\right)$, where $\sigma_{t}$ is the dispersion of the hypothetical distribution.
 Even in such a special environment, $\phi_{t}^{(2)}$, the transmission parameter of tail shocks, is separately identified from $\phi_{t}^{(1)}$.
-Identification does not require the shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks.
+Identification does not require shocks to be non-Gaussian, but skewness and kurtosis (both defined also in the Gaussian case) are still crucial for identifying the pass-through of tail shocks.
 Despite normality, shocks in the (normal) tails are associated with a different pass-through from that of average shocks.
@@
 B.4 Targeted moments
-There is a large number of over-identifying moments that we target in the estimation of the various specifications of the model.
-Table B.1 lists the income moments targeted in each case, while table B.2 lists the moments targeted in the different specifications of the consumption model.
-For simplicity, we illustrate the next statements assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters.
+The estimation targets several over-identifying moments.
+Table B.1 lists the target income moments, while table B.2 lists the moments targeted in the various specifications of the consumption model.
+For simplicity, we illustrate these moments assuming stationarity in the distributions of income shocks, taste heterogeneity, and measurement error, and time-invariance of the partial insurance parameters.
 These are not identifying assumptions but they make the illustration more compact by allowing us to pool common terms together.
@@
 Moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
-1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$
-2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$
-3 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$
-4 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$
-5 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$
-6 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$
-7 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$
-8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$
-9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$
+1 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$
+2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$
+3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$
+4 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$
+5 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$
+6 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$
+7 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})$
+8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$
+9 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}_{\rho}y_{it-2})$
+10 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{3})$
 Notes: The table lists the moments targeted in the estimation of the various specifications of the income process.
-$\Delta^{2}y_{t}=\Delta y_{t}+\Delta y_{t-1}=y_{t}-y_{t-2}$ indicates a first difference of $y$ between periods $t$ and $t-2$.
+$\Delta^{2}y_{t}=y_{t}-y_{t-2}$ indicates a first difference between periods $t$ and t2t-2; $\Delta^{2}_{\rho}y_{t}=y_{t}-\rho^{2}y_{t-2}$ indicates the analogous quasi-difference.
+Note that $\Delta^{2}y_{t}=\Delta^{2}_{\rho}y_{t}$ if $\rho=1$ (baseline permanent-transitory process).
 B.4.1 Moments of income
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2})$ $=2\sigma_{\zeta}^{2}+2\sigma_{v}^{2}+2\sigma_{{u}^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it-2})$ $=-\sigma_{v}^{2}-\sigma_{{u}^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3})$ $=2\gamma_{\zeta}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}y_{it-2})$ $=\gamma_{v}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{2})$ $=-\gamma_{v}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{4})$ $=2\kappa_{\zeta}+2\kappa_{v}+6(\sigma_{\zeta}^{2})^{2}+6(\sigma_{v}^{2})^{2}+12(\sigma_{u^{y}}^{2})^{2}+24\sigma_{\zeta}^{2}\sigma_{v}^{2}+24\sigma_{v}^{2}\sigma_{u^{y}}^{2}+24\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}y_{it-2})^{2})$ $=\kappa_{v}+4(\sigma_{\zeta}^{2})^{2}+3(\sigma_{v}^{2})^{2}+6(\sigma_{u^{y}}^{2})^{2}+8\sigma_{\zeta}^{2}\sigma_{v}^{2}+8\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}y_{it-2})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}y_{it-2})^{3})$ $=-\kappa_{v}-3(\sigma_{v}^{2})^{2}-6(\sigma_{u^{y}}^{2})^{2}-6\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-12\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2})$ $=(1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})\sigma_{v}^{2}+(1+\rho^{4})\sigma_{{u}^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it-2})$ $=-\rho^{2}\sigma_{v}^{2}-\rho^{2}\sigma_{{u}^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}y_{it-4})$ $=0$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3})$ $=(1+\rho^{3})\gamma_{\zeta}+(1-\rho^{6})\gamma_{v}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}_{\rho}y_{it-2})$ $=\rho^{4}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ $=-\rho^{2}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{4})$ $=(1+\rho^{4})\kappa_{\zeta}+(1+\rho^{8})\kappa_{v}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}+6\rho^{4}(\sigma_{v}^{2})^{2}+3(1+\rho^{4})^{2}(\sigma_{u^{y}}^{2})^{2}$
+$+6(1+\rho^{4})\Big((1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}_{\rho}y_{it-2})^{2})$ $=\rho^{4}\kappa_{v}+(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(1+\rho^{4}+\rho^{8})(\sigma_{v}^{2})^{2}+(1+4\rho^{4}+\rho^{8})(\sigma_{u^{y}}^{2})^{2}$
+$+2(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+(2(1+\rho^{4})^{2}+4\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}_{\rho}y_{it-2})$ $=-\rho^{6}\kappa_{v}-3\rho^{2}(\sigma_{v}^{2})^{2}-3\rho^{2}(1+\rho^{4})(\sigma_{u^{y}}^{2})^{2}$
+$-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-6\rho^{2}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}_{\rho}y_{it-2})^{3})$ $=-\rho^{2}\kappa_{v}-3\rho^{6}(\sigma_{v}^{2})^{2}-3\rho^{2}(1+\rho^{4})(\sigma_{u^{y}}^{2})^{2}$
+$-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}-6\rho^{2}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
 Table B.2: Targeted joint moments of consumption and income
@@
 2 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$
-3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$
-4 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$
+3 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$
+4 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$
 5 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$
@@
 7 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$
-8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$
-9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$
-10 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$
-11 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$
-12 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$
-13 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$
-14 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$
-15 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$
+8 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$
+9 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$
+10 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$
+11 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta c_{it-2})^{2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{2})$
+12 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$
+13 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$
+14 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$
+15 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$
 16 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$
@@
 19 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$
-20 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$
-21 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$
-22 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$
-23 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
-24 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$
-25 ${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$
-26 ${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$
+20 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$
+21 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it})$
+22 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{3})$
+23 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$
+24 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$
+25 ${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it-2})$
+26 ${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{3})$
 Notes: The table lists the moments targeted in the estimation of the different specifications of the consumption process.
-In all cases, the notation $\Delta^{2}x_{t}=\Delta x_{t}+\Delta x_{t-1}=x_{t}-x_{t-2}$ indicates the first difference of the generic variable $x$ between periods $t$ and $t-2$.
+The notation $\Delta^{2}x_{t}=x_{t}-x_{t-2}$ indicates a first difference of the generic variable $x$ between periods $t$ and t2t-2; $\Delta^{2}_{\rho}x_{t}=x_{t}-\rho^{2}x_{t-2}$ indicates the analogous quasi-difference.
 B.4.2 Moments of consumption – linear function
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=-\sigma_{{u}^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ $=2\gamma_{\xi}+2(\phi^{(1)})^{3}\gamma_{\zeta}+(\psi^{(1)})^{3}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{2}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3})$ $=2\gamma_{\xi}+(\phi^{(1)})^{3}(1+\rho^{3})\gamma_{\zeta}+(\psi^{(1)})^{3}\gamma_{v}$
 ${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\gamma_{u^{c}}$
 ${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-\gamma_{u^{c}}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{2})$ $=2(\phi^{(1)})^{2}\gamma_{\zeta}+(\psi^{(1)})^{2}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-(\psi^{(1)})^{2}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=0$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=0$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=0$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$ $=2\kappa_{\xi}+2(\phi^{(1)})^{4}\kappa_{\zeta}+(\psi^{(1)})^{4}\kappa_{v}+2\kappa_{u^{c}}+6(\sigma_{\xi}^{2})^{2}+6(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+6(\sigma_{u^{c}}^{2})^{2}$
-$+24(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+12(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+24\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}$
-$+12(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+24(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+8(\phi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+4(\phi^{(1)})^{4}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}$
-$+4(\phi^{(1)})^{2}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+8(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-6(\phi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\kappa_{\zeta}+6(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}+8\phi^{(1)}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\kappa_{\zeta}+6\phi^{(1)}(\sigma_{\zeta}^{2})^{2}+\psi^{(1)}\kappa_{v}+3\psi^{(1)}(\sigma_{v}^{2})^{2}$
-$+6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it})^{3})$ $=2(\phi^{(1)})^{3}\kappa_{\zeta}+6(\phi^{(1)})^{3}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{3}\kappa_{v}+12\phi^{(1)}\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+6\psi^{(1)}\sigma_{\xi}^{2}\sigma_{v}^{2}$
-$+6\phi^{(1)}(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}+12\phi^{(1)}\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$ $=4\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}(\sigma_{\zeta}^{2})^{2}+2(\psi^{(1)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+4\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{v}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+4\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
-$+4\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+4(\phi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+2(\psi^{(1)})^{2}\sigma_{u^{y}}^{2}\sigma_{v}^{2}+4\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\kappa_{v}-3\psi^{(1)}(\sigma_{v}^{2})^{2}-6\psi^{(1)}\sigma_{\zeta}^{2}\sigma_{v}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-(\psi^{(1)})^{3}\kappa_{v}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{\xi}^{2}-6(\phi^{(1)})^{2}\psi^{(1)}\sigma_{v}^{2}\sigma_{\zeta}^{2}-6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{2})$ $=(\phi^{(1)})^{2}(1+\rho^{3})\gamma_{\zeta}+(\psi^{(1)})^{2}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\rho^{4}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{2})$ $=-(\psi^{(1)})^{2}\rho^{2}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=0$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\rho^{2}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ $=0$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ $=0$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{4})$ $=2\kappa_{\xi}+(\phi^{(1)})^{4}(1+\rho^{4})\kappa_{\zeta}+(\psi^{(1)})^{4}\kappa_{v}+2\kappa_{u^{c}}+6(\sigma_{\xi}^{2})^{2}$
+$+6(\phi^{(1)})^{4}\rho^{2}(\sigma_{\zeta}^{2})^{2}+6(\sigma_{u^{c}}^{2})^{2}+12(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}$
+$+12(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+24\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}+6(\phi^{(1)})^{2}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+12(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+12(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=4(\sigma_{\xi}^{2})^{2}+4(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+4(\psi^{(1)})^{2}\sigma_{\xi}^{2}\sigma_{v}^{2}+8\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}$
+$+(\phi^{(1)})^{4}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+3(\sigma_{u^{c}}^{2})^{2}+2(\phi^{(1)})^{2}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+4(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+(\psi^{(1)})^{4}(\sigma_{v}^{2})^{2}+4(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}+\kappa_{u^{c}}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-3(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-\kappa_{u^{c}}-3(\sigma_{u^{c}}^{2})^{2}-6\sigma_{\xi}^{2}\sigma_{u^{c}}^{2}-3(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}-3(\psi^{(1)})^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{4})\kappa_{\zeta}+6(\phi^{(1)})^{2}\rho^{2}(\sigma_{\zeta}^{2})^{2}$
+$+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}$
+$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\kappa_{v}+(\psi^{(1)})^{2}\rho^{4}(\sigma_{v}^{2})^{2}$
+$+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}$
+$+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}+4\phi^{(1)}\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{4})\kappa_{\zeta}+6\phi^{(1)}\rho^{2}(\sigma_{\zeta}^{2})^{2}+\psi^{(1)}\kappa_{v}+3\psi^{(1)}\rho^{4}(\sigma_{v}^{2})^{2}$
+$+3\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+3\phi^{(1)}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+3\phi^{(1)}(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2}+3\psi^{(1)}(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it})^{3})$ $=(\phi^{(1)})^{3}(1+\rho^{4})\kappa_{\zeta}+6(\phi^{(1)})^{3}\rho^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{3}\kappa_{v}$
+$+6\phi^{(1)}(1+\rho^{2})\sigma_{\xi}^{2}\sigma_{\zeta}^{2}+6\psi^{(1)}\sigma_{\xi}^{2}\sigma_{v}^{2}+3\phi^{(1)}(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+3(\phi^{(1)})^{2}\psi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+6\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+6\psi^{(1)}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times(\Delta^{2}c_{it-2})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}$
+$+(\psi^{(1)})^{2}\rho^{4}\kappa_{v}+(\psi^{(1)})^{2}(\sigma_{v}^{2})^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}$
+$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times(\Delta^{2}c_{it})^{2})$ $=2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+(\psi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+$+2(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{v}^{2}\sigma_{\zeta}^{2}$
+$+(\psi^{(1)})^{2}(1+\rho^{4})(\sigma_{v}^{2})^{2}+2(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{c}}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\xi}^{2}$
+$+(\phi^{(1)})^{2}(1+\rho^{2})(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{v}^{2}+2(1+\rho^{4})\sigma_{u^{y}}^{2}\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{3}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{6}\kappa_{v}-3\psi^{(1)}\rho^{2}(\sigma_{v}^{2})^{2}-3\psi^{(1)}(1+\rho^{2})\rho^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}-3\psi^{(1)}(1+\rho^{4})\rho^{2}\sigma_{v}^{2}\sigma_{u^{y}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times(\Delta^{2}c_{it-2})^{3})$ $=-(\psi^{(1)})^{3}\rho^{2}\kappa_{v}-6\psi^{(1)}\rho^{2}\sigma_{v}^{2}\sigma_{\xi}^{2}-3(\phi^{(1)})^{2}\psi^{(1)}(1+\rho^{2})\rho^{2}\sigma_{v}^{2}\sigma_{\zeta}^{2}-6\psi^{(1)}\rho^{2}\sigma_{v}^{2}\sigma_{u^{c}}^{2}$
 B.4.3 Moments of consumption – quadratic function
-${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+2(\phi^{(1)})^{2}\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+2(\phi^{(2)})^{2}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2})+(\psi^{(2)})^{2}\kappa_{v}$
-$+2(\omega^{(22)})^{2}\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}+4\phi^{(1)}\phi^{(2)}\gamma_{\zeta}+2\psi^{(1)}\psi^{(2)}\gamma_{v}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=4(\phi^{(2)})^{2}(\sigma_{\zeta}^{2})^{2}+4\phi^{(2)}\psi^{(2)}\sigma_{\zeta}^{2}\sigma^{2}_{v}+(\psi^{(2)})^{2}(\sigma^{2}_{v})^{2}-\sigma_{u^{c}}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}+2\phi^{(2)}\gamma_{\zeta}+\psi^{(2)}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\sigma_{v}^{2}-\psi^{(2)}\gamma_{v}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it})$ $=2\phi^{(1)}\gamma_{\zeta}+\psi^{(1)}\gamma_{v}+2\phi^{(2)}(\kappa_{\zeta}+3(\sigma_{\zeta}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{\zeta}^{2}\sigma_{u^{y}}^{2})$
-$+\psi^{(2)}(\kappa_{v}+(\sigma_{v}^{2})^{2}+2\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{v}^{2}\sigma_{u^{y}}^{2})+4\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\gamma_{v}+4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}(2\sigma_{\zeta}^{2}+\sigma_{v}^{2}+2\sigma_{u^{y}}^{2}))$
-${\operatorname*{\mathbb{E}}}((\Delta^{2}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=4\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})+2\psi^{(2)}\sigma_{v}^{2}(\sigma_{\zeta}^{2}+\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\gamma_{v}-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}(\kappa_{v}+\sigma_{v}^{2}\sigma_{u^{y}}^{2})-2\omega^{(22)}\sigma_{\zeta}^{2}\sigma_{v}^{2}$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it-2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
-${\operatorname*{\mathbb{E}}}(\Delta^{2}y_{it}\times\Delta^{2}y_{it+2}\times\Delta^{2}c_{it+2})$ $=-2\phi^{(2)}\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}c_{it})^{2})$ $=2\sigma_{\xi}^{2}+(\phi^{(1)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}+(\psi^{(1)})^{2}\sigma_{v}^{2}+(\phi^{(2)})^{2}\Big((1+\rho^{4})\kappa_{\zeta}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}\Big)$
+$+(\psi^{(2)})^{2}\kappa_{v}+(\omega^{(22)})^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+2\sigma_{u^{c}}^{2}$
+$+2\phi^{(1)}\phi^{(2)}(1+\rho^{3})\gamma_{\zeta}+2\psi^{(1)}\psi^{(2)}\gamma_{v}+2\phi^{(2)}\psi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}c_{it}\times\Delta^{2}c_{it-2})$ $=(\phi^{(2)})^{2}(1+\rho^{2})^{2}(\sigma_{\zeta}^{2})^{2}+2\phi^{(2)}\psi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma^{2}_{v}+(\psi^{(2)})^{2}(\sigma^{2}_{v})^{2}-\sigma_{u^{c}}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{2})\sigma_{\zeta}^{2}+\psi^{(1)}\sigma_{v}^{2}+\phi^{(2)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(2)}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}c_{it-2})$ $=-\psi^{(1)}\rho^{2}\sigma_{v}^{2}-\psi^{(2)}\rho^{2}\gamma_{v}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(1)}(1+\rho^{3})\gamma_{\zeta}+\psi^{(1)}\gamma_{v}$
+$+\phi^{(2)}\Big((1+\rho^{4})\kappa_{\zeta}+6\rho^{2}(\sigma_{\zeta}^{2})^{2}+(1+\rho^{2})(1+\rho^{4})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
+$+\psi^{(2)}\Big(\kappa_{v}+\rho^{4}(\sigma_{v}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)+2\omega^{(22)}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it})^{2}\times\Delta^{2}c_{it-2})$ $=\psi^{(1)}\rho^{4}\gamma_{v}+\phi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
+$+\psi^{(2)}\Big(\rho^{4}\kappa_{v}+(\sigma_{v}^{2})^{2}+(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}+(1+\rho^{4})\sigma_{v}^{2}\sigma_{u^{y}}^{2}\Big)$
+${\operatorname*{\mathbb{E}}}((\Delta^{2}_{\rho}y_{it-2})^{2}\times\Delta^{2}c_{it})$ $=\phi^{(2)}(1+\rho^{2})\sigma_{\zeta}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
+$+\psi^{(2)}\sigma_{v}^{2}\Big((1+\rho^{2})\sigma_{\zeta}^{2}+(1+\rho^{4})(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})\Big)$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it})$ $=-\psi^{(1)}\rho^{2}\gamma_{v}-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
+$-\psi^{(2)}\rho^{2}(\kappa_{v}+\sigma_{v}^{2}\sigma_{u^{y}}^{2})-\omega^{(22)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}\sigma_{v}^{2}$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it-2})$ $=-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\rho^{2}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
+${\operatorname*{\mathbb{E}}}(\Delta^{2}_{\rho}y_{it}\times\Delta^{2}_{\rho}y_{it+2}\times\Delta^{2}c_{it+2})$ $=-\phi^{(2)}\rho^{2}(1+\rho^{2})\sigma_{\zeta}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})-\psi^{(2)}\rho^{2}\sigma_{v}^{2}(\sigma_{v}^{2}+\sigma_{u^{y}}^{2})$
 Appendix C Empirical details and additional results
@@
 Observations 20,866 4,258 5,159 2,428 9,021
-Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 19991999199919992019201920192019.
+Notes: The table reports summary statistics for consumption, income, wealth, and demographics in our baseline sample in the PSID and in subsamples sorted by wealth and education, over 1999199920192019.
 Monetary amounts are expressed in \$2018.
@@
 ${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.271 (0.010) 0.323
-${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.065 (0.004)
-${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.569 (0.128) -1.039
+${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.065 (0.004)
+${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.569 (0.128) -1.039
 ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.087 (0.010)
@@
 ${\operatorname*{\mathrm{Var}}}(\Delta y_{it})$ 0.233 (0.009)
-${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$0.056 (0.004)
-${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$0.728 (0.142)
+${\operatorname*{\mathrm{Cov}}}(\Delta y_{it},\Delta y_{it+1})$ -0.056 (0.004)
+${\operatorname*{\mathrm{Skew}}}(\Delta y_{it})$ -0.728 (0.142)
 ${\operatorname*{\mathrm{Cov}}}((\Delta y_{it})^{2},\Delta y_{it+1})$ 0.045 (0.008)
@@
 Bottom 10% Top 90%
-mean mean $p$-value $\Delta$mean
+[…]
 (1) (2) (3)
 % unemployed $t$ (male) 0.082 0.034 0.000
-% unemployed $t-1$ (male) 0.084 0.036 0.000
+% unemployed t1t-1 (male) 0.084 0.036 0.000
 % unemployed $t$ (female) 0.081 0.037 0.000
-% unemployed $t-1$ (female) 0.081 0.041 0.000
+% unemployed t1t-1 (female) 0.081 0.041 0.000
 % main job laid off (male) 0.054 0.039 0.003
@@
 This is something we have not allowed for in the paper, and neither have BPP.
-While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors).
+While $\phi_{t}^{(1)}$ is identified via ${\operatorname*{\mathbb{E}}}(\Delta c_{it}(\Delta y_{it-1}+\Delta y_{it}+\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$ in the baseline permanent-transitory income process, in practice BPP and virtually all follow-up papers estimate it through $({\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it})+{\operatorname*{\mathbb{E}}}(\Delta c_{it}\Delta y_{it+1}))/\sigma_{\zeta_{t}}^{2}$, assuming that $\sigma_{u_{t}}=0$ (uncorrelated errors).
 This is what we also do in our baseline estimation of $\phi_{t}^{(1)}$ in the linear specification.
@@
 One can show that, with imputed consumption, the moment used to identify $\phi_{t}^{(1)}$ in fact identifies $\phi_{t}^{(1)\text{imp*}}=\phi_{t}^{(1)}+\kappa_{t}\sigma_{u_{t}}/\sigma_{\zeta_{t}}^{2}$.
-As a result, $*\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
+As a result, $\phi_{t}^{(1)\text{imp*}}>\phi_{t}^{(1)\text{psid*}}$ and the imputation inflates the transmission parameter of permanent shocks vis-à-vis the true consumption series.
 C.5 Estimation using the original BPP data
@@
 The small unimportant numerical differences between column 1 herein and the original BPP estimates are likely due to different software and optimization algorithms (our code uses Matlab and fmincon while BPP use Gauss).
-#$\theta$ is the moving average parameter of the transitory shock.
+$\theta$ is the moving average parameter of the transitory shock.
 BPP assume that the transitory component of income is given by $v_{it}=\epsilon_{it}+\theta\epsilon_{it-1}$, where $\epsilon_{it}$ is the transitory shock.
@@
 C.6 Additional results
-Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix in the GMM estimation (optimal GMM).
+Tables C.6 and C.7 present the estimates of the parameters of the income process and the consumption function using the empirical covariance matrix of the target moments as weighting matrix (optimal GMM).
 As such, these tables accompany tables 2 and 3 in the paper, which report results from equally weighted GMM.
-The results are qualitatively (and most often numerically) similar to those reported in the paper.
-Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption data from the CEX.
-As such, this table accompanies table 3 in the main text, which reports results using consumption internally available in the PSID.
-In the linear specification, the transmission parameter of permanent shocks $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments.
+Table C.8 presents the estimates of the parameters of the consumption function using imputed consumption from the CEX.
+As such, this table accompanies table 3 in the paper, which reports results using consumption internally available in the PSID.
+In the linear specification, $\phi^{(1)}$ is consistently higher than that obtained from the consumption data internally available in the PSID, regardless of whether we target only second- or also higher-order moments.
 This echoes our earlier discussion about the imputation inflating the pass-through of shocks.
-We also find that the transmission parameter of transitory shocks $\psi^{(1)}$ is negative in all cases.
-Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude and negative ($-0.418$).
+We also find that $\psi^{(1)}$ is negative in all cases.
+Moreover, when up to fourth-order moments are targeted (column 3), the point estimate is particularly large in magnitude (0.418-0.418).
 This is counterintuitive, raising further concerns about the reliability of the imputation procedure, particularly when higher-order moments are involved.
 We draw similar conclusions about $\phi^{(1)}$ and $\psi^{(1)}$ in the context of the quadratic function.
-$\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID ($-0.036$ versus $-0.04$; albeit not significant).
+$\phi^{(2)}<0$ in that case, with a magnitude similar to our baseline results from the PSID (0.036-0.036 versus 0.04-0.04; albeit not significant).
 Once again, this is indicative of the asymmetric pass-through of negative versus positive permanent shocks.
-Similar to our baseline, we also find $\psi^{(2)}>0$.
-Table C.9 presents the estimates of the parameters of the income process by age and the household’s position in the income distribution.
-Skewness of all shocks becomes more negative, both economically and statistically, with age.
+Similar to our baseline, we find $\psi^{(2)}>0$.
+Table C.9 presents the estimates of the income process by age and the household’s position in the income distribution.
+Skewness of all shocks is more negative at older ages.
 Skewness of permanent shocks is weakly positive in the bottom income tertile but turns very negative in the top tertile.
@@
 Tables C.10 and C.11 present results from the linear consumption function by age and the household’s position in the income distribution, respectively.
-Parameter $\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65.
-A similar non-monotonic pattern appears in column 2, where second and third moments are targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted.
-This is again indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results.
+$\phi^{(1)}$ exhibits a non-monotonic pattern over the three age brackets – decreasing from $0.154$ over ages 30–40 to $0.126$ over 41–50 before increasing to $0.180$ over ages 51–65.
+A similar non-monotonic pattern appears in column 2, where third moments are also targeted, and is further accentuated in column 3 where all moments up to fourth-order are targeted.
+This is indicative of a misspecification in the linear function, in the sense that its theoretical restrictions (higher wealth insures income fluctuations better) are not supported by the empirical results.
 [^8] This conclusion is consistent with BPP, who find a decreasing pattern with age only when they impose a linear age trend on the transmission parameters.
@@
 By contrast, $\psi^{(1)}$ is statistically indistinguishable from zero in all three age brackets.
-Parameter $\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
+$\phi^{(1)}$ is uniformly lower in the top income tertile vis-à-vis the bottom tertile, suggesting that the income-rich, who are also wealthier, have more insurance against permanent income shocks than the poor.
 Table C.6: Estimates of the income process, optimal weighting
@@
 (0.028) (0.044) (0.025) (0.029) (0.030)
-$\psi^{(1)}$0.006 -0.109 -0.015 -0.004 -0.018
+$\psi^{(1)}$ -0.006 -0.109 -0.015 -0.004 -0.018
 (0.045) (0.078) (0.037) (0.041) (0.047)
-$\phi^{(2)}$0.037
-$\psi^{(2)}$0.007
+$\phi^{(2)}$ -0.037
+$\psi^{(2)}$ -0.007
 $\omega^{(22)}$ 0.740
@@
 (0.001) (0.004) (0.001) (0.001) (0.001)
-$\gamma_{\xi}$0.001 -0.001
+$\gamma_{\xi}$ -0.001 -0.001
 (0.001) (0.001)
@@
 (0.044) (0.074) (0.118) (0.043)
-$\psi^{(1)}$0.109 -0.155 -0.418 -0.108
+$\psi^{(1)}$ -0.109 -0.155 -0.418 -0.108
 (0.078) (0.098) (0.247) (0.086)
-$\phi^{(2)}$0.036
+$\phi^{(2)}$ -0.036
 $\psi^{(2)}$ 0.025
-$\omega^{(22)}$0.157
+$\omega^{(22)}$ -0.157
 $\sigma^{2}_{\xi}$ 0.019 0.018 0.000 0.019
 (0.004) (0.004) (0.004) (0.004)
-$\gamma_{\xi}$0.016 -0.016
+$\gamma_{\xi}$ -0.016 -0.016
 (0.007) (0.007)
@@
 (0.009) (0.009) (0.012) (0.009)
-$\gamma_{u_{c}}$0.185 -0.185
+$\gamma_{u_{c}}$ -0.185 -0.185
 (0.037) (0.037)
@@
 Table C.12 presents estimation results for the linear consumption function in the subsamples formed by wealth and education.
-As such, this table accompanies table 6 in the main text, which reported wealth- and education-specific results in the quadratic specification.
 The odd-numbered columns present parameter estimates when only second-order moments are targeted.
 $\phi^{(1)}$ becomes monotonically smaller as we move from group 1 (low wealth, no college) to group 4 (high wealth, some college), reflecting the insurance role of wealth, which increases monotonically across groups.
-This is similar to what we found in the quadratic specification.
-However, as soon as we target up to fourth-order moments in the linear specification, shown in the even-numbered columns, this monotonic pattern is lost.
+This is similar to what we find in the quadratic specification.
+However, as soon as we target up to fourth-order moments in the even-numbered columns, this monotonic pattern is lost.
 Transitory shocks transmit strongly in group 1, and with a negative coefficient.
+Table C.13 presents the estimates of the income process, when permanent income follows an AR(1) process, i.e., it features linear persistence.
+Table C.14 presents the associated estimates of the parameters of the linear and quadratic consumption functions.
 Table C.9: Estimates of the income process, by age and position in income distribution
@@
 (0.031) (0.032) (0.035) (0.036) (0.034) (0.041)
-$\psi^{(1)}$0.045 0.013 0.025 -0.004 0.101 0.071
+$\psi^{(1)}$ -0.045 0.013 0.025 -0.004 0.101 0.071
 (0.086) (0.038) (0.054) (0.093) (0.048) (0.073)
@@
 (0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
-$\gamma_{\xi}$0.001 -0.001 0.002 0.002
+$\gamma_{\xi}$ -0.001 -0.001 0.002 0.002
 (0.001) (0.001) (0.001) (0.001)
@@
 (0.053) (0.042) (0.053) (0.056) (0.071) (0.089) (0.041) (0.055)
-$\psi^{(1)}$0.178 -0.164 0.015 0.073 0.056 0.142 0.003 0.004
+$\psi^{(1)}$ -0.178 -0.164 0.015 0.073 0.056 0.142 0.003 0.004
 (0.107) (0.077) (0.089) (0.091) (0.089) (0.131) (0.059) (0.071)
@@
 (0.003) (0.003) (0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
-$\gamma_{\xi}$0.003 -0.001 -0.003 0.000
+$\gamma_{\xi}$ -0.003 -0.001 -0.003 0.000
 (0.002) (0.001) (0.002) (0.001)
@@
 (0.003) (0.003) (0.003) (0.003) (0.004) (0.004) (0.002) (0.002)
-$\gamma_{u_{c}}$0.001 0.004 0.010 0.002
+$\gamma_{u_{c}}$ -0.001 0.004 0.010 0.002
 (0.003) (0.002) (0.005) (0.001)
@@
 We do not report results when up to third-order moments of income and consumption are targeted, but these are available upon request.
-Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 19991999199919992019201920192019.
+Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth in the sample over 1999199920192019.
 No college versus some college is defined on the basis of the highest level of education attained by the male spouse (cutoff is 12 years of schooling).
@@
 This table accompanies the quadratic specification estimates in table 6 in the paper.
+Table C.13: Estimates of the income process, income persistence
+Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th
+Income data: PSID PSID PSID
+(1) (2) (3)
+Panel A. Permanent shocks
+Variance $\sigma_{\zeta}^{2}$ 0.058 (0.016) 0.056 (0.016) 0.036 (0.006)
+Skewness $\gamma_{\zeta}$ central -0.007 (0.003) -0.006 (0.003)
+standardized -0.503 (0.337) -0.910 (0.512)
+Kurtosis $\kappa_{\zeta}$ central 0.039 (0.007)
+standardized 29.891 (11.459)
+Panel B. Transitory shocks
+Variance $\sigma_{v}^{2}$ 0.011 (0.015) 0.014 (0.014) 0.030 (0.005)
+Skewness $\gamma_{v}$ central -0.005 (0.003) -0.006 (0.003)
+standardized -3.272 (5.415) -1.178 (0.694)
+Kurtosis $\kappa_{v}$ central 0.047 (0.006)
+standardized 52.091 (19.426)
+Panel C. Linear persistence
+AR(1) $\rho$ 0.846 (0.038) 0.854 (0.038) 0.931 (0.028)
+Notes: The table presents the estimates of the parameters of the AR(1) income process.
+Column 1 targets only the second moments of income growth; column 2 targets also third moments; column 3 targets all moments up to fourth-order.
+Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
+This table accompanies table 7 in the paper, as well as table C.14 that follows.
+Table C.14: Estimates of the consumption function, income persistence
+Consumption fn.: Linear Quadratic
+Income moments: 2nd 2nd, 3rd 2nd, 3rd, 4th 2nd, 3rd, 4th
+Consumption data: PSID PSID PSID PSID
+(1) (2) (3) (4)
+$\phi^{(1)}$ 0.089 0.090 0.110 0.112
+(0.029) (0.029) (0.041) (0.024)
+$\psi^{(1)}$ -0.139 -0.077 -0.013 -0.010
+(0.265) (0.083) (0.077) (0.066)
+$\phi^{(2)}$ -0.049
+$\psi^{(2)}$ 0.019
+$\omega^{(22)}$ 0.430
+$\sigma^{2}_{\xi}$ 0.019 0.019 0.020 0.019
+(0.001) (0.001) (0.001) (0.001)
+$\gamma_{\xi}$ -0.001 -0.001
+(0.001) (0.001)
+$\kappa_{\xi}$ 0.031
+$\sigma^{2}_{u_{c}}$ 0.044 0.044 0.044 0.044
+(0.002) (0.002) (0.001) (0.002)
+$\gamma_{u_{c}}$ 0.002 0.002
+(0.001) (0.001)
+$\kappa_{u_{c}}$ 0.013
+Notes: The table presents the estimates of the parameters of the consumption function, when income features linear persistence.
+Columns 1-3 present parameter estimates in the linear function, while column 4 presents estimates in the quadratic case; the order of moments targeted in each case is shown at the top of the table.
+Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses.
+This table accompanies table 7 in the paper.
 Appendix D Some welfare costs of tail income risk
@@
 Specifically, we draw a sequence of income shocks over the lifetime, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare based on those consumption responses.
-Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear and nonlinear transmission.
+Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear versus nonlinear transmission.
 Normal mixtures offer a flexible way of modeling non-Gaussian distributions and have by now become a standard choice in quantitative income dynamics (Guvenen et al., 2021, 2024).
 We draw permanent and transitory income shocks for 50,000 households from a mixture of two normal distributions, namely
-$\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob.
-}p_{\zeta}\\ {\cal N}(\mu_{\zeta,2},\sigma_{\zeta,2})\text{ with prob.
-}1-p_{\zeta}\\ \end{cases}\text{ and }~{}v_{it}\sim\begin{cases}{\cal N}(\mu_{v,1},\sigma_{v,1})\text{ with prob.
-}p_{v}\\ {\cal N}(\mu_{v,2},\sigma_{v,2})\text{ with prob.
-}1-p_{v}.\\ \end{cases}$
+[…]
 There are five parameters per shock $(p,\mu_{1},\sigma_{1},\mu_{2},\sigma_{2})$, which are estimated using the method of simulated moments to match the first four moments of shocks in the PSID (column 3, table 2).
@@
 In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data.
-We simulate consumption over a working life of 35 years, starting at age 30.888We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
+We simulate consumption over a working life of 35 years, starting at age 30.999We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$.
 We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample.
@@
 The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+#% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
 First, the unconditional cost is small but negative in both the first (linear) and second (nonlinear) scenarios, suggesting households may in fact prefer tail over Gaussian risk.
@@
 The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+#% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
 Table D.4: Welfare costs of non-Gaussian relative to Gaussian income risk, by wealth
@@
 The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+#% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
 Table D.5: Welfare costs of non-Gaussian relative to Gaussian income risk, varying $\gamma$
@@
 Panel B. $\boldsymbol{\gamma=3}$
-All households 100% -0.81 -0.88 0.66
-$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
-$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
-$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
+All households 100% -0.76 -0.81 0.66
+$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.68 6.32 7.69
+$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.22 8.59 9.93
+$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.40 11.89 13.17
 Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B).
@@
 The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience.
-#of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
+#% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.